A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1975–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.14

Title word cross-reference

#11812 [1953]. #14473 [1557]. #17627 [1946]. #8879 [1787].

(2 + 1) [1265, 1286]. (2 × 2) [1461]. (λϕ^{4} + σϕ^{2})_{2} [421]. (λϕ^{4} − μϕ)_{3} [181]. (λϕ^{4} − σϕ^{2} − μϕ)_{3} [808]. (Sp(2N); 2N) [1368, 1609]. ∗X ⊗ ∗X [52]. −Δ [936]. 1 [1198, 956, 811, 184, 209, 939, 1140]. 1/2 [1279, 1140]. 1/r^{2} [130, 131, 132, 102]. 2 [510]. 2 + 1 [498]. 2 + ϵ [632, 940]. {2^{N/2−S}, 1^{2S}} [1396]. 3 [1827, 1785, 1950, 1317, 1731, 660, 1197, 316, 317, 1464, 595]. 3/2 [1140]. 3j [1615]. 3V [1085]. 4 [1534]. 405 × 56 → 56, 70 [476]. 5/2 [1259]. 6 [750, 1894, 316, 317, 1443, 1390]. 9 [1780, 480]. ^{∗} [1409, 1869]. ^{−} [1098]. ^{2} [1521]. ^{3} [1521]. ^{6} [481]. ^{p} [721]. _{2} [32, 886]. _{4} [883]. _{7} [1207]. _{8} [628]. _{p} [252]. xy [541]. α [647]. B [33]. βϵ [326]. C [496, 490, 1447, 1888, 1874]. C∗ [845, 1005, 353]. C^{∗} [1906]. c^{−1} [1249]. c^{−4} [1431]. C^{∞} [1430]. H [1832]. D [619, 1833, 1690, 221, 379, 1602]. D(1, 0) [976]. δ [614]. δ = 2 [582, 412]. E [132]. ϵ [976]. F_{2} [669]. f_{tt} − f_{ss} = γ^{2}f [255, 6]. G [360]. G_{2} ⊇ A_{2} [712]. g_{ij} [21]. gl(n, C) [1667]. H [1641, 514, 867, 1201, 1202].

H = (−1/2)d^{2}/dx^{2} + x^{2}/2 + λ/x^{2} [232]. h → 0 [691]. I [132]. IO(n) [140, 1596, 1644]. IU(n) [140, 1596, 1644]. iU_{t} + Δ_{2}U = 0 [96]. J [1259, 83, 1780, 66, 660, 750, 1197, 1894, 316, 317, 480, 1443, 1464, 595, 1390]. J_{1} [1322]. J_{2} [1322]. J_{3} [1322]. K [1129, 1405, 1676, 1295, 1007, 1069]. K(z − z_{0}) [1729]. L [204]. l = 1 [892]. L^{2} [1294]. L_{2} [496]. L_{p} [494, 188]. λ [923, 1833]. λ: Φ_{4}^{4}: [45]. λϕ_{3}^{4} [351, 579]. λx^{2}/(1 + gx^{2}) [1455]. m ≥ 0 [820]. N [150, 890, 917, 1706, 336, 1100, 1386, 1896, 1419, 727, 538, 616, 1704, 1904, 1230, 1341, 599, 802, 803, 1696, 1697, 1768, 1060, 133, 315, 485, 1025, 1329, 1750, 1245, 1710, 1089, 66, 796, 1176, 1223, 528, 1295, 1048, 1047, 1453, 1135]. N/D [306]. n^{2} [1660]. N → ∞ [336]. O(2, 1) [1628]. O(2, 1) [434, 863, 435]. O(3) [1466]. O(3, 1) [434]. O(4, 1) [214]. O(4, 2) [192]. O(5) ⊆ O(3) [1411]. O(N) [1721, 1319, 1926]. O(n, 1) [140]. O_{n}^{5} [52]. O_{h}^{7} [52]. OSp(1/4) [1281]. P [279, 1301, 234]. p = ω [13]. P_{μ} = 0 [524]. ∂^{2}z/∂x^{1}∂x^{2} = f(z) [1504]. ∂_{tt}ψ = Δ_{2}ψ [459]. ϕ^{3} [1370]. P n3n [1776, 1649]. pp [1457]. Ψ^{2N} [1543]. Ψ_{tt} − Δ_{2}Ψ = 0 [399, 458]. ψ_{tt} − Δ_{3}ψ = 0 [1012, 1332]. R [458, 459, 1012, 1332, 1429, 1428]. r^{−2} [1480]. r^{a} [431]. R^{l_{ijk}} [21]. S [235, 234, 1579, 1676, 767, 307, 1287, 652, 466, 1393, 1313]. S(Q, ω) [561]. S^{2} [14]. S_{n} [1002, 1556]. S · π [1140]. σ [1926, 1924]. SL(2, C) [1767]. SL(2, C) [334]. sl(2, R) [250]. SL(3, R) [1779, 1949, 1474]. SL(3, R) [815, 1204, 64]. SO(1, 1) [1397]. SO(2) [1397]. SO(2, 1) [553, 1397]. SO(3) [1189, 1840]. SO(4) [553]. SO(4, 1) [82]. SO(5) [524]. SO(N) [629, 916, 1418, 70, 379, 474, 1236]. SO(n − 1, 1) [1458, 379]. Sp(2M) [629]. Sp(2n) [1721]. Sp(8, R) [1365]. s → ∞ [1188]. SU(1, 1) [305, 198, 644, 1285]. SU(2) [1322, 14, 305, 1681, 384]. SU(2)_{Λ}R^{4} [1770]. SU(2) ⊗ SU(2) [613]. SU(2) × SU(2) [638]. SU(2, 1) [1694]. SU(2, 2) [1024]. SU(2, 2/1) [1281]. SU(3) [696, 1034, 1371, 1840]. SU(4) [1692, 1083, 716, 638, 398]. SU(4) ⊆ SU(2) × SU(2) [925]. SU(4) ⊃ SU(2) ⊗ SU(2) [1205]. SU(6) [476]. SU(6) ⊇ SU(3) ⊗ SU(2) [1583, 1782]. SU(N) [1460, 1535, 1194, 597]. SU_{2} [119, 1172, 1549]. SU_{2} ⊃ ⋯ ⊃ G′′ ⊃ G′ ⊃ G [545]. SU_{2} × SU_{2} [214]. SU_{3}^{c} [1207]. SU_{6} [1207]. T [1676, 675, 8, 172, 1518, 1559, 241, 892]. × [1207]. U(1) [1880]. U(2) [364]. U(3) [387]. U(3): U(2) [387]. U(5) ⊆ O(5) ⊆ O(3) [516]. U(M) [629]. U(N) [629, 597, 1031, 1242, 364, 100, 1396, 133, 315, 648, 1390, 1893]. U(n, 1) [140]. U(p) × U(q) [1838, 1839]. U(p + q) [1838, 1839]. U(p, q) [1838]. U(p, q) [1839]. U_{1}(M) ⊗ U_{2}(M) ⊗ U_{3}(M) [995]. V (X) = a∥X∥ [605]. φ^{2} [1826, 1668]. φ^{4} [1826]. ϑ [775, 352, 544]. W^{∗} [1789]. W_{t}^{2} − W_{x}^{2} − W_{y}^{2} = 0 [1156]. W_{t}^{2} − W_{x}^{2}W_{y}^{2} = 0 [1157]. x = 1 [582]. x^{−2} [764]. x^{2α} [620]. y = ±1 [582]. Z [498]. Z_{2} ⊗ Z_{2} [1483].

-algebra [1005, 845]. -algebras [353, 1906]. -analytic [431]. -band [890]. -bell [923]. -body [1025, 1329, 1750, 528, 917, 1706, 336, 1100, 1386, 1904, 1696, 485]. -bubble [796, 1176]. -coefficients [1615, 316, 317]. -correlated [614]. -dimensional [1089, 1048, 1047]. -dimensions [221, 1245]. -electron [1710]. -equation [1641]. -equations [867, 1202]. -Fd3m [52]. -fibre [1880]. -Fm3m [52].

-function [1201]. -generalized [1007]. -harmonics [1129, 1405]. -Hilbert [360]. -level [727, 538, 616]. -like [764, 1480]. -matrix [307, 1069, 83]. -metric [490, 1447, 1874]. -mode [1341]. -model [1926]. -momentum [1731]. -parametric [647]. -plet [1660]. -point [1223]. -product [1869]. -quantum [1060]. -representability [802, 803]. -representation [279]. -representations [1409]. -separable [458, 459, 1012, 1332]. -site [33]. -soliton [1896]. -space [1827, 494]. -spaces [496, 514]. -sphere [1317]. -state [1313]. -structures [1461]. -system [1295]. -variable [150]. -vector [1534]. -wave [1785, 1950, 234, 1676].

15015b [1951]. 16 [624, 775]. 17 [1787]. 18 [1953, 1451, 1091, 1065, 915]. 19 [1686, 1887, 1888, 1943, 1941, 1591, 1557, 1590, 1946, 1605, 1815, 1726, 1814].

20 [1942, 1935, 1949, 1944, 1950, 1952, 1889, 1951, 1945, 1890, 1940, 1947, 1936, 1938, 1939].

55 [1953, 1787]. 56 [1946]. 58 [1557].

80b [1940]. 80f [1952]. 80g [1945]. 80h [1889]. 80i [1949]. 80k [1944]. 81038b [1451]. 81039a [1889]. 81d [1950]. 81e [1451]. 81j [1951]. 82045a [1940].

9 [225].

= [541].

Abelian [1761, 1923, 1192, 1289]. above [573]. Absence [1252, 1737]. absorbing [623]. abstract [736, 112]. accelerated [1758]. acceleration [1108, 107]. accidental [1046, 1048, 28, 304]. Acoustic [84, 853, 1210, 1277, 1531, 907]. acoustics [778, 1953]. across [1757]. acting [884, 1301]. action [1620, 1286, 1817, 974]. actions [1569]. activity [1633, 522, 561]. adapted [1400]. adapting [133, 315]. Addendum [624, 1590, 225, 1726, 1556, 915]. addition [704]. Adiabatic [156, 1713, 1640, 1948, 1280, 1478, 57]. adiabats [1226]. adjoint [353, 372]. adjointness [1865]. admit [1462, 1154]. admittance [467]. admitting [1832, 737, 943, 756, 944, 454, 1209]. advanced [10, 202]. Affine [1006, 441, 1537]. affinity [1830]. Airy [1863]. AKNS [831]. Albedo [1927, 1552]. algebra [163, 290, 1323, 1368, 1609, 250, 347, 276, 845, 1005, 1473, 1366, 468, 1454, 613, 1369, 1889, 1614, 1492, 1849, 375, 374, 384, 1680, 779, 712, 1748, 638, 515, 654, 655, 1497, 1365, 718]. Algebraic [100, 146, 939, 626, 1434, 964, 1666, 47, 1537, 1898, 341, 342, 1301, 1697]. Algebraically [514, 170, 410, 852, 1209]. algebras [1620, 1861, 987, 464, 88, 785, 1837, 1609, 1229, 1282, 276, 702, 319, 1634, 770, 1086, 1384, 1789, 297, 1391, 1853, 1091, 919, 821, 1468, 1612, 548, 353, 1906, 1109, 1639, 330, 505, 564, 965, 571, 1561, 1446, 1483, 654, 655, 768, 769, 1652, 1052]. algorithm [1720]. algorithms [966]. aligned [1121, 1542, 1159]. alloys [736]. Almost [1461, 547, 1372, 1461]. almost- [1372]. Almost-Hermite- [1461]. almost-Killing [1372]. almost-product- [1461]. Almost-structures [1461]. along [292, 560, 293]. alternative [1190]. Altman [1562]. among [1320, 100, 1653]. Ampère [1175]. Amplitude [1, 1719, 1516, 601, 225, 833, 1044]. amplitudes [1550, 307, 835, 1276, 971]. analogs [1380, 1557]. Analysis [1855, 230, 624, 1120, 1540, 569, 1886, 46, 1846, 1466, 835, 442, 1050, 287, 1890, 1619, 1940, 1777, 1489, 1338, 810, 1365, 133, 315]. Analytic [14, 1101, 50, 1201, 1544, 1527, 1293, 241, 492, 916, 1707, 1633, 434, 1628, 1835, 1361, 741, 543, 498, 815, 1204, 1667, 435, 431, 746, 531, 669, 607, 1533, 3, 44, 405, 1349]. analytical [1643, 1298, 1943, 1550]. analyticity [684]. analyzing [1377]. angle [1560, 1816]. angle-dependent [1560]. Angular [586, 1740, 235, 66, 1029, 1859, 1951, 1688, 686, 316, 317, 534, 83]. angular-spectrum [534]. anharmonic [1538, 1361, 481, 1174, 1292, 394, 313, 620, 442, 1574, 1668, 1670, 1745, 176]. anharmonic-oscillator [1670]. anharmonicity [313, 620]. anholonomic [868]. anisotropic [1855, 600, 1119, 358, 787, 1055, 1545, 872, 685, 1475, 1046, 1048, 1045]. anisotropy [1150]. annular [1120]. anticommuting [1139, 1360]. antifermion [29]. antinormal [1788]. any [1322, 1133, 1941, 1651, 1755, 1261]. Appell [669]. applicability [1817, 1080]. Application [481, 46, 1174, 1292, 1562, 1745, 5, 43, 1707, 865, 1784, 1609, 1361, 731, 1030, 1438, 1741, 1420, 1859, 1951, 424, 1865, 1147, 1824, 1903, 1708, 1252, 1123, 1453, 1135, 1432, 736, 91, 1927, 188, 1608]. Applications [1666, 1935, 1632, 668, 873, 674, 205, 1804, 1581, 233, 867, 1201, 213, 127, 1928, 1868, 752, 1028, 969, 881, 774]. applied [991, 1013, 1345, 1899]. Approach [444, 1088, 1934, 587, 336, 1884, 1886, 1151, 73, 124, 437, 438, 155, 1662, 895, 1203, 1794, 341, 1900, 512, 455, 28, 649, 146, 1697, 1768, 1548, 1885, 1190, 1568, 193, 1920]. approaches [556, 1661, 826]. approximant [499, 517]. approximants [150, 143, 162, 570, 58, 542, 1516]. Approximate [582, 375, 948, 1343, 1656, 969]. approximated [1222, 1431]. approximately [639]. Approximating [823, 968]. approximation [592, 529, 748, 1129, 849, 1572, 1952, 609, 151, 301, 1169, 1856, 1431, 1564, 181, 1060, 1106]. approximations [370, 91, 1405, 848, 1676, 549, 1312, 560, 808, 812, 621, 317, 534, 1393, 605]. arbitrary [243, 961, 1486, 1602, 1412, 1912, 1698, 132, 870, 1824, 311, 820, 83, 117]. arising [1108, 920]. array [114, 113]. Ashkin [842]. aspects [1387, 335, 767, 138, 1203, 789, 802, 803, 1042]. assembly [106, 1167]. associated [987, 1682, 1710, 1136, 1645, 1064, 334, 103, 622, 817, 1056, 1612, 440, 1389]. assumptions [730]. Asymmetric [879, 1275, 760, 1705]. Asymptotic [1543, 1273, 640, 1238, 448, 1754, 1364, 986, 902, 812, 1714, 1936, 525, 526, 534, 1291, 1487, 1376, 1550, 1386, 1168, 767, 1010, 151, 1459, 753, 328, 1863, 1395, 1605, 1285, 565]. asymptotically [1800, 68, 1912, 1799]. atmospheres [1927]. atom [592, 173, 665, 1304, 1298, 1943, 1032]. atomic [1250, 1659, 1682, 1710, 1298, 1943]. atomlike [1442]. atoms [1682, 1710, 760]. attempt [1675]. attractive [1739, 102]. augmented [1006, 1562]. autocatalysis [1934]. autocorrelation [54]. automorphism [1906]. Automorphisms [1597, 718, 1537, 353]. Average [366, 1456]. averaged [641]. averages [388]. averaging [359]. avoiding [1130]. axial [964, 40]. axially [1738, 942, 9, 171, 1103, 1587, 1674]. axially-symmetric [1103]. Axiomatic [928, 453, 975]. axiomatics [978]. axis [1536, 689, 687, 53]. Axisymmetric [795, 743, 1612, 147, 1728, 1417, 294]. axisymmetry [1715]. azimuthal [1290, 1590]. Azzarelli [1044].

Bach [1142]. background [48]. backgrounds [636, 1602]. Bäcklund [541, 238, 382, 1139, 1847, 1504]. bad [936]. balance [1731]. Banach [36, 1474]. band [890, 706]. bands [549]. Bargmann [358]. barotropic [1240]. barriers [573]. baryon [964, 1259]. baryonium [1692]. baryons [1692]. base [1049]. based [584, 1707, 91, 518, 1340, 1686]. bases [1397, 588, 589, 1400]. basic [1100, 739, 810, 1191, 437]. basis [1630, 1547, 434, 1925, 498, 79, 1838, 1839, 1218, 1840, 1396, 928]. Baxter [195]. Baym [1520]. BCH [208]. Beam [1536, 199, 48, 1798, 31, 1078]. behavior [1452, 1225, 1370, 1796, 391, 986, 902, 1276, 1863, 1395, 1605, 1291]. behaviour [1168]. bell [923]. belonging [943]. below [271, 1032]. Benjamin [1685, 1093]. Bergmann [1514]. Bernoulli [872]. Bertotti [1177]. Bertrand [1425]. Bessel [1138]. Bethe [361, 524, 29]. between [507, 470, 1143, 1155, 1259, 1433, 621, 1788, 1293]. BGK [810]. Bianchi [575, 1597]. Bifurcate [185, 1736, 98, 99]. Bifurcation [653, 705, 1406]. Bilinear [450, 548, 644, 1167]. Binary [663, 138, 1607]. binding [1570]. biorthogonal [1008]. Biot [1180]. Birkhoff [1231]. Birman [1819]. Bivariational [570]. Bivector [973]. Black [413, 1260, 433, 1022, 606, 732, 1759, 519]. blowing [1021]. Blume [1795]. bodies [671, 672]. body [1268, 249, 1025, 1329, 1750, 917, 1706, 336, 182, 1100, 1386, 901, 1705, 198, 155, 23, 1904, 528, 1725, 142, 859, 1254, 1336, 1786, 1945, 1906, 256, 1696, 485, 1346]. body-centred [142]. Bogoliubov [690, 1500, 847, 828]. Bogolubov [647]. Bogolyubov [1166]. Bogolyubov-boson [1166]. Bohr [1420]. Boltzmann [242, 1163, 1555, 631, 899, 284, 1248]. Bondi [816, 1872]. bonds [121]. Bonnor [1915]. Boost [1766, 1017, 379]. Bopp [788]. Borel [1872, 1292, 309]. Born [517, 621]. Bose [1889, 690, 1086, 1722, 1614, 1794, 752, 38, 1137, 1470, 1918]. Bose-like [1137]. Boson [1565, 814, 221, 729, 1606, 813, 590, 847, 76, 515, 1788, 1365, 1166, 550, 589]. boson-fermion [515]. bosons [846, 1167, 1877]. both [337]. Bound [1480, 592, 1421, 1298, 1943, 1098, 1870, 421, 1169, 478, 1482, 993, 1395, 1605, 479, 910, 1560, 531, 969, 893, 1479, 310]. bound-bound [1298, 1943]. bound-state [1870]. Boundaries [94, 1072, 287]. Boundary [1335, 603, 1119, 1277, 1645, 887, 1636, 191, 971]. bounded [91, 279, 86, 1010, 1011, 1807, 512, 418]. Boundedness [271, 252, 1867]. Bounds [467, 971, 570, 499, 865, 729, 1442, 1039, 1255, 1696, 579, 590, 1234, 1663, 531, 969, 1819, 1188]. box [1134, 351]. bracket [718]. brackets [674, 1166]. Branching [1239, 628]. Brans [1357, 1762, 1828, 1193, 1288, 795, 1563, 1875]. breaking [755, 637, 995, 994]. Breit [260]. Broken [970, 1921, 1523, 1489, 523]. Bronzan [1060]. Brownian [889]. bubble [796, 1176]. bulk [1026]. bundle [163, 952, 887, 1020, 1489]. bundles [360, 1429, 1428, 1880, 1266, 1691, 742, 222, 1786, 1945, 752, 1306]. Bures [1834].

calculate [657]. calculated [605]. calculating [549]. Calculation [967, 1066, 79, 354, 1250, 1280, 804, 888, 1313, 134, 90, 716, 1870]. calculations [1656, 1298, 1943, 229]. calculus [953, 1414, 429, 711, 1013, 1463, 1330, 1729]. Calogero [484, 126]. Canonical [250, 283, 483, 1577, 660, 750, 28, 304, 588, 589, 251, 639, 1604, 506, 1429, 598, 733, 1004, 1116, 728, 1409, 901, 941, 604, 1233, 1155, 991, 387, 1868, 277, 30, 323, 511, 246, 588]. carriers [916, 508, 509]. Cartan [952, 77]. Cartesian [257]. case [1718, 1942, 814, 280, 1363, 1087, 813, 177, 1126, 430, 188]. cases [1130, 887, 759]. Casimir [262, 25, 1368, 1853, 651, 957, 1109, 1488]. Cauchy [1379, 189, 1021, 1022, 1624, 117, 118]. Causal [791, 432]. Causality [1502]. Causally [985, 1107]. caused [877, 800]. cavity [877, 800]. Cayley [121]. CCR [452, 332, 1849]. cell [289, 504, 471]. censor [749]. Center [1750, 1900]. centered [717]. centers [901]. central [276, 1164]. centred [142]. certain [1827, 1388, 770, 1444, 364, 684, 738, 605, 1533, 223, 302]. Chaba [448]. chain [1411, 1804, 545, 1205, 712, 1803, 1190]. chains [427, 1321, 5]. Chandrasekhar [1641]. Change [129, 1022]. channel [235, 234, 1846, 601, 1663, 652]. channels [466]. Chapman [106]. Characteristic [1801, 1937, 320, 1757, 617, 1853, 1859, 1951, 539, 118]. Characterization [1388, 852, 566, 1149]. characterizations [1707]. Charge [1235, 511, 1025, 1329, 1143, 710, 301, 34, 756, 944, 1852, 1224, 11, 703]. Charged [1913, 861, 1750, 1280, 640, 839, 869, 1874, 217, 1084, 606, 1673, 828, 846, 1348]. charges [586, 1730, 1731, 1489]. charmonium [1824]. Chebyshev [1361]. chemical [930, 945]. chiral [1810, 214, 754]. CIR [728]. circle [1601]. Circular [1705, 1111, 1009]. class [917, 1072, 1525, 243, 305, 635, 1598, 1444, 851, 1409, 333, 340, 1842, 1164, 1003, 1227, 184, 209, 1463, 562, 959, 1110, 1665, 486, 515, 1252, 1615, 994, 999, 1337, 506, 906]. Classes [1033, 1248]. Classic [555]. Classical [1923, 1374, 1730, 1731, 1928, 880, 1106, 1797, 1097, 987, 464, 246, 190, 1845, 1835, 1198, 1465, 849, 946, 580, 335, 1384, 481, 941, 796, 1176, 1677, 1623, 758, 103, 817, 301, 219, 1868, 1564, 1592, 1850, 1673, 1468, 1708, 1109, 571, 1655, 1876, 1105, 1070, 915, 761, 656, 797, 192, 974, 950, 1185, 57, 223, 302, 861]. Classification [1599, 1609, 1922, 1944, 654, 655, 1515, 1832, 1885]. Clebsch [725, 1556, 1692, 553, 51, 52, 1646, 1648, 1649, 716, 1781, 435, 628, 384, 398, 397, 1354, 477, 756, 1002, 1017, 1841]. Clifford [1282, 1086, 1473, 341]. Closed [668, 50, 1307, 1582, 641, 1107, 847, 405]. Closure [787]. Cluster [587, 1311, 182, 438, 609, 643, 1160, 1293]. cnoidal [602]. coated [1144]. coaxial [160]. coefficient [507, 660, 750, 254, 1443, 1464]. coefficients [1692, 553, 51, 52, 725, 1319, 305, 1776, 1773, 1775, 1646, 1648, 1649, 1878, 578, 716, 495, 1781, 213, 1764, 1371, 384, 398, 397, 76, 1354, 477, 1615, 1002, 1556, 316, 317, 1583, 1782, 1400, 1399, 114, 550, 595, 648, 1017, 1390, 1236, 1893, 1841]. coherence [1768, 907, 1178]. Coherent [1503, 1064, 1308, 1121, 1159, 492, 190, 1538, 450, 481, 369, 183, 1470, 895]. Coherent-state [895]. cohesion [1350]. cohomology [1392]. cold [424]. collapse [1765]. collapsed [881]. Collas [1044]. collective [882, 1625, 1049]. Colliding [1578]. Collineations [862, 943]. collision [1101, 1120]. collisionless [1732, 1733]. collisions [1250]. color [1366]. colored [995]. combinations [1322]. combinatorial [285]. combinatorics [1171]. combined [1059]. Comment [1780, 775, 504, 480, 289]. Comments [1827, 1451, 1952, 1065]. common [1810]. communicability [978]. commutation [1409]. commutator [1492]. commuting [925]. compact [722, 850, 1472]. compactified [1681]. Compactness [1682, 1710]. comparison [1843, 826]. compatibility [122]. compatible [1487]. competing [656]. complementary [25]. Complete [955, 244, 687, 1487, 1904, 1674]. Completely [1569, 538, 1062, 1895]. completeness [1603]. Complex [1266, 1457, 1774, 9, 171, 1322, 916, 570, 416, 222, 104, 1255, 431, 386, 385]. complex-coordinate [1255]. complexion [72]. component [1871, 632, 940, 796, 1176, 391, 472]. components [686]. Composite [922, 1793, 871, 308, 1546, 461, 608]. composition [187]. compound [1542]. compressible [982]. Computation [1615, 70, 379, 474, 1458, 1651, 1353, 1354, 381]. computing [1361, 1358]. Comultiplicator [488]. concerning [1160, 1126]. concomitant [1435]. condensed [1362]. condition [169, 456, 1800, 767, 757, 1420, 1074, 184, 209, 1335, 1105, 1560]. Conditional [918, 1403, 273]. Conditioning [513, 1508]. Conditions [1476, 1119, 955, 1277, 1741, 1419, 1700, 1421, 1914, 1645, 1681, 802, 803, 562, 1192, 1638, 603, 633]. conducting [34, 164]. conduction [350, 167, 1180]. conductivity [736, 683, 462, 1813]. conductor [206]. cone [916, 396, 1426]. cones [528]. Configuration [250, 1659, 1682, 1710, 962, 178, 1293]. configuration-interaction [1293]. configurations [641, 858, 933]. confined [424, 1918]. confluent [1301]. conform [412]. Conformal [917, 1618, 468, 510, 1199, 1917, 175, 1595, 1018, 1522, 585, 1477, 611, 1243, 1298, 1943, 543, 65, 498, 432, 157, 822, 601, 407, 1684, 1375, 1372]. Conformally [145, 1037, 1240, 1747, 1563, 1886, 291, 1885]. congruences [921]. conjecture [80, 426, 749, 299]. conjugate [1779, 1949, 1034]. conjugation [1774, 1143]. connected [598, 236, 377, 1223, 558]. connecting [1397]. connection [952, 1143, 719, 1437, 1293]. connections [897, 1306]. consequence [1180]. Consequences [567, 341]. Conservation [1623, 710, 1506, 1235, 1448, 914, 1435, 789, 1155, 680, 1299, 1726, 525, 526]. Conservative [793, 22, 430, 562]. Conserved [1718, 1942, 201, 122, 1376, 1664, 1720, 406, 1499]. Consistency [105]. constant [1117, 1594, 1690, 1381, 523, 798]. constants [1148, 103, 622, 817, 934, 642, 227, 1468, 1644]. constitutive [672]. Constrained [1621]. Constraint [1542, 1505, 118]. Constraints [1394, 1514, 532, 806, 1289, 1638]. construct [1720]. constructed [774]. Constructing [1507]. Construction [1837, 234, 56, 886, 445, 1609, 1457, 1035, 1278, 45, 1237, 1036, 655]. Constructive [1707]. contact [727, 616]. containing [117]. content [1698]. continua [1324, 1325, 1326, 1327]. continuation [1550, 1201, 1667, 379, 746, 669]. continuations [1527]. continued [542, 1650]. Continuous [264, 265, 1094, 890, 1628, 1134, 47, 1064, 1107, 1013, 959, 304, 447]. Continuum [1932, 711, 1013, 1463, 1330, 1393, 671, 672]. contour [405]. contracted [943]. contraction [952, 1919]. contractions [1635, 1391]. contractor [1562]. Convergence [143, 336, 1579, 1108, 808, 1659]. Convergent [766, 425, 233, 1478]. converging [370]. Convolution [867, 1138]. Cook [128]. coordinate [1750, 1534, 1150, 458, 184, 209, 612, 1255, 1453, 1135]. coordinates [97, 371, 399, 792, 1629, 1830, 752]. Coping [1200, 1887]. core [559, 877, 800]. Cores [339]. corner [206]. Correction [162, 806, 570, 67]. corrections [501, 580]. correlated [614, 1121, 1159]. Correlation [996, 295, 1178, 580, 662, 1215, 1700, 1320, 561, 1295, 1456, 530, 1271, 1811, 215, 1114]. Correlations [1341, 1043, 272, 1737]. Correspondence [503, 355]. corresponding [1906, 652, 466, 1401]. Corrigendum [1414]. coset [133, 315, 278]. cosine [1283]. cosmic [749]. cosmological [636, 1073, 787, 1690, 659, 523]. cosmologies [1884, 1886, 1739, 1375, 999, 1000, 1885, 852]. Coulomb [1954, 487, 320, 1297, 1054, 263, 1471, 1905, 1257, 215, 1638, 649, 812, 1697, 735, 83, 485, 912, 1153, 1518, 241, 892, 893, 1479, 1349, 1703, 1910]. Coulomb-like [1471, 1703, 1910]. Coulomb-projected [1518]. Coulombian [565]. Coulter [510]. counting [1543, 1077]. Coupled [606, 1758, 367, 1538, 481, 231, 156, 1713, 914, 394, 101, 713, 714, 715, 1617, 1293]. coupled-cluster [1293]. Coupling [1776, 1773, 1775, 556, 1779, 1949, 235, 1606, 222, 1859, 1951, 829, 1594, 1115, 316, 317, 115, 1367, 1628]. covariance [32]. Covariant [953, 1414, 1660, 1602, 901, 402, 457, 1136, 1287, 383, 1324, 1325, 1326, 1327, 1924]. cover [806]. covering [65, 644, 1285]. criterion [878]. Critical [493, 676, 556, 1784, 958, 1257, 1745, 1445]. critical-coupling [556]. Critical-exponent [493]. Criticality [1530]. Critique [591, 1661]. Cross [207, 1, 621, 627, 120, 1117, 1731, 1456, 292, 560, 225, 1696, 531, 1381, 1797, 1188, 293]. cross-section [1117, 292, 1797]. crossed [601]. crossed-channel [601]. Crossing [1908, 1909]. crystal [1784, 51, 725, 197, 463, 442, 471, 697]. crystalline [1108, 705]. crystallographic [489]. crystals [1932, 1252, 1586]. cubic [1302, 142, 373, 282]. cumulant [63, 591, 74]. cumulants [1902]. curiosities [954]. curl [201]. Current [1732, 1733, 613, 1122, 375, 374]. currents [1136, 1287, 1921, 824, 146]. Currie [1249]. Curvature [896, 174, 719, 1309]. curve [1272]. curved [543, 1074, 432, 1079, 1141, 1907]. curves [441, 1786, 1945, 997, 959]. cut [971]. cutoff [1112, 1714, 1936]. cycle [1934]. Cyclic [1866]. cylinder [1111, 239]. cylinders [292, 560, 937, 293]. cylindrical [1829, 248, 1848, 596]. Cylindrically [1357, 199, 400, 1762, 92, 27, 424, 392]. cylindrically-symmetric [27].

D [1198]. D. [80]. damping [1307, 111, 231, 1661, 1823]. data [1800, 1339, 1801, 1937, 594]. Debye [1579, 1700, 1393]. Decay [272, 1737, 592, 1452, 1225, 439, 1870]. decays [964]. declension [1769]. decomposition [319, 364, 686, 1160, 1001, 1533, 372, 1553]. Decompositions [258, 1634, 1907]. defined [1384, 486]. Definition [1279, 125, 782, 1799, 1197, 1894]. deformable [671, 672, 164]. deformation [572, 877, 800]. Deformations [1873, 674, 1851, 103, 140]. deformed [1469]. Degeneracies [61, 1124, 1638]. degeneracy [1133, 1941, 1046, 1048, 922, 28, 304, 417]. degree [1368, 1299, 1726]. degrees [900, 1522]. delay [1706, 256]. delta [1253, 562, 1264, 1441]. democratic [517]. denominators [310, 425, 1478]. Dense [1133, 1941]. densities [1718, 1942, 1827, 848, 1720, 1499, 201]. density [935, 667, 182, 956, 1741, 1313, 641, 155, 958, 1122, 260, 1271, 1608, 471, 1811, 1877, 1179]. dependence [1751, 1290, 1590, 288]. dependent [736, 637, 587, 276, 1116, 189, 731, 1090, 1253, 894, 1427, 1055, 837, 1769, 528, 103, 622, 817, 934, 78, 1823, 1038, 991, 1573, 1574, 931, 1868, 346, 834, 1184, 1844, 1560, 1338, 1745, 810, 934]. depolarization [1542]. depth [1475]. Derivation [667, 561, 992, 805, 522, 929, 1147, 847, 1142, 1625, 16, 1398]. derivations [1620, 353]. derivative [1251, 1115]. derivatives [1897, 898, 1863, 1663]. derived [269, 643, 1564]. describing [77]. description [1487, 639, 161]. descriptions [250, 945, 342]. Details [1098]. detection [125]. determinantal [359]. determinants [666, 731, 1627]. Determination [498, 1, 225, 417, 1784, 319, 1634, 1174, 334, 21, 1257, 374]. determines [959]. determining [1712]. development [713]. deviation [291, 497]. diagonalizability [259]. Diagonalization [1167, 1617, 798]. diagonalizations [1858, 1859, 1951]. diagrams [714]. diamond [52]. Dicke [1357, 1762, 1828, 653, 745, 1193, 1288, 795, 1563, 1875]. Dicke-like [745]. dielectric [583]. difference [108, 569, 499, 578, 306, 1028]. different [1200, 1887, 826, 631]. Differential [731, 348, 1, 108, 569, 1707, 600, 913, 635, 1895, 1030, 1438, 552, 156, 1713, 1016, 1383, 432, 1842, 1265, 1227, 158, 724, 1407, 1867, 225, 938, 1378, 567, 1748, 812, 1924, 152, 1820, 1883, 506]. differential-difference [108, 569]. differentiation [457, 1013]. Diffraction [580, 689, 1274, 801, 1111, 1356, 1854]. diffractive [963]. diffusion [1072, 1525, 1934, 139, 151, 1003, 1148, 1157, 237, 562, 759, 603, 1097, 1398, 1821]. diffusion-reaction [1934]. Dilatations [296]. Dilations [911]. dimension [917, 1253, 1206, 215, 573, 564, 928, 1877]. dimensional [93, 1402, 1515, 1609, 1181, 82, 1738, 1589, 195, 324, 1054, 1089, 173, 1304, 1412, 1138, 1421, 1869, 210, 1650, 495, 75, 263, 1391, 859, 1050, 685, 1667, 1651, 1755, 1048, 1047, 300, 923, 1124, 1300, 875, 780, 1585, 295, 177, 1092, 1709, 123, 1303, 965, 1926, 1930, 1360, 1763, 365, 54, 605, 268, 933, 663, 417, 1398, 706]. dimensionality [1128, 838, 61, 178]. dimensionally [1719]. dimensions [221, 1245, 243, 1448, 382, 1588, 1785, 1950, 632, 940, 198, 498, 1265, 1286, 802, 803, 227, 694, 794, 1114, 1697, 1261, 1480]. Diophantine [1633]. dipole [245]. dipoles [1345]. Dirac [1526, 487, 1394, 449, 851, 333, 674, 1623, 1514, 105, 1059, 104, 1865, 119, 1882, 735, 1717, 135, 1172]. Direct [319, 1634, 1413, 1552, 384, 1344, 1081, 555, 1433, 1160, 268]. direct-interaction [1160]. Direct-inverse [1413, 1552]. directed [784, 1575]. direction [1475]. directions [681]. Discontinuities [1757, 1933]. Discrete [1088, 869, 1380, 1557, 215, 20, 286, 855, 1529, 771]. Disequilibrium [1899]. disordered [736, 1321, 1190, 5]. dispersion [1855, 230, 624, 507, 3, 24, 44, 405, 706]. dispersionless [1812]. Dispersive [946, 1584, 789, 1763]. Disproof [80]. dissipation [1662]. Dissipative [1876, 321, 175, 1655, 148, 799]. dissipatively [1432]. distance [1834, 639]. distinguishable [1124]. distinguished [689]. distorted [889]. distortion [1424]. distribution [557, 1762, 1273, 1550, 609, 958, 1532, 309, 1244, 1222, 1345, 1026]. Distributional [1749]. distributions [1357, 1846, 1463, 328, 546, 822, 314, 756, 944, 1121, 1159]. disturbances [1752]. divergence [1827, 1541, 1486, 973]. divergence-free [1827, 1541, 973]. divergent [1516, 1745]. diverging [91]. divided [306]. divided-difference [306]. divisible [772]. Does [1906, 1818]. domain [1277]. domains [1807]. domino [972]. Donder [184, 209]. Double [133, 315, 1571, 938, 1816, 278]. double-hypergeometric [938]. doubly [178, 113]. doubly-excited [178]. down [1117, 1878, 1540]. Dual [15, 290, 1535, 1008, 454]. Duality [472, 1761, 169, 456, 1805, 1847]. dually [839]. duals [1051, 1744]. Dubourdieu [1462]. due [805, 34, 1746, 1947]. dumbbells [1124]. Dunn [67, 1563]. Dust [1762, 1357, 1068, 937, 294, 392]. Dynamic [390]. Dynamical [465, 250, 679, 1742, 198, 538, 1769, 1131, 103, 622, 817, 934, 1671, 374, 1553, 1328, 1416, 568]. Dynamics [1134, 537, 1534, 174, 1226, 1862, 665, 1067, 116, 512, 1431, 765, 312]. dynamo [406]. dyon [758].

Eckart [821]. Ecker [1676, 1393]. edge [1356]. Edward [1952]. Effect [111, 71, 1152, 1650, 1500]. Effective [1523, 1572, 1952, 74, 75, 289, 504, 826, 1113, 455, 892, 74]. Effective-medium [74]. effective-potential [455]. effectively [1128, 838]. effects [139, 1758, 193]. eigenfrequencies [1550]. Eigenfrequency [935, 1273]. eigenfunction [430]. eigenfunctions [1395, 1605]. eigenmode [877, 800]. eigensolutions [1754]. eigenstate [1482]. Eigenstates [1322]. Eigenvalue [865, 1361, 1532, 1491, 1521, 1687, 1825, 865]. Eigenvalues [597, 651, 1140, 211, 1151, 1154, 1358, 746, 957, 1721, 1596]. Einstein [1059, 566, 874, 790, 611, 402, 955, 743, 92, 1388, 1715, 1657, 942, 783, 68, 1342, 1917, 610, 1211, 1911, 1912, 710, 1231, 1801, 1654, 1937, 1288, 1435, 980, 981, 1436, 1459, 35, 684, 973, 77, 1674, 1679, 145, 368, 737, 165, 95, 699, 385, 786, 687, 338, 129, 1417, 454, 303, 1498, 1385]. Einstein-massless-scalar [1435]. Einsteinian [1127]. elastic [1268, 230, 624, 961, 560, 1476, 1560, 833, 293, 1326]. elastically [357]. electric [1025, 1329, 586, 1345, 756, 944]. Electrical [1813, 736, 74, 75, 289, 504, 826]. electrically [869]. electrodynamical [1049]. electrodynamics [1382, 1917, 683, 1730, 1731, 337, 1324, 1325, 1326, 1327, 288, 824, 846, 771, 992, 1626]. Electromagnetic [1260, 1288, 1673, 1510, 877, 800, 199, 1229, 634, 1845, 845, 1005, 983, 1666, 1409, 1932, 998, 850, 1654, 1471, 1254, 424, 606, 1356, 17, 1221, 1032, 145, 737, 1612, 943, 707, 72, 756, 944, 1279, 1875, 49, 389, 1415, 1177, 1159, 932, 164]. electromagnetics [778, 1953]. Electromagnetism [1880]. electron [1659, 1682, 1710, 580, 641, 470, 1133, 1941]. electron-gas [1133, 1941]. electronic [1108]. electronic-structure [1108]. electrons [284]. Electrostatic [62, 160, 1856]. electrovac [451, 1342, 1833, 1690]. electrovacs [1694]. electrovacuum [98, 99]. element [1220]. Elementary [275, 1237, 1497, 1592]. elements [964, 702, 888, 1298, 1943, 134, 90, 470, 1838, 1839, 1781, 70, 474, 1458, 1651, 1755, 1353, 1766, 1892, 278, 140, 1017]. ellipsoids [160, 1542]. elliptic [1707, 97, 1283]. embedding [698, 876]. emission [853, 1210, 173, 1304]. empty [410]. endomorphisms [802, 803]. ends [1124]. Energetically [1494]. energies [1467]. Energy [889, 1435, 1731, 944, 310, 729, 1589, 1297, 549, 1796, 231, 1055, 1442, 61, 269, 56, 426, 749, 1315, 1591, 421, 1039, 409, 1289, 1276, 993, 131, 346, 834, 1184, 1491, 1625, 160, 1191, 735, 1610, 1611, 605, 912, 313, 620]. energy-dependent [1055]. Energy-momentum [1435, 944, 269]. ensemble [273, 941]. ensembles [149, 574]. Enskog [391]. Entangled [1505]. entire [343]. Entropy [1404, 1295, 1328, 273, 242, 941, 205, 1834, 149, 1529, 1440]. enveloping [1620, 638]. EPD [460]. Epifanio [1787]. equal [1401]. equality [652, 466]. equals [1116]. equation [36, 487, 667, 58, 1641, 96, 637, 905, 1156, 361, 913, 541, 1685, 238, 1598, 1394, 1363, 1588, 349, 153, 380, 449, 1822, 990, 602, 851, 1836, 507, 333, 524, 1099, 641, 578, 1075, 1163, 1545, 224, 41, 721, 522, 343, 1795, 1352, 1383, 326, 1165, 1314, 1157, 1492, 306, 1093, 6, 255, 458, 459, 460, 744, 792, 1012, 1331, 1332, 1118, 325, 1251, 1056, 521, 1050, 1359, 612, 1335, 929, 1283, 13, 1555, 1316, 478, 1482, 631, 132, 1499, 500, 899, 1079, 794, 177, 418, 1092, 1709, 479, 1824, 1903, 1358, 1520, 212, 1378, 1114]. equation [1491, 45, 404, 812, 621, 1476, 1882, 284, 1524, 525, 526, 1540, 1504, 1381, 878, 607, 1372, 1248, 144, 1097, 356, 1498, 568, 1385, 261, 29, 1672, 1920]. Equations [1752, 1718, 1942, 108, 367, 569, 1481, 1891, 1112, 143, 1526, 1707, 791, 106, 611, 1072, 1525, 668, 1835, 975, 670, 1702, 402, 955, 1535, 743, 1161, 189, 1100, 635, 329, 1277, 518, 92, 1388, 855, 1682, 1710, 1895, 1041, 731, 1030, 1438, 1785, 1950, 1087, 1519, 1896, 1715, 552, 1757, 1657, 1602, 942, 783, 1419, 615, 307, 156, 1713, 1720, 510, 1722, 1021, 1623, 1016, 982, 1645, 138, 1010, 1011, 391, 1103, 1076, 914, 1059, 1864, 39, 266, 1342, 244, 395, 1842, 610, 1211, 1911, 1912, 710, 1231, 1801, 1937, 758, 1003, 334, 67]. equations [1227, 233, 867, 1202, 1859, 1951, 980, 981, 1436, 1459, 35, 393, 789, 1155, 158, 724, 1407, 856, 1830, 986, 1058, 1312, 322, 973, 1033, 593, 1528, 960, 1674, 119, 1019, 671, 1324, 1325, 1326, 1327, 145, 368, 737, 165, 1132, 95, 699, 1867, 1079, 1141, 694, 794, 830, 1843, 1562, 938, 926, 1378, 1506, 331, 1008, 639, 567, 1272, 1028, 385, 1844, 687, 1924, 152, 1693, 1729, 969, 129, 1797, 1717, 37, 303, 12, 55, 633, 1498, 506, 261, 118, 194]. Equatorial [1009]. equilibria [1890, 1619, 1940]. Equilibrium [38, 809, 173, 1304, 580, 219, 532, 806, 1876, 1070, 713, 714, 715]. Equivalence [305, 960, 584, 195, 1054, 1444, 945, 1186, 77, 846]. equivalent [652, 466, 1401]. Ergodic [357, 854, 109, 1446]. Ergodicity [109, 378]. Erratum [1942, 428, 1935, 625, 1329, 1004, 1948, 1452, 725, 1953, 202, 726, 1005, 1949, 1944, 1950, 349, 350, 940, 1128, 1686, 171, 1887, 1888, 452, 1176, 1943, 224, 1941, 504, 1937, 1889, 1591, 1066, 1954, 255, 1951, 473, 1557, 1945, 1091, 1330, 1890, 1940, 474, 277, 1946, 1605, 1947, 1815, 505, 1210, 977, 87, 1936, 652, 1275, 1415, 776, 877, 703, 315, 1453, 1814, 1938, 1939, 43, 44, 172, 832, 544]. Error [1255, 969, 1656, 804, 134, 812]. especially [892]. essential [203, 1865]. estimates [1784, 1656, 437, 812]. Euclidean [1943, 755, 82, 748, 698, 700, 437, 438, 1298, 65, 1826, 1831, 1147, 351, 1014, 445, 1015]. Euler [1935, 1632, 1044, 261]. evaluate [1381]. Evaluation [243, 210, 1212, 1782, 1101, 775, 308, 126, 1463, 1727, 316, 1390, 352, 544]. evanescent [101]. Even [1942, 1914, 1273, 1161, 1550, 1718]. even-parity [1273]. events [491]. Everett [1095]. Evolution [816, 1072, 1525, 926, 1718, 1942, 367, 1481, 855, 1041, 615, 641, 1720, 914, 960, 1132, 1305, 694, 794, 102, 1000, 37, 12]. evolutional [55, 633]. Exact [1198, 665, 763, 641, 198, 1298, 1943, 1320, 1321, 325, 78, 1823, 1032, 923, 165, 780, 1585, 316, 1531, 1270, 798, 1566, 1939, 1385, 1526, 667, 801, 516, 333, 1904, 340, 758, 21, 561, 1251, 132, 54, 1000, 1248, 1498, 29, 299]. exactly [917, 855, 1130, 131]. example [438, 131]. Examples [1343, 849, 416, 277, 30]. Exceptional [1600, 1639, 1690, 655, 1126]. exchange [470]. Excitation [556, 963, 812]. excited [178, 263, 1928]. exclusive [531]. exhibiting [49]. exist [1906]. Existence [1379, 899, 10, 202, 189, 1808, 754, 1132]. Exotic [1593, 1692]. Expanding [1150, 1019]. Expansion [704, 697, 1616, 1465, 1311, 1579, 437, 438, 591, 430, 74, 561, 127, 1147, 601, 870, 1670, 1594, 714, 1607, 1220, 471, 1467, 860, 1881]. Expansions [373, 1830, 14, 587, 182, 766, 1486, 156, 1713, 1516, 1258, 1581, 151, 257, 1441, 1788, 820, 1249]. expectation [804, 714]. expectations [918]. experiment [1381]. Experimental [144]. experiments [484]. Explicit [69, 1870, 1824, 1191, 1533, 194, 1390]. exploding [1022]. explosive [55]. exponent [493]. Exponential [592, 814, 1810, 813]. exponents [1745]. expression [76]. expressions [1486, 1298, 1943, 1721, 201, 1349]. Extended [682, 1076, 1735, 818, 1060]. extended-source [1060]. Extension [1613, 188, 1688, 532, 806, 276, 1582, 1915, 1132, 687, 844]. Extensions [1052, 1040, 362, 1199, 1098, 83]. exterior [1277]. external [1753, 1792, 1809, 1916, 1039, 1032, 1748, 146, 886, 861, 1279]. extra [173, 1304]. extrapolation [1631, 594]. extremum [90].

F [541]. face [717]. factor [726, 576, 707, 489, 1542]. Factored [747, 1723]. Factorization [154]. Factorization-method [154]. factors [476]. family [783, 943, 1624, 881, 1125, 1348, 1424]. FCC [763]. Fd3m [52]. Feenberg [736, 112]. Feller [759, 1335]. Fermi [983, 1642, 1086, 760, 1369, 1830, 752, 713, 714, 715, 1607, 1608]. fermion [271, 587, 1264, 589, 515, 29]. fermion-antifermion [29]. fermions [1589, 1311]. ferrimagnetic [671, 672]. ferromagnet [1402, 701]. ferromagnetic [1633, 1232, 314]. ferromagnets [840, 933]. few [1028]. Feynman [1330, 1814, 1719, 543, 371, 711, 1013, 1463, 408, 902, 1276, 1212, 691, 1408]. fiber [952, 1786, 1945, 462]. Fibre [1020, 1489, 1428, 1880]. Fibre-bundle [1020]. Field [1419, 1429, 1428, 93, 1753, 1792, 221, 246, 1543, 1229, 1370, 169, 801, 975, 112, 748, 723, 845, 1005, 983, 1666, 743, 1808, 698, 1280, 1511, 245, 1757, 1657, 173, 1304, 450, 927, 1809, 1916, 412, 1409, 942, 9, 171, 783, 1857, 457, 510, 958, 1722, 1082, 1103, 1826, 1321, 475, 1059, 253, 1342, 1230, 604, 1911, 1806, 710, 1231, 1801, 1654, 1937, 1471, 758, 1889, 1614, 67, 1435, 980, 981, 1436, 1380, 1557, 35, 683, 1155, 1817, 1039, 973, 1032, 1668, 1564, 1345, 32, 737, 165, 1147, 1673, 979, 1079, 1141, 1746, 1947, 1562]. field [1493, 1815, 351, 181, 579, 590, 808, 122, 72, 486, 886, 1744, 1852, 128, 1745, 1279, 1213, 389, 1415, 11, 703, 1811, 1797, 37, 303, 1121, 932, 949, 1185, 1261, 798, 1179, 1104, 164, 194, 975, 671]. Fields [34, 1599, 85, 1622, 790, 1828, 1871, 1477, 1761, 456, 1845, 587, 203, 1802, 1829, 1922, 1944, 1041, 1752, 380, 839, 836, 1392, 1789, 1932, 873, 1419, 1507, 217, 1228, 1035, 850, 1587, 1654, 371, 21, 1794, 1006, 7, 473, 1084, 537, 536, 287, 1221, 643, 159, 1192, 157, 145, 1612, 257, 795, 943, 1306, 407, 1037, 147, 1728, 1621, 861, 756, 944, 1875, 1417, 950, 820, 1014, 1172, 445, 1036, 1432, 1015]. Finite [1095, 678, 1609, 48, 91, 614, 1860, 1752, 173, 350, 167, 439, 1545, 1454, 1098, 1552, 1218, 1380, 1557, 802, 803, 219, 1667, 1426, 1289, 431, 1530, 181, 1028, 1803, 488, 365, 1191, 1049, 735, 423, 1841]. finite-difference [1028]. finite-dimensional [1609, 1667]. First [1656, 1857, 668, 1129, 1501, 1148, 299, 1924, 110]. first- [668]. first-order [1501, 1148]. First-stage [1857]. five [1738, 210, 495, 938]. five-dimensional [1738, 210, 495]. five-parametric [938]. fixed [1072, 56]. flat [1800, 1886, 68, 1912, 145, 291, 1037, 1240, 1747, 1563, 1885]. flatness [1487, 1459, 1595, 1375]. flow [396]. flows [865, 1102, 1425]. Floyd [1952]. Fluctuation [1145, 182, 1449, 16, 1026]. fluctuations [1099, 945]. Fluid [943, 4, 428, 865, 1273, 1161, 195, 839, 390, 1084, 35, 1636, 1913, 1043, 1240, 999, 881, 568]. fluid-dynamical [568]. fluids [557, 958, 376, 343, 141, 924, 13, 248, 1848, 680, 596, 713, 714, 715, 809, 1325]. flux [586, 1078]. Fm3m [52]. Fock [1507, 233, 1264, 128]. focusing [1536]. Foldy [1698]. force [1823, 1456]. forced [60]. forces [859, 1675, 799, 229]. form [785, 1534, 1465, 1929, 582, 510, 1155, 1528, 1506, 707, 568]. Formal [1667, 1570, 677, 1223, 713]. formalism [773, 246, 816, 731, 1427, 498, 1567, 1113, 1217, 128]. formation [1692]. Forms [1154, 889, 1265, 548, 621, 1007, 1714, 1936, 644]. Formula [70, 474, 1539, 208, 1411, 1509, 805, 561, 328, 379, 1334, 847, 1816, 1123, 595, 1398]. formulas [1660, 704, 188, 1458, 1651, 1456, 1208]. Formulation [1636, 36, 1753, 441, 600, 1181, 964, 1282, 182, 845, 1005, 692, 1715, 836, 9, 171, 1200, 1887, 901, 105, 622, 301, 1006, 1308, 824, 765, 639, 2, 87, 1621, 135, 771]. formulations [1831]. forward [833, 3, 44]. Foundation [1584]. Foundations [274, 987, 1592]. founded [1683]. Four [1298, 1943, 277, 30, 1515, 82, 1355, 1391, 884, 965]. four-dimensional [1515, 82, 1391, 965]. four-parameter [884]. four-wave [1355]. Fourier [569, 1577, 714]. Fourth [1368, 1754, 212, 1115]. fractions [542, 1650]. frame [40, 1758, 819]. frames [1923, 682]. framework [1793, 1087, 1519, 342, 1730]. Frankel [508, 509]. Fredholm [666, 731, 1729, 969, 356]. Free [1067, 850, 1827, 1541, 1229, 748, 1589, 1409, 376, 1765, 1454, 151, 973, 159, 786, 820, 648, 1390, 1893]. freedom [900, 1299, 1726, 1522]. frequencies [1190, 1032]. frequency [230, 624, 736, 766, 1079, 1141, 292, 560, 1907, 115, 293, 5]. fronted [1820]. Full [556, 188, 419]. function [1855, 434, 1628, 733, 1004, 208, 320, 580, 1253, 507, 609, 657, 1051, 1302, 263, 33, 142, 967, 1066, 966, 561, 1905, 1954, 1056, 1201, 366, 562, 299, 1264, 435, 1345, 853, 1210, 675, 1930, 531, 54, 1527, 892]. Functional [58, 429, 1076, 1360, 143, 370, 941, 1897, 343, 1364, 354, 1013, 486, 1768, 607]. functionals [1040, 518, 1223]. functionals-positive [1040]. functions [557, 166, 625, 1171, 492, 120, 916, 553, 10, 202, 1757, 1486, 1810, 741, 704, 1029, 1700, 1138, 457, 1241, 775, 470, 65, 1320, 1223, 823, 1244, 717, 373, 996, 1870, 867, 1218, 1113, 1283, 387, 643, 119, 282, 1863, 908, 288, 215, 1301, 1709, 1562, 957, 938, 590, 530, 260, 1271, 1714, 1936, 1768, 1441, 20, 381, 49, 1586, 735, 1811, 1146, 1178, 906, 1104, 1026, 352, 544, 117, 1349, 1703]. Fundamental [598, 264, 265, 1094]. Further [873, 502]. fuzzy [584, 781, 491, 661].

G [883]. G. [1787]. Galilei [163, 1247, 1699, 1412, 1206, 1199, 558]. Galileian [1282]. gap [421]. gas [690, 1054, 580, 1133, 1941, 138, 1264, 396, 215, 899, 38, 1930, 879, 1275, 1918]. gases [557, 810]. Gauge [1845, 332, 452, 457, 688, 288, 659, 1852, 846, 1599, 1622, 1158, 1083, 1535, 1311, 324, 1922, 1944, 1923, 1809, 1756, 1873, 1724, 1382, 1600, 1523, 1485, 622, 843, 1380, 1557, 658, 1192, 1289, 1110, 824, 1183, 1638, 1306, 1065, 1389, 1744, 1489, 994, 403, 1172]. Gauge-covariant [457]. gauges [1831]. Gaussian [1453, 1099, 1145, 1222, 1849, 1675, 31, 1175, 1135]. Gaussian-approximated [1222]. Gaussians [1146]. Gegenbauer [704, 1816]. Gel′fand [1667]. Gel’fand [791, 1151, 1242, 498, 364, 1118, 177, 1092, 1843]. Gel’fand-Shilov [791]. General [773, 1434, 1486, 1646, 214, 857, 394, 1730, 1619, 1453, 262, 4, 428, 501, 1584, 71, 1376, 1664, 1025, 1329, 1119, 1538, 170, 1598, 1200, 1887, 217, 1914, 23, 1055, 858, 1466, 426, 1315, 1591, 1853, 1492, 1890, 1940, 730, 409, 275, 1671, 292, 560, 1680, 1913, 1028, 1693, 1624, 1400, 1617, 1167, 771, 293, 819, 392, 117, 118, 1522, 1753, 1773, 1010, 1859, 1951, 1732, 1324, 264, 251, 1135]. Generalization [733, 1004, 1030, 1438, 768, 1172, 1539, 736, 112, 1427, 831, 1092, 147, 1440, 1044]. Generalizations [1519]. Generalized [557, 1513, 1447, 1888, 279, 837, 1701, 159, 404, 1652, 1875, 860, 1881, 492, 218, 1294, 1702, 1277, 855, 1588, 918, 63, 591, 1099, 1420, 448, 1532, 1231, 1801, 1654, 1937, 759, 1335, 408, 1867, 693, 794, 1139, 988, 951, 1425, 1604, 1197, 1894, 1389, 1234, 1007, 1653, 1166, 308]. generate [1741, 1912, 384]. generated [1673, 1748]. Generating [855, 1241, 957, 464, 422, 967, 1066, 366, 1007, 1797]. Generation [1417, 1760, 1342, 1694, 610, 129]. generator [612]. Generators [1081, 214, 1136, 1853, 100, 1840, 1655, 140]. geodesic [1830, 291, 497, 386]. Geodesics [196, 1009, 1771]. Geometric [1042, 1625, 771, 441, 520, 1570, 1224]. Geometrical [501, 1635, 1501, 1584, 335, 424, 1203]. geometries [1693, 1918]. Geometrization [962]. Geometrodynamics [989]. Geometry [1268, 1558, 535, 752, 1683, 1488, 1009, 1771, 742, 1055, 1545, 383, 1898, 734, 1120, 348, 676, 834, 1184, 1749, 1693, 1213, 1820]. Ghost [26, 1382]. Gibbs [1802, 314]. given [1720, 823, 334, 897]. Glauber [1899]. Gleason [762]. glimpse [539]. Global [189, 1073, 1691, 1215, 820, 1934]. go [1347, 1160]. Goldberg [386]. Goldstone [423]. good [1339]. Gordan [725, 1556, 1692, 553, 51, 52, 1646, 1648, 1649, 716, 1781, 435, 628, 384, 398, 397, 1354, 477, 1002, 1017, 1841]. Gordon [1623, 1331, 393, 929, 1079, 1378, 1114, 1115]. Goursat [212]. Graded [768, 785, 1368, 1609, 1281, 1853, 752, 821, 330, 505, 1561, 1237, 1483, 654, 655, 769, 1052]. gradient [556, 1539, 1411]. gradients [958]. grand [733, 1004, 941, 657, 1930]. graph [1543, 971]. graphs [784, 1575, 972, 1212]. gravidynamics [1725]. gravitating [862, 1740, 248, 1848, 596, 1240]. gravitation [1828, 1750, 324, 1193, 1288, 337, 1747, 1563, 400, 403, 1873].

Gravitational [17, 1847, 1235, 1622, 1845, 174, 1550, 839, 1810, 1916, 412, 873, 9, 171, 1419, 1765, 50, 1103, 619, 1578, 1230, 21, 67, 7, 473, 1084, 606, 1758, 1831, 1221, 258, 1078, 1506, 1510, 147, 1728, 645, 160, 1213, 1522]. gravitational-wave [50]. graviton [1457]. gravity [1226, 340, 774]. Green [202, 1954, 1210, 208, 10, 1757, 1810, 457, 65, 1302, 717, 263, 33, 142, 1905, 373, 1056, 1113, 282, 288, 1709, 1562, 853, 20, 49]. Gribov [1638]. Ground [1402, 1039, 302]. Group [987, 1565, 882, 952, 789, 1698, 1592, 388, 1777, 1743, 93, 88, 1323, 840, 434, 1397, 1628, 1793, 755, 1477, 82, 422, 296, 358, 1239, 1410, 726, 576, 1872, 1087, 1519, 956, 1699, 1774, 1647, 214, 1741, 198, 1298, 65, 498, 696, 1199, 850, 178, 841, 1162, 744, 1331, 1735, 539, 213, 393, 1278, 1194, 159, 70, 474, 1651, 1755, 1047, 1045, 435, 119, 123, 264, 265, 709, 563, 1094, 977, 251, 639, 407, 1803, 1766, 1389, 1684, 644, 1285, 1001, 1892, 558, 447, 1400, 1399, 1309, 268, 204, 1291, 844, 1017, 1353]. group-theoretic [178, 1353]. Group-theoretical [1592]. groups [25, 598, 464, 1635, 678, 722, 51, 725, 1318, 1282, 726, 576, 236, 1860, 1411, 1247, 1195, 81, 1344, 1571, 1775, 1648, 1151, 1537, 1412, 1873, 1418, 1723, 1943, 876, 1281, 1462, 1597, 1064, 1600, 827, 967, 1066, 1838, 1839, 1781, 1218, 1081, 393, 1155, 884, 1110, 631, 401, 1362, 1906, 651, 957, 183, 264, 265, 1094, 1512, 1237, 488, 477, 489, 270, 1469, 948, 1343, 1533, 211, 1841]. growth [1933]. guiding [1900]. Gurtin [440]. Gurtin-type [440]. gyrotropic [689].

H [1943, 1298, 1098]. Haag [788]. hadronic [1204, 1665]. hadrons [1076, 1735]. Half [810, 801, 689, 188, 418, 420]. half-plane [801, 689]. Half-range [420]. Half-space [810, 418]. Hamilton [1155, 1513, 905, 1156, 244, 395, 175]. Hamiltonian [200, 1845, 587, 280, 103, 301, 1629, 232, 1900, 1316, 1727, 180, 765, 1425, 1007, 1621, 1401, 968, 1617, 1167]. Hamiltonians [1605, 652, 889, 347, 1659, 1682, 1710, 1862, 1929, 653, 1082, 1186, 1038, 991, 1194, 1573, 180, 1468, 870, 1395, 764, 466, 1270, 798, 223, 1219]. Hankel [1627]. hard [1804, 1608, 559, 54]. hard-core [559]. hard-sphere [1804]. Harmonic [208, 239, 1539, 14, 1181, 97, 427, 873, 837, 1769, 1904, 79, 934, 78, 1823, 1038, 991, 1283, 931, 1046, 1048, 1047, 1045, 1617, 1626]. harmonics [1630, 1129, 1405, 700, 720, 524, 1096, 1317, 1521, 1783, 1549]. Hartree [592, 189, 233, 1264, 593]. Hawking [1500]. Heat [350, 167, 1119, 1097, 164]. heavenly [502]. Heavens [904]. Heisenberg [1402, 1169, 701, 270, 302]. helical [1272]. helicity [1630, 1162, 1351]. helium [592, 178]. Helmholtz [1165, 1359, 686]. Hermite [1461]. Hermitian [169, 768]. Hertz [1510]. Hertzian [619]. heterogeneous [230, 624, 74, 75, 289, 504, 826, 1013]. HH [1246]. hierarchy [1553]. Higgs [85, 1679, 1189, 995, 1014]. High [1257, 1276, 560, 836, 1796, 683, 292, 293]. High-energy [1276, 1796]. High-precision [1257]. high-spin [836]. Higher [1935, 1632, 324, 788, 1572, 1952, 1091, 709, 977, 223, 501, 1536, 1405, 830, 1178]. Higher-dimensional [324]. higher-order [1536]. Highest [919, 297, 1091]. highly [881]. Hilbert [87, 360, 163, 492, 570, 736, 91, 112, 698, 1100, 1386, 762, 696, 903, 1423, 979, 2, 55, 633]. Hill [1467]. Hinshelwood [1148]. Hodge [1748]. Hoffman [1679]. hole [606, 732, 519]. holes [1260, 413, 433, 1022, 1759]. holonomy [1462]. homogeneity [1187]. Homogeneous [1226, 93, 1502, 168, 1150, 439, 84, 227, 226, 1493, 1815, 1766, 1658, 1375, 1017, 1733]. homographic [1323]. Homothetic [611, 1558, 1704]. homotopy [1362]. Hooft [1168, 1217, 1879]. horizons [1618, 98, 99, 1736]. horizontally [556]. hot [1273, 1161, 284]. Hove [1346]. Husimi [1806]. Huygens [1074]. hybrid [1256]. Hydrodynamic [1449, 668, 259]. hydrogen [599]. hydrogenic [1032]. Hydrostatic [958]. hyperbolic [1186]. hyperboloid [1096]. hyperboloids [1471]. hypergeometric [938, 669]. hyperheavens [740, 998]. hyperspace [537, 535, 536]. hypersurfaces [185, 884, 1624]. hypothesis [749, 109].

I. [80]. i.e [1091]. Ideal [1448, 38]. identical [1434, 1267]. Identification [841, 1162]. identities [553, 436, 66, 186, 1853, 1217, 1879, 100, 122]. identity [1639, 1142]. II [814, 1792, 509, 1525, 420, 402, 240, 698, 1386, 1682, 1438, 720, 463, 1508, 615, 124, 438, 727, 1713, 1581, 90, 1546, 1011, 1321, 924, 266, 1230, 395, 410, 620, 1286, 1203, 817, 682, 1859, 1951, 981, 1839, 1182, 536, 803, 158, 1731, 1733, 1013, 1848, 159, 1048, 435, 1564, 1592, 131, 672, 1325, 31, 699, 1141, 794, 479, 649, 265, 589, 639, 497, 677, 76, 1354, 714, 1608, 1554, 655, 128, 526, 1885, 1611, 5, 43, 1881, 118, 1920]. III [1402, 1628, 250, 1710, 1886, 205, 934, 1436, 537, 724, 1463, 322, 1482, 132, 1326, 28, 1094, 715, 1624]. illustration [1519]. image [1631, 594]. imaginary [1162]. Imbedding [42, 1693, 1624, 356]. Impedance [993]. impenetrable [1877]. implementability [658]. imply [1700]. important [1101, 480]. imprimitivity [781, 1027]. impurity [641, 1586]. Ince [97]. incident [34]. including [943, 798]. Inclusions [961]. inclusive [531]. incoherent [369]. incomplete [723]. incorporates [543, 432]. indecomposable [1454]. Independence [1589, 1902, 1403]. independent [1087, 1218, 618, 819]. indeterminacy [1292]. index [401]. index-3 [401]. indices [709, 977]. induce [802, 803]. Induced [1063, 1430, 678, 1152, 1024]. inelastically [1178, 1179]. Inequalities [1196, 493, 662, 1215, 141, 924, 158, 724, 1407, 295, 314, 1114, 701, 1551]. inequality [690]. Inequivalent [925, 1409]. inertia [1750]. inertial [682, 1758]. Infeld [1679]. infinite [1402, 1871, 1095, 1682, 1710, 762, 427, 1304, 197, 463, 1253, 272, 914, 1081, 38, 808, 1114, 1876, 1105, 1070, 1931, 1291, 5, 43, 179]. infinite-range [272]. infinitely [772, 475, 926, 880, 900]. infinitesimal [1838, 1839, 100]. infinity [1487, 1376, 946, 1339, 1213]. influence [231, 1078]. infrared [146]. infrared-problem [146]. ingoing [1550]. inherently [1320]. Inhomogeneous [1724, 1000, 276, 1884, 1886, 349, 153, 380, 673, 136, 1227, 1927, 237, 1890, 1619, 1940, 1271, 948, 1343, 1375, 907, 999, 1885, 852, 1672]. initial [427, 1339, 1801, 1937, 562, 1638, 1490, 118]. initial-data [1339]. initial-value [427, 118]. inner [1173]. Inönü [1391]. Inönü-Wigner [1391].

Instability [890, 602, 633, 1011]. Instantaneous [1022, 1379]. instantons [1245, 1015]. integer [1453]. integrability [955, 1419, 1674]. integrable [1569, 1895, 1316]. Integral [1719, 1119, 1161, 1227, 1696, 670, 904, 1598, 1438, 1363, 1101, 549, 1572, 888, 1421, 1952, 126, 1352, 371, 298, 306, 1629, 612, 1013, 986, 1058, 1312, 1033, 1727, 500, 177, 1562, 1551, 252, 1768, 1560, 531, 605, 1408, 1814, 1467, 356, 906, 774]. Integrals [245, 370, 503, 1678, 990, 695, 283, 483, 1505, 1227, 1053, 354, 711, 1463, 1330, 1356, 408, 902, 738, 1120, 355, 1170, 1946, 1360, 735, 691, 1291]. Integration [595, 58, 1702, 1460, 1795, 859, 938, 486]. integro [600, 1383]. integro-differential [600, 1383]. integrodifferential [986, 1312]. intelligent [708, 1353, 1354]. interacting [310, 221, 1565, 198, 1320, 639, 1621, 1432, 1918]. Interaction [1185, 911, 1659, 1682, 1710, 111, 1051, 520, 1254, 226, 555, 1433, 1221, 1249, 1455, 1160, 160, 663, 1293, 164]. interactions [200, 487, 1355, 637, 1090, 1606, 272, 439, 86, 1564, 1906, 988, 656, 1819]. interchange [461]. interest [1138]. interior [1595, 881]. intermembrane [555]. intermolecular [1598]. Internal [1699, 1925, 1728, 327, 1262, 1086, 415, 348]. Internal-labeling [327]. interpolation [518, 1269]. interpolative [1174]. Interpretation [411, 1095, 1554, 454]. interval [179]. Intrinsic [1884, 734, 1886, 740, 1885]. introduction [166, 625]. Invariance [1681, 393, 1299, 1726, 276, 411, 1223, 827, 175, 1155, 228, 1180, 631, 1183, 1272, 1901]. Invariant [1869, 1223, 1003, 1053, 1693, 356, 1596, 840, 1477, 1116, 1280, 1460, 741, 1243, 1741, 65, 1442, 1904, 1681, 622, 1574, 824, 601, 548, 1721, 644, 1406, 1885, 1820, 211]. invariantly [728]. Invariants [1620, 1038, 931, 564, 262, 1861, 597, 1429, 1547, 1151, 837, 967, 1066, 1853, 1194, 1478, 1109, 563, 779, 1844, 896, 774]. Inverse [673, 623, 431, 1453, 1135, 179, 953, 1414, 249, 778, 1953, 1061, 552, 551, 1373, 1791, 318, 1796, 583, 1164, 1413, 1552, 856, 1679, 1290, 1590, 277, 30, 693, 1472, 677, 1517, 1548, 607, 1568, 8, 172]. inverse-power [1796]. inverses [218]. inversion [731, 1030, 1509, 1484]. inversionlike [1438, 1363]. investigation [1387, 1462, 883]. Investigations [884]. involutional [1858]. involving [1008, 1399, 485, 1219]. ion [1098, 1531]. ionic [197, 463]. ionization [1559]. ionized [899]. ions [760]. irradiance [609]. Irreducible [545, 257, 769, 1609, 1409, 747, 1723, 1241, 1242, 554, 841, 1162, 1218, 1512, 712, 1396, 64]. irregular [1586]. irreps [1207]. irrotational [1240]. Ising [1633, 1321, 123, 1284, 1803, 1931]. Isolating [227]. Isometries [1487, 1377, 816, 1342]. isoscalar [476]. isotropic [1226, 1047, 1493, 1815, 701, 1416]. isotropically [1927]. Isotropy [1189]. IST-solvable [1132]. iteration [1641, 1202]. iterations [233]. Iterative [593, 308, 191]. IV [210, 74, 1459, 989, 1407, 1327, 304, 1586, 405]. Iwasawa [319].

J [1942, 1935, 624, 1953, 1451, 1949, 1944, 1950, 1686, 1787, 1887, 1888, 1952, 1943, 775, 1941, 1889, 1591, 1954, 1951, 1557, 1945, 1091, 1890, 1940, 1590, 1946, 1605, 1947, 225, 1815, 1065, 1936, 1726, 915, 1814, 1938, 1939]. Jacobi [1513, 905, 1156, 280, 244, 395, 1157, 175]. Jacobian [775, 544, 352]. Jauch [777]. jet [657]. jets [1428]. Jin [910]. Jordan [1738, 1249]. Josephson [1524]. Jost [1936, 1709, 1714, 675, 1349, 1703]. junction [1914].

K.M.S. [1105]. Kac [1010, 1011]. Kadanoff [1520]. Kähler [753]. Källén [404]. Kantowski [1073]. Kato [411]. KdV [789]. Kepler [320, 1869, 872, 465, 192]. kernel [1089, 306, 177]. kernels [1090, 1729]. Kerr [137, 921, 1009, 851, 1836, 433, 1669, 383, 17, 18, 1372, 1125, 1348, 1424]. Kiehn [1065]. Killing [785, 1041, 1618, 1558, 1246, 1832, 739, 976, 141, 924, 39, 266, 617, 1154, 410, 734, 1485, 1736, 1372, 1018, 1209]. kind [908, 381]. kinematic [406]. kinematical [1387, 1546]. Kinematics [536, 1495, 1496]. kinetic [529, 1181, 138, 390, 391, 899, 1520, 1625, 810, 1883]. kinetics [109, 854]. King [972]. Kinks [362, 1115]. Kirkwood [522, 1010, 1011, 247, 521]. Klein [1623, 1331, 929, 1079, 1115]. KMS [946, 1688, 467, 1070]. Kolmogorov [106, 1440]. Komar [122]. Konhauser [1008]. Koornwinder [1684]. Korteweg [602, 1118, 1499, 830, 418, 1653, 261].

L [721, 252]. label [618, 925]. labeling [629, 327]. laboratory [163]. Ladder [1892, 890, 971]. Lagrange [645, 261]. Lagrangian [1726, 1429, 865, 540, 836, 214, 827, 1523, 622, 639, 122, 1007, 1299, 1821]. Lagrangians [1296, 415, 1278]. Lamé [1352]. laminae [1144]. Lanczos [1142]. Landau [401]. Langevin [1099]. Langlands [1634]. Langmuir [1772, 1148, 1337]. languages [1200, 1887]. Laplacian [1819]. Large [1452, 1225, 963, 1148, 1305, 1023, 886]. laser [1784, 1383, 1341]. Lattice [33, 142, 181, 114, 557, 935, 1761, 1503, 748, 484, 243, 1311, 238, 763, 1804, 1242, 775, 210, 1051, 448, 1108, 1302, 717, 811, 373, 1232, 923, 1300, 808, 1930, 447, 115, 663, 417, 352, 544]. Lattices [1128, 838, 197, 463, 282, 780, 1585]. Lauricella [1527]. law [1934, 1148, 1175]. laws [1448, 1623, 914, 1435, 789, 1155, 680, 1299, 1726, 525, 526]. Lax [1784, 1132]. layered [1711]. leading [680]. least [854]. Lee [733, 1004, 1085, 1060, 1716]. length [1257, 676]. level [665, 727, 538, 616, 216]. levels [61, 313, 620, 1039, 605]. Levinson [988, 951, 818]. Levitan [1118, 177, 1092, 1843]. Lewis [1760]. Lie [255, 505, 1620, 1861, 598, 987, 464, 785, 1515, 1323, 1837, 1793, 1368, 96, 97, 236, 702, 319, 81, 1634, 1087, 770, 1519, 741, 1384, 297, 1281, 1597, 850, 1391, 1853, 6, 399, 458, 459, 460, 744, 792, 1012, 1331, 1332, 1781, 1380, 1557, 342, 1091, 919, 752, 1478, 821, 1612, 1301, 548, 1109, 1639, 183, 330, 564, 965, 712, 571, 1561, 1237, 1483, 654, 655, 768, 769, 1652, 270, 1469, 1497, 1365, 1052, 718]. Lie-graded [1052]. Light [1426, 1078, 444]. Light-cone [1426]. like [745, 1471, 1093, 1182, 1137, 639, 764, 147, 1480, 1703, 1910]. limit [221, 1934, 336, 1802, 1606, 1384, 1558, 1677, 1826, 1148, 683, 1698, 1668, 1818, 649, 808, 1114, 1655, 1023, 809, 1291, 1346]. limiting [500, 1170, 1946]. limits [1386, 829]. line [801, 1266, 222, 301, 1901]. line-source [801]. Linear [1651, 550, 1346, 1920, 1474, 1088, 773, 1171, 1891, 1294, 1322, 143, 518, 1682, 1710, 81, 1582, 231, 578, 156, 1713, 970, 297, 1846, 1853, 1227, 1794, 1056, 719, 991, 1032, 1755, 1028, 1491, 76, 252, 1553, 1400, 1187, 1185, 1190, 107, 5, 43]. linearised [1385]. Linearization [874, 1622, 95, 699]. Linearized [1890, 1619, 1940, 402, 1163, 391, 1636, 1831]. Linked [182]. Liouville [776, 529, 1296, 414]. liquid [1932, 1252]. liquids [1252]. liquids-application [1252]. list [1462]. Local [402, 1802, 1265, 1286, 1229, 1752, 1724, 340, 658, 1110, 375, 374, 824, 1737, 11, 703, 1104]. locality [203]. localizable [643]. localization [1122]. Localized [932, 949, 1243, 754, 658, 950, 1185]. locally [1624]. logarithmic [1903]. logic [777]. logics [446, 1367]. Long [197, 463, 391, 1258, 1581, 528, 1675, 455, 559, 1825, 577, 832, 1219]. long-range [528, 455, 559, 577, 832, 1219]. Long-wavelength [197, 463]. longitudinal [1855, 1280]. look [1626]. Lorentz [1262, 1323, 434, 1397, 1628, 1394, 1344, 1925, 383, 1382, 744, 682, 885, 1182, 1149, 1651, 1755, 435, 32, 1766, 1389, 1882, 1365, 1790, 204, 1017]. Lorentz-type [1262]. Lorentzian [1461]. loss [735]. Low [115, 1918, 230, 624, 766, 564, 1608]. low-frequency [766]. Low-temperature [1918]. Lower [1442, 1188, 1256, 1039, 910, 1560, 531]. Lowering [1644, 1031, 1318]. LSZ [767]. Lubkin [805]. lumps [1763]. Lund [1519].

M [1065]. machine [1929]. Machlup [1398, 1821]. Macrocausal [767]. Macroscopic [369, 945]. Magnetic [1083, 586, 487, 1829, 1280, 1168, 245, 1756, 413, 433, 1321, 74, 75, 289, 504, 826, 1231, 1039, 488, 477, 1777, 756, 944, 798, 1841]. magnetization [1857]. magnetofluids [862, 857]. magnetohydrodynamics [730]. magnetoplasma [1531]. magnon [478, 1482, 479]. magnons [763]. Main [118]. major [1661]. Majorana [482]. Mandelstam [1217]. manifold [1374, 850, 465, 1437]. manifolds [441, 904, 723, 1461, 1057, 239, 873, 1514, 843, 1065]. many [182, 1659, 1682, 1710, 1895, 235, 234, 198, 1336, 926, 880, 900, 1663, 148, 1049, 472, 1346]. many-body [182, 198, 1336, 1346]. many-channel [235, 234, 1663]. many-component [472]. many-electron [1659, 1682, 1710]. many-particle [1895]. many-species [148]. mappings [239, 873, 622, 817, 943, 680, 1745]. maps [741, 1408, 1814]. Marchenko [1663]. Marginally [306]. Markov [1027, 1444, 1071, 1789, 1035, 1529, 797, 1036]. Markovian [679, 1439, 1919, 109, 854]. Martin [910]. Mass [1751, 1793, 361, 1660, 412, 901, 510, 1098, 1162, 1351, 1113, 902, 365, 1224, 820, 1125]. Massive [554, 1719, 1382, 895, 1141, 1139, 1495]. Massless [1114, 1229, 748, 942, 554, 1435, 407]. Master [329, 36, 434, 1628, 435, 1920]. matching [1586]. material [289, 504]. materials [74, 75, 289, 504, 826, 462, 461, 1813]. Math [1942, 1935, 624, 1953, 1451, 1949, 1944, 1950, 1686, 1787, 1887, 1888, 1952, 1943, 775, 1941, 1889, 1591, 1954, 1951, 1557, 1945, 1091, 1890, 1940, 1590, 1946, 1605, 1947, 225, 1815, 1065, 1936, 1726, 915, 1814, 1938, 1939]. Mathematical [138, 954]. mathematics [508, 509]. Mathieu [166, 625, 1314]. matrices [218, 1294, 1397, 1676, 956, 1741, 379, 180, 574, 1840, 1518, 241, 1488]. Matrix [964, 470, 898, 1838, 1839, 1708, 1220, 1883, 140, 1616, 670, 702, 235, 234, 552, 1579, 767, 664, 1582, 1787, 307, 888, 1287, 155, 1298, 1943, 134, 90, 1858, 1859, 1951, 1169, 1781, 70, 474, 1458, 1651, 1755, 259, 1353, 1766, 812, 675, 1892, 1811, 278, 1069, 1877, 1017, 83, 1559, 892, 1858]. matter [729, 73, 124, 1362, 1665]. maximal [571]. maximum [470, 1927, 1865, 149]. Maxwell [874, 790, 975, 743, 1829, 92, 1657, 942, 1623, 68, 1059, 1342, 1211, 1911, 1912, 710, 1231, 1801, 1654, 1937, 980, 981, 1436, 1459, 759, 17, 18, 973, 1528, 1019, 145, 368, 737, 331, 786, 1417, 454]. Mayer [1465, 1722, 77]. mean [151, 996, 683, 854]. means [1294, 1283, 812, 605]. measurable [1188]. Measure [613, 1223, 512]. Measure-theoretic [613]. measurement [1095]. measurements [584, 1764]. measures [762, 1613, 359, 903, 1843, 491]. Mechanical [1490, 1181, 91, 1247, 551, 662, 321, 1603, 561, 1045, 378, 979, 1520, 1137, 1553, 1554, 316, 317, 1191, 1825]. mechanics [584, 781, 987, 464, 755, 1095, 370, 1296, 1845, 1198, 1465, 646, 207, 692, 1422, 1752, 788, 1444, 335, 168, 782, 139, 46, 1507, 437, 438, 318, 941, 86, 1662, 390, 1576, 895, 1203, 622, 719, 1169, 1233, 1308, 532, 806, 1431, 1564, 1592, 323, 870, 1708, 765, 1901, 339, 799, 135, 797, 691, 974, 1097, 286, 16, 992]. mechanism [1679]. media [1584, 230, 624, 1536, 1678, 1812, 609, 1643, 583, 623, 237, 337, 49, 1711, 107, 1327]. media-cluster [609]. medium [48, 600, 667, 1117, 961, 430, 689, 1552, 1475, 77, 331, 19, 74]. membrane [1286]. membranes [1753, 1792, 1433]. memory [614]. Menten [1148]. merits [342]. Meromorphic [1895, 235, 234]. merons [1851]. meson [1435, 1037]. Mesonic [1392]. Meta [825]. Metastability [746]. Metastable [1931, 1857]. method [150, 1250, 112, 1361, 983, 1129, 1405, 518, 1659, 1682, 1710, 591, 1804, 1712, 74, 267, 909, 1314, 154, 298, 545, 1858, 1859, 1951, 1856, 424, 856, 1817, 1675, 1255, 277, 30, 1132, 1358, 191, 264, 268, 1214, 83, 860, 1881]. methods [1707, 930, 253, 60, 1167]. metric [1235, 137, 174, 490, 1447, 1888, 1874, 1669, 1830, 1683, 497, 1695, 1938]. metric-torsion [1235]. metrics [170, 502, 1704, 1154, 410, 1690, 948, 1343, 454, 294, 1209, 117]. Metzner [1872]. Michaelis [1148, 1157]. microcanonical [273]. Microscopic [1607, 1608, 369]. Micu [1861]. Micu-type [1861]. migration [1857]. Mills [1547, 1835, 1226, 742, 457, 1681, 758, 1528, 1015]. Minimal [222, 184, 209]. minimum [224, 41]. Minkowski [720, 1681, 1694, 1096, 1831]. Minkowskian [1799]. Mirror [1102, 1929]. misalignments [1536]. missing [242, 618, 925]. Miura [1689]. mixed [1585]. mixings [178]. mixture [663]. mixtures [138]. modal [1711]. mode [197, 463, 1341, 813]. model [1633, 668, 748, 271, 983, 1565, 882, 1511, 1519, 788, 665, 1822, 350, 167, 1243, 760, 1734, 1340, 1686, 657, 871, 126, 138, 475, 1597, 1806, 1364, 1148, 1157, 421, 1544, 1636, 1316, 1139, 1080, 123, 1284, 579, 1926, 1085, 1060, 128, 365, 810, 1381, 1580, 656, 1931, 1261, 472, 842, 1899, 392, 1716, 194]. modelocking [1383]. Models [200, 508, 509, 290, 1226, 1073, 1805, 787, 1722, 745, 876, 996, 1380, 1557, 1665, 590, 1924, 148, 1658, 1625, 1497, 302]. modes [1550, 1603, 101]. Modified [1711, 602, 1572, 1952, 1352, 136]. Modulational [602]. molecular [557, 545]. molecule [422]. Molien [1066, 967, 966]. Moment [848, 1804, 314, 1633, 487, 668, 91, 1040, 1292, 212]. Moment-method [1804]. Moment-theory [848]. momenta [316, 317]. moments [150, 1536, 91, 1912, 1244, 79]. momentum [963, 1740, 586, 235, 66, 269, 1435, 1859, 1951, 1731, 1688, 1456, 686, 944, 1453, 1135, 83]. Monopole [864, 1168, 1231, 1172, 1549]. Monopoles [742, 1880, 1083, 1756, 1015]. Monotonic [370, 1062]. Monotonicity [530]. Monroe [522]. Morse [154]. most [1914]. Motion [23, 1268, 1448, 245, 901, 1705, 103, 622, 817, 934, 227, 1431, 1564, 1612, 639, 1272, 1844, 1209]. motions [862, 611, 1704, 1786, 1945, 884]. mound [1075]. moving [357, 997, 1746, 1947, 49]. MR [1451, 1949, 1944, 1950, 1952, 1889, 1951, 1945, 1940, 1953, 1787, 1557, 1946]. Muhlerin [80]. Multi [1093]. Multi-soliton-like [1093]. Multichannel [2, 87, 1616, 577, 832]. Multidimensional [1222, 1363, 1919, 1337]. multifrequency [1768]. Multigroup [419, 420, 507, 429, 721, 1202, 685, 1807]. Multimode [814]. multiparticle [366, 1909, 1401]. Multiple [1148, 1157, 442, 1271, 113, 634, 666, 1427, 470, 1244, 1050, 738, 907, 751]. multiple-scales [1050]. multiple-scatter [907]. Multiplet [554]. Multiplicative [608, 63, 205, 1566, 1939]. Multiplicities [1647]. Multiplicity [1454, 1390, 470, 366, 648, 1893]. Multiplicity-free [1390, 648, 1893]. multipliers [1872, 1842, 645]. multiplying [106]. multipoint [518]. Multipole [471, 1586, 380, 1486, 1912, 1441, 1416]. multipoles [700, 720]. Multiquark [876]. multirooted [1575]. multishell [630]. multispecies [1546]. multispinor [1278].

Naimark [498]. Nambu [765]. Natural [1306, 1603]. nature [1687]. Navier [982]. near [1260, 206, 582, 958, 1043]. nearest [121, 922, 923, 663, 417]. necessary [1800, 1074, 802]. negative [231, 1079, 1141, 1907]. neighbor [763, 121, 922, 923, 1124, 663, 417]. neighbor-pair [1124]. nematic [1932, 1252]. net [389, 1415]. network [856]. networks [345, 646]. Neumann [224, 776, 1229, 1277, 1582, 41, 414, 135]. neutral [649]. Neutrino [1829, 1587]. Neutrinos [27, 575, 26]. Neutron [496, 1117, 1550, 1055, 237, 363, 1260, 106, 419, 420, 793, 1273, 1161, 867, 1201, 1202, 685, 1807, 300, 346, 1530, 1184, 1540, 1381]. neutron-transport [1381]. Newman [137, 1009, 851, 1836, 1669]. Newtonian [1329, 1025, 1750, 196, 719, 1102, 639]. Newtonian-like [639]. next [922, 663]. nilpotent [1380, 1557]. No [1249, 1953, 1451, 1949, 1944, 1950, 1787, 1952, 1347, 1889, 1591, 1351, 1951, 1557, 520, 1945, 1890, 1940, 1946, 1160, 1814, 1938, 1939]. no-go [1347, 1160]. No-interaction [1249, 520]. nodal [490]. Noether [1429, 1844]. noise [1846]. noises [614]. non [1761, 1923, 1439, 1919, 1193, 342, 1192, 1289, 1123, 1185]. non-Abelian [1761, 1923, 1192, 1289]. non-Lie [342]. non-linear [1185]. non-Markovian [1439, 1919]. non-standard [1123]. non-static [1193]. Nonbijective [1233]. Noncentral [951]. noncommutative [1789, 532, 806]. noncommutativity [1902]. noncompact [319, 1634, 1064]. Noncompletely [1389]. nonconservative [335, 1868]. nonconstant [1154]. Noncontinuous [1756]. noncovariant [1921]. nondegenerate [1173]. nondiverging [185, 1833]. nondynamic [1320]. nonequilibrium [469, 797, 205]. Nonexistence [760, 148]. nonhomogeneous [48]. noninteger [1852, 928, 1453, 1135]. nonintegral [1128, 838, 1424]. noninteracting [836, 880]. noninteraction [1564]. noninvariance [1047, 1520]. Nonlinear [108, 569, 1344, 323, 1310, 713, 714, 1568, 1718, 1942, 367, 1481, 246, 1707, 1784, 791, 1072, 1525, 668, 190, 1934, 670, 1641, 855, 1895, 1030, 1438, 1588, 1785, 1950, 1457, 1767, 425, 615, 111, 439, 1021, 914, 754, 60, 267, 909, 371, 325, 1251, 789, 1155, 856, 1573, 322, 960, 1868, 1132, 694, 794, 1562, 1493, 1815, 567, 1926, 45, 812, 621, 1844, 1763, 525, 526, 1717, 1248, 37, 932, 949, 950, 12, 55, 633, 1337, 1398, 261]. nonlinearly [231]. Nonlocal [988, 1610, 1611, 3, 8, 24, 44, 172, 405]. nonlocalizable [203]. nonmetric [1053]. Nonmetrical [312]. nonnegative [848]. Nonorthogonal [459]. nonperturbative [1734]. nonpolynomial [415]. nonrelativistic [584, 755, 276, 782, 308, 1546, 1698, 659, 974]. nonrigid [1762]. nonrotating [1161]. nonseparable [1637]. nonsimply [377]. nonsingular [737]. nonspin [1677]. Nonspreading [635, 349, 153]. Nonstandard [897, 46, 649]. Nonstationary [751, 667, 126]. nonstrictly [643]. nonsymmorphic [726, 576]. nonsynchronous [1150]. Nontranslationally [1136, 1287]. nontrivial [1880]. nontwisting [1833]. Nonuniqueness [778, 1953, 1373, 540]. nonunitary [161]. nonvacuum [684]. nonzero [1351]. Nordström [1771, 732]. norm [1867, 1334]. Normal [1929, 371, 197, 463, 332, 452, 1830, 1426, 1788, 1190]. Normalization [1319, 507, 1870, 1300, 1644]. Note [863, 379, 1755, 1529, 1177, 1695, 1938, 1689, 1641, 729, 1369, 738, 1371, 1080, 183, 1854, 1790, 135]. Notes [1016]. notion [203, 289, 504]. notions [1071]. Novel [66, 1174]. nuclear [930, 1098]. Nuclearity [1430]. nucleus [1565, 516, 882]. Null [1339, 1654, 291, 497, 386, 1487, 1376, 185, 1200, 1887, 884, 145, 1624, 1365, 681]. null-plane [1365]. Number [618, 914, 1023, 1261, 893]. numbers [104, 527]. Numerical [484, 1581, 1693]. NUT [147]. NUT-like [147].

object [1631, 868]. objects [84, 881]. observable [1382, 109]. Observables [104, 163, 946, 987]. observations [978, 1778]. observer [1758]. obstacles [113]. obtained [791]. occupation [557, 780, 1585]. occurs [1383]. octahedral [1754]. Octonionic [696]. Off [675, 1610, 1611, 1910, 1676, 424, 912, 1349, 1703]. Off-energy-shell [1610, 1611, 912]. Off-shell [675, 1910, 1676, 1349, 1703]. One [632, 940, 263, 1124, 1877, 1402, 917, 1589, 195, 1054, 173, 1304, 1253, 1206, 1919, 1421, 1869, 796, 1176, 1650, 1098, 859, 1050, 1817, 685, 923, 1300, 215, 875, 573, 177, 1092, 123, 1360, 1299, 1726, 1381, 54, 605, 663, 417, 1398, 1480, 75, 1930]. One- [75]. One-component [632, 940, 796, 1176]. One-dimensional [263, 1124, 1402, 1589, 195, 1054, 173, 1304, 1421, 1869, 1650, 859, 1050, 685, 923, 1300, 875, 177, 1092, 123, 1360, 54, 605, 663, 417, 1398, 1930]. one-field [1817]. one-speed [1381]. only [1098, 1889, 1614]. Ono [1685, 1093]. Onsager [259, 1398, 1821, 1488]. operational [1729]. Operator [290, 1122, 1368, 1534, 1613, 1902, 641, 578, 69, 429, 498, 841, 1162, 1351, 867, 1201, 1006, 247, 1807, 322, 256, 388, 1446, 1663, 414, 776, 820, 1883, 1179, 8, 172, 860, 1881, 802, 803]. operator-valued [1613, 578, 867, 1201]. operators [814, 25, 1935, 1632, 1294, 88, 91, 1031, 1319, 1318, 964, 1062, 690, 496, 664, 1582, 1787, 747, 1723, 279, 1754, 136, 104, 1466, 1492, 1838, 1839, 1908, 1909, 1865, 1034, 813, 875, 821, 1301, 353, 651, 957, 1721, 1818, 1109, 618, 1604, 1876, 925, 252, 652, 466, 559, 1892, 339, 1788, 327, 211, 1596, 1644, 577, 832, 372, 608]. Optical [220, 1338, 1410, 642, 287, 594]. optics [501, 1584, 1784, 1631, 89, 424, 1811]. Optimal [726, 576]. orbit [56, 1164]. orbitals [1453, 1135]. Orbits [956, 57, 193]. order [1718, 1942, 1429, 1428, 501, 668, 1536, 1280, 1656, 804, 1572, 1952, 578, 156, 1713, 134, 1501, 1383, 39, 266, 326, 1148, 599, 299, 228, 1431, 830, 938, 1115, 1924, 110, 1768, 1777, 1745, 1178, 5, 43]. ordered [1404, 1902]. Ordering [814, 813, 695, 1252]. orders [1732, 1733]. ordinary [578, 1842, 1227]. origin [1313]. origins [1822]. Ornstein [493, 195, 1080]. Orthocomplemented [1173]. Orthogonal [240, 458, 1318, 1151, 1601, 651, 957, 1553, 844, 211]. Orthonormality [278]. orthorhombic [717]. orthosymplectic [1853]. oscillations [935, 60, 267, 1126]. oscillator [1397, 1640, 1948, 208, 1181, 190, 1361, 358, 97, 1174, 837, 1769, 1904, 154, 79, 934, 78, 1823, 1038, 991, 1574, 931, 1046, 1048, 1047, 1045, 1668, 1670, 1137, 1745, 1626]. oscillator-type [1038]. oscillators [1538, 427, 481, 1292, 394, 313, 620, 1928, 1617, 176]. oscillatory [1516]. osp [769]. other [1138]. outer [1399].

Padé [570, 143, 162, 499, 58, 542, 1516]. Painlevé [908]. pair [86, 1124, 1121, 1159, 580]. pair-correlated [1121, 1159]. Pairing [1023]. Palatini [1214]. Para [1470, 1369, 1889, 1614]. Para-Bose [1470, 1889, 1614]. para-Fermi [1369]. parabolic [1707, 1634, 1186, 506]. parabose [69]. Paraboson [900]. paradox [866]. parallel [287, 288, 1274]. parameter [731, 50, 884, 1687, 1188, 1424]. parameters [660, 750, 331]. parametric [647, 439, 938]. parametric-decay [439]. parametrically [1928]. parametrization [269]. Parasiuk [647]. parity [1273, 1161, 1914]. Partial [552, 835, 1173, 1895, 1030, 1438, 1516, 1258, 1581, 1378, 567, 714, 152, 395]. Partial-wave [835]. partially [1404]. Particle [1606, 630, 200, 1434, 1534, 1280, 1895, 1699, 1925, 807, 1340, 1686, 1067, 86, 1082, 1382, 357, 103, 622, 817, 1039, 1673, 1903, 661, 102, 1279, 1877]. Particle-boson [1606]. particle-field [1082]. particlelike [1621]. Particles [59, 1134, 1793, 1267, 1802, 640, 161, 956, 869, 1798, 1243, 198, 871, 308, 1546, 1567, 895, 1433, 902, 512, 923, 780, 1585, 527, 639, 1495, 880, 1085, 1023, 828, 861, 995, 1497, 1580, 1106]. particular [1896, 1227]. partition [733, 1004, 657, 1930]. Path [1678, 695, 1057, 126, 298, 1505, 859, 1356, 1727, 1088, 503, 283, 483, 543, 371, 1629, 711, 1013, 1463, 1330, 408, 355, 500, 1170, 1946, 1768, 691, 1087]. Path-integral [126, 1727]. Pathria [448]. paths [889, 151]. patterns [1934]. Pauli [105, 404]. PCAC [613]. penetrable [766, 84]. Percolation [784, 1575, 121, 1805, 1445]. Percus [1598, 1075, 326]. Perfect [680, 1273, 1161, 839, 376, 1084, 35, 999]. perfect-fluid [1273]. periodic [230, 624, 1505, 875, 1425, 351, 113, 261]. Periods [843, 1065]. permittivity [556]. permutation [630]. perpendicular [689]. Perturbation [1075, 1113, 1153, 166, 625, 736, 1181, 1616, 112, 425, 416, 1174, 1501, 216, 599, 78, 732, 20, 915, 761, 5, 43]. perturbation-Theoretical [166, 625]. perturbations [1934, 1273, 1161, 1914, 606, 258, 877, 800]. Perturbative [91, 1598, 326, 336, 1823]. Perturbative-variational [91]. perturbed [1862, 1602, 154, 1050]. Petrie [218]. Petrov [955, 1704, 883, 129]. Phase [120, 549, 888, 895, 1203, 1170, 1946, 781, 1743, 1134, 250, 692, 1656, 1439, 1404, 483, 1572, 1421, 1796, 1952, 745, 1011, 281, 56, 247, 521, 1737, 1358, 765, 491, 661, 1495, 1496, 1043, 1777, 461, 605]. Phase-integral [888, 1421, 605]. Phase-space [895, 1203, 692]. phenomena [1049]. phenomenological [444]. photoelectric [1032]. photoelectrons [1586]. photon [173, 1304, 1298, 1943, 1496, 1049]. Phys [1942, 1935, 624, 1953, 1451, 1949, 1944, 1950, 1686, 1787, 1887, 1888, 1952, 1943, 775, 1941, 1889, 1591, 1954, 1951, 1557, 1945, 1091, 1890, 1940, 1590, 1946, 1605, 1947, 225, 1815, 1065, 1936, 1726, 915, 1814, 1938, 1939]. Physical [205, 1138, 1804, 89, 1482]. physics [508, 509, 1752, 545, 920, 1928, 1362, 978, 264, 265, 1094]. picture [702, 1507]. piezoelectric [1933]. pinch [1636]. pion [214, 707]. Piron [777]. planar [115, 114, 113]. Plane [1828, 786, 790, 1866, 801, 748, 1387, 26, 1246, 1055, 1578, 689, 424, 34, 1274, 1365, 1820, 1518]. plane-fronted [1820]. plane-symmetric [26]. planes [1102]. planetary [193]. plasma [1855, 865, 668, 1829, 632, 940, 439, 796, 1176, 1732, 1733, 1890, 1619, 1940, 1928, 1126, 877, 800, 607]. plasmas [1355, 424]. plate [853, 1210]. plates [287]. plet [1660]. plethysms [1399]. plus [936, 241, 892, 893]. Pochhammer [1044]. Poincaré [1206, 1743, 93, 296, 1247, 161, 1412, 696, 841, 1162, 1331, 1735, 213, 264, 563, 251, 639, 1389, 1820]. Poincaré-invariant [1820]. point [1268, 125, 320, 1534, 1571, 1223, 1673, 191, 1512, 489, 1819]. point-particle [1673]. points [913, 958]. Pointwise [1446]. Poisson [674, 1208, 718]. polar [1629]. polarizabilities [1234]. Polarization [963, 1347, 714, 1279]. polarized [556]. pole [1370]. poles [1250, 582, 130, 131, 132]. Polish [1195]. poloidal [686]. polycrystalline [1813]. polygon [357]. Polynomial [1547, 1571, 1512, 1361, 1241, 1573, 1467]. polynomials [97, 240, 1242, 1601, 972, 1008, 1684, 1816, 718]. polyomino [972]. population [106]. position [1351]. Positive [1907, 346, 834, 1040, 231, 538, 313, 426, 749, 1079, 1141]. Positivity [1700, 1315, 1591]. possessing [250, 926]. possible [1880, 342, 337, 1381, 881]. Post [1025, 1329, 1750]. post- [1750]. Post-Newtonian [1329, 1025]. potential [1171, 1526, 494, 913, 1598, 1676, 206, 1798, 9, 171, 894, 757, 1075, 1010, 1011, 326, 1603, 1257, 1039, 1831, 455, 936, 875, 177, 1903, 1491, 102, 675, 1338, 1191, 969, 1069, 1393, 1610, 1611, 605, 949, 984, 893, 1479]. potentials [166, 625, 249, 1640, 1948, 1835, 975, 731, 449, 849, 1253, 1756, 139, 1572, 1420, 1421, 1796, 1952, 1258, 1581, 89, 619, 976, 528, 56, 1164, 21, 455, 130, 131, 132, 431, 1843, 1824, 603, 1358, 951, 649, 1189, 1510, 1472, 1714, 1936, 756, 559, 547, 627, 1825, 3, 8, 24, 44, 172, 405, 577, 832, 485, 1219, 241, 892, 1480, 1910]. Potts [1511, 1805]. power [1796, 1824, 1594]. precision [1257]. prediction [178]. Predictive [1576, 1730, 1731, 1431, 1564, 1901]. Preliminary [36]. presence [173, 1304]. preserve [1459]. pressures [579]. Presymplectic [1514]. primitive [1645]. Principal [681]. principle [280, 692, 411, 86, 90, 224, 41, 1074, 344, 1118, 1333, 1927, 1865, 469, 1180, 284]. principles [1428, 518, 1196, 186]. probabilities [557, 1403, 1445]. Probability [491, 453, 918, 1789, 1313, 1120, 274]. problem [953, 1414, 249, 1750, 1633, 865, 91, 778, 1953, 320, 1934, 336, 182, 516, 189, 1808, 1297, 1054, 692, 552, 551, 1089, 1373, 1791, 9, 171, 901, 1705, 73, 124, 318, 1869, 1804, 1021, 583, 1292, 1662, 872, 1904, 281, 1164, 1122, 306, 1413, 1552, 1927, 301, 859, 802, 803, 920, 612, 1544, 285, 435, 1679, 300, 1290, 1590, 431, 693, 28, 212, 191, 346, 146, 1360, 1854, 1504, 1548, 1624, 1825, 607, 144, 192, 1490, 842, 8, 172, 117, 118]. problems [1025, 1329, 917, 1631, 668, 1361, 1040, 427, 206, 416, 629, 909, 623, 1413, 1552, 1359, 188, 1356, 228, 677, 1517, 363, 1521, 1687, 969, 49, 697, 939, 304]. procedure [1516, 500, 1170, 1946]. process [1152, 1919, 759, 676, 797, 1398, 1821]. processes [125, 614, 329, 220, 1444, 63, 205, 1145, 562, 759, 1849, 1077, 1566, 1939]. produced [1796]. Product [990, 1173, 1461, 1344, 1869, 328, 127, 476, 1001, 1123, 820]. production [1529, 1328]. products [762, 364, 1081, 1426, 1708, 1285, 1533]. profile [1072, 1525]. Projected [1396, 1518]. projection [322]. Projective [1770, 1323, 236, 1860, 1063, 1131, 387, 1818, 489]. prolate [381]. Prolongation [1772, 615, 693, 694, 794, 12, 1786, 1945, 831, 1139]. Proof [1098, 11, 703, 1682, 1710, 757, 1351, 520, 686, 1659]. propagating [101, 560]. Propagation [600, 292, 19, 293, 1536, 1752, 673, 1643, 101, 1544, 440, 1208, 388, 107]. Propagator [715, 78, 1727, 1670, 1594]. propagators [1500, 1270]. proper [1758]. Properties [1535, 1685, 297, 1107, 407, 917, 592, 91, 1062, 728, 1659, 1682, 1710, 946, 580, 1691, 807, 1851, 653, 1546, 739, 448, 68, 1223, 1898, 357, 74, 75, 289, 504, 826, 1908, 1909, 1039, 1034, 387, 631, 1120, 1665, 899, 1301, 440, 1197, 1894, 809, 278, 607, 1453, 1135, 302]. property [1810, 1071, 643, 1160]. proposition [116]. proposition-state [116]. propositions [811]. Pseudo [614, 548, 1309]. pseudo-Riemannian [1309]. Pseudoparticle [933]. pseudopotentials [533, 770]. Pseudoscalar [1259]. pulse [1643, 1544]. pulses [1451, 1111, 1862, 947, 1795]. pump [1812]. Punctual [1516]. pure [1673]. pure-radiation [1673]. pursuit [1866]. Pyramidal [187].

Quadratic [1194, 814, 1727, 813, 1270, 1617, 223]. quadripotential [1673]. quadrupole [516, 193]. quanta [1351]. quantal [888, 1420, 1823, 812]. quantisation [757]. quantities [1376, 1664, 406]. Quantization [626, 572, 1296, 692, 1228, 1437, 246, 845, 1005, 1666, 1642, 1734, 1420, 1421, 1869, 1742, 308, 1546, 1471, 371, 325, 843, 1637, 511, 465, 1065, 176, 1261, 1432]. quantizations [402]. quantized [1691, 871]. quantizing [983]. Quantum [1871, 1451, 646, 723, 1862, 1511, 168, 662, 947, 313, 620, 903, 1423, 1850, 978, 527, 661, 1077, 1440, 1416, 310, 36, 584, 781, 987, 221, 1134, 755, 453, 125, 1095, 1181, 370, 91, 456, 975, 112, 250, 1534, 276, 207, 983, 1247, 1808, 1100, 1386, 329, 679, 692, 1422, 551, 173, 1444, 1304, 450, 927, 446, 1809, 1916, 782, 1384, 139, 46, 1507, 198, 1677, 1661, 1662, 1795, 1035, 1382, 394, 253, 321, 1846, 604, 61, 1603, 1600, 561, 895, 1203, 1233, 1849, 1308, 1045, 1592, 378, 32, 288, 323, 824, 1147, 979, 1305, 866, 870]. quantum [1520, 1562, 351, 1114, 1495, 1655, 777, 1060, 339, 128, 316, 317, 1191, 1328, 274, 1403, 1778, 1049, 1825, 11, 703, 37, 135, 797, 691, 974, 286, 1178, 1179, 423, 445, 1036, 992, 1626, 774]. quantum-mechanical [561, 1520, 316, 317, 1191]. quantum-statistical [173]. quark [1570]. Quartic [1639, 313]. Quasiclassical [573]. quasilinear [1378]. Quasiprobability [1846]. quasispin [495]. Quasisteady [1645]. question [1796].

R [1952, 1065]. Racah [213, 387, 254, 1893]. Radial [735, 702, 1905, 1954, 1714, 1936, 381]. Radiating [1669]. Radiation [1307, 636, 1260, 199, 640, 173, 1304, 1730, 1731, 1222, 17, 34, 1673, 229, 1178, 1179]. radiative [1119, 216, 1927, 751]. Radon [1509]. Raising [1318, 1644]. Raman [1812]. Random [1439, 299, 1187, 48, 600, 1536, 1181, 1451, 1802, 1678, 1791, 1130, 947, 1795, 74, 75, 289, 504, 826, 1806, 1364, 1475, 180, 574, 331, 1121, 1159, 19]. random-axis [1536]. random-walk [1364]. randomly [1050, 907, 1190]. randomness [1095]. range [1526, 419, 272, 1258, 1581, 528, 1675, 455, 431, 1080, 559, 1825, 423, 577, 832, 1219, 892, 420]. rank [1754, 133, 315, 1291]. Rarita [1836]. rational [1481, 150, 241, 892]. Ray [287, 55]. Ray-optical [287]. reaction [1934, 1148, 1157, 229]. reaction-diffusion [1148, 1157]. reactions [930, 945]. real [1634, 1384, 1391, 564, 965, 571, 1252]. realization [1397, 1365]. Realizations [276, 598, 1344, 1767, 251, 639]. reciprocal [906]. Reciprocity [833]. Recoupling [1400, 1399, 1371]. rectangular [115, 114]. recurrence [854]. recursion [1088, 773, 1803]. recursive [316]. Rederivation [910]. reduced [1138, 263, 1905, 1954, 1859, 1951, 1877, 1179]. reducibility [878]. reducible [1389, 1018, 159]. Reduction [1926, 1671, 1285, 480, 434, 261]. reductive [654, 655]. reference [1923, 1758]. reflection [1878, 1587, 1764]. reflections [885, 1182]. reflectors [1530]. reformulated [1263]. Regge [1370, 1250, 1519, 894, 307, 1204, 130, 131, 132]. region [1878, 512, 968]. regions [556]. Regular [1501]. regularity [592]. Regularization [1112, 791, 1835, 1516]. regularized [1719]. reinforced [462]. Reissner [732]. related [1640, 1948, 767, 244, 395]. relation [195, 1199, 686, 3, 44, 405, 706]. Relations [1653, 1088, 773, 195, 1486, 1409, 1320, 1803, 833, 774]. relationship [507, 1155, 461]. relative [1834, 1867]. Relativistic [634, 868, 1263, 1495, 1580, 1584, 246, 1273, 698, 1374, 857, 901, 1705, 217, 1576, 334, 1122, 895, 1203, 1735, 730, 1555, 1431, 1564, 1324, 1325, 1326, 1327, 639, 407, 45, 828, 1778, 11, 703, 878, 798, 881, 193, 1716]. Relativistically [937, 141, 924]. Relativity [1469, 4, 428, 501, 71, 1376, 1664, 1025, 1329, 1266, 1200, 1887, 23, 858, 426, 1315, 1591, 409, 275, 1680, 1913, 1624, 771, 819, 117, 118, 1522]. relaxation [173, 1304, 1919, 1416]. relaxing [728]. remark [1761, 1035, 1036]. Remarks [585, 1168, 1787, 804, 1629, 282, 1061, 102, 698]. Removal [490]. Renormalization [840, 123, 447, 539, 1803, 194, 608]. renormalizations [365]. Renormalized [912, 1518, 475, 322, 1276, 1879]. reobserved [232]. representability [1741, 802, 803]. Representation [1860, 941, 65, 1781, 1294, 1402, 722, 1322, 553, 863, 647, 1719, 82, 1538, 517, 1195, 47, 481, 664, 1582, 1787, 581, 578, 279, 871, 308, 1546, 334, 1369, 841, 1162, 895, 1909, 1233, 226, 70, 474, 1458, 978, 1137, 1553, 1554, 1396, 1811, 1883, 844, 1044, 53, 372]. Representations [1889, 1614, 1735, 374, 644, 1744, 1497, 268, 781, 1743, 1474, 1430, 1635, 678, 1770, 916, 434, 1628, 1609, 236, 1860, 1040, 161, 1063, 1774, 1647, 1409, 1418, 1242, 69, 297, 1454, 498, 815, 1204, 613, 364, 554, 1227, 1218, 1667, 1024, 159, 1651, 1755, 1045, 323, 375, 183, 709, 977, 712, 1389, 1684, 489, 1285, 768, 769, 1001, 534, 64, 1453, 1135, 1533]. Repulsive [1739, 1796]. reservoir [727, 616, 1295]. Resistance [701]. Resolution [1729]. Resolvent [1702, 968, 860, 1881]. resonance [290, 1152, 22]. Resonances [1650, 245, 1338]. resonant [344, 1478]. Resonantly [367]. resonator [877, 800]. respective [140]. response [1133, 1941, 1032, 713, 714, 1711, 1490, 1346, 1920]. responses [1732, 1733]. rest [40]. Restoration [1897, 1631]. restricted [1918]. restrictions [170]. result [733, 1004, 1794]. resulting [1560]. results [36, 763, 1537, 198, 1232, 1867, 1191, 1610, 1611]. resummation [15, 714]. Retarded [380, 10, 202, 1524]. retrieval [281]. reversal [1520]. reverse [662]. Reversible [116]. revisited [299, 1626, 1346, 1920]. Riccati [812]. Ricci [943]. Rice [1044]. Riemann [436, 551, 49, 774]. Riemannian [1374, 239, 873, 386, 1309]. Rigid [1786, 1945, 1357]. Rigorous [522, 1232, 531]. Robinson [1177]. rocksalt [52]. rod [676]. rods [54]. role [1821]. roots [1488]. rotated [968, 29]. Rotating [858, 217, 141, 924, 35, 1758, 1820, 937, 294]. rotation [1357, 1762, 868, 956, 268]. rotational [49, 1467]. rotationally [1741, 1406]. rotator [748]. Rowlinson [1930]. Roy [1112]. Ruch [469]. Ruelle [643]. rule [11, 703]. rules [1239, 187, 47, 503, 695, 1409, 642, 628, 1840].

S [1521]. Sachs [1872, 1073, 386]. Sagnac [71]. Salpeter [361, 524, 29]. Salsburg [1010, 1011, 247, 521]. satisfied [731]. satisfying [1909]. Sato [582, 412, 581, 687, 1125, 1348, 1424]. saturation [1934, 1148]. Saxon [675, 1069]. Scalar [1245, 449, 1809, 1916, 1037, 1229, 169, 1829, 349, 153, 1896, 1486, 1216, 942, 475, 895, 1435, 1006, 257, 1493, 1815, 786, 1747, 519, 400, 338, 1121, 949, 1185]. scalar-tensor [1747, 400, 338]. Scalar-vector [1245]. scalars [638, 1207]. scale [1826, 442]. scales [1050]. Scaling [1272, 1668, 567, 57]. scatter [907]. scattered [1178, 1179]. scatterers [766, 1121, 1159]. Scattering [166, 625, 48, 1798, 1601, 1475, 519, 1219, 179, 1526, 1640, 1948, 1434, 634, 667, 494, 1119, 1706, 1616, 345, 1061, 1100, 1386, 59, 552, 1089, 1656, 1373, 411, 782, 1253, 1812, 1516, 1258, 1581, 1427, 1055, 1545, 136, 528, 1603, 1164, 835, 623, 1927, 84, 1908, 1909, 424, 1336, 856, 1731, 685, 455, 1764, 601, 573, 693, 177, 1160, 1906, 1358, 1472, 677, 2, 87, 661, 812, 1697, 1517, 652, 466, 1401, 1360, 968, 969, 1580, 1610, 1611, 114, 113, 444, 971, 751, 984, 1568, 1044, 83, 577, 832, 485, 912, 1153, 1672, 1910, 1402]. scheme [1174]. Schild [921, 383]. Schrödinger [1526, 58, 637, 913, 347, 702, 1588, 1822, 990, 1021, 105, 1795, 244, 395, 1199, 325, 1251, 789, 1751, 132, 180, 875, 500, 794, 1092, 1709, 1358, 102, 621, 525, 526, 1097, 568, 1672]. Schwartz [1604]. Schwartzschild [497]. Schwarzschild [1938, 1216, 1595, 519, 1695]. Schwinger [1040, 1836, 1864, 590, 1819]. screened [1257, 627]. screening [1700, 1257, 62]. screw [1636]. Search [1425]. Second [246, 134, 871, 308, 1546, 668, 1280, 763, 121, 804, 578, 156, 1713, 1383, 561, 228, 401, 938, 1777, 381, 1536]. Second- [1536]. second-neighbor [763]. second-order [668, 156, 1713, 1383, 228, 1777]. Second-quantization [308, 1546]. Second-quantized [871]. Section [1, 120, 1117, 1456, 292, 560, 225, 1381, 1797, 1188, 293]. sections [1307, 207, 1731, 1696, 531, 627]. secular [425]. Segal [42, 1742, 1794]. Segrè [617]. Self [221, 1130, 248, 1848, 353, 596, 1240, 862, 1535, 1779, 1949, 1152, 1865, 1034, 1115, 906, 372]. Self-adjoint [353, 372]. self-adjointness [1865]. Self-avoiding [1130]. self-conjugate [1779, 1949, 1034]. self-coupling [1115]. self-dual [1535]. Self-gravitating [248, 1848, 596, 1240, 862]. self-induced [1152]. Self-interacting [221]. self-reciprocal [906]. semi [91, 1374, 437, 438, 179]. semi-bounded [91]. semi-Euclidean [437, 438].

Semi-infinite [179]. semi-Riemannian [1374]. Semiclassical [1616, 1637, 317, 1261, 870, 1670, 809]. semiconductor [284]. semiconductors [1933, 164]. semigroup [161]. semigroups [538, 1328, 1416]. Semisimple [81, 330, 505, 1368, 319, 1634, 297, 1781, 1091, 919, 270]. Semisubgroup [399, 792]. semiunitary [1860]. Sen [67, 1563]. separability [1378]. separable [458, 459, 1012, 1332, 241, 892]. separate [1675]. separating [204]. Separation [244, 395, 358, 96, 97, 1156, 1148, 6, 255, 399, 458, 459, 460, 1127, 506]. Sequences [1082, 1483, 1095, 1204, 872, 1188]. Series [542, 499, 741, 796, 1176, 628, 1008, 1594, 1745, 669, 24]. set [508, 509, 1720, 1396]. sets [1404, 1218, 925]. setting [262, 1671]. Settling [1796]. seven [93]. seven-dimensional [93]. several [200, 416]. several-particle [200]. shadow [287]. shape [961]. sharp [584, 1636]. sharp-boundary [1636]. Shear [1765, 376]. Shear-free [1765, 376]. shearfree [1150]. sheet [1222]. shell [1676, 675, 363, 1400, 1610, 1611, 912, 1349, 1703, 1910]. shells [630]. shift [1319, 1466]. shifted [1479]. shifts [120, 1656, 1796, 56, 1358]. Shilov [791]. Shock [730]. Short [805, 1526, 1675]. Siegert [344]. signal [350, 167]. signals [907]. similar [1378]. Similarity [396, 506]. similitude [265, 779]. Simple [443, 1351, 76, 1349, 1861, 195, 59, 533, 66, 1148, 1157, 373, 1316, 686, 548, 1212, 1109, 847, 1043, 654, 655, 1398]. simplicial [389, 1415]. simplicity [1238]. Simplified [229]. simultaneous [1891]. sine [1283, 1114, 393, 1378]. sine-cosine [1283]. sine-Gordon [1114, 393, 1378]. single [1268, 757, 1420, 1421, 69, 394, 1889, 1614, 780, 1585, 1395, 1605, 652, 466, 813]. single-variable [1395]. single-well [757, 1420, 1421]. Singular [1835, 1934, 1783, 166, 625, 586, 1296, 913, 139, 727, 1796, 616, 306, 78, 806, 603, 1358, 829, 547, 1104]. Singularities [377, 13, 1739, 1852, 896]. singularity [490, 823, 764, 302]. site [33]. Sitter [1067, 527, 1094]. sixth [1754]. skew [1344]. skew-symmetric [1344]. Skorohod [1463]. slab [1857, 1545, 834, 1530, 1184]. Slater [1453, 1135]. Slater-type [1453, 1135]. Slavnov [1217]. Slavnov-’t [1217]. slices [611, 1886, 1885]. slits [1274]. slope [1188]. Slow [1431, 1564]. slowing [1117, 1878, 1540]. Small [310, 425, 522, 151]. smooth [1112, 1421, 543]. smooth-path [543]. smoothing [299, 110]. soft [662]. solenoidal [686]. solids [1108, 1326]. solitary [1701, 1531, 1256]. Soliton [830, 1717, 1087, 1519, 1896, 807, 1093, 1850, 1272]. Solitonlike [1283]. Solitons [1481, 533, 770, 997, 1264, 666, 1772, 1340, 1686, 1056, 1817, 1786, 1945, 891, 1303, 1561]. Soluble [656, 855, 1439, 1130, 706]. Solution [249, 487, 427, 206, 721, 1352, 1383, 1202, 1903, 1803, 1825, 1372, 1171, 1891, 143, 1526, 801, 670, 516, 1598, 518, 1896, 551, 582, 581, 307, 1075, 583, 1904, 340, 326, 281, 1364, 1251, 1823, 612, 1544, 396, 1032, 593, 132, 368, 737, 418, 1824, 123, 346, 45, 1595, 1747, 687, 1476, 1560, 1693, 54, 303, 1498]. Solutions [1588, 1785, 1950, 1836, 856, 478, 1482, 479, 338, 392, 1481, 556, 791, 611, 1361, 1738, 189, 635, 238, 92, 1388, 1895, 349, 153, 602, 760, 851, 942, 783, 333, 156, 1713, 1163, 1021, 982, 1193, 50, 1010, 391, 343, 1103, 1321, 1681, 1342, 754, 610, 1231, 758, 1003, 1288, 1093, 459, 1012, 1332, 35, 1335, 1283, 299, 1058, 1289, 145, 165, 500, 899, 1079, 1141, 830, 1709, 834, 385, 1833, 677, 621, 705, 786, 1563, 148, 1621, 525, 526, 400, 969, 129, 1797, 1000, 1717, 1760, 1248, 932, 949, 950, 1185, 55, 939, 881, 1566, 1939, 506]. solutions [1385, 1125, 1348, 1424, 261, 1293, 29, 1672]. solvable [917, 1033, 131, 1665, 1132, 571]. solving [773, 1356, 191, 49]. Some [553, 170, 1061, 728, 954, 849, 1537, 416, 1193, 1103, 739, 68, 35, 802, 803, 1289, 1868, 574, 385, 1690, 102, 1248, 1935, 1632, 1743, 1430, 1198, 1387, 427, 335, 887, 341, 1817, 1038, 1356, 969, 37, 1188, 16, 939, 1566, 1939, 1386, 1130]. something [242]. Sommerfeld [1420, 1854]. sound [1111, 1746, 1947]. source [778, 1953, 801, 217, 1222, 1746, 1947, 786, 1060]. source-free [786]. sourceless [1306]. sources [1791, 989]. Space [161, 359, 699, 1658, 36, 1827, 1743, 1474, 93, 570, 917, 585, 1477, 1307, 51, 725, 736, 91, 494, 82, 112, 358, 726, 576, 1387, 698, 1923, 692, 1344, 700, 720, 1216, 1774, 1775, 1648, 1647, 962, 851, 1836, 1206, 239, 451, 1067, 483, 1507, 1723, 377, 1914, 543, 1074, 40, 619, 68, 688, 1681, 1462, 617, 1694, 432, 410, 1587, 1096, 967, 1066, 895, 1203, 744, 792, 1012, 1331, 1332, 753, 1050, 184, 209, 885, 1182, 903, 1786, 1945, 997, 188, 884, 1831, 684, 1423, 1727, 157, 822, 979, 1170, 1946, 1079, 1141, 418, 943, 1493, 1815]. space [765, 1114, 527, 1799, 291, 2, 87, 491, 661, 1496, 1037, 1240, 477, 810, 681, 896, 286, 771, 639]. Spacelike [1771, 611, 1471, 1683, 1213]. Spaces [916, 1079, 1141, 781, 1173, 492, 1430, 1294, 863, 453, 496, 1100, 1386, 1502, 762, 168, 1412, 514, 1246, 1832, 721, 696, 1053, 17, 18, 1604, 386, 1495, 252, 928, 1309, 55, 633, 1018]. Spacetime [95, 72, 1800, 1392, 1501, 1154, 1917, 959, 1683, 1907, 1372]. spacetimes [26, 27, 94, 1724, 98, 99, 1342, 1912, 985, 852]. Spatial [1375, 907, 1376, 1588, 1785, 1950, 1807, 694, 794, 1261]. spatially [1150]. special [1419, 438, 514, 410, 1853, 887, 354, 759, 852, 1209]. species [148]. specify [1800]. spectra [304]. Spectral [1640, 1948, 1062, 1277, 1517, 372, 667, 91, 1659, 1682, 1710, 848, 578, 1851]. spectroscopy [1824, 1903]. Spectrum [464, 422, 247, 890, 1793, 869, 1064, 1807, 988, 951, 534, 414, 776, 5]. speed [350, 167, 1381]. sphere [4, 428, 1804, 1317, 1913, 1608]. spheres [881]. Spherical [1441, 1539, 700, 720, 524, 1263, 1096, 996, 1222, 363, 1521, 1693, 877, 800]. Spherically [1563, 760, 1914, 340]. spheroidal [1029, 381]. Spin [1029, 819, 708, 1282, 1642, 788, 956, 836, 510, 470, 1321, 56, 1164, 811, 1492, 1232, 1259, 1698, 1327, 295, 311, 1279, 1220, 1106, 656, 54, 1416, 939, 1783, 844, 423, 472, 1899, 1282]. spin- [956, 510, 811, 1259, 939, 1259]. Spin-frame [819]. spin-orbit [56, 1164]. spin-spin [54]. Spin-weighted [1029]. spinel [33]. spinless [1534, 337]. spinning [640, 1243, 412, 1567, 1286, 77, 337, 1125, 53]. Spinor [489, 93, 791, 82, 361, 524, 1228, 1165, 897, 1334]. spinor-spinor [361].

Spinorial [998, 740]. spinors [566, 739, 1898, 1593, 681]. spins [1802, 447, 1140]. spiraling [1280]. spl [769]. Splitting [1195]. Spontaneous [995]. Spontaneously [523, 1921, 1523, 1489]. springs [662]. square [249, 772, 1075, 326, 1191]. square-mound [1075]. square-well [326, 1191]. Stability [439, 982, 1867, 1284, 1070, 633, 874, 1622, 22, 1215, 73, 124, 745, 390, 909, 219, 95, 699, 1185, 1695, 1938]. Stable [1621, 1072, 1525, 1494, 1238, 1010, 1011, 1105]. stage [1857]. standard [1123]. star [1748]. Stark [890, 1650, 1219]. stars [1260, 1273, 1161, 1550, 954, 1263, 1102]. state [1402, 592, 190, 1538, 481, 1313, 858, 116, 629, 1098, 1364, 263, 1870, 1039, 13, 478, 1482, 854, 479, 1818, 183, 20, 969, 1393, 1498, 302, 895]. states [310, 708, 492, 1503, 788, 1925, 450, 513, 446, 1508, 332, 452, 1421, 1382, 364, 1064, 344, 1148, 1157, 1169, 219, 658, 1308, 1688, 1020, 993, 38, 467, 1105, 1353, 1354, 515, 777, 1401, 1406, 1470, 1191, 471, 1931, 630, 565, 893, 1480, 1479]. Static [92, 954, 99, 1747, 1828, 1657, 1193, 68, 1257, 1141, 1563, 1060, 1875]. Stationary [839, 942, 1874, 98, 1084, 18, 950, 485, 1740, 199, 1738, 743, 1388, 1715, 1558, 239, 451, 40, 1103, 1211, 1911, 1912, 980, 981, 1436, 1459, 7, 473, 684, 1674, 455, 1079, 795, 2, 87, 1553, 1417, 294, 1821, 392]. Statistical [253, 74, 75, 289, 504, 826, 370, 1198, 1465, 173, 1304, 437, 438, 318, 941, 86, 532, 806, 1665, 1708, 1403, 1811, 797, 444, 16, 1902]. statistics [36, 1143, 469, 923, 1300, 780, 1585]. Steady [685, 858, 1364, 1148, 1157]. steady-state [1364]. steep [1420, 958]. stellar [881]. step [1301]. steps [437]. Stiefel [1306]. Stieltjes [499, 542, 1292]. stiff [1498]. stimulated [1812]. Stochastic [529, 1496, 1551, 799, 1432, 668, 329, 788, 1822, 63, 205, 1145, 1712, 158, 724, 1407, 929, 1849, 322, 1308, 31, 1867, 1495, 1566, 1939, 992, 1626, 1554]. stochastically [101]. stochasticity [1812]. Stokes [982, 1842, 331]. straight [301]. straight-line [301]. stratified [583, 1102]. strength [152]. stress [409]. stress-energy [409]. stretched [1345]. string [1792, 864]. stringlike [1340, 1686]. Strings [1183, 1753, 1792, 1505, 1433, 1570]. strong [340]. strongly [1796, 713, 714, 715]. Structural [745, 39, 266, 1034, 387]. Structure [1811, 1044, 1541, 1377, 1376, 840, 1633, 1226, 1772, 1073, 446, 740, 543, 1347, 116, 1108, 554, 753, 1336, 1020, 401, 831, 1139, 548, 988, 951, 148, 1309, 648, 1522, 706]. structures [1461, 1057, 615, 998, 1501, 976, 432, 1786, 1945, 693, 694, 794, 1127, 1777, 12]. Studies [137, 1108]. study [501, 1865, 1353]. sub [1094, 1127]. sub-group [1094]. sub-structures [1127]. Subalgebras [965, 779, 1634, 571]. subclass [1209]. Subcriticality [1184, 834]. subdynamics [36, 679]. subgroup [1206, 1365]. Subgroups [755, 563, 1477, 1239, 1410, 81, 401, 1189, 264, 265, 270, 558, 133, 315]. subject [1451, 1932, 947, 1795]. subspaces [1173, 782]. Successive [1778]. Sudarshan [1249]. sufficient [1800, 803]. Sum [642, 1840, 210, 1816]. summability [1292, 309]. Summation [1258, 1581, 1745]. sums [935, 1503, 243, 775, 210, 448, 1108, 114, 352, 544]. Super [1473, 1478, 1186].

Super-convergent [1478]. super-Hamiltonians [1186]. superalgebras [1515, 1837, 1110, 1483]. superconducting [1857]. superconductivity [1593]. supercriticality [1184]. supereigenvalue [612]. Superfields [844, 298, 341]. superfluid [440, 1208, 1746, 1947, 1607, 1608]. superfluorescent [1452, 1225]. supergauge [468, 1042]. supergravity [752, 1265]. superluminal [49]. Superposition [446, 344]. superpropagator [415]. superpropagators [898]. superselection [47, 1409, 11, 703]. superspace [468, 1110, 1305]. superspaces [752]. supersymmetric [835]. supersymmetries [443]. supersymmetry [1265, 1286, 554, 1053, 1485]. supervariational [186]. support [1472]. Surface [345, 1245, 763, 1339, 292, 560, 697, 293]. surfaces [586, 1379, 1757, 185, 1558, 1022, 1801, 1937, 386]. surroundings [1416]. susceptibility [1701]. symbol [1780, 480]. symbols [66, 1197, 1894, 1390]. symmetric [1357, 1762, 199, 790, 1828, 916, 422, 1738, 1387, 92, 1344, 26, 27, 942, 9, 171, 1914, 429, 1103, 340, 1587, 1278, 1674, 1563, 1001, 400, 1450, 1399, 985, 392, 1479]. symmetrical [416, 269]. Symmetries [1158, 905, 1448, 927, 1911, 1912, 980, 981, 1436, 1459, 1443, 1464, 1429, 1376, 1884, 1886, 161, 1086, 502, 1287, 1921, 1016, 1769, 688, 1131, 103, 934, 1435, 95, 699, 926, 254, 72, 1728, 659, 1885]. symmetrization [588, 589]. Symmetry [637, 1156, 758, 149, 1192, 631, 755, 358, 1116, 1779, 1949, 1262, 239, 747, 1723, 1136, 50, 40, 415, 103, 622, 817, 342, 1194, 248, 1848, 643, 259, 348, 943, 680, 705, 596, 687, 995, 994, 49, 948, 1343, 133, 315, 192, 523, 630, 223, 1488]. symmetry-equals-invariant [1116]. symplectic [1569, 276, 741, 1151, 651, 957, 765, 1506, 1533, 211]. Synchronized [891]. system [874, 1134, 1451, 1682, 1710, 1699, 173, 1304, 869, 1439, 871, 308, 1546, 616, 947, 1795, 1681, 1231, 811, 1295, 1529, 831, 1818, 1070, 1608, 1924, 1553, 1554, 152, 447, 54, 261]. systematic [883]. Systems [781, 158, 724, 1407, 1569, 310, 1891, 1740, 865, 586, 91, 1296, 1198, 250, 276, 1589, 1779, 1949, 679, 1027, 728, 1895, 1642, 946, 42, 1494, 1691, 22, 1067, 727, 1742, 198, 86, 1677, 1016, 538, 1320, 1442, 321, 61, 216, 75, 1131, 458, 1232, 103, 817, 175, 1637, 219, 1868, 275, 1671, 1727, 1592, 378, 1305, 295, 1395, 1605, 649, 834, 880, 1876, 1105, 1271, 1551, 1607, 1608, 1007, 489, 828, 1401, 1553, 1554, 1299, 1726, 1763, 1653, 1187, 423, 5, 43, 485]. Szego [1627]. Szekeres [852].

’t [1168, 1217, 1879]. tableau [384]. tachyon [1068, 1450]. Tachyonic [1216]. tachyons [311]. Tani [308, 1546]. Taylor [441]. technique [1377, 1061, 855, 1108, 322, 697]. techniques [1543, 494, 1702, 63, 394, 359, 1562, 715]. Teller [842]. temperature [1877, 1918]. tempered [649]. Temperley [1806]. temporal [1812, 1187]. Temporally [136]. Tensor [1630, 88, 1477, 700, 720, 1317, 1521, 1827, 566, 722, 916, 975, 436, 762, 1422, 1344, 412, 451, 376, 340, 1754, 269, 1435, 545, 537, 536, 989, 1034, 157, 821, 257, 786, 1747, 1001, 400, 338, 1533, 445]. tensors [1541, 1571, 1241, 39, 266, 617, 1154, 409, 1512, 944, 1744, 133, 315, 1018, 774]. term [1352, 561]. terms [566, 1171, 162, 570, 517, 1245, 425, 21, 119, 177]. ternary [1837]. test [1392, 1243]. tetrad [1200, 1887]. th [1429, 1428, 599, 1768, 133, 315]. th-order [1768]. their [1935, 1632, 1707, 904, 1886, 450, 1067, 502, 1199, 233, 1233, 1928, 643, 1651, 1755, 323, 1468, 548, 563, 201, 779, 644, 1000, 1820, 971, 1596, 1644, 774, 374]. Theorem [1418, 963, 441, 647, 1802, 1116, 1422, 921, 760, 1787, 1627, 1347, 1231, 469, 127, 409, 1249, 1564, 401, 821, 978, 686, 1160, 988, 951, 1425, 818, 386, 1844, 57, 423]. theorems [1613, 220, 704, 15, 520, 833, 16, 1026]. theoretic [613, 178, 789, 1553, 1554, 1261, 1353]. Theoretical [166, 987, 625, 1722, 512, 1592, 1777]. theories [1235, 1513, 1158, 1083, 271, 174, 1311, 324, 1923, 450, 927, 1384, 1145, 1449, 105, 340, 688, 604, 754, 827, 1523, 973, 1289, 77, 1160, 122, 786, 1563, 486, 659, 1489, 994, 1400, 403]. Theory [1753, 1792, 1144, 1812, 609, 130, 131, 132, 311, 36, 1429, 1428, 493, 93, 1635, 784, 1575, 1402, 221, 246, 1357, 1762, 1828, 1871, 453, 1640, 1948, 1434, 1543, 508, 509, 1370, 736, 529, 572, 1119, 1845, 975, 670, 793, 112, 96, 97, 1535, 1738, 723, 983, 1565, 882, 1808, 698, 1100, 1386, 1277, 864, 1511, 59, 848, 552, 918, 1101, 1773, 1646, 411, 1253, 1384, 66, 1409, 1789, 425, 1582, 1741, 1932, 139, 63, 468, 796, 1176, 1643, 1601, 1193, 1514, 1501, 1661, 65, 1010, 390, 343, 1826, 151, 1382, 253, 136, 104, 313]. theory [216, 620, 495, 74, 75, 289, 504, 826, 1654, 1600, 306, 1288, 6, 255, 399, 458, 459, 460, 744, 792, 1012, 1331, 1332, 1413, 1552, 1333, 867, 1201, 599, 175, 1859, 1951, 1380, 1557, 1254, 1113, 903, 1732, 991, 299, 406, 555, 1433, 658, 1890, 1619, 1940, 1264, 1042, 228, 1423, 1564, 455, 671, 672, 32, 1290, 1590, 875, 1147, 866, 1080, 603, 1906, 1562, 1493, 1815, 351, 181, 808, 251, 1028, 2, 87, 388, 45, 1446, 786, 1747, 1607, 1608, 110, 886, 20, 861, 756, 944, 645, 128, 915, 761, 1745, 1224, 810, 400, 338, 1875]. theory [274, 1403, 1778, 11, 703, 907, 37, 1711, 444, 1185, 984, 1883, 1445, 5, 43, 1104, 1821, 1899, 1153, 1219, 1672, 1346, 1920]. there [1906]. thermal [1846, 74, 75, 289, 504, 826, 396, 1180]. thermodynamic [221, 1020, 649]. Thermodynamical [1665, 1215]. Thermodynamics [672, 650, 390, 879, 1275, 1918]. thermopiezoelectric [1144]. thickness [1784]. third [908]. Thirring [1139]. Thomas [760]. those [305]. Three [1355, 1303, 249, 1181, 1152, 1785, 1950, 220, 665, 1138, 111, 859, 227, 478, 1482, 300, 1585, 1709, 479, 256, 965, 268, 706, 1480, 774]. Three- [1355, 965]. three-body [249, 859, 256]. Three-dimensional [1303, 1181, 1138, 300, 1585, 1709, 268, 706]. three-level [665]. three-magnon [1482]. three-to-three [220]. three-wave [1152, 111]. threshold [1032]. Time [1706, 1116, 1253, 1902, 103, 817, 934, 1520, 1477, 1452, 1225, 637, 587, 276, 280, 189, 161, 1216, 1705, 1150, 1427, 543, 1074, 391, 837, 1769, 688, 617, 432, 528, 744, 792, 1012, 1331, 1332, 622, 78, 1823, 442, 1050, 184, 209, 885, 1182, 1038, 991, 1573, 1574, 1222, 931, 1868, 1727, 854, 1305, 699, 1906, 1493, 1815, 256, 346, 1114, 1844, 1658, 810, 1450, 896, 771, 934, 1182, 639]. time-asymmetric [1705]. time-delay [256]. Time-dependent [1253, 103, 817, 934, 189, 837, 1769, 528, 622, 78, 1823, 1038, 991, 1573, 1574, 931, 1868, 1844, 810, 934]. Time-reversal [1520]. time-symmetric [1450]. time-varying [280]. timelike [959, 1739]. times [1387, 239, 451, 377, 1914, 40, 619, 68, 1462, 1587, 884, 684, 1079, 1141, 943, 1799, 291, 681]. Toda [238]. Tomimatsu [1125, 1348, 1424, 582, 412, 581, 687]. top [422]. Topological [1891, 1350, 1561, 345, 646, 920, 1489, 1852]. topologically [1307]. Topology [85, 1015, 543, 1022, 432, 959, 204, 1014]. Toroidal [1759, 686, 303]. torsion [1235, 1724, 1658]. total [1696, 1188]. trace [1639]. tracks [543]. trajectories [196, 1558, 894, 872, 734]. transfer [963, 1119, 1927, 1456, 751]. transform [179, 1061, 546, 1484, 1472]. Transformation [1788, 1166, 1689, 1397, 238, 382, 1915, 510, 308, 1546, 1858, 1698, 180, 277, 30, 1139, 28, 76, 1663, 1790, 550, 482, 472]. transformations [1323, 1750, 541, 1116, 283, 483, 332, 452, 1724, 970, 1082, 1382, 682, 1858, 1859, 1951, 1233, 885, 1182, 393, 1500, 991, 1573, 1149, 658, 960, 1868, 323, 304, 1847, 847, 828, 1166, 1912, 1459]. Transforming [1146]. transforms [250, 1577, 1478, 1453, 1135, 506, 906]. transient [1711]. transients [19]. Transition [453, 1298, 1943, 745, 1712, 1259, 1043]. transitions [1134, 1404, 1011, 247, 521, 1737, 1777]. Translation [1090, 1089, 1442, 1904, 288]. translation-invariant [1442, 1904]. translational [1223]. translationally [840]. Transmission [1878, 1359, 1764]. transparency [1152]. transparent [849]. Transport [1764, 331, 667, 419, 420, 670, 793, 1702, 496, 1117, 507, 1055, 1545, 429, 721, 151, 1413, 1552, 1333, 867, 1201, 1202, 237, 685, 1807, 300, 1290, 1590, 346, 676, 834, 1530, 1184, 363, 1381]. trapped [185]. treated [128]. treatment [1376, 383, 321, 154]. tree [121]. trees [15]. triad [226]. triad-interaction [226]. trilinear [644]. Trotter [1123]. tunnelling [1310]. turbulence [668, 1054, 1856, 227, 226]. turbulent [609, 1643]. twist [1230]. twisted [88, 328, 127]. twisting [1419]. twistor [753]. Twistors [1024]. Two [320, 216, 1725, 213, 1763, 1025, 1329, 499, 913, 1100, 1386, 1448, 382, 1682, 1588, 1344, 1699, 1925, 901, 1705, 155, 1298, 1943, 50, 391, 340, 39, 266, 75, 1148, 301, 1254, 1730, 1731, 227, 287, 780, 1585, 295, 694, 794, 1906, 1274, 191, 1114, 639, 1926, 2, 87, 661, 365, 933, 1899, 1522, 630]. two- [1585]. Two-body [1725, 1025, 1329, 901, 1705, 155, 1254, 1906]. two-charge [301]. Two-dimensional [1763, 75, 780, 295, 1926, 365, 933]. two-electron [1682]. two-Hilbert [87, 2]. Two-level [216]. two-parameter [50]. two-particle [1699, 1925, 661]. two-photon [1298]. Two-point [320, 191]. two-spin [1899]. two-structure [1522]. two-tensor [340]. Type [1230, 1861, 955, 1262, 1602, 807, 1419, 1919, 1857, 619, 1704, 789, 1038, 991, 1455, 440, 122, 1833, 1690, 129, 1453, 1135]. type- [1602]. types [883].

ultralocal [927]. unbounded [690, 1802, 664, 1787]. uncertainties [144]. uncertainty [1196]. uncorrelated [657]. undergoing [107]. Unequal [361]. uniaxial [230, 624]. Unification [403, 337]. unifications [324]. Unified [1433, 1376, 1738, 1509, 1380, 1557, 28, 135]. Uniform [151, 1312, 579, 590, 1039, 107, 1804]. Uniformly [1058]. unique [486]. Uniqueness [1163, 1143, 732, 1794, 899, 900, 1476, 1560]. unit [1601, 1096]. Unitarily [1401, 652]. Unitarity [1879, 1, 1909, 225, 466, 1476, 1560, 144]. Unitary [815, 1204, 1769, 25, 1239, 236, 1151, 498, 554, 841, 1162, 1081, 588, 589, 64, 133, 315]. Universal [1104, 1086, 65, 644, 1285, 1376]. universe [1307, 1829, 413, 433, 1150, 1019, 825, 1068, 1450, 1177]. universes [575, 1150, 50, 1107]. unstable [161]. upon [584, 187]. Upper [1663, 1645, 421]. Ursell [285]. use [888, 1278, 384]. Useful [90]. using [143, 1507, 63, 498, 1889, 1614, 824].

V [1911, 75]. vacuum [170, 239, 1704, 1739, 795, 681, 303, 1385, 1209]. valid [1058]. validity [1074, 1381]. value [427, 1638, 191, 1490, 118]. valued [1613, 578, 867, 1201]. values [804, 714]. values-polarization [714]. Variable [178, 150, 61, 1162, 1395, 1605, 1078]. variables [358, 96, 97, 1156, 369, 416, 244, 395, 6, 255, 399, 458, 459, 460, 1360, 819, 506]. Variation [909, 1817, 1222, 1214]. Variational [1428, 930, 186, 1567, 60, 267, 1314, 1269, 1687, 370, 91, 518, 692, 804, 155, 86, 134, 90, 1118, 1333, 1927, 228, 1180, 1142, 1234, 284, 621]. variations [953, 1414]. Varieties [994]. various [1831]. varying [280]. Vector [1429, 89, 1096, 157, 1468, 14, 964, 1534, 1245, 1752, 1832, 897, 295, 1141, 1109, 201]. vectors [1430, 1246, 410, 1485, 973, 686, 122]. velocities [49]. velocity [1090, 894, 841, 1162, 1338]. velocity-dependent [1338]. verifiable [811]. version [1175]. versus [299]. vertex [842]. very [1420]. Vestigial [139]. VI [1912, 289, 504]. via [1691, 1858, 1859, 1951, 20]. vibrations [516, 197, 463, 1144]. viewpoint [626]. VII [173, 826]. VIII [865, 1304]. virial [1422, 1147]. viscous [396]. Visual [1898]. Vlasov [668, 391, 1856, 1636]. Vlasov-fluid [1636]. Vlasov-plasma [668]. Volterra [1058, 1033, 148]. volume [181, 808, 1114]. vortex [1448]. vortices [742]. vorticity [1374]. Vries [602, 1118, 1499, 830, 418, 1653, 261].

Waals [1080]. Walfisz [935]. walk [1364]. walks [1130]. wall [418]. walls [1420]. Wannier [1146]. Wave [101, 559, 107, 577, 832, 556, 791, 529, 1355, 1550, 1277, 1152, 1785, 1950, 349, 153, 380, 1896, 234, 1676, 673, 1439, 602, 1602, 111, 1643, 1516, 1258, 1581, 50, 136, 334, 835, 458, 459, 744, 792, 1012, 1331, 1332, 1870, 1050, 1701, 1222, 1933, 1824, 1903, 440, 1208, 1137, 45, 786, 878, 1811, 1432, 1518]. wave-kinetic [529]. wave-mechanical [1137]. wave-projected [1518]. wavefunction [1336, 993, 1220]. Wavefunctions [1267, 592, 888, 186, 1174, 1038, 1751, 1714, 1936, 1586, 1393]. waveguide [673]. waveguides [907]. wavelength [197, 463]. Wavepacket [984]. waves [1855, 556, 230, 624, 634, 1457, 1678, 1216, 231, 1578, 267, 424, 789, 730, 292, 560, 519, 1531, 113, 1159, 1820, 833, 1337, 293, 1256]. Weak [1734, 1606, 340, 1463, 829]. Weakly [649, 10, 202, 899, 1918]. Wegner [748]. weight [297]. weighted [1029, 252]. weights [1091, 919]. Weinberg [510]. Weinless [1794]. Weiss [1284]. Weisskopf [173, 1304]. Weizel [1676, 1393]. well [757, 1420, 1421, 326, 486, 1191, 605]. well-defined [486]. Wentzel [757]. Weyl [412, 451, 376, 546, 1484, 159, 119, 355, 588, 1497, 1760, 176, 1390, 1236]. Wheeler [1467]. where [556]. which [1384, 543, 1383, 1462, 1154, 432, 1912, 1459, 802, 803]. white [614]. whole [239]. whose [654, 655]. Wick [772, 29]. Wick-rotated [29]. Widom [1930]. widths [1870]. Wien [759]. Wiener [1814, 408, 1408]. Wightman [643, 822]. Wigner [305, 1780, 173, 1304, 495, 1391, 821, 260, 1583, 1782, 595, 648, 1893]. without [1880, 864, 1756, 425, 888, 78, 500, 1170, 1946, 681]. WKB [1169, 621]. WKBJ [1106]. Woods [675, 1069]. World [1901, 1226]. World-line [1901]. Wouthuysen [1698]. Wronskian [1061, 855]. Wronskians [187].

x [541, 481].

Yamaguchi [893]. Yang [1547, 733, 1004, 1835, 1226, 742, 457, 1681, 758, 1528, 1015]. Yevick [1598, 1075, 326]. yielding [1401]. York [1800]. Young [384]. Yukawa [32, 913, 1676, 886, 365].

z [541]. Zeno [866]. Zermelo [508, 509]. Zernike [493, 195, 1080]. zero [556, 1297, 510, 1162, 1351, 902, 993, 131, 1877]. zero-energy [1297]. zeroes [707]. Zeros [733, 1004, 1930, 366]. Zinnes [80]. Zorn [1680]. Zucker [775].

References

Newton:1968:DAD

[1] Roger G. Newton. Determination of the amplitude from the differential cross section by unitarity. Journal of Mathematical Physics, 9(12):2050– 2055, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

      7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
      jmapaq/v9/i12/p2050_s1. See addendum [225].

Prugovecki:1973:MSS

[2] E. Prugovečki. Multichannel stationary scattering theory in two-Hilbert space formulation. Journal of Mathematical Physics, 14(7):957–962, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

      tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i7/p957_s1. See erratum [87].

Yao:1973:ANP

[3] Te Hai Yao. On analytic nonlocal potentials. I. A forward dispersion relation. Journal of Mathematical Physics, 14(8):1141–1151, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i8/p1141_s1. See erratum [44].

Adler:1974:FSG

[4] Ronald J. Adler. A fluid sphere in general relativity. Journal of Mathematical Physics, 15(6):727–729, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i6/p727_s1. See erratum [428].

Wu:1974:AIOb

[5] Shi yu Wu. Application of infinite order perturbation theory in linear systems. II. The frequency spectrum of disordered chains. Journal of Mathematical Physics, 15(7):947–949, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p947_s1. See erratum [43].

Kalnins:1974:LTSa

[6] E. G. Kalnins and W. Miller, Jr. Lie theory and separation of variables. 3. The equation f_{tt} −f_{ss} = γ^{2}f. Journal of Mathematical Physics, 15(7): 1025–1032, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i7/p1025_s1. See erratum [255].

Kloster:1974:SGF

[7] S. Kloster, M. M. Som, and A. Das. On the stationary gravitational fields. Journal of Mathematical Physics, 15(7):1096–1102, July 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p1096_

s1. See erratum [473].

Yao:1974:OIP

[8] Te Hai Yao. The T operator and an inverse problem for nonlocal potentials. Journal of Mathematical Physics, 15(8):1227–1231, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v15/i8/p1227_s1. See erratum [172].

Ernst:1974:CPF

[9] Frederick J. Ernst. Complex potential formulation of the axially symmetric gravitational field problem. Journal of Mathematical Physics, 15 (9):1409–1412, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i9/p1409_s1. See erratum [171].

Capri:1974:EWR

[10] Anton Z. Capri. On the existence of weakly retarded and advanced Green’s functions. Journal of Mathematical Physics, 15(12):2093–2095, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2093_s1. See erratum [202].

Strocchi:1974:PCS

[11] F. Strocchi and A. S. Wightman. Proof of the charge superselection rule in local relativistic quantum field theory. Journal of Mathematical Physics, 15(12):2198–2224, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i12/p2198_s1. See erratum [703].

Wahlquist:1975:PSN

[12] H. D. Wahlquist and F. B. Estabrook. Prolongation structures of nonlinear evolution equations. Journal of Mathematical Physics, 16(1):1– 7, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i1/p1_s1.

Letelier:1975:SFE

[13] P. Letelier and R. Tabensky. Singularities for fluids with p = ω equation of state. Journal of Mathematical Physics, 16(1):8–10, January 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i1/p8_s1.

Beers:1975:AVH

[14] Brian Lee Beers and Richard S. Millman. Analytic vector harmonic expansions on SU(2) and S^{2}. Journal of Mathematical Physics, 16(1): 11–19, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i1/p11_s1.

Friedberg:1975:DTR

[15] R. Friedberg. Dual trees and resummation theorems. Journal of Mathematical Physics, 16(1):20–30, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i1/p20_s1.

Vezzetti:1975:NDS

[16] David J. Vezzetti. A new derivation of some fluctuation theorems in statistical mechanics. Journal of Mathematical Physics, 16(1):31–33, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p31_s1.

Lind:1975:GER

[17] Robert W. Lind. Gravitational and electromagnetic radiation in Kerr– Maxwell spaces. Journal of Mathematical Physics, 16(1):34–38, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p34_s1.

Lind:1975:SKM

[18] Robert W. Lind. Stationary Kerr–Maxwell spaces. Journal of Mathematical Physics, 16(1):39–43, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i1/p39_s1.

Wenzel:1975:PTR

[19] Alan R. Wenzel. Propagation of transients in a random medium. Journal of Mathematical Physics, 16(1):44–49, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i1/p44_s1.

Rubinson:1975:DSP

[20] William Rubinson. Discrete state perturbation theory via Green’s functions. Journal of Mathematical Physics, 16(1):50–53, January 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i1/p50_s1.

Ihrig:1975:EDG

[21] Edwin Ihrig. An exact determination of the gravitational potentials g_{ij} in terms of the gravitational fields R^{l_{ijk}}. Journal of Mathematical Physics, 16(1):54–55, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i1/p54_s1.

Dysthe:1975:RSC

[22] K. B. Dysthe and O. T. Gudmestad. On resonance and stability of conservative systems. Journal of Mathematical Physics, 16(1):56–64, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p56_s1.

Geroch:1975:MBG

[23] Robert Geroch and Pong Soo Jang. Motion of a body in general relativity. Journal of Mathematical Physics, 16(1):65–67, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i1/p65_s1.

Yao:1975:DSN

[24] T. H. Yao. A dispersion series for nonlocal potentials. Journal of Mathematical Physics, 16(1):68–71, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i1/p68_s1.

Aguilera-Navarro:1975:COC

[25] M. C. K. Aguilera-Navarro and V. C. Aguilera-Navarro. Casimir operators of complementary unitary groups. Journal of Mathematical Physics, 16(1):72–74, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i1/p72_s1.

Davis:1975:GNP

[26] Talmadge M. Davis and John R. Ray. Ghost neutrinos in planesymmetric spacetimes. Journal of Mathematical Physics, 16(1):75–79, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i1/p75_s1.

Davis:1975:NCS

[27] Talmadge M. Davis and John R. Ray. Neutrinos in cylindricallysymmetric spacetimes. Journal of Mathematical Physics, 16(1):80–81, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i1/p80_s1.

Moshinsky:1975:CTAa

[28] M. Moshinsky, J. Patera, and P. Winternitz. Canonical transformation and accidental degeneracy. III. A unified approach to the problem. Journal of Mathematical Physics, 16(1):82–92, January 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i1/p82_s1.

ZumWinkel:1975:ESW

[29] D. Zum Winkel. On the exact solutions of the Wick-rotated fermionantifermion Bethe–Salpeter equation. Journal of Mathematical Physics, 16(1):93–95, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i1/p93_s1.

McLaughlin:1975:FEI

[30] D. W. McLaughlin. Four examples of the inverse method as a canonical transformation. Journal of Mathematical Physics, 16(1):96–99, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v16/i1/p96_s1. See erratum [277].

McLaughlin:1975:SGB

[31] D. W. McLaughlin. A stochastic Gaussian beam. II. Journal of Mathematical Physics, 16(1):100–103, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i1/p100_s1.

McBryan:1975:LCY

[32] Oliver A. McBryan and Yong Moon Park. Lorentz covariance of the yukawa_{2} quantum field theory. Journal of Mathematical Physics, 16(1):104–110, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i1/p104_s1.

Inoue:1975:LGFa

[33] Michiko Inoue. Lattice Green’s function for B-site lattice in spinel. Journal of Mathematical Physics, 16(1):111–115, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i1/p111_s1.

Maresca:1975:FRD

[34] Neal J. Maresca and Richard L. Liboff. Fields and radiation due to a charge incident on a conducting plane. Journal of Mathematical Physics, 16(1):116–124, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i1/p116_s1.

Krasinski:1975:SSE

[35] Andrzej Krasiński. Some solutions of the Einstein field equations for a rotating perfect fluid. Journal of Mathematical Physics, 16(1):125– 131, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i1/p125_s1.

Alberti:1975:PRB

[36] M. A. Alberti, P. Cotta-Ramusino, and G. Ramella. Preliminary results in the Banach space formulation of master equation and subdynamics theory in quantum statistics. Journal of Mathematical Physics, 16(1):132–139, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i1/p132_s1.

Tesei:1975:SNE

[37] A. Tesei. On some nonlinear evolution equations in quantum field theory. Journal of Mathematical Physics, 16(1):140–146, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i1/p140_s1.

Moya:1975:ESI

[38] R. P. Moya. Equilibrium states for the infinite ideal Bose gas. Journal of Mathematical Physics, 16(1):147–149, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i1/p147_s1.

Hauser:1975:SEKa

[39] I. Hauser and R. J. Malhiot. Structural equations for Killing tensors of order two. I. Journal of Mathematical Physics, 16(1):150–152, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p150_s1.

Greene:1975:RFS

[40] Richard D. Greene, Engelbert L. Schucking, and C. V. Vishveshwara. The rest frame in stationary space–times with axial symmetry. Journal of Mathematical Physics, 16(1):153–157, January 1975. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i1/p153_s1.

Golden:1975:MPN

[41] Sidney Golden. A minimum principle for von Neumann’s equation. Journal of Mathematical Physics, 16(1):158–162, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i1/p158_s1. See

erratum [224].

Deliyannis:1975:ISS

[42] Platon C. Deliyannis. Imbedding of Segal systems. Journal of Mathematical Physics, 16(1):163–170, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i1/p163_s1.

Wu:1975:EAI

[43] Shi yu Wu. Erratum: “Application of infinite order perturbation theory in linear systems. II”. Journal of Mathematical Physics, 16(1):171, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p171_s1. See [5].

Yao:1975:EAN

[44] Te Hai Yao. Erratum: “On analytic nonlocal potentials. I. A forward dispersion relation”. Journal of Mathematical Physics, 16(1):171, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p171_s2. See [3].

Raczka:1975:CSN

[45] R. Raczka. The construction of solution of nonlinear relativistic wave equation in λ: Φ_{4}^{4}: theory. Journal of Mathematical Physics, 16(2):173– 176, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p173_s1.

Farrukh:1975:ANA

[46] M. O. Farrukh. Application of nonstandard analysis to quantum mechanics. Journal of Mathematical Physics, 16(2):177–200, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p177_
s1.

Cirelli:1975:ARC

[47] Renzo Cirelli, Franco Gallone, and Bella Gubbay. An algebraic representation of continuous superselection rules. Journal of Mathematical Physics, 16(2):201–213, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i2/p201_s1.

Beran:1975:SFB

[48] M. J. Beran and A. M. Whitman. Scattering of a finite beam in a random medium with a nonhomogeneous background. Journal of Mathematical Physics, 16(2):214–217, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i2/p214_s1.

Solimeno:1975:RGF

[49] Salvatore Solimeno. Riemann–Green’s functions for solving electromagnetic problems exhibiting rotational symmetry in media moving with superluminal velocities. Journal of Mathematical Physics, 16(2):218– 223, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p218_s1.

Gowdy:1975:CGW

[50] Robert H. Gowdy. Closed gravitational-wave universes: Analytic solutions with two-parameter symmetry. Journal of Mathematical Physics, 16(2):224–226, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i2/p224_s1.

Berenson:1975:CGCa

[51] Rhoda Berenson and Joseph L. Birman. Clebsch–Gordan coefficients for crystal space groups. Journal of Mathematical Physics, 16(2):227– 235, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p227_s1. See erratum [725].

Berenson:1975:CGCb

[52] Rhoda Berenson, Irving Itzkan, and Joseph L. Birman. Clebsch–Gordan coefficients for ∗X ⊗ ∗X in diamond O_{h}^{7}-Fd3m and rocksalt O_{n}^{5}-Fm3m. Journal of Mathematical Physics, 16(2):236–242, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p236_
s1.

deAzcarraga:1975:SAR

[53] J. A. de Azcárraga and L. Navarro. On the spinning axis representation. Journal of Mathematical Physics, 16(2):243–246, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p243_
s1.

Subramanian:1975:ESS

[54] G. Subramanian, D. Levitt, and H. T. Davis. An exact solution of the spin-spin autocorrelation function for a one-dimensional system of hard rods. Journal of Mathematical Physics, 16(2):247–250, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i2/p247_s1.

Wang:1975:RES

[55] P. K. C. Wang. Ray and explosive solutions of nonlinear evolutional equations in Hilbert spaces. Journal of Mathematical Physics, 16(2):251– 256, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p251_s1.

Hooshyar:1975:CSO

[56] M. A. Hooshyar. Construction of spin-orbit potentials from the phase shifts at fixed energy. Journal of Mathematical Physics, 16(2):257–262, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p257_s1.

Witten:1975:TAS

[57] T. A. Witten, Jr. A theorem on the adiabatic scaling of classical orbits. Journal of Mathematical Physics, 16(2):263–267, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p263_
s1.

Bove:1975:FIP

[58] A. Bove, G. Fano, G. Turchetti, and A. G. Teolis. Functional integration, Padé approximants, and the Schrödinger equation. Journal of Mathematical Physics, 16(2):268–270, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i2/p268_s1.

Corbett:1975:PSS

[59] J. V. Corbett. Particles and simple scattering theory. Journal of Mathematical Physics, 16(2):271–274, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i2/p271_s1.

Hsieh:1975:VMNa

[60] D. Y. Hsieh. Variational methods and nonlinear forced oscillations. Journal of Mathematical Physics, 16(2):275–280, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p275_s1.

Herrick:1975:DEL

[61] D. R. Herrick. Degeneracies in energy levels of quantum systems of variable dimensionality. Journal of Mathematical Physics, 16(2):281– 283, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p281_s1.

Rauch:1975:ES

[62] Jeffrey Rauch and Michael Taylor. Electrostatic screening. Journal of Mathematical Physics, 16(2):284–288, February 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i2/p284_s1.

Fox:1975:GTM

[63] Ronald Forrest Fox. A generalized theory of multiplicative stochastic processes using cumulant techniques. Journal of Mathematical Physics, 16(2):289–297, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i2/p289_s1.

Sijacki:1975:UIR

[64] Dj. Sijacki. The unitary irreducible representations of S̄L(3, R). Journal of Mathematical Physics, 16(2):298–311, February 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i2/p298_s1.

Grensing:1975:RTU

[65] G. Grensing. Representation theory of the universal covering of the Euclidean conformal group and conformal invariant Green’s functions. Journal of Mathematical Physics, 16(2):312–317, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p312_s1.

Dunlap:1975:NIS

[66] B. I. Dunlap and B. R. Judd. Novel identities for simple n-j symbols (in angular momentum theory). Journal of Mathematical Physics, 16(2):318–319, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i2/p318_s1.

Jeavons:1975:CSD

[67] J. S. Jeavons, C. B. G. McIntosh, and D. K. Sen. A correction to the Sen and Dunn gravitational field equations. Journal of Mathematical Physics, 16(2):320–321, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i2/p320_s1.

Hallidy:1975:SPA

[68] William Hallidy. Some properties of asymptotically flat, static, Einstein– Maxwell space–times. Journal of Mathematical Physics, 16(2):322–325, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p322_s1.

Gray:1975:ERS

[69] D. A. Gray and C. A. Hurst. Explicit representations of a single parabose operator. Journal of Mathematical Physics, 16(2):326–333, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i2/p326_s1.

Maekawa:1975:FCR

[70] Takayoshi Maekawa. Formula for the computation of the representation matrix elements of the group SO(n). Journal of Mathematical Physics, 16(2):334–340, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i2/p334_s1. See erratum [474].

Ashtekar:1975:SEG

[71] Abhay Ashtekar and Anne Magnon. The Sagnac effect in general relativity. Journal of Mathematical Physics, 16(2):341–344, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p341_
s1.

Ray:1975:SSC

[72] John R. Ray and Everette L. Thompson. Spacetime symmetries and the complexion of the electromagnetic field. Journal of Mathematical Physics, 16(2):345–346, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i2/p345_s1.

Federbush:1975:NASa

[73] Paul Federbush. A new approach to the stability of matter problem. I. Journal of Mathematical Physics, 16(2):347–351, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p347_
s1.

Hori:1975:STEa

[74] Motoo Hori and Fumiko Yonezawa. Statistical theory of effective electrical, thermal, and magnetic properties of random heterogeneous materials. IV. Effective-medium theory and cumulant expansion method. Journal of Mathematical Physics, 16(2):352–364, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p352_s1. See

erratum [504].

Hori:1975:STEb

[75] Motoo Hori and Fumiko Yonezawa. Statistical theory of effective electrical, thermal, and magnetic properties of random heterogeneous materials. V. One- and two-dimensional systems. Journal of Mathematical Physics, 16(2):365–377, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i2/p365_s1.

Rashid:1975:SEL

[76] M. A. Rashid. Simple expression for the linear boson transformation coefficients. II. Journal of Mathematical Physics, 16(2):378–380, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i2/p378_s1.

Mahanta:1975:EEM

[77] M. N. Mahanta and A. Sudhakar Rao. On the equivalence of the Einstein–Mayer and Einstein–Cartan theories for describing a spinning medium. Journal of Mathematical Physics, 16(2):381–383, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i2/p381_s1.

Khandekar:1975:EPT

[78] D. C. Khandekar and S. V. Lawande. Exact propagator for a timedependent harmonic oscillator with and without a singular perturbation. Journal of Mathematical Physics, 16(2):384–388, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p384_
s1.

Joshi:1975:CMH

[79] P. C. Joshi. Calculation of moments in the harmonic oscillator basis. Journal of Mathematical Physics, 16(2):389–390, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p389_
s1.

Dollard:1975:DCD

[80] John D. Dollard. Disproof of a conjecture of D. Muhlerin and I. Zinnes. Journal of Mathematical Physics, 16(2):391–393, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p391_
s1.

Cornwell:1975:SSL

[81] J. F. Cornwell and J. M. Ekins. Semisimple subgroups of linear semisimple Lie groups. Journal of Mathematical Physics, 16(2):394–399, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i2/p394_s1.

Bogdanovic:1975:RGF

[82] R. Bogdanović and M. A. Whitehead. The representation of the SO(4, 1) group in four-dimensional Euclidean and spinor space. Journal of Mathematical Physics, 16(2):400–409, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i2/p400_s1.

Yamani:1975:MME

[83] Hashim A. Yamani and Louis Fishman. J-matrix method: Extensions to arbitrary angular momentum and to Coulomb scattering. Journal of Mathematical Physics, 16(2):410–420, February 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i2/p410_s1.

Kittappa:1975:ASP

[84] R. Kittappa and R. E. Kleinman. Acoustic scattering by penetrable homogeneous objects. Journal of Mathematical Physics, 16(2):421–432, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p421_s1.

Arafune:1975:THF

[85] J. Arafune, P. G. O. Freund, and C. J. Goebel. Topology of Higgs fields. Journal of Mathematical Physics, 16(2):433–437, February 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i2/p433_
s1.

Gerardi:1975:VPS

[86] A. Gerardi, N. Ianiro, C. Maffei, and C. Marchioro. A variational principle in statistical mechanics for particle systems with bounded pair interactions. Journal of Mathematical Physics, 16(2):438–441, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i2/p438_s1.

Prugovecki:1975:EMS

[87] E. Prugovečki. Erratum: “Multichannel stationary scattering theory in two-Hilbert space formulation”. Journal of Mathematical Physics, 16(2): 442, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p442_s1. See [2].

Backhouse:1975:TOT

[88] N. B. Backhouse. Tensor operators and twisted group algebras. Journal of Mathematical Physics, 16(3):443–447, March 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i3/p443_s1.

Gordon:1975:VPP

[89] William B. Gordon. Vector potentials and physical optics. Journal of Mathematical Physics, 16(3):448–454, March 1975. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i3/p448_s1.

Gerjuoy:1975:UEP

[90] Edward Gerjuoy, L. Rosenberg, and Larry Spruch. Useful extremum principle for the variational calculation of matrix elements. II. Journal of Mathematical Physics, 16(3):455–461, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i3/p455_s1.

Bessis:1975:PVA

[91] D. Bessis and M. Villani. Perturbative-variational approximations to the spectral properties of semi-bounded Hilbert space operators, based on the moment problem with finite or diverging moments. Application to quantum mechanical systems. Journal of Mathematical Physics, 16(3): 462–474, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p462_s1.

Chitre:1975:SCS

[92] D. M. Chitre, R. Güven, and Y. Nutku. Static cylindrically symmetric solutions of the Einstein–Maxwell equations. Journal of Mathematical Physics, 16(3):475–477, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i3/p475_s1.

Ardalan:1975:SFT

[93] Farhad Ardalan and G. N. Fleming. A spinor field theory on a sevendimensional homogeneous space of the Poincaré group. Journal of Mathematical Physics, 16(3):478–484, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i3/p478_s1.

Duncan:1975:BS

[94] D. P. Duncan and L. C. Shepley. Boundaries of spacetimes. Journal of Mathematical Physics, 16(3):485–492, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i3/p485_s1.

Moncrief:1975:SSL

[95] Vincent Moncrief. Spacetime symmetries and linearization stability of the Einstein equations. I. Journal of Mathematical Physics, 16(3):493– 498, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p493_s1.

Boyer:1975:LTSa

[96] C. P. Boyer, E. G. Kalnins, and W. Miller, Jr. Lie theory and separation of variables. 6. The equation iU_{t} + Δ_{2}U = 0. Journal of Mathematical Physics, 16(3):499–511, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i3/p499_s1.

Boyer:1975:LTSb

[97] C. P. Boyer, E. G. Kalnins, and W. Miller, Jr. Lie theory and separation of variables. 7. the harmonic oscillator in elliptic coordinates and Ince polynomials. Journal of Mathematical Physics, 16(3):512–517, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p512_s1.

Hajicek:1975:SESa

[98] P. Hájícek. Stationary electrovacuum spacetimes with bifurcate horizons. Journal of Mathematical Physics, 16(3):518–522, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p518_s1.

Hajicek:1975:SESb

[99] P. Hájícek. Static electrovacuum spacetimes with bifurcate horizons. Journal of Mathematical Physics, 16(3):523–527, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p523_s1.

Okubo:1975:AIA

[100] Susumu Okubo. Algebraic identities among U(n) infinitesimal generators. Journal of Mathematical Physics, 16(3):528–535, March 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p528_
s1.

Kohler:1975:WPS

[101] Werner E. Kohler. Wave propagation with stochastically coupled propagating and evanescent modes. Journal of Mathematical Physics, 16(3): 536–543, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p536_s1.

Radin:1975:SRE

[102] Charles Radin. Some remarks on the evolution of a Schrödinger particle in an attractive 1/r^{2} potential. Journal of Mathematical Physics, 16(3): 544–547, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p544_s1.

Katzin:1975:TDD

[103] Gerald H. Katzin and Jack Levine. Time-dependent dynamical symmetries, associated constants of motion, and symmetry deformations of the Hamiltonian in classical particle systems. Journal of Mathematical Physics, 16(3):548–555, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i3/p548_s1.

Hestenes:1975:OOC

[104] David Hestenes. Observables, operators, and complex numbers in the Dirac theory. Journal of Mathematical Physics, 16(3):556–572, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p556_s1.

Gurtler:1975:CFD

[105] R. Gurtler and D. Hestenes. Consistency in the formulation of the Dirac, Pauli, and Schrödinger theories. Journal of Mathematical Physics, 16(3): 573–584, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p573_s1.

Belleni-Morante:1975:CKE

[106] A. Belleni-Morante. On Chapman–Kolmogorov equations for neutron population in a multiplying assembly. Journal of Mathematical Physics, 16(3):585–589, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i3/p585_s1.

Whicker:1975:WPM

[107] Donald Whicker and George Lianis. Wave propagation in media undergoing uniform linear acceleration. Journal of Mathematical Physics, 16(3): 590–597, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p590_s1.

Ablowitz:1975:NDD

[108] M. J. Ablowitz and J. F. Ladik. Nonlinear differential-difference equations. Journal of Mathematical Physics, 16(3):598–603, March 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p598_
s1.

Matsuno:1975:EOE

[109] Koichiro Matsuno. Ergodicity of observable and ergodic hypothesis in Markovian kinetics. Journal of Mathematical Physics, 16(3):604– 608, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p604_s1.

Roberts:1975:FOS

[110] P. H. Roberts and A. M. Soward. On first order smoothing theory. Journal of Mathematical Physics, 16(3):609–615, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p609_s1.

Fuchs:1975:EDN

[111] V. Fuchs and G. Beaudry. Effect of damping on nonlinear three-wave interaction. Journal of Mathematical Physics, 16(3):616–619, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p616_s1.

Bowen:1975:AHS

[112] Samuel P. Bowen. The abstract Hilbert space generalization of Feenberg’s perturbation theory—A new method of quantum field theory. Journal of Mathematical Physics, 16(3):620–632, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i3/p620_s1.

Twersky:1975:MSW

[113] Victor Twersky. Multiple scattering of waves by the doubly periodic planar array of obstacles. Journal of Mathematical Physics, 16(3):633– 643, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p633_s1.

Twersky:1975:LSS

[114] Victor Twersky. Lattice sums and scattering coefficients for the rectangular planar array. Journal of Mathematical Physics, 16(3):644–657, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p644_s1.

Twersky:1975:LFC

[115] Victor Twersky. Low frequency coupling in the planar rectangular lattice. Journal of Mathematical Physics, 16(3):658–666, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p658_s1.

Gorini:1975:RDP

[116] Vittorio Gorini and Antonio Zecca. Reversible dynamics in a propositionstate structure. Journal of Mathematical Physics, 16(3):667–669, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p667_s1.

dInverno:1975:CPGa

[117] R. A. d’Inverno. The Cauchy problem in general relativity. I. Metrics containing arbitrary functions. Journal of Mathematical Physics, 16(3): 670–673, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p670_s1.

dInverno:1975:CPGb

[118] R. A. d’Inverno. The Cauchy problem in general relativity. II. Main and constraint equations in the characteristic initial-value problem. Journal of Mathematical Physics, 16(3):674–678, March 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i3/p674_s1.

Malin:1975:WDE

[119] S. Malin. The Weyl and Dirac equations in terms of functions over the group SU_{2}. Journal of Mathematical Physics, 16(3):679–684, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p679_s1.

Atkinson:1975:PSF

[120] D. Atkinson, M. Kaekebeke, and M. de Roo. Phase shifts as functions of the cross section. Journal of Mathematical Physics, 16(3):685–693, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p685_s1.

Eggarter:1975:PCT

[121] T. P. Eggarter. Percolation on a Cayley tree with second nearest neighbor bonds. Journal of Mathematical Physics, 16(3):694–695, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p694_s1.

Pavelle:1975:CVK

[122] Richard Pavelle. Conserved vectors of the Komar type and compatibility identities in Lagrangian field theories. Journal of Mathematical Physics, 16(3):696–702, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i3/p696_s1.

Nauenberg:1975:RGS

[123] M. Nauenberg. Renormalization group solution of the one-dimensional Ising model. Journal of Mathematical Physics, 16(3):703–705, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p703_s1.

Federbush:1975:NASb

[124] Paul Federbush. A new approach to the stability of matter problem. II. Journal of Mathematical Physics, 16(3):706–709, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/p706_s1.

Benard:1975:DQD

[125] C. Bénard. A definition of quantum detection point processes. Journal of Mathematical Physics, 16(3):710–719, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i3/p710_s1.

Goovaerts:1975:PIE

[126] M. J. Goovaerts. Path-integral evaluation of a nonstationary Calogero model. Journal of Mathematical Physics, 16(3):720–723, March 1975.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i3/
p720_s1.

Liu:1975:ETT

[127] Kuang Chi Liu. An expansion theorem for the twisted product with applications. Journal of Mathematical Physics, 16(3):724–728, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i3/p724_s1.

Schroeck:1975:MQF

[128] Franklin E. Schroeck, Jr. A model of quantum field theory treated in the Fock–Cook formalism. II. Journal of Mathematical Physics, 16(3): 729–739, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p729_s1.

Sneddon:1975:CPT

[129] G. E. Sneddon. Change of Petrov type under generation of solutions of Einstein’s equations. Journal of Mathematical Physics, 16(3):740– 742, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i3/p740_s1.

Mastalir:1975:TRPa

[130] Robert O. Mastalir. Theory of Regge poles for 1/r^{2} potentials. I. Journal of Mathematical Physics, 16(4):743–748, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p743_s1.

Mastalir:1975:TRPb

[131] Robert O. Mastalir. Theory of Regge poles for 1/r^{2} potentials. II. An exactly solvable example at zero energy. Journal of Mathematical Physics, 16(4):749–751, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p749_s1.

Mastalir:1975:TRPc

[132] Robert O. Mastalir. Theory of Regge poles for 1/r^{2} potentials. III. An exact solution of Schrödinger’s equation for arbitrary I and E. Journal of Mathematical Physics, 16(4):752–755, April 1975. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p752_s1.

Sullivan:1975:DCA

[133] John J. Sullivan. Double coset analysis for symmetry adapting Nth rank tensors of U(n) to its unitary subgroups. Journal of Mathematical Physics, 16(4):756–760, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p756_s1. See erratum [315].

Gerjuoy:1975:SOE

[134] Edward Gerjuoy. Second order error in variational calculation of matrix elements. Journal of Mathematical Physics, 16(4):761–765, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p761_s1.

Tjostheim:1975:NUD

[135] D. Tjøstheim. A note on the unified Dirac–von Neumann formulation of quantum mechanics. Journal of Mathematical Physics, 16(4):766– 767, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p766_s1.

Hendrickson:1975:TIS

[136] Joseph H. Hendrickson. Temporally inhomogeneous scattering theory for modified wave operators. Journal of Mathematical Physics, 16(4): 768–771, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p768_s1.

Bose:1975:SKN

[137] S. K. Bose. Studies in the Kerr–Newman metric. Journal of Mathematical Physics, 16(4):772–775, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p772_s1.

Greene:1975:MAK

[138] David E. Greene. Mathematical aspects of kinetic model equations for binary gas mixtures. Journal of Mathematical Physics, 16(4):776–782, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p776_s1.

Ezawa:1975:VES

[139] H. Ezawa, J. R. Klauder, and L. A. Shepp. Vestigial effects of singular potentials in diffusion theory and quantum mechanics. Journal of Mathematical Physics, 16(4):783–799, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i4/p783_s1.

Wong:1975:MEG

[140] M. K. F. Wong and Hsin-Yang Yeh. Matrix elements of the generators of IU(n) and IO(n) and respective deformations to U(n, 1) and O(n, 1). Journal of Mathematical Physics, 16(4):800–803, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i4/p800_s1.

Hansen:1975:KIR

[141] R. O. Hansen and Jeffrey Winicour. Killing inequalities for relativistically rotating fluids. Journal of Mathematical Physics, 16(4):804–808, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p804_s1.

Inoue:1975:LGFb

[142] Michiko Inoue. Lattice Green’s function for the body-centred cubic lattice. Journal of Mathematical Physics, 16(4):809–812, April 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i4/p809_
s1.

Baker:1975:CPA

[143] George A. Baker, Jr. Convergence of Padé approximants using the solution of linear functional equations. Journal of Mathematical Physics, 16(4):813–822, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p813_s1.

Tortorella:1975:EUP

[144] Michael Tortorella. Experimental uncertainties in the problem of the unitarity equation. Journal of Mathematical Physics, 16(4):823–828, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p823_s1.

McLenaghan:1975:CFS

[145] R. G. McLenaghan, N. Tariq, and B. O. J. Tupper. Conformally flat solutions of the Einstein–Maxwell equations for null electromagnetic fields. Journal of Mathematical Physics, 16(4):829–831, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i4/p829_s1.

Provost:1975:AAI

[146] J. P. Provost, F. Rocca, and G. Vallée. Algebraic approach of the infrared-problem for external currents. Journal of Mathematical Physics, 16(4):832–833, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p832_s1.

Reina:1975:NLG

[147] C. Reina and A. Treves. NUT-like generalization of axisymmetric gravitational fields. Journal of Mathematical Physics, 16(4):834–835, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p834_s1.

Rosen:1975:NDS

[148] Gerald Rosen. Nonexistence of dissipative structure solutions to Volterra many-species models. Journal of Mathematical Physics, 16(4):836, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p836_s1.

Lubkin:1975:SEM

[149] Elihu Lubkin. Symmetry of ensembles of maximum entropy. Journal of Mathematical Physics, 16(4):837–839, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p837_s1.

Alabiso:1975:VRA

[150] C. Alabiso and P. Butera. N-variable rational approximants and method of moments. Journal of Mathematical Physics, 16(4):840–845, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p840_s1.

Habetler:1975:UAE

[151] G. J. Habetler and B. J. Matkowsky. Uniform asymptotic expansions in transport theory with small mean free paths, and the diffusion approximation. Journal of Mathematical Physics, 16(4):846–854, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p846_s1.

Schutz:1975:SSP

[152] Bernard F. Schutz. On the strength of a system of partial differential equations. Journal of Mathematical Physics, 16(4):855–856, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p855_s1.

Couch:1975:NSI

[153] W. E. Couch and R. J. Torrence. Nonspreading solutions of the inhomogeneous scalar wave equation. Journal of Mathematical Physics, 16(4): 857–861, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p857_s1. See erratum [349].

Huffaker:1975:FMT

[154] J. N. Huffaker and P. H. Dwivedi. Factorization-method treatment of the perturbed Morse oscillator. Journal of Mathematical Physics, 16(4): 862–867, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p862_s1.

Garrod:1975:VAT

[155] C. Garrod, M. V. Mihailović, and M. Rosina. The variational approach to the two-body density matrix. Journal of Mathematical Physics, 16(4): 868–874, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p868_s1.

Fulling:1975:AES

[156] S. A. Fulling. Adiabatic expansions of solutions of coupled second-order linear differential equations. I. Journal of Mathematical Physics, 16(4): 875–883, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p875_s1.

Mayer:1975:VTF

[157] D. H. Mayer. Vector and tensor fields on conformal space. Journal of Mathematical Physics, 16(4):884–893, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p884_s1.

Ladde:1975:SDI

[158] G. S. Ladde. Systems of differential inequalities and stochastic differential equations. II. Journal of Mathematical Physics, 16(4):894–900, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p894_s1.

Lukierski:1975:GFF

[159] J. Lukierski and W. Sienkiewicz. Generalized free fields and the representations of Weyl group. II. Reducible representations. Journal of Mathematical Physics, 16(4):901–904, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p901_s1.

Schultheis:1975:EGI

[160] H. Schultheis and R. Schultheis. Electrostatic or gravitational interaction energy of coaxial ellipsoids. Journal of Mathematical Physics, 16(4): 905–909, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p905_s1.

Comi:1975:STS

[161] M. Comi, L. Lanz, L. A. Lugiato, and G. Ramella. Space–time symmetries and nonunitary representations of the Poincaré semigroup: Description of unstable particles. Journal of Mathematical Physics, 16(4): 910–917, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p910_s1.

Barnsley:1975:CTP

[162] M. F. Barnsley. Correction terms for Padé approximants. Journal of Mathematical Physics, 16(4):918–928, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p918_s1.

Abbati:1975:LGH

[163] Maria C. Abbati, Renzo Cirelli, and Franco Gallone. From a “laboratory” Galilei–Hilbert bundle to an algebra of observables. Journal of Mathematical Physics, 16(4):929–937, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p929_s1.

deLorenzi:1975:IEF

[164] H. G. de Lorenzi and H. F. Tiersten. On the interaction of the electromagnetic field with heat conducting deformable semiconductors. Journal of Mathematical Physics, 16(4):938–957, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i4/p938_s1.

Melnick:1975:ESE

[165] Jorge Melnick and Romualdo Tabensky. Exact solutions to Einstein field equations. Journal of Mathematical Physics, 16(4):958–960, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p958_s1.

Aly:1975:SSP

[166] H. H. Aly, H. J. W. Müller-Kirsten, and N. Vahedi-Faridi. Scattering by singular potentials with a perturbation-theoretical introduction to Mathieu functions. Journal of Mathematical Physics, 16(4):961–970, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v16/i4/p961_s1. See erratum [625].

DeFacio:1975:HCM

[167] B. DeFacio. Heat conduction model with finite signal speed. Journal of Mathematical Physics, 16(4):971–974, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v16/i4/p971_s1. See erratum [350].

Doebner:1975:QMH

[168] H. D. Doebner and J. Tolar. Quantum mechanics on homogeneous spaces. Journal of Mathematical Physics, 16(4):975–984, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i4/p975_s1.

Bisognano:1975:DCH

[169] Joseph J. Bisognano and Eyvind H. Wichmann. On the duality condition for a Hermitian scalar field. Journal of Mathematical Physics, 16(4): 985–1007, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i4/p985_s1.

Brans:1975:SRA

[170] Carl H. Brans. Some restrictions on algebraically general vacuum metrics. Journal of Mathematical Physics, 16(4):1008–1010, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i4/p1008_s1.

Ernst:1975:ECP

[171] Frederick J. Ernst. Erratum: “Complex potential formulation of the axially symmetric gravitational field problem”. Journal of Mathematical Physics, 16(4):1011, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p1011_s1. See [9].

Yao:1975:EOI

[172] Te Hai Yao. Erratum: “The T operator and an inverse problem for nonlocal potentials”. Journal of Mathematical Physics, 16(4):1011, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p1011_s2. See [8].

Davidson:1975:RQS

[173] Russell Davidson and John J. Kozak. On the relaxation to quantumstatistical equilibrium of the Wigner–Weisskopf atom in a onedimensional radiation field. VII. Emission in a finite system in the presence of an extra photon. Journal of Mathematical Physics, 16(5):1013– 1022, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i5/p1013_s1.

Camenzind:1975:CDM

[174] M. Camenzind. On the curvature dynamics for metric gravitational theories. Journal of Mathematical Physics, 16(5):1023–1031, May 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/p1023_
s1.

Kiehn:1975:CIH

[175] R. M. Kiehn. Conformal invariance and Hamilton Jacobi theory for dissipative systems. Journal of Mathematical Physics, 16(5):1032–1033, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1032_s1.

Truong:1975:WQA

[176] T. T. Truong. Weyl quantization of anharmonic oscillators. Journal of Mathematical Physics, 16(5):1034–1043, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i5/p1034_s1.

Moses:1975:IEG

[177] H. E. Moses. An integral equation for the Gel’fand–Levitan kernel in terms of the scattering potential in the one-dimensional case. Journal of Mathematical Physics, 16(5):1044–1046, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i5/p1044_s1.

Herrick:1975:VDG

[178] David R. Herrick. Variable dimensionality in the group-theoretic prediction of configuration mixings for doubly-excited helium. Journal of Mathematical Physics, 16(5):1047–1053, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i5/p1047_s1.

Ablowitz:1975:IST

[179] M. J. Ablowitz and H. Segur. The inverse scattering transform: Semiinfinite interval. Journal of Mathematical Physics, 16(5):1054–1056, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1054_s1.

McDonald:1975:TRS

[180] J. F. McDonald and L. D. Favro. On the transformation from random Schrödinger Hamiltonians to random Hamiltonian matrices. Journal of Mathematical Physics, 16(5):1057–1064, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i5/p1057_s1.

Park:1975:LAF

[181] Yong Moon Park. Lattice approximation of the (λϕ^{4} − μϕ)_{3} field theory in a finite volume. Journal of Mathematical Physics, 16(5):1065– 1075, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i5/p1065_s1.

Campbell:1975:LCE

[182] C. E. Campbell. Linked cluster expansions in the density fluctuation formulation of the many-body problem. Journal of Mathematical Physics, 16(5):1076–1086, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i5/p1076_s1.

Onofri:1975:NCS

[183] Enrico Onofri. A note on coherent state representations of Lie groups. Journal of Mathematical Physics, 16(5):1087–1089, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/p1087_s1.

Krause:1975:DCCa

[184] Jorge Krause. The de Donder coordinate condition and minimal class 1 space–time. Journal of Mathematical Physics, 16(5):1090–1092, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1090_s1.

Demmie:1975:BNN

[185] Paul N. Demmie. Bifurcate nondiverging null hypersurfaces and trapped surfaces. Journal of Mathematical Physics, 16(5):1093–1103, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/
p1093_s1.

Gerjuoy:1975:VPV

[186] E. Gerjuoy, A. R. P. Rau, L. Rosenberg, and Larry Spruch. Variational principles, variational identities, and supervariational principles for wavefunctions. Journal of Mathematical Physics, 16(5):1104–1111, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1104_s1.

Brunet:1975:PCR

[187] Robert C. Brunet. Pyramidal composition rules for Wronskians upon Wronskians. Journal of Mathematical Physics, 16(5):1112–1116, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1112_s1.

Larsen:1975:ECF

[188] E. W. Larsen, S. Sancaktar, and P. F. Zweifel. Extension of the Case formulas to L_{p}. Application to half and full space problems. Journal of Mathematical Physics, 16(5):1117–1121, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i5/p1117_s1.

Chadam:1975:GES

[189] J. M. Chadam and R. T. Glassey. Global existence of solutions to the Cauchy problem for time-dependent Hartree equations. Journal of Mathematical Physics, 16(5):1122–1130, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i5/p1122_s1.

Bhaumik:1975:CNO

[190] K. Bhaumik and B. Dutta-Roy. The classical nonlinear oscillator and the coherent state. Journal of Mathematical Physics, 16(5):1131–1133, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1131_s1.

Pao:1975:IMS

[191] C. V. Pao. An iterative method for solving a two-point boundary value problem. Journal of Mathematical Physics, 16(5):1134–1138, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/
p1134_s1.

Tripathy:1975:SCK

[192] K. C. Tripathy, Rakesh Gupta, and J. D. Anand. O(4, 2) symmetry and the classical Kepler problem. Journal of Mathematical Physics, 16(5): 1139–1144, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i5/p1139_s1.

Yaakobi:1975:RAQ

[193] Daniel Yaakobi. A relativistic approach to quadrupole effects on planetary orbits. Journal of Mathematical Physics, 16(5):1145–1147, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1145_s1.

deMottoni:1975:EMR

[194] P. de Mottoni and A. Tesei. An explicit model for the renormalization of field equations. Journal of Mathematical Physics, 16(5):1148– 1149, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i5/p1148_s1.

Chen:1975:EOZ

[195] M. Chen. On the equivalence of the Ornstein–Zernike relation and Baxter’s relations for a one-dimensional simple fluid. Journal of Mathematical Physics, 16(5):1150–1151, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i5/p1150_s1.

Cohn:1975:GNT

[196] Jack Cohn. Geodesics and Newtonian trajectories. Journal of Mathematical Physics, 16(5):1152–1155, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i5/p1152_s1.

Davies:1975:LWN

[197] John A. Davies and Carol L. Mainville. Long-wavelength normal mode vibrations of infinite, ionic crystal lattices. Journal of Mathematical Physics, 16(5):1156–1171, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i5/p1156_s1.

Gambardella:1975:ERQ

[198] Pascal J. Gambardella. Exact results in quantum many-body systems of interacting particles in many dimensions with S̄U(1, 1) as the dynamical group. Journal of Mathematical Physics, 16(5):1172–1187, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1172_s1.

Banerjee:1975:CSS

[199] A. Banerjee. Cylindrically symmetric stationary beam of electromagnetic radiation. Journal of Mathematical Physics, 16(5):1188–1190, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i5/p1188_s1.

Arens:1975:MHS

[200] Richard Arens. Models of Hamiltonian several-particle interactions. Journal of Mathematical Physics, 16(5):1191–1198, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/p1191_s1.

Pavelle:1975:CVD

[201] Richard Pavelle. Conserved vector densities and their curl expressions. Journal of Mathematical Physics, 16(5):1199–1200, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/p1199_s1.

Capri:1975:EEW

[202] A. Z. Capri. Erratum: “On the existence of weakly retarded and advanced Green’s functions”. Journal of Mathematical Physics, 16(5): 1201, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i5/p1201_s1. See [10].

Bummerstede:1975:NEL

[203] J. Bümmerstede and W. Lücke. The notion of essential locality for nonlocalizable fields. Journal of Mathematical Physics, 16(6):1203–1209, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i6/p1203_s1.

Vroegindeweij:1975:STL

[204] P. G. Vroegindeweij. The separating topology for the Lorentz group L. Journal of Mathematical Physics, 16(6):1210–1213, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/p1210_s1.

Fox:1975:PAM

[205] Ronald Forrest Fox. Physical applications of multiplicative stochastic processes. III. Nonequilibrium entropy. Journal of Mathematical Physics, 16(6):1214–1218, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i6/p1214_s1.

Davis:1975:SPP

[206] L. C. Davis and J. R. Reitz. Solution of potential problems near the corner of a conductor. Journal of Mathematical Physics, 16(6):1219– 1226, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i6/p1219_s1.

Campesino-Romeo:1975:CSQ

[207] Enrique Campesino-Romeo and John R. Taylor. Cross sections in quantum mechanics. Journal of Mathematical Physics, 16(6):1227–1230, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i6/p1227_s1.

Berendt:1975:HOG

[208] G. Berendt, E. Weimar, and R. Gilmore. Harmonic oscillator Green’s function from a BCH formula. Journal of Mathematical Physics, 16(6): 1231–1233, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i6/p1231_s1.

Krause:1975:DCCb

[209] Jorge Krause. The de Donder coordinate condition and minimal class 1 space–time. Journal of Mathematical Physics, 16(6):1234–1236, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i6/p1234_s1.

Glasser:1975:ELS

[210] M. L. Glasser. Evaluation of lattice sums. IV. A five-dimensional sum. Journal of Mathematical Physics, 16(6):1237–1238, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/p1237_s1.

Wong:1975:EIO

[211] M. K. F. Wong and Hsin-Yang Yeh. Eigenvalues of the invariant operators of the orthogonal and symplectic groups. Journal of Mathematical Physics, 16(6):1239–1243, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i6/p1239_s1.

Oguztoreli:1975:GPF

[212] M. N. Oguztöreli. A Goursat problem for the fourth moment equation. Journal of Mathematical Physics, 16(6):1244–1246, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/p1244_s1.

Klink:1975:TAR

[213] W. H. Klink. Two applications of the Racah coefficients of the Poincaré group. Journal of Mathematical Physics, 16(6):1247–1252, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/
p1247_s1.

Ebrahim:1975:GCP

[214] Abdul Ebrahim. General chiral SU_{2} × SU_{2} pion Lagrangian and the generators of O(4, 1) group. Journal of Mathematical Physics, 16(6): 1253–1255, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i6/p1253_s1.

Mehta:1975:DCG

[215] M. L. Mehta and G. C. Mehta. Discrete Coulomb gas in one dimension: Correlation functions. Journal of Mathematical Physics, 16(6): 1256–1258, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i6/p1256_s1.

Hioe:1975:TLR

[216] F. T. Hioe and E. W. Montroll. Two-level radiative systems and perturbation theory. Journal of Mathematical Physics, 16(6):1259–1270, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i6/p1259_s1.

Finkelstein:1975:GRF

[217] Robert J. Finkelstein. The general relativistic fields of a charged rotating source. Journal of Mathematical Physics, 16(6):1271–1277, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i6/p1271_s1.

Aronowitz:1975:PMG

[218] S. Aronowitz and B. E. Eichinger. Petrie matrices and generalized inverses. Journal of Mathematical Physics, 16(6):1278–1283, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/
p1278_s1.

Lebowitz:1975:SES

[219] Joel L. Lebowitz, Michael Aizemann, and Sheldon Goldstein. On the stability of equilibrium states of finite classical systems. Journal of Mathematical Physics, 16(6):1284–1287, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i6/p1284_s1.

Coster:1975:OTT

[220] Joseph Coster and Henry P. Stapp. Optical theorems for three-to-three processes. Journal of Mathematical Physics, 16(6):1288–1323, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/
p1288_s1.

Baker:1975:SIB

[221] George A. Baker, Jr. Self-interacting, boson, quantum, field theory and the thermodynamic limit in d-dimensions. Journal of Mathematical Physics, 16(6):1324–1346, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i6/p1324_s1.

Greub:1975:MCC

[222] Werner Greub and Herbert-Rainer Petry. Minimal coupling and complex line bundles. Journal of Mathematical Physics, 16(6):1347–1351, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i6/p1347_s1.

Wollenberg:1975:HSC

[223] L. S. Wollenberg. Higher symmetry of certain classical quadratic Hamiltonians. Journal of Mathematical Physics, 16(6):1352–1357, June 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/
p1352_s1.

Golden:1975:EMP

[224] Sidney Golden. Erratum: “A minimum principle for von Neumann’s equation”. Journal of Mathematical Physics, 16(6):1358, June 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/p1358_

s2. See [41].

Newton:1975:ADA

[225] Roger G. Newton. Addendum: “Determination of the amplitude from the differential cross section by unitarity”, [J. Math. Phys. 9, 2050 (1968)]. Journal of Mathematical Physics, 16(6):1358, June 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i6/p1358_s1. See [1].

Lee:1975:TIR

[226] Jon Lee. The triad-interaction representation of homogeneous turbulence. Journal of Mathematical Physics, 16(7):1359–1366, July 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i7/p1359_
s1.

Lee:1975:ICM

[227] Jon Lee. Isolating constants of motion for the homogeneous turbulence of two and three dimensions. Journal of Mathematical Physics, 16(7): 1367–1373, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1367_s1.

Logan:1975:ITS

[228] J. D. Logan and J. S. Blakeslee. An invariance theory for second-order variational problems. Journal of Mathematical Physics, 16(7):1374–1379, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i7/p1374_s1.

Tabensky:1975:SCR

[229] R. Tabensky and D. Villarroel. Simplified calculations for radiation reaction forces. Journal of Mathematical Physics, 16(7):1380–1382, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i7/p1380_s1.

Balanis:1975:ADL

[230] George N. Balanis. Analysis of the dispersion of low frequency uniaxial waves in heterogeneous periodic elastic media. Journal of Mathematical Physics, 16(7):1383–1387, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i7/p1383_s1. See addendum [624].

Fuchs:1975:ILD

[231] V. Fuchs. The influence of linear damping on nonlinearly coupled positive and negative energy waves. Journal of Mathematical Physics, 16(7): 1388–1392, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1388_s1.

Lathouwers:1975:HR

[232] L. Lathouwers. The Hamiltonian H = (−1/2)d^{2}/dx^{2} + x^{2}/2 + λ/x^{2} reobserved. Journal of Mathematical Physics, 16(7):1393–1395, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i7/p1393_s1.

Kelemen:1975:CIT

[233] A. Kelemen. On convergent iterations and their applications to Hartree– Fock equations. Journal of Mathematical Physics, 16(7):1396–1401, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i7/p1396_s1.

Cox:1975:CMM

[234] Joseph R. Cox and Hugo R. Garcia. Construction of a meromorphic many-channel p-wave S matrix. Journal of Mathematical Physics, 16(7): 1402–1409, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1402_s1.

Cox:1975:AMC

[235] Joseph R. Cox. On angular momentum and channel coupling for a meromorphic many-channel S matrix. Journal of Mathematical Physics, 16(7):1410–1415, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i7/p1410_s1.

Carinena:1975:PUR

[236] Jose F. Cariñena and Mariano Santander. On the projective unitary representations of connected Lie groups. Journal of Mathematical Physics, 16(7):1416–1420, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i7/p1416_s1.

Larsen:1975:NTD

[237] Edward W. Larsen. Neutron transport and diffusion in inhomogeneous media. I. Journal of Mathematical Physics, 16(7):1421–1427, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i7/
p1421_s1.

Chen:1975:BTS

[238] Hsing-Hen Chen and Chuan-Sheng Liu. Bäcklund transformation solutions of the Toda lattice equation. Journal of Mathematical Physics, 16(7):1428–1430, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i7/p1428_s1.

Eris:1975:HMR

[239] Ahmet Eris and Yavuz Nutku. Harmonic mappings of Riemannian manifolds and stationary vacuum space–times with whole cylinder symmetry. Journal of Mathematical Physics, 16(7):1431–1434, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i7/p1431_s1.

Case:1975:OPI

[240] K. M. Case. Orthogonal polynomials. II. Journal of Mathematical Physics, 16(7):1435–1440, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i7/p1435_s1.

vanHaeringen:1975:AMC

[241] H. van Haeringen and R. van Wageningen. Analytic T matrices for Coulomb plus rational separable potentials. Journal of Mathematical Physics, 16(7):1441–1452, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i7/p1441_s1.

Forte:1975:SMB

[242] B. Forte and C. C. A. Sastri. Is something missing in the Boltzmann entropy? Journal of Mathematical Physics, 16(7):1453–1456, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i7/
p1453_s1.

Chaba:1975:ECL

[243] A. N. Chaba and R. K. Pathria. Evaluation of a class of lattice sums in arbitrary dimensions. Journal of Mathematical Physics, 16(7):1457– 1460, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1457_s1.

Havas:1975:SVHa

[244] Peter Havas. Separation of variables in the Hamilton–Jacobi, Schrödinger, and related equations. I. complete separation. Journal of Mathematical Physics, 16(7):1461–1468, July 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i7/p1461_s1.

Contopoulos:1975:IMR

[245] G. Contopoulos and L. Vlahos. Integrals of motion and resonances in a dipole magnetic field. Journal of Mathematical Physics, 16(7):1469– 1474, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1469_s1.

Balaban:1975:SQC

[246] T. Balaban and R. Raczka. Second quantization of classical nonlinear relativistic field theory. I. Canonical formalism. Journal of Mathematical Physics, 16(7):1475–1481, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v16/i7/p1475_s1.

Klein:1975:SKS

[247] William Klein. Spectrum of the Kirkwood–Salsburg operator and phase transitions. Journal of Mathematical Physics, 16(7):1482–1487, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i7/p1482_s1.

Letelier:1975:SGF

[248] P. S. Letelier. Self-gravitating fluids with cylindrical symmetry. Journal of Mathematical Physics, 16(7):1488–1490, July 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i7/p1488_s1.

Avishai:1975:STB

[249] Y. Avishai. Solution of the three-body problem with inverse square potentials. Journal of Mathematical Physics, 16(7):1491–1492, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i7/
p1491_s1.

Boyer:1975:CTI

[250] Charles P. Boyer and Kurt Bernardo Wolf. Canonical transforms. III. configuration and phase descriptions of quantum systems possessing an sl(2, R) dynamical algebra. Journal of Mathematical Physics, 16(7):1493– 1502, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1493_s1.

Pauri:1975:CRP

[251] M. Pauri and G. M. Prosperi. Canonical realizations of the Poincaré group. I. General theory. Journal of Mathematical Physics, 16(7):1503– 1521, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1503_s1.

Ronen:1975:BLI

[252] Yigal Ronen. Boundedness of linear integral operators in weighted L_{p} spaces. Journal of Mathematical Physics, 16(7):1522–1523, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i7/
p1522_s1.

Hanckowiak:1975:SMQ

[253] Jerzy Hańćkowiak. Statistical methods in quantum field theory. Journal of Mathematical Physics, 16(7):1524–1527, July 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i7/p1524_s1.

Rao:1975:SRC

[254] K. Srinivasa Rao, T. S. Santhanam, and K. Venkatesh. On the symmetries of the Racah coefficient. Journal of Mathematical Physics, 16(7): 1528–1530, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i7/p1528_s1.

Kalnins:1975:ELT

[255] E. G. Kalnins and W. Miller. Erratum: “Lie theory and separation of variables. 3. The equation f_{tt} − f_{ss} = γ^{2}f”. Journal of Mathematical Physics, 16(7):1531, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i7/p1531_s1. See [6].

Osborn:1975:TBT

[256] T. A. Osborn and D. Bollé. The three-body time-delay operator. Journal of Mathematical Physics, 16(8):1533–1549, August 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1533_s1.

Moraal:1975:ICT

[257] H. Moraal. Irreducible Cartesian tensor expansions of scalar fields. Journal of Mathematical Physics, 16(8):1550–1555, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1550_s1.

Moncrief:1975:DGP

[258] Vincent Moncrief. Decompositions of gravitational perturbations. Journal of Mathematical Physics, 16(8):1556–1560, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1556_s1.

Maltman:1975:OSD

[259] Kim Maltman and W. G. Laidlaw. Onsager symmetry and the diagonalizability of the hydrodynamic matrix. Journal of Mathematical Physics, 16(8):1561–1568, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i8/p1561_s1.

Perry:1975:DBW

[260] W. L. Perry and C. D. Luning. On the density of the Breit–Wigner functions. Journal of Mathematical Physics, 16(8):1569–1572, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i8/p1569_s1.

Zielke:1975:NSE

[261] Walter Zielke. A nonlinear system of Euler–Lagrange equations. Reduction to the Korteweg–de Vries equation and periodic solutions. Journal of Mathematical Physics, 16(8):1573–1579, August 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1573_s1.

Abellanas:1975:GSC

[262] L. Abellanas and L. Martinez Alonso. A general setting for Casimir invariants. Journal of Mathematical Physics, 16(8):1580–1584, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i8/p1580_s1.

Hostler:1975:ODE

[263] Levere Hostler. One-dimensional excited state reduced Coulomb Green’s function. Journal of Mathematical Physics, 16(8):1585–1596, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i8/p1585_s1.

Patera:1975:CSFa

[264] J. Patera, P. Winternitz, and H. Zassenhaus. Continuous subgroups of the fundamental groups of physics I. General method and the Poincaré group. Journal of Mathematical Physics, 16(8):1597–1614, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i8/
p1597_s1.

Patera:1975:CSFb

[265] J. Patera, P. Winternitz, and H. Zassenhaus. Continuous subgroups of the fundamental groups of physics II. The similitude group. Journal of Mathematical Physics, 16(8):1615–1624, August 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1615_s1.

Hauser:1975:SEKb

[266] I. Hauser and R. J. Malhiot. Structural equations for Killing tensors of order two. II. Journal of Mathematical Physics, 16(8):1625–1629, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i8/p1625_s1.

Hsieh:1975:VMNb

[267] Din-Yu Hsieh. Variational method and nonlinear oscillations and waves. Journal of Mathematical Physics, 16(8):1630–1636, August 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1630_
s1.

Torruella:1975:RTD

[268] A. J. Torruella. Representations of the three-dimensional rotation group by the direct method. Journal of Mathematical Physics, 16(8):1637– 1642, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i8/p1637_s1.

Hjalmars:1975:SEM

[269] Stig Hjalmars. The symmetrical energy-momentum tensor derived by parametrization. Journal of Mathematical Physics, 16(8):1643–1645, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i8/p1643_s1.

Scull:1975:HSS

[270] Sidney C. Scull. Heisenberg subgroups of semisimple Lie groups. Journal of Mathematical Physics, 16(8):1646–1648, August 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1646_s1.

Brydges:1975:BBF

[271] David Brydges. Boundedness below for fermion model theories. I. Journal of Mathematical Physics, 16(8):1649–1661, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1649_s1.

Duneau:1975:DCI

[272] M. Duneau, B. Souillard, and D. Iagolnitzer. Decay of correlations for infinite-range interactions. Journal of Mathematical Physics, 16(8):1662– 1666, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i8/p1662_s1.

Dietz:1975:CEM

[273] David Dietz and William Greenberg. The conditional entropy in the microcanonical ensemble. Journal of Mathematical Physics, 16(8):1667– 1671, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i8/p1667_s1.

Srinivas:1975:FQP

[274] M. D. Srinivas. Foundations of a quantum probability theory. Journal of Mathematical Physics, 16(8):1672–1685, August 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i8/p1672_s1.

Maia:1975:ESG

[275] M. D. Maia. Elementary systems in general relativity. Journal of Mathematical Physics, 16(8):1686–1691, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i8/p1686_s1.

Burdet:1975:RCE

[276] G. Burdet and M. Perrin. Realizations of the central extension of the inhomogeneous symplectic algebra as time dependent invariance algebras of nonrelativistic quantum systems. Journal of Mathematical Physics, 16(8):1692–1703, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i8/p1692_s1.

McLaughlin:1975:EFE

[277] D. W. McLaughlin. Erratum: “Four examples of the inverse method as a canonical transformation”. Journal of Mathematical Physics, 16(8): 1704, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i8/p1704_s1. See [30].

Sullivan:1975:OPD

[278] John J. Sullivan. Orthonormality properties of double coset matrix elements. Journal of Mathematical Physics, 16(9):1707–1709, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i9/p1707_s1.

Gautrin:1975:GRB

[279] Henri-François Gautrin. Generalized P -representation for bounded operators. Journal of Mathematical Physics, 16(9):1710–1713, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i9/p1710_s1.

Cawley:1975:JPC

[280] Robert G. Cawley. Jacobi’s principle for the case of time-varying Hamiltonian. Journal of Mathematical Physics, 16(9):1714–1718, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i9/p1714_s1.

Hoenders:1975:SPR

[281] B. J. Hoenders. On the solution of the phase retrieval problem. Journal of Mathematical Physics, 16(9):1719–1725, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1719_s1.

Mano:1975:RGF

[282] Koichi Mano. Remarks on the Green’s functions for the cubic lattices. Journal of Mathematical Physics, 16(9):1726–1728, September 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1726_
s1.

Fanelli:1975:CTP

[283] Robert Fanelli. Canonical transformations and path integrals. Journal of Mathematical Physics, 16(9):1729–1732, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1729_s1.

Schlup:1975:VPB

[284] W. A. Schlup. A variational principle for the Boltzmann equation for hot electrons in a semiconductor. Journal of Mathematical Physics, 16 (9):1733–1734, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i9/p1733_s1.

MacGillivray:1975:UCP

[285] A. D. MacGillivray. On Ursell’s combinatorial problem. Journal of Mathematical Physics, 16(9):1735–1738, September 1975. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i9/p1735_s1.

Vasholz:1975:DSQ

[286] David P. Vasholz. Discrete space quantum mechanics. Journal of Mathematical Physics, 16(9):1739–1745, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i9/p1739_s1.

Lee:1975:ROA

[287] S. W. Lee and J. Boersma. Ray-optical analysis of fields on shadow boundaries of two parallel plates. Journal of Mathematical Physics, 16 (9):1746–1764, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i9/p1746_s1.

Meetz:1975:GDG

[288] Kurt Meetz. Gauge dependence of Green’s functions in quantum electrodynamics from parallel translation. Journal of Mathematical Physics, 16(9):1765–1771, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i9/p1765_s1.

Hori:1975:STEc

[289] Motoo Hori. Statistical theory of effective electrical, thermal and magnetic properties of random heterogeneous materials. VI. Comment on the notion of a cell material. Journal of Mathematical Physics, 16(9):1772– 1775, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i9/p1772_s1.

Arik:1975:OAD

[290] M. Arik, D. D. Coon, and Yuk ming Lam. Operator algebra of dual resonance models. Journal of Mathematical Physics, 16(9):1776–1779, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i9/p1776_s1.

Peters:1975:NGD

[291] P. C. Peters. Null geodesic deviation. I. Conformally flat space–times. Journal of Mathematical Physics, 16(9):1780–1785, September 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1780_
s1.

Morrison:1975:PHF

[292] J. A. Morrison. Propagation of high frequency surface waves along cylinders of general cross-section. Journal of Mathematical Physics, 16 (9):1786–1794, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i9/p1786_s1.

Wilson:1975:PHF

[293] L. O. Wilson and J. A. Morrison. Propagation of high frequency elastic surface waves along cylinders of general cross section. Journal of Mathematical Physics, 16(9):1795–1805, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i9/p1795_s1.

Winicour:1975:ASA

[294] Jeffrey Winicour. All stationary axisymmetric rotating dust metrics. Journal of Mathematical Physics, 16(9):1806–1808, September 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1806_
s1.

Monroe:1975:CIT

[295] James L. Monroe. Correlation inequalities for two-dimensional vector spin systems. Journal of Mathematical Physics, 16(9):1809–1812, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i9/p1809_s1.

Boya:1975:DPG

[296] Luis J. Boya, Jose F. Cariñena, and Mariano Santander. Dilatations and the Poincaré group. Journal of Mathematical Physics, 16(9):1813–1815, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i9/p1813_s1.

Gruber:1975:PLR

[297] B. Gruber and A. U. Klimyk. Properties of linear representations with a highest weight for the semisimple Lie algebras. Journal of Mathematical Physics, 16(9):1816–1832, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1816_s1.

Huq:1975:PIM

[298] M. Huq. Path integral method for superfields. Journal of Mathematical Physics, 16(9):1833–1837, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1833_s1.

Lerche:1975:RFT

[299] I. Lerche and E. N. Parker. Random function theory revisited: Exact solutions versus the first order smoothing conjecture. Journal of Mathematical Physics, 16(9):1838–1839, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i9/p1838_s1.

Mangiarotti:1975:TDN

[300] Luigi Mangiarotti. A three-dimensional neutron transport problem. Journal of Mathematical Physics, 16(9):1840–1843, September 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1840_
s1.

Kennedy:1975:HFC

[301] Frederick James Kennedy. Hamiltonian formulation of the classical twocharge problem in straight-line approximation. Journal of Mathematical Physics, 16(9):1844–1856, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1844_s1.

Wreszinski:1975:PSG

[302] Walter F. Wreszinski. On properties of the singularity of the ground state in certain classical Heisenberg models. Journal of Mathematical Physics, 16(9):1857–1859, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i9/p1857_s1.

Thorne:1975:TSV

[303] Kip S. Thorne. A toroidal solution of the vacuum Einstein field equations. Journal of Mathematical Physics, 16(9):1860–1865, September 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1860_
s1.

Moshinsky:1975:CTAb

[304] M. Moshinsky and J. Patera. Canonical transformations and accidental degeneracy. IV. Problems with continuous spectra. Journal of Mathematical Physics, 16(9):1866–1875, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i9/p1866_s1.

Chacon:1975:ECW

[305] E. Chacón, D. Levi, and M. Moshinsky. Equivalence of a class of Wigner coefficients of SU(1, 1) with those of SU(2). Journal of Mathematical Physics, 16(9):1876–1879, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1876_s1.

Johnson:1975:MSI

[306] Porter W. Johnson and Robert L. Warnock. Marginally singular integral equation with divided-difference kernel: a problem in N/D theory. Journal of Mathematical Physics, 16(9):1880–1885, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1880_s1.

Frederiksen:1975:RAT

[307] Jørgen S. Frederiksen, Porter W. Johnson, and Robert L. Warnock. Regge amplitudes through solution of S-matrix equations. Journal of Mathematical Physics, 16(9):1886–1900, September 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1886_s1.

Girardeau:1975:SQR

[308] M. D. Girardeau. Second-quantization representation for a nonrelativistic system of composite particles. I. Generalized Tani transformation and its iterative evaluation. Journal of Mathematical Physics, 16 (9):1901–1919, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i9/p1901_s1.

Hillion:1975:BSD

[309] P. Hillion. Borel summability and distribution. Journal of Mathematical Physics, 16(9):1920–1927, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1920_s1.

Aks:1975:SED

[310] S. Ø. Aks, B. B. Varga, and J. J. Sienicki. Small energy denominators in interacting quantum systems: Bound states. Journal of Mathematical Physics, 16(9):1928–1933, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1928_s1.

Shay:1975:TTA

[311] D. Shay. Theory of tachyons for arbitrary spin. Journal of Mathematical Physics, 16(9):1934–1938, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1934_s1.

Pereira:1975:ND

[312] C. Martin Pereira. Nonmetrical dynamics. I. Journal of Mathematical Physics, 16(9):1939–1944, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i9/p1939_s1.

Hioe:1975:QTA

[313] F. T. Hioe and E. W. Montroll. Quantum theory of anharmonic oscillators. I. Energy levels of oscillators with positive quartic anharmonicity. Journal of Mathematical Physics, 16(9):1945–1955, September 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/p1945_
s1.

Newman:1975:MIF

[314] Charles M. Newman. Moment inequalities for ferromagnetic Gibbs distributions. Journal of Mathematical Physics, 16(9):1956–1959, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i9/p1956_s1.

Sullivan:1975:EDC

[315] John J. Sullivan. Erratum: “Double coset analysis for symmetry adapting Nth rank tensors of U(n) to its unitary subgroups”. Journal of Mathematical Physics, 16(9):1959, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i9/p1959_s1. See [133].

Schulten:1975:ERE

[316] Klaus Schulten and Roy G. Gordon. Exact recursive evaluation of 3j- and 6j-coefficients for quantum-mechanical coupling of angular momenta. Journal of Mathematical Physics, 16(10):1961–1970, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p1961_s1.

Schulten:1975:SAC

[317] Klaus Schulten and Roy G. Gordon. Semiclassical approximations to 3j- and 6j-coefficients for quantum-mechanical coupling of angular momenta. Journal of Mathematical Physics, 16(10):1971–1988, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p1971_s1.

Fesciyan:1975:IPS

[318] S. Fesciyan and H. L. Frisch. An inverse problem in statistical mechanics. Journal of Mathematical Physics, 16(10):1989–1991, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/
p1989_s1.

Cornwell:1975:DDI

[319] J. F. Cornwell. Direct determination of the Iwasawa decomposition for noncompact semisimple Lie algebras. Journal of Mathematical Physics, 16(10):1992–1999, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i10/p1992_s1.

Blinder:1975:TPC

[320] S. M. Blinder. Two-point characteristic function for the Kepler–Coulomb problem. Journal of Mathematical Physics, 16(10):2000–2004, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2000_s1.

Hasse:1975:QMT

[321] Rainer W. Hasse. On the quantum mechanical treatment of dissipative systems. Journal of Mathematical Physics, 16(10):2005–2011, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2005_s1.

LoDato:1975:RPO

[322] Vincent A. LoDato. The renormalized projection operator technique for nonlinear stochastic equations. III. Journal of Mathematical Physics, 16(10):2012–2016, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i10/p2012_s1.

Mello:1975:NCT

[323] P. A. Mello and M. Moshinsky. Nonlinear canonical transformations and their representations in quantum mechanics. Journal of Mathematical Physics, 16(10):2017–2028, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i10/p2017_s1.

Cho:1975:HDU

[324] Y. M. Cho. Higher-dimensional unifications of gravitation and gauge theories. Journal of Mathematical Physics, 16(10):2029–2035, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2029_s1.

Kaup:1975:EQN

[325] D. J. Kaup. Exact quantization of the nonlinear Schrödinger equation. Journal of Mathematical Physics, 16(10):2036–2041, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/
p2036_s1.

Henderson:1975:PSO

[326] Douglas Henderson and Masao Chen. Perturbative solution to order βϵ of the Percus–Yevick equation for the square-well potential. Journal of Mathematical Physics, 16(10):2042–2049, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2042_s1.

Sharp:1975:ILO

[327] R. T. Sharp. Internal-labeling operators. Journal of Mathematical Physics, 16(10):2050–2053, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i10/p2050_s1.

Liu:1975:AFT

[328] Kuang Chi Liu. An asymptotic formula for the twisted product of distributions. Journal of Mathematical Physics, 16(10):2054–2056, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2054_s1.

Chen:1975:MEQ

[329] Etang Chen. Master equations in quantum stochastic processes. Journal of Mathematical Physics, 16(10):2057–2061, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2057_s1.

Pais:1975:SGL

[330] A. Pais and V. Rittenberg. Semisimple graded Lie algebras. Journal of Mathematical Physics, 16(10):2062–2073, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2062_s1. See er-

ratum [505].

Papanicolaou:1975:TES

[331] G. C. Papanicolaou and R. Burridge. Transport equations for the Stokes parameters from Maxwell’s equations in a random medium. Journal of Mathematical Physics, 16(10):2074–2085, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2074_s1.

Fannes:1975:GTN

[332] M. Fannes and A. Verbeure. Gauge transformations and normal states of the CCR. Journal of Mathematical Physics, 16(10):2086–2088, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v16/i10/p2086_s1. See erratum [452].

Felber:1975:NCE

[333] Franklin S. Felber and John H. Marburger. New class of exact solutions of the Dirac equation. Journal of Mathematical Physics, 16(10):2089– 2092, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i10/p2089_s1.

Hurley:1975:DRW

[334] William J. Hurley and E. C. G. Sudarshan. On the determination of the relativistic wave equations associated with a given representation of SL(2, C). Journal of Mathematical Physics, 16(10):2093–2098, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2093_s1.

Djukic:1975:SGA

[335] Dj. Djukic and B. Vujanovic. On some geometrical aspects of classical nonconservative mechanics. Journal of Mathematical Physics, 16 (10):2099–2102, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i10/p2099_s1.

Calogero:1975:CPA

[336] F. Calogero and F. Palumbo. Convergence of the perturbative approach to the N-body problem in the N → ∞ limit. Journal of Mathematical Physics, 16(10):2103–2105, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i10/p2103_s1.

Mahanta:1975:PUG

[337] M. N. Mahanta. A possible unification of gravitation and electrodynamics both for spinless and spinning media. Journal of Mathematical Physics, 16(10):2106–2108, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i10/p2106_s1.

Singh:1975:SST

[338] T. Singh. Solutions of a scalar-tensor theory from Einstein’s theory. Journal of Mathematical Physics, 16(10):2109–2111, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2109_s1.

Schroeck:1975:COQ

[339] F. E. Schroeck, Jr. Cores for operators in quantum mechanics. Journal of Mathematical Physics, 16(10):2112–2113, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2112_s1.

Hammel:1975:ESS

[340] William C. Hammel. An exact spherically symmetric local solution in a class of “strong” and “weak” two-tensor gravity theories. Journal of Mathematical Physics, 16(10):2114–2122, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2114_s1.

Ktorides:1975:CAA

[341] C. N. Ktorides. A Clifford algebraic approach to superfields and some consequences. Journal of Mathematical Physics, 16(10):2123–2129, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2123_s1.

Ktorides:1975:NLA

[342] C. N. Ktorides. A non-Lie algebraic framework and its possible merits for symmetry descriptions. Journal of Mathematical Physics, 16(10):2130– 2141, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i10/p2130_s1.

Gross:1975:ESF

[343] Fred Gross, Charles F. Osgood, and Chung-Chun Yang. On the entire solutions of a functional equation in the theory of fluids. Journal of Mathematical Physics, 16(10):2142–2147, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2142_s1.

Huestis:1975:SPS

[344] David L. Huestis. A superposition principle for Siegert resonant states. Journal of Mathematical Physics, 16(10):2148–2152, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i10/p2148_s1.

Budgor:1975:SST

[345] Aaron B. Budgor. Surface scattering on topological networks. Journal of Mathematical Physics, 16(10):2153–2165, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2153_s1.

Pao:1975:PST

[346] C. V. Pao. Positive solution of a time and energy dependent neutron transport problem. Journal of Mathematical Physics, 16(10):2166– 2171, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i10/p2166_s1.

Burdet:1975:HSA

[347] G. Burdet and M. Perrin. The Hamiltonians of the Schrödinger algebra. Journal of Mathematical Physics, 16(10):2172–2176, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/
p2172_s1.

Mirman:1975:DGI

[348] R. Mirman. Differential geometry and internal symmetry. Journal of Mathematical Physics, 16(10):2177–2179, October 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i10/p2177_s1.

Couch:1975:ENS

[349] W. E. Couch and R. J. Torrence. Erratum: “Nonspreading solutions of the inhomogeneous scalar wave equation”. Journal of Mathematical Physics, 16(10):2180, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i10/p2180_s2. See [153].

DeFacio:1975:EHC

[350] B. DeFacio. Erratum: “Heat conduction model with finite signal speed”. Journal of Mathematical Physics, 16(10):2180, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i10/

p2180_s1. See [167].

Park:1975:EQF

[351] Yong Moon Park. The λϕ_{3}^{4} Euclidean quantum field theory in a periodic box. Journal of Mathematical Physics, 16(11):2183–2188, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2183_s1.

Zucker:1975:NJF

[352] I. J. Zucker. New Jacobian ϑ functions and the evaluation of lattice sums. Journal of Mathematical Physics, 16(11):2189–2191, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v16/i11/p2189_s1. See erratum [544] and comment [775].

Narnhofer:1975:SAO

[353] Heide Narnhofer. Self-adjoint operators, derivations and automorphisms on C∗-algebras. Journal of Mathematical Physics, 16(11):2192–2196, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i11/p2192_s1.

Kollmann:1975:CSF

[354] W. Kollmann. Calculation of special functional integrals. Journal of Mathematical Physics, 16(11):2197–2200, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/p2197_s1.

Mizrahi:1975:WCP

[355] Maurice M. Mizrahi. The Weyl correspondence and path integrals. Journal of Mathematical Physics, 16(11):2201–2206, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/
p2201_s1.

Wang:1975:IIF

[356] Alan P. Wang. Invariant imbedding and the Fredholm integral equation. Journal of Mathematical Physics, 16(11):2207–2209, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2207_s1.

Hobson:1975:EPP

[357] Arthur Hobson. Ergodic properties of a particle moving elastically inside a polygon. Journal of Mathematical Physics, 16(11):2210–2214, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2210_s1.

Boyer:1975:AOS

[358] Charles P. Boyer and Kurt Bernardo Wolf. The 2:1 anisotropic oscillator, separation of variables and symmetry group in Bargmann space. Journal of Mathematical Physics, 16(11):2215–2223, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/p2215_s1.

Kelson:1975:SAT

[359] I. Kelson and G. Shadmon. Space averaging techniques of determinantal measures. Journal of Mathematical Physics, 16(11):2224–2232, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2224_s1.

Abbati:1975:HB

[360] Maria C. Abbati, Renzo Cirelli, and Franco Gallone. G-Hilbert bundles. Journal of Mathematical Physics, 16(11):2233–2240, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2233_s1.

Brennan:1975:UMS

[361] B. J. Brennan. Unequal mass spinor-spinor Bethe–Salpeter equation. Journal of Mathematical Physics, 16(11):2241–2249, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/
p2241_s1.

Finkelstein:1975:KE

[362] David Finkelstein and Gin McCollum. Kinks and extensions. Journal of Mathematical Physics, 16(11):2250–2259, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/p2250_s1.

Sahni:1975:NTP

[363] D. C. Sahni. Neutron transport problems in a spherical shell. Journal of Mathematical Physics, 16(11):2260–2270, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/p2260_s1.

Henrich:1975:GSC

[364] Christopher J. Henrich. The Gel’fand states of certain representations of U(n) and the decomposition of products of representations of U(2).

Journal of Mathematical Physics, 16(11):2271–2288, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/
p2271_s1.

Seiler:1975:FMR

[365] Erhard Seiler and Barry Simon. On finite mass renormalizations in the two-dimensional Yukawa model. Journal of Mathematical Physics, 16 (11):2289–2293, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i11/p2289_s1.

Kinoshita:1975:AMZ

[366] Toichiro Kinoshita and Jnanadeva Maharana. Average multiplicity and the zeros of multiparticle generating function. Journal of Mathematical Physics, 16(11):2294–2300, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i11/p2294_s1.

Ablowitz:1975:RCN

[367] Mark J. Ablowitz and Richard Haberman. Resonantly coupled nonlinear evolution equations. Journal of Mathematical Physics, 16(11):2301– 2305, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i11/p2301_s1.

McLenaghan:1975:NSE

[368] Raymond G. McLenaghan and Nessim Tariq. A new solution of the Einstein–Maxwell equations. Journal of Mathematical Physics, 16(11): 2306–2312, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i11/p2306_s1.

Ekstein:1975:MMC

[369] H. Ekstein. Macroscopic and microscopic, coherent and incoherent variables. Journal of Mathematical Physics, 16(11):2313–2317, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2313_s1.

Bessis:1975:MCV

[370] D. Bessis, P. Moussa, and M. Villani. Monotonic converging variational approximations to the functional integrals in quantum statistical mechanics. Journal of Mathematical Physics, 16(11):2318–2325, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i11/p2318_s1.

Hsue:1975:NCQ

[371] C. S. Hsue. Normal coordinates and quantization of nonlinear fields by Feynman’s path integral. Journal of Mathematical Physics, 16(11):2326– 2327, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i11/p2326_s1.

deDormale:1975:SRD

[372] Bernard M. de Dormale and Henri-François Gautrin. Spectral representation and decomposition of self-adjoint operators. Journal of Mathematical Physics, 16(11):2328–2332, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i11/p2328_s1.

Jorna:1975:ESC

[373] S. Jorna. Expansions for simple cubic lattice Green’s functions. Journal of Mathematical Physics, 16(11):2333–2338, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i11/p2333_s1.

Menikoff:1975:RLC

[374] Ralph Menikoff and David H. Sharp. Representations of a local current algebra: Their dynamical determination. Journal of Mathematical Physics, 16(12):2341–2352, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i12/p2341_s1.

Menikoff:1975:ARL

[375] Ralph Menikoff and David H. Sharp. Approximate representations of a local current algebra. Journal of Mathematical Physics, 16(12):2353– 2360, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i12/p2353_s1.

Glass:1975:WTS

[376] E. N. Glass. The Weyl tensor and shear-free perfect fluids. Journal of Mathematical Physics, 16(12):2361–2363, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2361_s1.

Gannon:1975:SNC

[377] Dennis Gannon. Singularities in nonsimply connected space–times. Journal of Mathematical Physics, 16(12):2364–2367, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2364_s1.

Matsuno:1975:EQM

[378] Koichiro Matsuno. Ergodicity of quantum mechanical systems. Journal of Mathematical Physics, 16(12):2368–2371, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2368_s1.

Maekawa:1975:NCF

[379] Takayoshi Maekawa. Note on the computation formula of the boost matrices of SO(n−1, 1) and continuation to the d matrices of SO(n). Journal of Mathematical Physics, 16(12):2372–2374, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2372_s1.

Couch:1975:RMF

[380] W. E. Couch and R. J. Torrence. Retarded multipole fields and the inhomogeneous wave equation. Journal of Mathematical Physics, 16 (12):2375–2377, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i12/p2375_s1.

Sinha:1975:CPS

[381] B. P. Sinha and R. H. MacPhie. On the computation of the prolate spheroidal radial functions of the second kind. Journal of Mathematical Physics, 16(12):2378–2381, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i12/p2378_s1.

Chen:1975:BTT

[382] Hsing-Hen Chen. A Bäcklund transformation in two dimensions. Journal of Mathematical Physics, 16(12):2382–2384, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2382_s1.

Gurses:1975:LCT

[383] Metin Gürses and Feza Gürsey. Lorentz covariant treatment of the Kerr– Schild geometry. Journal of Mathematical Physics, 16(12):2385–2390, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i12/p2385_s1.

Nachamkin:1975:DUY

[384] Jack Nachamkin. Direct use of Young tableau algebra to generate the Clebsch–Gordan coefficients of SU(2). Journal of Mathematical Physics, 16(12):2391–2394, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i12/p2391_s1.

Plebanski:1975:SSC

[385] J. F. Plebañski. Some solutions of complex Einstein equations. Journal of Mathematical Physics, 16(12):2395–2402, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2395_s1.

Plebanski:1975:NGS

[386] J. F. Plebañski and S. Hacyan. Null geodesic surfaces and Goldberg– Sachs theorem in complex Riemannian spaces. Journal of Mathematical Physics, 16(12):2403–2407, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i12/p2403_s1.

Louck:1975:SPC

[387] J. D. Louck, M. A. Lohe, and L. C. Biedenharn. Structural properties of the canonical U(3) Racah functions and the U(3): U(2) projective functions. Journal of Mathematical Physics, 16(12):2408–2426, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i12/p2408_s1.

Quesne:1975:GTP

[388] C. Quesne. Group theory and propagation of operator averages. Journal of Mathematical Physics, 16(12):2427–2431, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2427_s1.

Sorkin:1975:EFS

[389] Rafael Sorkin. The electromagnetic field on a simplicial net. Journal of Mathematical Physics, 16(12):2432–2440, December 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/

p2432_s1. See erratum [1415].

Grmela:1975:DST

[390] Miroslav Grmela. Dynamic stability and thermodynamics in kinetic theory and fluid mechanics. Journal of Mathematical Physics, 16(12):2441– 2449, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i12/p2441_s1.

Grmela:1975:LTB

[391] Miroslav Grmela. Long time behavior of solutions to the linearized two component Enskog–Vlasov kinetic equations. Journal of Mathematical Physics, 16(12):2450–2457, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i12/p2450_s1.

Zimmerman:1975:SGC

[392] John Charles Zimmerman. Solutions for the general cylindrically symmetric stationary dust model. Journal of Mathematical Physics, 16 (12):2458–2460, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i12/p2458_s1.

Kumei:1975:ITI

[393] Sukeyuki Kumei. Invariance transformations, invariance group transformations and invariance groups of the Sine-Gordon equations. Journal of Mathematical Physics, 16(12):2461–2468, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v16/i12/p2461_s1.

Halpern:1975:GTS

[394] Francis R. Halpern. General techniques for single and coupled quantum anharmonic oscillators. Journal of Mathematical Physics, 16(12):2469– 2475, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i12/p2469_s1.

Havas:1975:SVHb

[395] Peter Havas. Separation of variables in the Hamilton–Jacobi, Schrödinger and related equations. II. Partial separation. Journal of Mathematical Physics, 16(12):2476–2489, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i12/p2476_s1.

Liron:1975:SSV

[396] N. Liron and H. E. Wilhelm. Similarity solution for viscous and thermal gas flow in a cone. Journal of Mathematical Physics, 16(12):2490– 2493, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i12/p2490_s1.

Rabl:1975:SCG

[397] Veronika Rabl, George Campbell, and Kameshwar C. Wali. SU(4) Clebsch–Gordan coefficients. Journal of Mathematical Physics, 16(12): 2494–2506, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i12/p2494_s1.

Rabl:1975:CGC

[398] V. Rabl, G. Campbell, Jr., and K. C. Wali. SU(4) Clebsch–Gordan coefficients. Journal of Mathematical Physics, 16(12):2495–2506, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1975:LTS

[399] E. G. Kalnins and Willard Miller, Jr. Lie theory and separation of variables. 8. Semisubgroup coordinates for Ψ_{tt}−Δ_{2}Ψ = 0. Journal of Mathematical Physics, 16(12):2507–2516, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v16/i12/p2507_s1.

Singh:1975:CSS

[400] T. Singh. Cylindrically symmetric solutions of a scalar-tensor theory of gravitation. Journal of Mathematical Physics, 16(12):2517–2519, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i12/p2517_s1.

Meisel:1975:SGI

[401] L. V. Meisel, D. M. Gray, and E. Brown. The structure of groups with index-3 subgroups and Landau’s second theorem. Journal of Mathematical Physics, 16(12):2520–2521, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v16/i12/p2520_s1.

Bracci:1975:LCQ

[402] L. Bracci and F. Strocchi. Local and covariant quantizations of linearized Einstein’s equations. II. Journal of Mathematical Physics, 16(12):2522– 2526, December 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i12/p2522_s1.

Tabensky:1976:UGG

[403] R. Tabensky. Unification of gauge and gravitation theories. Journal of Mathematical Physics, 17(1):1–5, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i1/p1_s1.

Raju:1976:GKP

[404] Cvavb Chandra Raju. Generalized Källén–Pauli equation. Journal of Mathematical Physics, 17(1):6–12, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i1/p6_s1.

Yao:1976:ANP

[405] Te Hai Yao. On analytic nonlocal potentials. IV. A closed contour dispersion relation. Journal of Mathematical Physics, 17(1):13–16, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p13_s1.

Lerche:1976:CQK

[406] I. Lerche. On conserved quantities in kinematic dynamo theory. Journal of Mathematical Physics, 17(1):17–23, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i1/p17_s1.

Post:1976:PMR

[407] G. Post. Properties of massless relativistic fields under the conformal group. Journal of Mathematical Physics, 17(1):24–32, January 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i1/p24_s1.

Maheshwari:1976:GWF

[408] Amar Maheshwari. The generalized Wiener–Feynman path integrals. Journal of Mathematical Physics, 17(1):33–36, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i1/p33_s1.

Lord:1976:TSE

[409] Eric A. Lord. A theorem on stress-energy tensors (in general relativity). Journal of Mathematical Physics, 17(1):37–38, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i1/p37_s1.

Held:1976:KVE

[410] A. Held. Killing vectors in empty space algebraically special metrics. II. Journal of Mathematical Physics, 17(1):39–45, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i1/p39_s1.

Dollard:1976:IKI

[411] John D. Dollard. Interpretation of Kato’s invariance principle in scattering theory. Journal of Mathematical Physics, 17(1):46–51, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p46_s1.

Economou:1976:WCT

[412] John E. Economou and Frederick J. Ernst. Weyl conform tensor of δ = 2 Tomimatsu–Sato spinning mass gravitational field. Journal of Mathematical Physics, 17(1):52–53, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i1/p52_s1.

Ernst:1976:BHM

[413] Frederick J. Ernst. Black holes in a magnetic universe. Journal of Mathematical Physics, 17(1):54–56, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i1/p54_s1.

Spohn:1976:SLN

[414] Herbert Spohn. The spectrum of the Liouville–von Neumann operator. Journal of Mathematical Physics, 17(1):57–60, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i1/p57_s1. See

[776].

Kapoor:1976:SNL

[415] A. K. Kapoor. On superpropagator for nonpolynomial Lagrangians with internal symmetry. Journal of Mathematical Physics, 17(1):61–62, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p61_s1.

Gherman:1976:SES

[416] O. Gherman and L. Méder. Some examples of symmetrical perturbation problems in several complex variables. Journal of Mathematical Physics, 17(1):63–68, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i1/p63_s1.

Yan:1976:DNN

[417] C. C. Yan. Determination of nearest neighbor degeneracy on a onedimensional lattice. Journal of Mathematical Physics, 17(1):69–72, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p69_s1.

Moses:1976:SKV

[418] H. E. Moses. A solution of the Korteweg–de Vries equation in a halfspace bounded by a wall. Journal of Mathematical Physics, 17(1):73– 75, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i1/p73_s1.

Bowden:1976:MNTa

[419] R. L. Bowden, S. Sancaktar, and P. F. Zweifel. Multigroup neutron transport. I. Full range. Journal of Mathematical Physics, 17(1):76– 81, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i1/p76_s1.

Bowden:1976:MNTb

[420] R. L. Bowden, S. Sancaktar, and P. F. Zweifel. Multigroup neutron transport. II. Half-range. Journal of Mathematical Physics, 17(1):82– 86, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i1/p82_s1.

Klein:1976:UBE

[421] Abel Klein and Lawrence J. Landau. An upper bound on the energy gap in the (λϕ^{4} + σϕ^{2})_{2} model. Journal of Mathematical Physics, 17(1): 87–93, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i1/p87_s1.

Bohm:1976:SGG

[422] Arno Bohm and R. B. Teese. Spectrum generating group of the symmetric top molecule. Journal of Mathematical Physics, 17(1):94–108, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p94_s1.

Wreszinski:1976:GTQ

[423] Walter F. Wreszinski. Goldstone’s theorem for quantum spin systems of finite range. Journal of Mathematical Physics, 17(1):109–111, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p109_s1.

Kriegsmann:1976:AMG

[424] Gregory A. Kriegsmann. An application of the method of geometrical optics to the scattering of plane electromagnetic waves off cylindrically confined cold plasmas. Journal of Mathematical Physics, 17(1):112– 120, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i1/p112_s1.

Eminhizer:1976:CNP

[425] Charles R. Eminhizer, Robert H. G. Helleman, and Elliott W. Montroll. On a convergent nonlinear perturbation theory without small denominators or secular terms. Journal of Mathematical Physics, 17(1):121– 140, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i1/p121_s1.

Jang:1976:PEC

[426] Pong Soo Jang. On the positive energy conjecture, (in general relativity). Journal of Mathematical Physics, 17(1):141–145, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i1/p141_s1.

Christian:1976:SIV

[427] W. G. Christian, A. G. Law, W. F. Martens, A. L. Mullikin, and M. B. Sledd. Solution of initial-value problems for some infinite chains of harmonic oscillators. Journal of Mathematical Physics, 17(1):146–158, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i1/p146_s1.

Adler:1976:EFS

[428] Ronald J. Adler. Erratum: “A fluid sphere in general relativity”. Journal of Mathematical Physics, 17(1):158, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i1/p158_s1. See [4].

Greenberg:1976:FCS

[429] William Greenberg. Functional calculus for the symmetric multigroup transport operator. Journal of Mathematical Physics, 17(2):159–162, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i2/p159_s1.

Greenberg:1976:CEE

[430] William Greenberg and Paul F. Zweifel. The Case eigenfunction expansion for a conservative medium. Journal of Mathematical Physics, 17(2):163–167, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i2/p163_s1.

Miodek:1976:IPC

[431] Ilia Miodek. Inverse problem for complex “r^{a}-analytic” potentials of finite range. Journal of Mathematical Physics, 17(2):168–173, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p168_s1.

Hawking:1976:NTC

[432] S. W. Hawking, A. R. King, and P. J. McCarthy. A new topology for curved space–time which incorporates the causal, differential, and conformal structures. Journal of Mathematical Physics, 17(2):174–181, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p174_s1.

Ernst:1976:KBH

[433] Frederick J. Ernst and Walter J. Wild. Kerr black holes in a magnetic universe. Journal of Mathematical Physics, 17(2):182–184, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p182_s1.

Basu:1976:MAF

[434] Debabrata Basu and S. Datta Majumdar. The master analytic function and the Lorentz group. I. Reduction of the representations of O(3, 1) in O(2, 1) basis. Journal of Mathematical Physics, 17(2):185–192, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p185_s1.

Majumdar:1976:MAF

[435] S. Datta Majumdar. The master analytic function and the Lorentz group. II. The Clebsch–Gordan problem for O(2, 1). Journal of Mathematical Physics, 17(2):193–197, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i2/p193_s1.

Cheng:1976:NIR

[436] Kuo shung Cheng. New identities on the Riemann tensor. Journal of Mathematical Physics, 17(2):198–199, February 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i2/p198_s1.

Federbush:1976:SEAa

[437] Paul Federbush. The semi-Euclidean approach in statistical mechanics. I. Basic expansion steps and estimates. Journal of Mathematical Physics, 17(2):200–203, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i2/p200_s1.

Federbush:1976:SEAb

[438] Paul Federbush. The semi-Euclidean approach in statistical mechanics. II. The cluster expansion, a special example. Journal of Mathematical Physics, 17(2):204–207, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i2/p204_s1.

Fuchs:1976:SNP

[439] V. Fuchs and G. Beaudry. Stability of nonlinear parametric-decay interactions in finite homogeneous plasma. Journal of Mathematical Physics, 17(2):208–213, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i2/p208_s1.

Murray:1976:GTP

[440] J. C. Murray. Gurtin-type properties associated with wave propagation in a superfluid. Journal of Mathematical Physics, 17(2):214–216, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p214_s1.

Bazanski:1976:GFT

[441] Stanislaw L. Bazański. A geometric formulation of the Taylor theorem for curves in affine manifolds. Journal of Mathematical Physics, 17(2):217– 221, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i2/p217_s1.

Khoo:1976:MTS

[442] I. C. Khoo and Y. K. Wang. Multiple time scale analysis of an anharmonic crystal. Journal of Mathematical Physics, 17(2):222–227, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p222_s1.

Freund:1976:SS

[443] Peter G. O. Freund and Irving Kaplansky. Simple supersymmetries. Journal of Mathematical Physics, 17(2):228–231, February 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i2/p228_
s1.

Varatharajulu:1976:AST

[444] V. Varatharajulu and D. J. Vezzetti. Approach of the statistical theory of light scattering to the phenomenological theory. Journal of Mathematical Physics, 17(2):232–240, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i2/p232_s1.

Yao:1976:CQF

[445] Te Hai Yao. Construction of quantum fields from Euclidean tensor fields. Journal of Mathematical Physics, 17(2):241–247, February 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i2/p241_
s1.

Deliyannis:1976:SSS

[446] Platon C. Deliyannis. Superposition of states and the structure of quantum logics. Journal of Mathematical Physics, 17(2):248–254, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p248_s1.

Subbarao:1976:RGS

[447] K. Subbarao. Renormalization group for a system of continuous spins on a lattice. Journal of Mathematical Physics, 17(2):255–258, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p255_s1.

Hall:1976:APG

[448] George L. Hall. Asymptotic properties of generalized Chaba and Pathria lattice sums. Journal of Mathematical Physics, 17(2):259–260, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p259_s1.

Critchfield:1976:SPD

[449] Charles L. Critchfield. Scalar potentials in the Dirac equation. Journal of Mathematical Physics, 17(2):261–266, February 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i2/p261_s1.

DeFacio:1976:BQF

[450] B. DeFacio and C. L. Hammer. Bilinear quantum field theories and their coherent states. Journal of Mathematical Physics, 17(2):267–281, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i2/p267_s1.

Esposito:1976:WTS

[451] F. Paul Esposito and E. N. Glass. The Weyl tensor and stationary electrovac space–times. Journal of Mathematical Physics, 17(2):282–283, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i2/p282_s1.

Fannes:1976:EGT

[452] M. Fannes and A. Verbeure. Erratum: “Gauge transformations and normal states of the CCR”. Journal of Mathematical Physics, 17(2): 284, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i2/p284_s1. See [332].

Belinfante:1976:TPS

[453] Johan G. F. Belinfante. Transition probability spaces (in axiomatic quantum theory). Journal of Mathematical Physics, 17(3):285–291, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p285_s1.

Tariq:1976:EMM

[454] N. Tariq and B. O. J. Tupper. Einstein–Maxwell metrics admitting a dual interpretation. Journal of Mathematical Physics, 17(3):292–296, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p292_s1.

Masson:1976:EPA

[455] D. Masson and E. Prugovečki. An effective-potential approach to stationary scattering theory for long-range potentials. Journal of Mathematical Physics, 17(3):297–302, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i3/p297_s1.

Bisognano:1976:DCQ

[456] Joseph J. Bisognano and Eyvind H. Wichmann. On the duality condition for quantum fields. Journal of Mathematical Physics, 17(3):303– 321, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i3/p303_s1.

Frey:1976:GCD

[457] Heinz Frey. Gauge-covariant differentiation and Green’s functions for the Yang–Mills field. Journal of Mathematical Physics, 17(3):322–330, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p322_s1.

Kalnins:1976:LTSa

[458] E. G. Kalnins and W. Miller, Jr. Lie theory and separation of variables. 9. orthogonal R-separable coordinate systems for the wave equation ψ_{tt} − Δ_{2}ψ = 0. Journal of Mathematical Physics, 17(3):331– 355, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i3/p331_s1.

Kalnins:1976:LTSb

[459] E. G. Kalnins and W. Miller, Jr. Lie theory and separation of variables. 10. Nonorthogonal R-separable solutions of the wave equation ∂_{tt}ψ = Δ_{2}ψ. Journal of Mathematical Physics, 17(3):356–368, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p356_s1.

Kalnins:1976:LTSc

[460] E. G. Kalnins and Willard Miller, Jr. Lie theory and separation of variables. 11. The EPD equation. Journal of Mathematical Physics, 17(3): 369–377, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i3/p369_s1.

Schulgasser:1976:PIR

[461] Kalman Schulgasser. On a phase interchange relationship for composite materials. Journal of Mathematical Physics, 17(3):378–381, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p378_s1.

Schulgasser:1976:CFR

[462] Kalman Schulgasser. On the conductivity of fiber reinforced materials. Journal of Mathematical Physics, 17(3):382–387, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i3/p382_s1.

Davies:1976:LWN

[463] John A. Davies. Long-wavelength normal mode vibrations of infinite, ionic crystal lattices. II. Journal of Mathematical Physics, 17(3):388– 393, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i3/p388_s1.

Andrie:1976:SGA

[464] M. Andrié. Spectrum generating algebras and Lie groups in classical mechanics. Journal of Mathematical Physics, 17(3):394–400, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i3/
p394_s1.

Onofri:1976:DQK

[465] Enrico Onofri. Dynamical quantization of the Kepler manifold. Journal of Mathematical Physics, 17(3):401–408, March 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i3/p401_s1.

Saenz:1976:EOC

[466] A. W. Sáenz and W. W. Zachary. On the equality of S operators corresponding to unitarity equivalent Hamiltonians in single channels scattering. Journal of Mathematical Physics, 17(3):409–418, March 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i3/p409_

s1. See [652].

Naudts:1976:BAK

[467] J. Naudts and A. Verbeure. Bounds on the admittance for KMS states. Journal of Mathematical Physics, 17(3):419–423, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i3/p419_s1.

Freund:1976:CAS

[468] Peter G. O. Freund. Conformal algebra in superspace and supergauge theory. Journal of Mathematical Physics, 17(3):424–426, March 1976.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i3/
p424_s1.

Lesche:1976:TRP

[469] Bernhard Lesche. A theorem on Ruch’s principle of nonequilibrium statistics. Journal of Mathematical Physics, 17(3):427–430, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i3/
p427_s1.

Girardeau:1976:MEM

[470] M. D. Girardeau. Matrix elements of multiple electron exchange between spin functions of maximum multiplicity. Journal of Mathematical Physics, 17(3):431–433, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i3/p431_s1.

Strinati:1976:MED

[471] G. Strinati and U. Fano. Multipole expansion of the density of states about a crystal cell. Journal of Mathematical Physics, 17(3):434– 438, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i3/p434_s1.

Wu:1976:DTM

[472] F. Y. Wu and Y. K. Wang. Duality transformation in a many-component spin model. Journal of Mathematical Physics, 17(3):439–440, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p439_s1.

Kloster:1976:ESG

[473] S. Kloster, M. M. Som, and A. Das. Erratum: “On the stationary gravitational fields”. Journal of Mathematical Physics, 17(3):441, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p441_s2. See [7].

Maekawa:1976:EFC

[474] Takayoshi Maekawa. Erratum: “Formula for the computation of the representation matrix elements of the group SO(n)”. Journal of Mathematical Physics, 17(3):441, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i3/p441_s1. See [70].

Hamilton:1976:IRF

[475] Eugene Hamilton. The infinitely renormalized field in the scalar field model. Journal of Mathematical Physics, 17(4):443–457, April 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p443_
s1.

Machacek:1976:IFP

[476] M. Machacek and Y. Tomozawa. SU(6) isoscalar factors for the product 405 × 56 → 56, 70. Journal of Mathematical Physics, 17(4):458– 462, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p458_s1.

Rudra:1976:CGC

[477] P. Rudra and M. K. Sikdar. Clebsch–Gordan coefficients of magnetic space groups. Journal of Mathematical Physics, 17(4):463–473, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i4/p463_s1.

Majumdar:1976:STM

[478] Chanchal K. Majumdar. Solutions of the three magnon bound state equation. I. Journal of Mathematical Physics, 17(4):474–477, April 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p474_
s1.

Mukhopadhyay:1976:STM

[479] S. K. Mukhopadhyay and C. K. Majumdar. Solutions of the three magnon bound state equation. II. Journal of Mathematical Physics, 17(4):478–479, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i4/p478_s1.

Stassis:1976:CRI

[480] C. Stassis and S. A. Williams. Comment on the reduction of an important 9-j symbol. Journal of Mathematical Physics, 17(4):480–481, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i4/p480_s1.

Dutt:1976:ACS

[481] R. Dutt and M. Lakshmanan. Application of coherent state representation to classical x^{6} and coupled anharmonic oscillators. Journal of Mathematical Physics, 17(4):482–484, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i4/p482_s1.

Weaver:1976:MT

[482] D. L. Weaver. On the Majorana transformation. Journal of Mathematical Physics, 17(4):485–489, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i4/p485_s1.

Fanelli:1976:CTP

[483] Robert Fanelli. Canonical transformations and phase space path integrals. Journal of Mathematical Physics, 17(4):490–493, April 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p490_
s1.

Casati:1976:NEC

[484] Giulio Casati and Joseph Ford. Numerical experiments on the Calogero lattice. Journal of Mathematical Physics, 17(4):494–497, April 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p494_
s1.

Zorbas:1976:SSB

[485] J. Zorbas. Stationary scattering for N-body systems involving Coulomb potentials. Journal of Mathematical Physics, 17(4):498–502, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/
p498_s1.

Ringwood:1976:CFT

[486] G. A. Ringwood. A class of field theories with unique well-defined functional integration. Journal of Mathematical Physics, 17(4):503–505, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i4/p503_s1.

Barut:1976:SDE

[487] A. O. Barut and J. Kraus. Solution of the Dirac equation with Coulomb and magnetic moment interactions. Journal of Mathematical Physics, 17(4):506–508, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i4/p506_s1.

Rudra:1976:CFM

[488] P. Rudra. Comultiplicator of finite magnetic groups. Journal of Mathematical Physics, 17(4):509–511, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i4/p509_s1.

Rudra:1976:SRP

[489] P. Rudra. Spinor representations and projective factor systems of crystallographic point groups. Journal of Mathematical Physics, 17(4):512– 514, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p512_s1.

Ernst:1976:RNS

[490] Frederick J. Ernst. Removal of the nodal singularity of the C-metric. Journal of Mathematical Physics, 17(4):515–516, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p515_s1.

Prugovecki:1976:PMF

[491] E. Prugovečki. Probability measures on fuzzy events in phase space. Journal of Mathematical Physics, 17(4):517–523, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p517_s1.

Arik:1976:HSA

[492] M. Arik and D. D. Coon. Hilbert spaces of analytic functions and generalized coherent states. Journal of Mathematical Physics, 17(4):524– 527, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p524_s1.

Alexanian:1976:CEI

[493] Moorad Alexanian. Critical-exponent inequalities in the Ornstein– Zernike theory. Journal of Mathematical Physics, 17(4):528–529, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i4/p528_s1.

Blaauw:1976:STP

[494] H. J. Blaauw and A. Tip. L_{p}-space techniques in potential scattering. Journal of Mathematical Physics, 17(4):530–534, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p530_s1.

Hongoh:1976:WCT

[495] M. Hongoh. Wigner coefficients for the theory of five-dimensional quasispin. Journal of Mathematical Physics, 17(4):535–541, April 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p535_
s1.

Busoni:1976:NTO

[496] G. Busoni, L. Mangiarotti, and G. Frosali. Neutron transport operators in C and L_{2}-spaces. Journal of Mathematical Physics, 17(4):542– 545, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p542_s1.

Peters:1976:NGD

[497] P. C. Peters. Null geodesic deviation. II. The Schwartzschild metric. Journal of Mathematical Physics, 17(4):546–553, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p546_s1.

Guler:1976:DUA

[498] Y. Güler. Determination of unitary, analytic representations of the conformal group in 2 + 1 dimensions using the operator formalism of the Gel’fand–Naimark Z basis. Journal of Mathematical Physics, 17(4):554– 558, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p554_s1.

Barnsley:1976:PAB

[499] M. F. Barnsley. Padé approximant bounds for the difference of two series of Stieltjes. Journal of Mathematical Physics, 17(4):559–565, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i4/p559_s1.

Mizrahi:1976:PIS

[500] Maurice M. Mizrahi. On path integral solutions of the Schrödinger equation, without limiting procedure. Journal of Mathematical Physics, 17(4): 566–575, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p566_s1.

Anile:1976:GOG

[501] A. M. Anile. Geometrical optics in general relativity: a study of the higher order corrections. Journal of Mathematical Physics, 17(4):576– 584, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i4/p576_s1.

Finley:1976:FHM

[502] J. D. Finley III and J. F. Plebański. Further heavenly metrics and their symmetries. Journal of Mathematical Physics, 17(4):585–596, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i4/p585_s1.

Cohen:1976:CRP

[503] Leon Cohen. Correspondence rules and path integrals. Journal of Mathematical Physics, 17(4):597–598, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i4/p597_s1.

Hori:1976:EST

[504] Motoo Hori. Erratum: “Statistical theory of effective electrical, thermal, and magnetic properties of random heterogeneous materials. VI. Comment on the notion of a cell material”. Journal of Mathematical Physics, 17(4):598, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i4/p598_s1. See [74].

Pais:1976:ESG

[505] A. Pais and V. Rittenberg. Erratum: “Semisimple [graded] Lie algebras”. Journal of Mathematical Physics, 17(4):598, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i4/p598_s2. See

[330].

Wolf:1976:CTS

[506] Kurt Bernardo Wolf. Canonical transforms, separation of variables, and similarity solutions for a class of parabolic differential equations. Journal of Mathematical Physics, 17(5):601–613, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i5/p601_s1.

Feigenbaum:1976:RBN

[507] Mitchell J. Feigenbaum. The relationship between the normalization coefficient and dispersion function for the multigroup transport equation. Journal of Mathematical Physics, 17(5):614–617, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p614_s1.

Benioff:1976:MZFa

[508] Paul A. Benioff. Models of Zermelo Frankel set theory as carriers for the mathematics of physics. I. Journal of Mathematical Physics, 17(5): 618–628, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p618_s1.

Benioff:1976:MZFb

[509] Paul A. Benioff. Models of Zermelo Frankel set theory as carriers for the mathematics of physics. II. Journal of Mathematical Physics, 17(5): 629–640, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p629_s1.

Galles:1976:CTC

[510] Carlos D. Galles and Oscar S. Zandron. Conformal transformation for the Coulter–Weinberg form of the equations for mass zero spin-2 field. Journal of Mathematical Physics, 17(5):641–642, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p641_s1.

Miller:1976:CQC

[511] J. G. Miller. Charge quantization and canonical quantization. Journal of Mathematical Physics, 17(5):643–646, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i5/p643_s1.

Marchioro:1976:DPB

[512] C. Marchioro, A. Pellegrinotti, E. Presutti, and M. Pulvirenti. On the dynamics of particles in a bounded region: A measure theoretical approach. Journal of Mathematical Physics, 17(5):647–652, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p647_s1.

Deliyannis:1976:CS

[513] Platon C. Deliyannis. Conditioning of states. Journal of Mathematical Physics, 17(5):653–659, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i5/p653_s1.

Fette:1976:ASS

[514] C. W. Fette, Allen I. Janis, and Ezra T. Newman. Algebraically special H-spaces. Journal of Mathematical Physics, 17(5):660–664, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i5/p660_s1.

Rocca:1976:CSB

[515] F. Rocca, P. N. M. Sisson, and A. Verbeure. A class of states on the boson-fermion algebra. Journal of Mathematical Physics, 17(5):665–667, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i5/p665_s1.

Chacon:1976:ESP

[516] E. Chacón, M. Moshinsky, and R. T. Sharp. U(5) ⊆ O(5) ⊆ O(3) and the exact solution for the problem of quadrupole vibrations of the nucleus. Journal of Mathematical Physics, 17(5):668–676, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p668_s1.

Brunet:1976:ADR

[517] Robert C. Brunet. An approximant for democratic representation of all Born terms. Journal of Mathematical Physics, 17(5):677–682, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i5/p677_s1.

Cheng:1976:MIM

[518] Edward T. Cheng and Robert W. Conn. A multipoint interpolation method based on variational principles for functionals of the solution to linear equations. Journal of Mathematical Physics, 17(5):683–687, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i5/p683_s1.

Sanchez:1976:SSW

[519] Norma G. Sanchez. Scattering of scalar waves from a Schwarzschild black hole. Journal of Mathematical Physics, 17(5):688–692, May 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p688_
s1.

Kracklauer:1976:GPN

[520] Aloysius F. Kracklauer. A geometric proof of no-interaction theorems. Journal of Mathematical Physics, 17(5):693–694, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p693_s1.

Klein:1976:KSE

[521] W. Klein. The Kirkwood–Salsburg equation and phase transitions. Journal of Mathematical Physics, 17(5):695–698, May 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i5/p695_s1.

Grewe:1976:RDK

[522] N. Grewe and W. Klein. Rigorous derivation of the Kirkwood–Monroe equation for small activity. Journal of Mathematical Physics, 17(5): 699–703, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p699_s1.

Wang:1976:SBS

[523] M. Y. Wang. Spontaneously broken symmetry and cosmological constant. Journal of Mathematical Physics, 17(5):704–705, May 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p704_
s1.

Floratos:1976:BSS

[524] E. G. Floratos. Bethe–Salpeter spinor equation at P_{μ} = 0 and SO(5) spinor spherical harmonics. Journal of Mathematical Physics, 17(5): 706–709, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p706_s1.

Segur:1976:ASCa

[525] Harvey Segur and Mark J. Ablowitz. Asymptotic solutions and conservation laws for the nonlinear Schrödinger equation. I. Journal of Mathematical Physics, 17(5):710–713, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i5/p710_s1.

Segur:1976:ASCb

[526] Harvey Segur. Asymptotic solutions and conservation laws for the nonlinear Schrödinger equation. II. Journal of Mathematical Physics, 17(5): 714–716, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p714_s1.

Patera:1976:QNP

[527] J. Patera, P. Winternitz, and H. Zassenhaus. Quantum numbers for particles in de Sitter space. Journal of Mathematical Physics, 17(5): 717–728, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p717_s1.

Hendrickson:1976:BSC

[528] Joseph H. Hendrickson. n-body scattering into cones with long-range time-dependent potentials. Journal of Mathematical Physics, 17(5):729– 733, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p729_s1.

Besieris:1976:SWK

[529] Ioannis M. Besieris and Frederick D. Tappert. Stochastic wavekinetic theory in the Liouville approximation. Journal of Mathematical Physics, 17(5):734–743, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i5/p734_s1.

Pearce:1976:MCF

[530] Paul A. Pearce. Monotonicity of correlation functions. Journal of Mathematical Physics, 17(5):744–746, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i5/p744_s1.

Scanio:1976:RBI

[531] J. J. G. Scanio and P. Suranyi. Rigorous bounds on inclusive and exclusive cross sections I. Lower bound on the integral of an analytic function. Journal of Mathematical Physics, 17(5):747–752, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p747_s1.

Lubkin:1976:ESM

[532] Elihu Lubkin. Extension of the statistical mechanics of equilibrium to noncommutative constraints. Journal of Mathematical Physics, 17(5): 753–755, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i5/p753_s1. See correction [806].

Corones:1976:SSP

[533] James Corones. Solitons and simple pseudopotentials. Journal of Mathematical Physics, 17(5):756–759, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i5/p756_s1.

Sherman:1976:AAA

[534] George C. Sherman, Jakob J. Stamnes, and Éamon Lalor. Asymptotic approximations to angular-spectrum representations. Journal of Mathematical Physics, 17(5):760–776, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i5/p760_s1.

Kuchar:1976:GH

[535] Karel Kuchař. Geometry of hyperspace. I. Journal of Mathematical Physics, 17(5):777–791, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i5/p777_s1.

Kuchar:1976:KTF

[536] Karel Kuchař. Kinematics of tensor fields in hyperspace. II. Journal of Mathematical Physics, 17(5):792–800, May 1976. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i5/p792_s1.

Kuchar:1976:DTF

[537] Karel Kuchař. Dynamics of tensor fields in hyperspace. III. Journal of Mathematical Physics, 17(5):801–820, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i5/p801_s1.

Gorini:1976:CPD

[538] Vittorio Gorini, Andrzej Kossakowski, and E. C. G. Sudarshan. Completely positive dynamical semigroups of N-level systems. Journal of Mathematical Physics, 17(5):821–825, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i5/p821_s1.

Klauder:1976:CGR

[539] John R. Klauder. A characteristic glimpse of the renormalization group. Journal of Mathematical Physics, 17(5):826–830, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p826_s1.

Browne:1976:NL

[540] Seán Browne. On the nonuniqueness of the Lagrangian. Journal of Mathematical Physics, 17(5):831–835, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i5/p831_s1.

Byrnes:1976:BTE

[541] S. G. Byrnes. Bäcklund transformations and the equation z_{xy} = F (x, y, z). Journal of Mathematical Physics, 17(5):836–842, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p836_s1.

Field:1976:SSP

[542] David A. Field. Series of Stieltjes, Padé approximants and continued fractions. Journal of Mathematical Physics, 17(5):843–844, May 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/p843_
s1.

Gobel:1976:SPT

[543] Rüdiger Göbel. The smooth-path topology for curved space–time which incorporates the conformal structure and analytic Feynman tracks. Journal of Mathematical Physics, 17(5):845–853, May 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i5/p845_s1.

Zucker:1976:ENJ

[544] I. J. Zucker. Erratum: “New Jacobian ϑ functions and the evaluation of lattice sums”. Journal of Mathematical Physics, 17(5):853, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i5/p853_s1. See [352, 775].

Kibler:1976:ITM

[545] Maurice R. Kibler. Irreducible tensor method for a chain SU_{2} ⊃ ⋯ ⊃ G′′ ⊃ G′ ⊃ G and molecular physics. Journal of Mathematical Physics, 17(6):855–858, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p855_s1.

Liu:1976:WTD

[546] Kuang Chi Liu. The Weyl transform of distributions. Journal of Mathematical Physics, 17(6):859–862, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i6/p859_s1.

Schwartz:1976:ASP

[547] Charles Schwartz. Almost singular potentials. Journal of Mathematical Physics, 17(6):863–867, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p863_s1.

Nahm:1976:SSP

[548] W. Nahm and M. Scheunert. On the structure of simple pseudo Lie algebras and their invariant bilinear forms. Journal of Mathematical Physics, 17(6):868–879, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p868_s1.

Floyd:1976:PIA

[549] Edward R. Floyd. Phase integral approximations for calculating energy bands. Journal of Mathematical Physics, 17(6):880–884, June 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i6/p880_
s1.

Vaidya:1976:LBT

[550] Arvind Narayan Vaidya. Linear boson transformation coefficients. Journal of Mathematical Physics, 17(6):885–887, June 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i6/p885_s1.

Coz:1976:RSI

[551] Marcel Coz and Christiane Coudray. The Riemann solution and the inverse quantum mechanical problem. Journal of Mathematical Physics, 17(6):888–893, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p888_s1.

Coz:1976:PDM

[552] M. Coz and P. Rochus. Partial differential matrix equations for the inverse problem of scattering theory. Journal of Mathematical Physics, 17(6):894–899, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p894_s1.

Barut:1976:SNI

[553] A. O. Barut and Raj Wilson. Some new identities of Clebsch–Gordan coefficients and representation functions of SO(2, 1) and SO(4). Journal of Mathematical Physics, 17(6):900–915, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i6/p900_s1.

Jarvis:1976:MMS

[554] P. D. Jarvis. Massive and massless supersymmetry: Multiplet structure and unitary irreducible representations. Journal of Mathematical Physics, 17(6):916–922, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p916_s1.

Letelier:1976:CTD

[555] Patricio S. Letelier. Classic theory of direct intermembrane interaction. Journal of Mathematical Physics, 17(6):923–928, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i6/p923_s1.

Bahar:1976:EHP

[556] E. Bahar. Excitation of horizontally polarized waves in critical-coupling regions where the permittivity gradient approaches zero — full wave solutions. Journal of Mathematical Physics, 17(6):929–936, June 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i6/p929_
s1.

Allnatt:1976:GMD

[557] A. R. Allnatt. Generalized molecular distribution functions for fluids and occupation probabilities for lattice gases. Journal of Mathematical Physics, 17(6):937–940, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p937_s1.

Sorba:1976:GGC

[558] Paul Sorba. The Galilei group and its connected subgroups. Journal of Mathematical Physics, 17(6):941–953, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i6/p941_s1.

Saenz:1976:WOL

[559] A. W. Sáenz and W. W. Zachary. Wave operators for long-range hardcore potentials. Journal of Mathematical Physics, 17(6):954–957, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i6/p954_s1.

Morrison:1976:HFA

[560] J. A. Morrison. High frequency approximations for elastic surface waves propagating along cylinders of general cross section. Journal of Mathematical Physics, 17(6):958–963, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i6/p958_s1.

Johnson:1976:DEQ

[561] R. E. Johnson and N. K. Pope. Derivation of an exact quantummechanical formula for the second term in the activity expansion of the correlation function S(Q, ω). Journal of Mathematical Physics, 17(6): 964–972, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i6/p964_s1.

Lehnigk:1976:CCD

[562] Siegfried H. Lehnigk. A class of conservative diffusion processes with delta function initial conditions. Journal of Mathematical Physics, 17(6): 973–976, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i6/p973_s1.

Patera:1976:SPG

[563] J. Patera, R. T. Sharp, P. Winternitz, and H. Zassenhaus. Subgroups of the Poincaré group and their invariants. Journal of Mathematical Physics, 17(6):977–985, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i6/p977_s1.

Patera:1976:IRL

[564] J. Patera, R. T. Sharp, P. Winternitz, and H. Zassenhaus. Invariants of real low dimension Lie algebras. Journal of Mathematical Physics, 17(6): 986–994, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i6/p986_s1.

vanHaeringen:1976:CAS

[565] H. van Haeringen. Coulombian asymptotic states. Journal of Mathematical Physics, 17(6):995–1000, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i6/p995_s1.

Anderson:1976:CET

[566] Ian M. Anderson and David Lovelock. A characterization of the Einstein tensor in terms of spinors. Journal of Mathematical Physics, 17(6):1001– 1003, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i6/p1001_s1.

Pereira:1976:CSN

[567] Nino R. Pereira. Consequences of scaling in nonlinear partial differential equations. Journal of Mathematical Physics, 17(6):1004–1007, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i6/p1004_s1.

Wong:1976:SEF

[568] C. Y. Wong. On the Schrödinger equation in fluid-dynamical form. Journal of Mathematical Physics, 17(6):1008–1010, June 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i6/p1008_s1.

Ablowitz:1976:NDD

[569] M. J. Ablowitz and J. F. Ladik. Nonlinear differential-difference equations and Fourier analysis. Journal of Mathematical Physics, 17(6):1011– 1018, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i6/p1011_s1.

Barnsley:1976:BBC

[570] M. F. Barnsley and George A. Baker, Jr. Bivariational bounds in a complex Hilbert space, and correction terms for padé approximants. Journal of Mathematical Physics, 17(6):1019–1027, June 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i6/p1019_s1.

Perroud:1976:MSS

[571] Marcel Perroud. The maximal solvable subalgebras of the real classical Lie algebras. Journal of Mathematical Physics, 17(6):1028–1033, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i6/p1028_s1.

Bloore:1976:QDT

[572] F. J. Bloore, M. Assimakopoulos, and I. R. Ghobrial. Quantization as deformation theory. Journal of Mathematical Physics, 17(6):1034–1038, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i6/p1034_s1.

Meyer:1976:QSA

[573] R. E. Meyer. Quasiclassical scattering above barriers in one dimension. Journal of Mathematical Physics, 17(6):1039–1041, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i6/p1039_s1.

McDonald:1976:SNE

[574] J. F. McDonald. Some new ensembles of random matrices. Journal of Mathematical Physics, 17(6):1042–1048, June 1976. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i6/p1042_s1.

Davis:1976:NBU

[575] T. M. Davis and J. R. Ray. Neutrinos and Bianchi I universes. Journal of Mathematical Physics, 17(6):1049–1050, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i6/p1049_s1.

Car:1976:OFG

[576] R. Car, G. Ciucci, and L. Quartapelle. Optimal factor group for nonsymmorphic space groups. Journal of Mathematical Physics, 17(6):1051– 1055, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i6/p1051_s1. See [726].

Zachary:1976:WOM

[577] W. W. Zachary. Wave operators for multichannel scattering by longrange potentials. Journal of Mathematical Physics, 17(6):1056–1063, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i6/p1056_s1. See [832].

Fukushima:1976:SRO

[578] Masahisa Fukushima. A spectral representation on ordinary linear difference equation with operator-valued coefficients of the second order. Journal of Mathematical Physics, 17(6):1064–1072, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i6/p1064_s1.

Park:1976:UBP

[579] Yong Moon Park. Uniform bounds of the pressures of the λϕ_{3}^{4} field model. Journal of Mathematical Physics, 17(6):1073–1075, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i6/p1073_s1.

Deutsch:1976:DCE

[580] C. Deutsch and M. M. Gombert. Diffraction corrections to the equilibrium properties of the classical electron gas. Pair correlation function. Journal of Mathematical Physics, 17(7):1077–1090, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1077_s1.

Ernst:1976:NRT

[581] Frederick J. Ernst. New representation of the Tomimatsu–Sato solution. Journal of Mathematical Physics, 17(7):1091–1094, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1091_s1.

Economou:1976:AFT

[582] John E. Economou. Approximate form of the Tomimatsu–Sato δ = 2 solution near the poles x = 1, y = ±1. Journal of Mathematical Physics, 17(7):1095–1098, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i7/p1095_s1.

Gopinath:1976:SIP

[583] B. Gopinath. The solution to an inverse problem in stratified dielectric media. Journal of Mathematical Physics, 17(7):1099–1104, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i7/p1099_s1.

Ali:1976:ENQ

[584] S. Twareque Ali and H. D. Doebner. On the equivalence of nonrelativistic quantum mechanics based upon sharp and fuzzy measurements. Journal of Mathematical Physics, 17(7):1105–1111, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1105_s1.

Bayen:1976:RCS

[585] F. Bayen and M. Flato. Remarks on conformal space. Journal of Mathematical Physics, 17(7):1112–1114, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i7/p1112_s1.

Barut:1976:AMS

[586] A. O. Barut and H. Schneider. Angular momentum of systems of electric and magnetic charges and of singular flux surfaces. Journal of Mathematical Physics, 17(7):1115–1117, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i7/p1115_s1.

Brydges:1976:CEF

[587] David C. Brydges. Cluster expansions for fermion fields by the time dependent Hamiltonian approach. Journal of Mathematical Physics, 17(7):

1118–1124, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1118_s1.

Patterson:1976:CSUa

[588] Chris W. Patterson and William G. Harter. Canonical symmetrization for unitary bases. I. Canonical Weyl bases. Journal of Mathematical Physics, 17(7):1125–1136, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i7/p1125_s1.

Patterson:1976:CSUb

[589] Chris W. Patterson and William G. Harter. Canonical symmetrization for the unitary bases. II. Boson and fermion bases. Journal of Mathematical Physics, 17(7):1137–1142, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i7/p1137_s1.

Park:1976:UBS

[590] Yong Moon Park. Uniform bounds of the Schwinger functions in boson field models. Journal of Mathematical Physics, 17(7):1143–1147, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i7/p1143_s1.

Fox:1976:CGC

[591] Ronald Forrest Fox. Critique of the generalized cumulant expansion method. Journal of Mathematical Physics, 17(7):1148–1153, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/
p1148_s1.

Benci:1976:EDR

[592] V. Benci, D. Fortunato, and F. Zirilli. Exponential decay and regularity properties of the Hartree approximation to the bound state wavefunctions of the helium atom. Journal of Mathematical Physics, 17(7): 1154–1155, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1154_s1.

Luning:1976:ISH

[593] C. D. Luning and W. L. Perry. Iterative solution of the Hartree equations. Journal of Mathematical Physics, 17(7):1156–1159, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1156_s1.

Viano:1976:EOI

[594] G. A. Viano. On the extrapolation of optical image data. Journal of Mathematical Physics, 17(7):1160–1165, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i7/p1160_s1.

Winch:1976:IFW

[595] D. E. Winch. Integration formula for Wigner 3-j coefficients. Journal of Mathematical Physics, 17(7):1166–1170, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i7/p1166_s1.

Ray:1976:SGF

[596] Dipankar Ray. Self-gravitating fluids with cylindrical symmetry. Journal of Mathematical Physics, 17(7):1171–1172, July 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1171_s1.

Aguilera-Navarro:1976:EI

[597] M. C. K. Aguilera-Navarro and V. C. Aguilera-Navarro. Eigenvalues of invariants of U(n) and Su(n). Journal of Mathematical Physics, 17(7): 1173–1176, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1173_s1.

Alonso:1976:FCR

[598] Luis Martinez Alonso. Fundamental canonical realizations of connected Lie groups. Journal of Mathematical Physics, 17(7):1177–1181, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i7/p1177_s1.

Kelsey:1976:TOP

[599] Edward J. Kelsey and Joseph Macek. Nth order perturbation theory for hydrogen. Journal of Mathematical Physics, 17(7):1182–1185, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i7/p1182_s1.

Beran:1976:PTA

[600] M. J. Beran and J. J. McCoy. Propagation through an anisotropic random medium. An integro-differential formulation. Journal of Mathematical Physics, 17(7):1186–1189, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i7/p1186_s1.

Meuldermans:1976:CCE

[601] R. Meuldermans. A crossed-channel expansion of a conformal invariant scattering amplitude. Journal of Mathematical Physics, 17(7):1190– 1195, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1190_s1.

Driscoll:1976:MIC

[602] C. F. Driscoll and T. M. O’Neil. Modulational instability of cnoidal wave solutions of the modified Korteweg–de Vries equation. Journal of Mathematical Physics, 17(7):1196–1200, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i7/p1196_s1.

Narnhofer:1976:BCS

[603] H. Narnhofer and J. R. Klauder. Boundary conditions and singular potentials in diffusion theory. Journal of Mathematical Physics, 17(7): 1201–1209, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1201_s1.

Herbst:1976:CQF

[604] Ira W. Herbst. On canonical quantum field theories. Journal of Mathematical Physics, 17(7):1210–1221, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i7/p1210_s1.

Thide:1976:ELO

[605] Bo Thidé. The energy levels of the one-dimensional potential well V (X) = a∥X∥ calculated by means of certain phase-integral approximations. Journal of Mathematical Physics, 17(7):1222–1225, July 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1222_
s1.

Lee:1976:CGE

[606] Chul Hoon Lee. Coupled gravitational and electromagnetic perturbations around a charged black hole. Journal of Mathematical Physics, 17(7): 1226–1235, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1226_s1.

Szu:1976:NFE

[607] H. H. Szu, C. E. Carroll, C. C. Yang, and S. Ahn. A new functional equation in the plasma inverse problem and its analytic properties. Journal of Mathematical Physics, 17(7):1236–1247, July 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1236_s1.

deVega:1976:MRC

[608] H. J. de Vega and F. Schaposnik. Multiplicative renormalization of composite operators. Journal of Mathematical Physics, 17(7):1248–1251, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i7/p1248_s1.

Furutsu:1976:TID

[609] K. Furutsu. Theory of irradiance distribution function in turbulent media-cluster approximation. Journal of Mathematical Physics, 17(7): 1252–1263, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1252_s1.

Hoenselaers:1976:GSE

[610] C. Hoenselaers. On generation of solutions of Einstein’s equations. Journal of Mathematical Physics, 17(7):1264–1267, July 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1264_s1.

Berger:1976:HCM

[611] Beverly K. Berger. Homothetic and conformal motions in spacelike slices of solutions of Einstein’s equations. Journal of Mathematical Physics, 17(7):1268–1273, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i7/p1268_s1.

Lathouwers:1976:SPS

[612] L. Lathouwers. A supereigenvalue problem for the solution of the generator coordinate integral equation. Journal of Mathematical Physics, 17(7): 1274–1275, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1274_s1.

Haba:1976:MTR

[613] Zbigniew Haba. Measure-theoretic representations of the SU(2) ⊗ SU(2) current algebra with PCAC. Journal of Mathematical Physics, 17(7): 1276–1285, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1276_s1.

Brissaud:1976:PWN

[614] A. Brissaud, J. Léorat, and A. Mazure. Pseudo white noises: δ-correlated processes with finite memory. Journal of Mathematical Physics, 17(7): 1286–1292, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1286_s1.

Estabrook:1976:PSN

[615] F. B. Estabrook and H. D. Wahlquist. Prolongation structures of nonlinear evolution equations. II. Journal of Mathematical Physics, 17(7): 1293–1297, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1293_s1.

Gorini:1976:LSC

[616] Vittorio Gorini and Andrzej Kossakowski. N-level system in contact with a singular reservoir. Journal of Mathematical Physics, 17(7):1298– 1305, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1298_s1.

Hauser:1976:STK

[617] I. Hauser and R. J. Malhiot. On space–time Killing tensors with a Segrè characteristic ((11)(11)). Journal of Mathematical Physics, 17(7):1306– 1312, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1306_s1.

Peccia:1976:NIM

[618] A. Peccia and R. T. Sharp. Number of independent missing label operators. Journal of Mathematical Physics, 17(7):1313–1314, July 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1313_
s1.

Guven:1976:HGP

[619] R. Güven. Hertzian gravitational potentials for type D space–times. Journal of Mathematical Physics, 17(7):1315–1319, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1315_s1.

Hioe:1976:QTA

[620] F. T. Hioe, Don MacMillen, and E. W. Montroll. Quantum theory of anharmonic oscillators. II. Energy levels of oscillators with x^{2α} anharmonicity. Journal of Mathematical Physics, 17(7):1320–1337, July 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1320_
s1.

Rau:1976:CBB

[621] A. R. P. Rau. Cross between Born and WKB approximations: Variational solutions of nonlinear forms of the Schrödinger equation. Journal of Mathematical Physics, 17(7):1338–1344, July 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i7/p1338_s1.

Katzin:1976:GIF

[622] Gerald H. Katzin and Jack Levine. A gauge invariant formulation of time-dependent dynamical symmetry mappings and associated constants of motion for Lagrangian particle mechanics. I. Journal of Mathematical Physics, 17(7):1345–1350, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i7/p1345_s1.

Jaulent:1976:ISP

[623] M. Jaulent. Inverse scattering problems in absorbing media. Journal of Mathematical Physics, 17(7):1351–1360, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i7/p1351_s1.

Balanis:1976:AAD

[624] George N. Balanis. Addendum: “Analysis of the dispersion of low frequency uniaxial waves in heterogeneous periodic elastic media [J. Math. Phys. 16, 1383 (1975)]”. Journal of Mathematical Physics, 17(7):1360, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i7/p1360_s1. See [230].

Aly:1976:ESS

[625] H. H. Aly, H. J. W. Müller-Kirsten, and N. Vahedi-Faridi. Erratum: “Scattering by singular potentials with a perturbation-theoretical introduction to Mathieu functions”. Journal of Mathematical Physics, 17(7): 1361, July 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i7/p1361_s1. See [166].

Abellanas:1976:QAV

[626] L. Abellanas and L. Martinez Alonso. Quantization from the algebraic viewpoint. Journal of Mathematical Physics, 17(8):1363–1365, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i8/p1363_s1.

Semon:1976:CSS

[627] Mark D. Semon and John R. Taylor. Cross sections for screened potentials. Journal of Mathematical Physics, 17(8):1366–1370, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i8/p1366_s1.

McKay:1976:BRC

[628] W. McKay, J. Patera, and R. T. Sharp. Branching rules and Clebsch– Gordan series for E_{8}. Journal of Mathematical Physics, 17(8):1371– 1375, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i8/p1371_s1.

Green:1976:SLP

[629] H. S. Green, C. A. Hurst, and Y. Ilamed. The state labeling problems for SO(N) in U(N) and U(M) in Sp(2M). Journal of Mathematical Physics, 17(8):1376–1382, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i8/p1376_s1.

Williams:1976:PPS

[630] S. A. Williams and D. L. Pursey. Particle permutation symmetry of multishell states. I. Two shells. Journal of Mathematical Physics, 17(8):1383– 1391, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i8/p1383_s1.

Massot:1976:SIP

[631] J. N. Massot and R. Bacis. Symmetry and invariance properties of the Boltzmann equation on different groups. Journal of Mathematical Physics, 17(8):1392–1403, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i8/p1392_s1.

Deutsch:1976:OCP

[632] C. Deutsch. One-component plasma in 2+ϵ dimensions. Journal of Mathematical Physics, 17(8):1404–1413, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i8/p1404_s1. See [940].

Wang:1976:SIC

[633] P. K. C. Wang. Stability and instability conditions for nonlinear evolutional equations in Hilbert spaces. Journal of Mathematical Physics, 17(8):1414–1420, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i8/p1414_s1.

Berntsen:1976:RMS

[634] Svend Berntsen and Günther Johannsen. Relativistic multiple scattering of electromagnetic waves. Journal of Mathematical Physics, 17(8):1421– 1428, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i8/p1421_s1.

Chang:1976:NSC

[635] S. C. Chang and Allen I. Janis. Nonspreading solutions to a class of differential equations. Journal of Mathematical Physics, 17(8):1429– 1431, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i8/p1429_s1.

Chang:1976:RCB

[636] S. C. Chang and Allen I. Janis. Radiation in cosmological backgrounds. Journal of Mathematical Physics, 17(8):1432–1438, August 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i8/p1432_
s1.

Boyer:1976:SBI

[637] C. P. Boyer, R. T. Sharp, and P. Winternitz. Symmetry breaking interactions for the time dependent Schrödinger equation. Journal of Mathematical Physics, 17(8):1439–1451, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i8/p1439_s1.

Quesne:1976:SEA

[638] C. Quesne. SU(2) × SU(2) scalars in the enveloping algebra of SU(4). Journal of Mathematical Physics, 17(8):1452–1467, August 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i8/p1452_
s1.

Pauri:1976:CRP

[639] M. Pauri and G. M. Prosperi. Canonical realizations of the Poincaré group. II. Space–time description of two particles interacting at a distance, Newtonian-like equations of motion and approximately relativistic Lagrangian formulation. Journal of Mathematical Physics, 17(8):1468– 1495, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i8/p1468_s1.

Cohn:1976:ARS

[640] J. Cohn and H. Wiebe. Asymptotic radiation from spinning charged particles. Journal of Mathematical Physics, 17(8):1496–1500, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i8/p1496_s1.

Fujita:1976:ECE

[641] Shigeji Fujita. Exact closed evolution equation for the electron density operator averaged over impurity configurations. Journal of Mathematical Physics, 17(8):1501–1508, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v17/i8/p1501_s1.

King:1976:SRO

[642] Frederick W. King. Sum rules for the optical constants. Journal of Mathematical Physics, 17(8):1509–1514, August 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i8/p1509_s1.

Lucke:1976:RCP

[643] W. Lücke. Ruelle’s cluster property for nonstrictly localizable fields derived from symmetry of their Wightman functions. Journal of Mathematical Physics, 17(8):1515–1520, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i8/p1515_s1.

Ruhl:1976:RUC

[644] W. Rühl and B. C. Yunn. Representations of the universal covering group of SU(1, 1) and their bilinear and trilinear invariant forms. Journal of Mathematical Physics, 17(8):1521–1530, August 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i8/p1521_s1.

Safko:1976:LMG

[645] J. L. Safko and F. Elston. Lagrange multipliers and gravitational theory. Journal of Mathematical Physics, 17(8):1531–1537, August 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i8/p1531_
s1.

Budgor:1976:QMT

[646] Aaron B. Budgor. Quantum mechanics on topological networks. Journal of Mathematical Physics, 17(8):1538–1545, August 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i8/p1538_s1.

Bergere:1976:BPT

[647] M. C. Bergere and Yuk-Ming P. Lam. Bogolubov–Parasiuk theorem in the α-parametric representation. Journal of Mathematical Physics, 17(8):1546–1557, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i8/p1546_s1.

Wong:1976:SMF

[648] M. K. F. Wong. On the structure of the multiplicity-free Wigner coefficients of U(n). Journal of Mathematical Physics, 17(8):1558–1569, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i8/p1558_s1.

Ostebee:1976:NAT

[649] A. Ostebee, P. Gambardella, and M. Dresden. A “nonstandard” approach to the thermodynamic limit. II. weakly tempered potentials and neutral Coulomb systems. Journal of Mathematical Physics, 17(8):1570– 1578, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i8/p1570_s1.

Zeleznik:1976:T

[650] Frank J. Zeleznik. Thermodynamics. Journal of Mathematical Physics, 17(8):1579–1610, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i8/p1579_s1.

Nwachuku:1976:ECO

[651] C. O. Nwachuku and M. A. Rashid. Eigenvalues of the Casimir operators of the orthogonal and symplectic groups. Journal of Mathematical Physics, 17(8):1611–1616, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i8/p1611_s1.

Saenz:1976:EEO

[652] A. W. Sáenz and W. W. Zachary. Erratum: “On the equality of S operators corresponding to unitarily equivalent Hamiltonians in single channel scattering”. Journal of Mathematical Physics, 17(8):1616, August 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i8/p1616_s1. See [466].

Gilmore:1976:BPD

[653] R. Gilmore and C. M. Bowden. Bifurcation properties of Dicke Hamiltonians. Journal of Mathematical Physics, 17(9):1617–1625, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i9/p1617_s1.

Scheunert:1976:CASa

[654] M. Scheunert, W. Nahm, and V. Rittenberg. Classification of all simple graded Lie algebras whose Lie algebra is reductive. I. Journal of Mathematical Physics, 17(9):1626–1639, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i9/p1626_s1.

Scheunert:1976:CASb

[655] M. Scheunert, W. Nahm, and V. Rittenberg. Classification of all simple graded Lie algebras whose Lie algebra is reductive. II. Construction of the exceptional algebras. Journal of Mathematical Physics, 17(9):1640– 1644, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i9/p1640_s1.

Stephenson:1976:SCS

[656] John Stephenson. Soluble classical spin model with competing interactions. Journal of Mathematical Physics, 17(9):1645–1654, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i9/p1645_s1.

Gasser:1976:HCG

[657] J. Gasser. How to calculate the grand partition function in the uncorrelated jet model. Journal of Mathematical Physics, 17(9):1655–1661, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i9/p1655_s1.

Leveille:1976:ILG

[658] J. P. Leveille and P. Roman. On the implementability of local gauge transformations in a theory with localized states. Journal of Mathematical Physics, 17(9):1662–1663, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i9/p1662_s1.

Roman:1976:GTN

[659] P. Roman and J. Haavisto. Gauge theories and nonrelativistic cosmological symmetries. Journal of Mathematical Physics, 17(9):1664–1670, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i9/p1664_s1.

Lockwood:1976:CPC

[660] Loren A. Lockwood. Canonical parameters of the 3j coefficient. Journal of Mathematical Physics, 17(9):1671–1672, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i9/p1671_s1.

Prugovecki:1976:QTP

[661] E. Prugovečki. Quantum two-particle scattering in fuzzy phase space. Journal of Mathematical Physics, 17(9):1673–1681, September 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i9/p1673_
s1.

Ellis:1976:QMS

[662] Richard S. Ellis and Charles M. Newman. Quantum mechanical soft springs and reverse correlation inequalities. Journal of Mathematical Physics, 17(9):1682–1683, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i9/p1682_s1.

Yan:1976:BMN

[663] C. C. Yan. Binary mixture with nearest and next nearest neighbor interaction on a one-dimensional lattice. Journal of Mathematical Physics, 17(9):1684–1687, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i9/p1684_s1.

Epifanio:1976:MRU

[664] G. Epifanio. On the matrix representation of unbounded operators. Journal of Mathematical Physics, 17(9):1688–1691, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i9/p1688_s1.

Davidson:1976:EDM

[665] Russell Davidson and John J. Kozak. Exact dynamics of a model for a three-level atom. Journal of Mathematical Physics, 17(9):1692–1702, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i9/p1692_s1.

Case:1976:FDM

[666] K. M. Case. Fredholm determinants and multiple solitons. Journal of Mathematical Physics, 17(9):1703–1706, September 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i9/p1703_s1.

Besieris:1976:DES

[667] Ioannis M. Besieris. Derivation of an exact spectral density transport equation for a nonstationary scattering medium. Journal of Mathematical Physics, 17(9):1707–1710, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i9/p1707_s1.

Besieris:1976:CFS

[668] Ioannis M. Besieris and Wojciech B. Stasiak. Closed first- and secondorder moment equations for stochastic nonlinear problems with applications to model hydrodynamic and Vlasov-plasma turbulence. Journal of Mathematical Physics, 17(9):1711–1718, September 1976. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v17/i9/p1711_s1.

Sud:1976:NAC

[669] Krishan Sud and L. E. Wright. A new analytic continuation of Appell’s hypergeometric series F_{2}. Journal of Mathematical Physics, 17(9):1719– 1721, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i9/p1719_s1.

Bowden:1976:SNM

[670] R. L. Bowden, P. F. Zweifel, and R. Menikoff. On the solution of nonlinear matrix integral equations in transport theory. Journal of Mathematical Physics, 17(9):1722–1726, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i9/p1722_s1.

Maugin:1976:CTDa

[671] Gérard A. Maugin. A continuum theory of deformable ferrimagnetic bodies. I. Field equations. Journal of Mathematical Physics, 17(9):1727– 1738, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i9/p1727_s1.

Maugin:1976:CTDb

[672] Gérard A. Maugin. A continuum theory of deformable ferrimagnetic bodies. II. thermodynamics, constitutive theory. Journal of Mathematical Physics, 17(9):1739–1751, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i9/p1739_s1.

DeSanto:1976:IWP

[673] John A. DeSanto. Inverse wave propagation in an inhomogeneous waveguide. Journal of Mathematical Physics, 17(9):1752–1753, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i9/p1752_s1.

Flato:1976:DPB

[674] M. Flato, A. Lichnerowicz, and D. Sternheimer. Deformations of Poisson brackets, Dirac brackets and applications. Journal of Mathematical Physics, 17(9):1754–1762, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i9/p1754_s1.

Roy:1976:SJF

[675] M. N. Sinha Roy, B. Talukdar, and D. Chattarji. Off-shell Jost function and T matrix for the Woods–Saxon potential. Journal of Mathematical Physics, 17(9):1763–1766, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i9/p1763_s1.

Pao:1976:CLT

[676] C. V. Pao. Critical length of a transport process in rod geometry. Journal of Mathematical Physics, 17(9):1767–1772, September 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i9/p1767_s1.

Prosser:1976:FSI

[677] Reese T. Prosser. Formal solutions of inverse scattering problems. II. Journal of Mathematical Physics, 17(10):1775–1779, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i10/
p1775_s1.

Backhouse:1976:IRF

[678] Nigel B. Backhouse and Patricia Gard. On induced representations for finite groups. Journal of Mathematical Physics, 17(10):1780–1784, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1780_s1.

Chen:1976:MSQ

[679] Etang Chen. Markovian subdynamics in quantum dynamical systems. Journal of Mathematical Physics, 17(10):1785–1789, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i10/
p1785_s1.

Oliver:1976:PFS

[680] D. R. Oliver, Jr. and W. R. Davis. Perfect fluids and symmetry mappings leading to conservation laws. Journal of Mathematical Physics, 17(10):1790–1792, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i10/p1790_s1.

Stehney:1976:PND

[681] Ann K. Stehney. Principal null directions without spinors (in vacuum space times). Journal of Mathematical Physics, 17(10):1793–1796, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1793_s1.

Kerner:1976:EIF

[682] Edward H. Kerner. Extended inertial frames and Lorentz transformations. II. Journal of Mathematical Physics, 17(10):1797–1807, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1797_s1.

Krause:1976:HCL

[683] F. Krause and P. H. Roberts. The high conductivity limit in mean field electrodynamics. Journal of Mathematical Physics, 17(10):1808– 1809, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i10/p1808_s1.

Lindblom:1976:ACS

[684] Lee Lindblom. On the analyticity of certain stationary nonvacuum Einstein space–times. Journal of Mathematical Physics, 17(10):1810– 1811, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i10/p1810_s1.

Larsen:1976:SOD

[685] Edward W. Larsen and Paul F. Zweifel. Steady, one-dimensional multigroup neutron transport with anisotropic scattering. Journal of Mathematical Physics, 17(10):1812–1820, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i10/p1812_s1.

Moses:1976:SPA

[686] H. E. Moses. A simple proof of the angular momentum Helmholtz theorem and the relation of the theorem to the decomposition of solenoidal vectors into poloidal and toroidal components. Journal of Mathematical Physics, 17(10):1821–1823, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i10/p1821_s1.

Reina:1976:CES

[687] C. Reina. Complete extension of the symmetry axis of the Tomimatsu– Sato solution of the Einstein equations. Journal of Mathematical Physics, 17(10):1824–1826, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i10/p1824_s1.

Harnad:1976:GTS

[688] J. P. Harnad and R. B. Pettitt. Gauge theories for space–time symmetries. I. Journal of Mathematical Physics, 17(10):1827–1837, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1827_s1.

Hurd:1976:DHP

[689] R. A. Hurd and S. Przeździecki. Diffraction by a half-plane perpendicular to the distinguished axis of a gyrotropic medium. Journal of Mathematical Physics, 17(10):1838–1847, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i10/p1838_s1.

Bouziane:1976:BIU

[690] M. Bouziane and Ph. A. Martin. Bogoliubov inequality for unbounded operators and the Bose gas. Journal of Mathematical Physics, 17 (10):1848–1851, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i10/p1848_s1.

Truman:1976:FPI

[691] Aubrey Truman. Feynman path integrals and quantum mechanics as h → 0. Journal of Mathematical Physics, 17(10):1852–1862, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1852_s1.

Cohen:1976:QPV

[692] Leon Cohen. Quantization problem and variational principle in the phase-space formulation of quantum mechanics. Journal of Mathematical Physics, 17(10):1863–1866, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i10/p1863_s1.

Morris:1976:PSG

[693] Hedley C. Morris. Prolongation structures and a generalized inverse scattering problem. Journal of Mathematical Physics, 17(10):1867–1869, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1867_s1.

Morris:1976:PSN

[694] H. C. Morris. Prolongation structures and nonlinear evolution equations in two spatial dimensions. Journal of Mathematical Physics, 17(10):1870– 1872, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i10/p1870_s1.

Dowker:1976:PIO

[695] J. S. Dowker. Path integrals and ordering rules. Journal of Mathematical Physics, 17(10):1873–1874, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i10/p1873_s1.

Gunaydin:1976:OHS

[696] Murat Günaydin. Octonionic Hilbert spaces, the Poincaré group and SU(3). Journal of Mathematical Physics, 17(10):1875–1883, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1875_s1.

Trullinger:1976:ETC

[697] S. E. Trullinger. Expansion technique for crystal surface problems. Journal of Mathematical Physics, 17(10):1884–1888, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i10/p1884_s1.

Challifour:1976:EFT

[698] John L. Challifour. Euclidean field theory. II. Remarks on embedding the relativistic Hilbert space. Journal of Mathematical Physics, 17(10):1889– 1892, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i10/p1889_s1.

Moncrief:1976:STS

[699] Vincent Moncrief. Space–time symmetries and linearization stability of the Einstein equations. II. Journal of Mathematical Physics, 17(10):1893– 1902, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i10/p1893_s1.

Daumens:1976:TSHa

[700] M. Daumens and P. Minnaert. Tensor spherical harmonics and tensor multipoles. I. Euclidean space. Journal of Mathematical Physics, 17(10):1903–1909, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i10/p1903_s1.

Powers:1976:RII

[701] Robert T. Powers. Resistance inequalities for the isotropic Heisenberg ferromagnet. Journal of Mathematical Physics, 17(10):1910–1918, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i10/p1910_s1.

Chacon:1976:LAS

[702] E. Chacón, D. Levi, and M. Moshinsky. Lie algebras in the Schrödinger picture and radial matrix elements. Journal of Mathematical Physics, 17(10):1919–1929, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i10/p1919_s1.

Strocchi:1976:EPC

[703] F. Strocchi and A. S. Wightman. Erratum: “Proof of the charge superselection rule in local relativistic quantum field theory”. Journal of Mathematical Physics, 17(10):1930–1931, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i10/p1930_s1. See [11].

Durand:1976:EFA

[704] Loyal Durand, Paul M. Fishbane, and L. M. Simmons, Jr. Expansion formulas and addition theorems for Gegenbauer functions. Journal of Mathematical Physics, 17(11):1933–1948, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i11/p1933_s1.

Raveche:1976:BSC

[705] Harold J. Raveché and C. A. Stuart. Bifurcation of solutions with crystalline symmetry. Journal of Mathematical Physics, 17(11):1949–1953, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p1949_s1.

daCosta:1976:SDR

[706] R. C. T. da Costa and R. Lobo. A soluble dispersion relation for a three-dimensional band structure. Journal of Mathematical Physics, 17 (11):1954–1956, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p1954_s1.

Raszillier:1976:ZPE

[707] I. Raszillier, W. Schmidt, and I.-S. Stefanescu. On zeroes of the pion electromagnetic form factor. Journal of Mathematical Physics, 17(11):1957– 1962, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p1957_s1.

Aragone:1976:ISS

[708] C. Aragone, E. Chalbaud, and S. Salamó. On intelligent spin states. Journal of Mathematical Physics, 17(11):1963–1971, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i11/
p1963_s1.

Patera:1976:HIG

[709] J. Patera, R. T. Sharp, and P. Winternitz. Higher indices of group representations. Journal of Mathematical Physics, 17(11):1972–1979, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i11/p1972_s1. See [977].

Horndeski:1976:CCE

[710] Gregory Walter Horndeski. Conservation of charge and the Einstein– Maxwell field equations. Journal of Mathematical Physics, 17(11):1980– 1987, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p1980_s1.

Lee:1976:CCF

[711] L. L. Lee. Continuum calculus and Feynman’s path integrals. Journal of Mathematical Physics, 17(11):1988–1997, November 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i11/

p1988_s1. See [1330].

Perroud:1976:IRL

[712] M. Perroud. On the irreducible representations of the Lie algebra chain G_{2} ⊇ A_{2}. Journal of Mathematical Physics, 17(11):1998–2006, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i11/p1998_s1.

Reichl:1976:NREa

[713] L. E. Reichl. Nonlinear response of equilibrium strongly coupled Fermi fluids. I. Formal development. Journal of Mathematical Physics, 17 (11):2007–2022, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2007_s1.

Reichl:1976:NREb

[714] L. E. Reichl. Nonlinear response of equilibrium strongly coupled Fermi fluids. II. Fourier expansion and partial resummation of expectation values-polarization diagrams. Journal of Mathematical Physics, 17 (11):2023–2033, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2023_s1.

Reichl:1976:PTE

[715] L. E. Reichl. Propagator techniques for equilibrium strongly coupled Fermi fluids. III. Journal of Mathematical Physics, 17(11):2034–2040, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p2034_s1.

Haacke:1976:CCG

[716] E. M. Haacke, J. W. Moffat, and P. Savaria. A calculation of SU(4) Clebsch–Gordan coefficients. Journal of Mathematical Physics, 17(11): 2041–2066, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2041_s1.

Hoshen:1976:GFF

[717] J. Hoshen and R. Kopelman. Green’s functions for a face centered orthorhombic lattice. Journal of Mathematical Physics, 17(11):2067–2073, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p2067_s1.

Wollenberg:1976:ALA

[718] L. S. Wollenberg. Automorphisms of the Lie algebra of polynomials under Poisson bracket. Journal of Mathematical Physics, 17(11):2074– 2081, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p2074_s1.

Keskinen:1976:LCC

[719] R. Keskinen and M. Lehtinen. On the linear connection and curvature in Newtonian mechanics. Journal of Mathematical Physics, 17(11):2082– 2084, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p2082_s1.

Daumens:1976:TSHb

[720] M. Daumens and P. Minnaert. Tensor spherical harmonics and tensor multipoles. II. Minkowski space. Journal of Mathematical Physics, 17 (11):2085–2091, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2085_s1.

Greenberg:1976:SMT

[721] William Greenberg and Selim Sancaktar. Solution of the multigroup transport equation in L^{p} spaces. Journal of Mathematical Physics, 17 (11):2092–2097, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2092_s1.

Backhouse:1976:TRC

[722] Nigel Backhouse and Patricia Gard. On the tensor representation for compact groups. Journal of Mathematical Physics, 17(11):2098–2100, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p2098_s1.

Candelas:1976:QFT

[723] P. Candelas and D. J. Raine. Quantum field theory on incomplete manifolds. Journal of Mathematical Physics, 17(11):2101–2112, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i11/p2101_s1.

Ladde:1976:SDI

[724] G. S. Ladde. Systems of differential inequalities and stochastic differential equations. III. Journal of Mathematical Physics, 17(11):2113–2120, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i11/p2113_s1.

Berenson:1976:ECG

[725] Rhoda Berenson and Joseph L. Birman. Erratum: “Clebsch–Gordan coefficients for crystal space groups”. Journal of Mathematical Physics, 17(11):2121, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2121_s1. See [51].

Car:1976:EOF

[726] R. Car, G. Ciucci, and L. Quartapelle. Erratum: “Optimal factor group for nonsymmorphic space groups”. Journal of Mathematical Physics, 17(11):2121, November 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i11/p2121_s2. See [576].

Frigerio:1976:LSC

[727] Alberto Frigerio and Vittorio Gorini. N-level systems in contact with a singular reservoir. II. Journal of Mathematical Physics, 17(12):2123– 2127, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i12/p2123_s1.

Chong:1976:SPC

[728] Tak-Yuen Chong. Some properties of canonical invariantly relaxing (CIR) systems. Journal of Mathematical Physics, 17(12):2128–2132, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i12/p2128_s1.

Brydges:1976:NEB

[729] David Brydges and Paul Federbush. A note on energy bounds for boson matter. Journal of Mathematical Physics, 17(12):2133–2134, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i12/p2133_s1.

Lichnerowicz:1976:SWR

[730] André Lichnerowicz. Shock waves in relativistic magnetohydrodynamics under general assumptions. Journal of Mathematical Physics, 17 (12):2135–2142, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i12/p2135_s1.

Cornille:1976:DES

[731] Henri Cornille. Differential equations satisfied by Fredholm determinants and application to the inversion formalism for parameter dependent potentials. Journal of Mathematical Physics, 17(12):2143–2158, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i12/p2143_s1.

Lee:1976:UPR

[732] Chul Hoon Lee and W. D. McGlinn. Uniqueness of perturbation of a Reissner–Nordström black hole. Journal of Mathematical Physics, 17 (12):2159–2165, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i12/p2159_s1.

Baur:1976:ZGC

[733] M. E. Baur and R. M. Sinder. Zeros of the grand canonical partition function: Generalization of a result of Lee and Yang. Journal of Mathematical Physics, 17(12):2166–2168, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i12/p2166_s1. See [1004].

Honig:1976:IGK

[734] E. Honig. Intrinsic geometry of Killing trajectories. Journal of Mathematical Physics, 17(12):2169–2174, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i12/p2169_s1.

Sud:1976:RIF

[735] Krishan Sud, L. E. Wright, and D. S. Onley. Radial integrals with finite energy loss for Dirac–Coulomb functions. Journal of Mathematical Physics, 17(12):2175–2181, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i12/p2175_s1.

Berryman:1976:FDE

[736] James G. Berryman and Samuel P. Bowen. The frequency dependent electrical conductivity for disordered alloys: Application of an abstract Hilbert space generalization of Feenberg’s perturbation theory. Journal of Mathematical Physics, 17(12):2182–2191, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v17/i12/p2182_s1.

McLenaghan:1976:SEM

[737] Raymond G. McLenaghan and Nessim Tariq. On a solution of the Einstein–Maxwell equations admitting a nonsingular electromagnetic field. Journal of Mathematical Physics, 17(12):2192–2197, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i12/p2192_s1.

Mehta:1976:NCM

[738] M. L. Mehta. A note on certain multiple integrals. Journal of Mathematical Physics, 17(12):2198–2202, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i12/p2198_s1.

Hacyan:1976:SBP

[739] Shahen Hacyan and Jerzy Plebański. Some basic properties of Killing spinors. Journal of Mathematical Physics, 17(12):2203–2206, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i12/p2203_s1.

Finley:1976:ISS

[740] J. D. Finley III and J. F. Plebański. The intrinsic spinorial structure of hyperheavens. Journal of Mathematical Physics, 17(12):2207–2214, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i12/p2207_s1.

Dragt:1976:LSI

[741] Alex J. Dragt and John M. Finn. Lie series and invariant functions for analytic symplectic maps. Journal of Mathematical Physics, 17(12):2215– 2227, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i12/p2215_s1.

Ezawa:1976:MVG

[742] Z. F. Ezawa and H. C. Tze. Monopoles, vortices and the geometry of the Yang–Mills bundles. Journal of Mathematical Physics, 17(12):2228– 2231, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v17/i12/p2228_s1.

Catenacci:1976:SAE

[743] Roberto Catenacci and Joaquin Diaz Alonso. On the stationary axisymmetric Einstein–Maxwell field equations. Journal of Mathematical Physics, 17(12):2232–2235, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i12/p2232_s1.

Kalnins:1977:LTWa

[744] E. G. Kalnins and W. Miller, Jr. Lie theory and the wave equation in space–time. I. The Lorentz group. Journal of Mathematical Physics, 18(1):1–16, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p1_s1.

Gilmore:1977:SSP

[745] R. Gilmore. Structural stability of the phase transition in Dicke-like models. Journal of Mathematical Physics, 18(1):17–22, January 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i1/p17_s1.

Newman:1977:MAC

[746] C. M. Newman and L. S. Schulman. Metastability and the analytic continuation of eigenvalues. Journal of Mathematical Physics, 18(1): 23–30, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i1/p23_s1.

Folland:1977:FIS

[747] N. O. Folland. Factored irreducible symmetry operators. Journal of Mathematical Physics, 18(1):31–36, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i1/p31_s1.

Bricmont:1977:WAP

[748] J. Bricmont and F. Debacker-Mathot. The Wegner approximation of the plane rotator model as a massless, free, lattice, Euclidean field. Journal of Mathematical Physics, 18(1):37–40, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i1/p37_s1.

Jang:1977:PEC

[749] Pong Soo Jang and Robert M. Wald. The positive energy conjecture and the cosmic censor hypothesis. Journal of Mathematical Physics, 18(1): 41–44, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i1/p41_s1.

Lockwood:1977:CPC

[750] Loren A. Lockwood. Canonical parameters of the 6j coefficient. Journal of Mathematical Physics, 18(1):45–46, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i1/p45_s1.

Wang:1977:NMS

[751] A. P. Wang. Nonstationary multiple scattering (in radiative transfer). Journal of Mathematical Physics, 18(1):47–51, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i1/p47_s1.

Mansouri:1977:GSB

[752] Freydoon Mansouri. Geometry of superspaces with Bose and Fermi coordinates and applications to graded Lie bundles and supergravity. Journal of Mathematical Physics, 18(1):52–57, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i1/p52_s1.

Ko:1977:KSA

[753] M. Ko, E. T. Newman, and R. Penrose. The Kähler structure of asymptotic twistor space. Journal of Mathematical Physics, 18(1):58–64, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p58_s1.

Hietarinta:1977:ELS

[754] Jarmo Hietarinta. On the existence of localized solutions in nonlinear chiral theories. Journal of Mathematical Physics, 18(1):65–71, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p65_s1.

Beckers:1977:SEG

[755] J. Beckers, J. Patera, M. Perroud, and P. Winternitz. Subgroups of the Euclidean group and symmetry breaking in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 18(1):72–83, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p72_s1.

Rund:1977:CPT

[756] Hanno Rund. Clebsch potentials in the theory of electromagnetic fields admitting electric and magnetic charge distributions. Journal of Mathematical Physics, 18(1):84–95, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i1/p84_s1.

Froman:1977:WPQ

[757] Nanny Fröman and Per Olof Fröman. On Wentzel’s proof of the quantisation condition for a single-well potential. Journal of Mathematical Physics, 18(1):96–99, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p96_s1.

Hsu:1977:SED

[758] J. P. Hsu and E. Mac. Symmetry and exact dyon solutions for classical Yang–Mills field equations. Journal of Mathematical Physics, 18(1):100– 103, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i1/p100_s1.

Lehnigk:1977:MWP

[759] Siegfried H. Lehnigk. Maxwell and Wien processes as special cases of the generalized Feller diffusion process. Journal of Mathematical Physics, 18(1):104–105, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p104_s1.

Eimerl:1977:NTS

[760] David Eimerl. Nonexistence theorem for spherically asymmetric solutions of the Fermi–Thomas model of atoms and ions. Journal of Mathematical Physics, 18(1):106–109, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i1/p106_s1.

Schwartz:1977:CPT

[761] Charles Schwartz. A classical perturbation theory. Journal of Mathematical Physics, 18(1):110–112, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i1/p110_s1. See addendum [915].

Christensen:1977:GMI

[762] Mark J. Christensen. Gleason measures on infinite tensor products of Hilbert spaces. Journal of Mathematical Physics, 18(1):113–115, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p113_s1.

Dobson:1977:ERS

[763] J. Dobson. Exact results for second-neighbor surface magnons in an FCC lattice. Journal of Mathematical Physics, 18(1):116–117, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p116_s1.

Pick:1977:HLS

[764] Stepán Pick. Hamiltonians with x^{−2}-like singularity. Journal of Mathematical Physics, 18(1):118–119, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i1/p118_s1.

Oliveira:1977:HDS

[765] Colber G. Oliveira. Hamiltonian dynamics in symplectic phase space and Nambu’s formulation of mechanics. Journal of Mathematical Physics, 18(1):120–125, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p120_s1.

Dassios:1977:CLF

[766] George Dassios. Convergent low-frequency expansions for penetrable scatterers. Journal of Mathematical Physics, 18(1):126–137, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p126_s1.

Doebner:1977:MAM

[767] H. D. Doebner and W. Lücke. Macrocausal aspects of the S matrix related to the LSZ asymptotic condition. Journal of Mathematical Physics, 18(1):138–145, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p138_s1.

Scheunert:1977:GLA

[768] M. Scheunert, W. Nahm, and V. Rittenberg. Graded Lie algebras: Generalization of Hermitian representations. Journal of Mathematical Physics, 18(1):146–154, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i1/p146_s1.

Scheunert:1977:IRO

[769] M. Scheunert, W. Nahm, and V. Rittenberg. Irreducible representations of the osp(2,1) and spl(2,1) graded Lie algebras. Journal of Mathematical Physics, 18(1):155–162, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i1/p155_s1.

Corones:1977:SPC

[770] James Corones. Solitons, pseudopotentials, and certain Lie algebras. Journal of Mathematical Physics, 18(1):163–164, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i1/p163_s1.

Weingarten:1977:GFE

[771] Don Weingarten. Geometric formulation of electrodynamics and general relativity in discrete space–time. Journal of Mathematical Physics, 18(1):165–170, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p165_s1.

Baumann:1977:WSI

[772] Klaus Baumann and Gerhard C. Hegerfeldt. Is the Wick square infinitely divisible? Journal of Mathematical Physics, 18(1):171–172, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p171_s1.

Antippa:1977:GFS

[773] Adel F. Antippa and Alain J. Phares. General formalism solving linear recursion relations. Journal of Mathematical Physics, 18(1):173– 181, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i1/p173_s1.

vanNieuwenhuizen:1977:IRI

[774] P. van Nieuwenhuizen and C. C. Wu. On integral relations for invariants constructed from three Riemann tensors and their applications in quantum gravity. Journal of Mathematical Physics, 18(1):182–186, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p182_s1.

Geller:1977:CNJ

[775] Murray Geller. Comment on “New Jacobian ϑ functions and the evaluation of lattice sums”, by I. J. Zucker (J. Math. Phys. 16, 2189 (1975)). Journal of Mathematical Physics, 18(1):187, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i1/p187_s1. See

[352, 544].

Spohn:1977:ESL

[776] Herbert Spohn. Erratum: “The spectrum of the Liouville–von Neumann operator”. Journal of Mathematical Physics, 18(1):188, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i1/p188_s1. See [414].

Ruttimann:1977:JPS

[777] Gottfried T. Rüttimann. Jauch–Piron states (in quantum logic). Journal of Mathematical Physics, 18(2):189–193, February 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i2/p189_s1.

Bleistein:1977:NIS

[778] Norman Bleistein and Jack K. Cohen. Nonuniqueness in the inverse source problem in acoustics and electromagnetics. Journal of Mathematical Physics, 18(2):194–201, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i2/p194_s1. See erratum [1953].

Peccia:1977:SSA

[779] Anthony Peccia. Subalgebras of the similitude algebra and their invariants. Journal of Mathematical Physics, 18(2):202–214, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i2/p202_s1.

Miyazaki:1977:EOS

[780] Eizo Miyazaki and Iwao Yasumori. Exact occupation statistics for twodimensional lattices of single particles. Journal of Mathematical Physics, 18(2):215–218, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i2/p215_s1.

Ali:1977:SIR

[781] S. Twareque Ali and E. Prugovečki. Systems of imprimitivity and representations of quantum mechanics on fuzzy phase spaces. Journal of Mathematical Physics, 18(2):219–228, February 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i2/p219_s1.

Dollard:1977:DSS

[782] John D. Dollard. On the definition of scattering subspaces in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 18(2):229– 232, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i2/p229_s1.

Ernst:1977:NFS

[783] Frederick J. Ernst. A new family of solutions of the Einstein field equations. Journal of Mathematical Physics, 18(2):233–234, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i2/p233_s1.

Arrowsmith:1977:PTD

[784] D. K. Arrowsmith and J. W. Essam. Percolation theory on directed graphs. Journal of Mathematical Physics, 18(2):235–238, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i2/
p235_s1.

Backhouse:1977:KFG

[785] Nigel B. Backhouse. The Killing form for graded Lie algebras. Journal of Mathematical Physics, 18(2):239–244, February 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i2/p239_s1.

Ray:1977:PWS

[786] Dipankar Ray. Plane wave solutions in scalar tensor theories and solutions of source-free Einstein–Maxwell theory. Journal of Mathematical Physics, 18(2):245–249, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i2/p245_s1.

Galloway:1977:CAC

[787] Gregory J. Galloway. Closure in anisotropic cosmological models. Journal of Mathematical Physics, 18(2):250–252, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i2/p250_s1.

Dankel:1977:HSS

[788] Thad Dankel, Jr. Higher spin states in the stochastic mechanics of the Bopp–Haag spin model. Journal of Mathematical Physics, 18(2):253– 255, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i2/p253_s1.

Kumei:1977:GTA

[789] Sukeyuki Kumei. Group theoretic aspects of conservation laws of nonlinear dispersive waves: KdV type equations and nonlinear Schrödinger equations. Journal of Mathematical Physics, 18(2):256–264, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i2/p256_s1.

Banerjee:1977:PSE

[790] A. Banerjee and N. Chakrabarty. On plane symmetric Einstein–Maxwell fields. Journal of Mathematical Physics, 18(2):265–266, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i2/
p265_s1.

Baumgarten:1977:CSN

[791] D. Baumgarten, G. Braunss, and O. Wagner. Causal solutions of nonlinear wave and spinor equations obtained by Gel’fand-Shilov regularization. Journal of Mathematical Physics, 18(2):267–270, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i2/
p267_s1.

Kalnins:1977:LTWb

[792] E. G. Kalnins and W. Miller, Jr. Lie theory and the wave equation in space–time. 3. Semisubgroup coordinates. Journal of Mathematical Physics, 18(2):271–280, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i2/p271_s1.

Bowden:1977:CNT

[793] R. L. Bowden, W. L. Cameron, and P. F. Zweifel. Conservative neutron transport theory. Journal of Mathematical Physics, 18(2):281–284, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i2/p281_s1.

Morris:1977:PSN

[794] H. C. Morris. Prolongation structures and nonlinear evolution equations in two spatial dimensions. II. A generalized nonlinear Schrödinger equation. Journal of Mathematical Physics, 18(2):285–288, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i2/p285_s1.

Nayak:1977:ASB

[795] B. K. Nayak and R. N. Tiwari. Axisymmetric stationary Brans–Dicke vacuum fields. Journal of Mathematical Physics, 18(2):289–291, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i2/p289_s1.

Furutani:1977:BST

[796] Y. Furutani and C. Deutsch. The n-bubble series in the theory of the classical one-component plasma. Journal of Mathematical Physics, 18(2):292–300, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i2/p292_s1. See [1176].

Torres:1977:CMP

[797] P. L. Torres. A classical Markov process in nonequilibrium quantum statistical mechanics. Journal of Mathematical Physics, 18(2):301–305, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i2/p301_s1.

Weaver:1977:EDR

[798] D. L. Weaver. Exact diagonalization of relativistic Hamiltonians including a constant magnetic field. Journal of Mathematical Physics, 18(2):306–307, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i2/p306_s1.

Skagerstam:1977:SMD

[799] Bo-Sture K. Skagerstam. Stochastic mechanics and dissipative forces. Journal of Mathematical Physics, 18(2):308–311, February 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i2/p308_
s1.

Steiger:1977:EEP

[800] Arno D. Steiger. Electromagnetic eigenmode perturbations caused by deformation of a plasma core in a spherical cavity resonator. Journal of Mathematical Physics, 18(2):312–320, February 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i2/p312_s1. See [877].

Boersma:1977:ESD

[801] J. Boersma and S. W. Lee. An exact solution for diffraction of a linesource field by a half-plane. Journal of Mathematical Physics, 18(2):321– 328, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i2/p321_s1.

Kummer:1977:SARa

[802] H. Kummer, I. Absar, and A. J. Coleman. Some aspects of the Nrepresentability problem in finite dimensions. I. Operator endomorphisms which induce necessary conditions. Journal of Mathematical Physics, 18(2):329–334, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i2/p329_s1.

Kummer:1977:SARb

[803] H. Kummer and I. Absar. Some aspects of the N-representability problem in finite dimensions. II. Operator endomorphisms which induce sufficient conditions. Journal of Mathematical Physics, 18(2):335–341, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i2/p335_s1.

Epstein:1977:RSO

[804] Saul T. Epstein. Remarks on the second order error in the variational calculation of expectation values. Journal of Mathematical Physics, 18(2):342–343, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i2/p342_s1.

Epstein:1977:SDF

[805] Saul T. Epstein. Short derivation of a formula due to Lubkin. Journal of Mathematical Physics, 18(2):344, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i2/p344_s1.

Lubkin:1977:CES

[806] Elihu Lubkin. Correction of “Extension of the statistical mechanics of equilibrium to noncommutative constraints” to cover singular constraints. Journal of Mathematical Physics, 18(2):345–346, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i2/

p345_s1. See [532].

Enz:1977:NTS

[807] U. Enz. A new type of soliton with particle properties. Journal of Mathematical Physics, 18(3):347–353, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p347_s1.

Park:1977:CLA

[808] Yong Moon Park. Convergence of lattice approximations and infinite volume limit in the (λϕ^{4} − σϕ^{2} − μϕ)_{3} field theory. Journal of Mathematical Physics, 18(3):354–366, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i3/p354_s1.

Sinha:1977:EPF

[809] S. K. Sinha and Y. Singh. Equilibrium properties of fluids in the semiclassical limit. Journal of Mathematical Physics, 18(3):367–375, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i3/p367_s1.

Siewert:1977:HSA

[810] C. E. Siewert and E. E. Burniston. Half-space analysis basic to the time-dependent BGK model in the kinetic theory of gases. Journal of Mathematical Physics, 18(3):376–380, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p376_s1.

Hultgren:1977:LVP

[811] Bror O. Hultgren III and Abner Shimony. The lattice of verifiable propositions of the spin-1 system. Journal of Mathematical Physics, 18(3): 381–394, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p381_s1.

Rasmussen:1977:AAE

[812] C. H. Rasmussen and G. H. Rawitscher. Asymptotic approximations, with error estimates, of the scattering matrix for quantal Coulomb excitation by means of a nonlinear (riccati) matrix differential equation. Journal of Mathematical Physics, 18(3):395–403, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i3/p395_s1.

Mehta:1977:OEQ

[813] C. L. Mehta. Ordering of the exponential of a quadratic in boson operators. I. Single mode case. Journal of Mathematical Physics, 18(3): 404–407, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p404_s1.

Agrawal:1977:OEQ

[814] G. P. Agrawal and C. L. Mehta. Ordering of the exponential of a quadratic in boson operators. II. Multimode case. Journal of Mathematical Physics, 18(3):408–412, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i3/p408_s1.

Guler:1977:UAR

[815] Y. Güler. Unitary analytic representations of SL(3, R). Journal of Mathematical Physics, 18(3):413–417, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i3/p413_s1.

Berezdivin:1977:EIB

[816] R. Berezdivin and L. Herrera. Evolution of isometries in the Bondi formalism. Journal of Mathematical Physics, 18(3):418–423, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i3/
p418_s1.

Katzin:1977:TDDa

[817] Gerald H. Katzin, Jack Levine, and Robert N. Sane. Time-dependent dynamical symmetry mappings and associated constants of motion for classical particle systems. II. Journal of Mathematical Physics, 18(3): 424–431, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p424_s1.

Osborn:1977:ELT

[818] T. A. Osborn and D. Bollé. An extended Levinson’s theorem. Journal of Mathematical Physics, 18(3):432–440, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p432_s1.

Wodkiewicz:1977:SFI

[819] K. Wódkiewicz. Spin-frame independent variables in general relativity. Journal of Mathematical Physics, 18(3):441–444, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i3/p441_s1.

Volkel:1977:GOP

[820] A. H. Völkel. Global operator product expansions for free fields of arbitrary mass m ≥ 0. Journal of Mathematical Physics, 18(3):445– 452, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p445_s1.

Mezincescu:1977:WET

[821] L. Mezincescu. Wigner–Eckart theorem for tensor operators of graded Lie algebras. Journal of Mathematical Physics, 18(3):453–455, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i3/p453_s1.

Mayer:1977:WDC

[822] D. H. Mayer. Wightman distributions on conformal space. Journal of Mathematical Physics, 18(3):456–464, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p456_s1.

Hillion:1977:AFG

[823] Pierre Hillion. Approximating functions with a given singularity. Journal of Mathematical Physics, 18(3):465–470, March 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i3/p465_s1.

Menikoff:1977:GIF

[824] R. Menikoff and D. H. Sharp. A gauge invariant formulation of quantum electrodynamics using local currents. Journal of Mathematical Physics, 18(3):471–482, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i3/p471_s1.

Srivastava:1977:MU

[825] Sushil K. Srivastava and Mahendra P. Pathak. Meta universe. Journal of Mathematical Physics, 18(3):483–486, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p483_s1.

Hori:1977:STE

[826] Motoo Hori. Statistical theory of effective electrical, thermal, and magnetic properties of random heterogeneous materials. VII. Comparison of different approaches. Journal of Mathematical Physics, 18(3):487– 501, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p487_s1.

Houard:1977:IGL

[827] J. C. Houard. On invariance groups and Lagrangian theories. Journal of Mathematical Physics, 18(3):502–516, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p502_s1.

Ruijsenaars:1977:BTS

[828] S. N. M. Ruijsenaars. On Bogoliubov transformations for systems of relativistic charged particles. Journal of Mathematical Physics, 18(3): 517–526, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p517_s1.

Palmer:1977:SCW

[829] P. F. Palmer. The singular coupling and weak coupling limits. Journal of Mathematical Physics, 18(3):527–529, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i3/p527_s1.

Morris:1977:SSH

[830] H. C. Morris. Soliton solutions and the higher order Korteweg–de Vries equations. Journal of Mathematical Physics, 18(3):530–532, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i3/
p530_s1.

Morris:1977:PSA

[831] H. C. Morris. A prolongation structure for the AKNS system and its generalization. Journal of Mathematical Physics, 18(3):533–536, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i3/p533_s1.

Zachary:1977:EWO

[832] W. W. Zachary. Erratum: “Wave operators for multichannel scattering by long-range potentials”. Journal of Mathematical Physics, 18(3): 536, March 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i3/p536_s1. See [577].

Varatharajulu:1977:RRF

[833] V. Varatharajulu. Reciprocity relations and forward amplitude theorems for elastic waves. Journal of Mathematical Physics, 18(4):537–543, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p537_s1.

Pao:1977:PSS

[834] C. V. Pao. Positive solutions and subcriticality of energy dependent transport systems in slab geometry. Journal of Mathematical Physics, 18(4):544–550, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i4/p544_s1.

Jarvis:1977:PWA

[835] P. D. Jarvis. Partial-wave analysis for supersymmetric scattering amplitudes. Journal of Mathematical Physics, 18(4):551–563, April 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/p551_
s1.

Doria:1977:LFN

[836] Francisco Antonio Doria. A Lagrangian formulation for noninteracting high-spin fields. Journal of Mathematical Physics, 18(4):564–571, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p564_s1.

Gunther:1977:GIT

[837] N. J. Günther and P. G. L. Leach. Generalized invariants for the timedependent harmonic oscillator. Journal of Mathematical Physics, 18(4): 572–576, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p572_s1.

Dhar:1977:LEN

[838] Deepak Dhar. Lattices of effectively nonintegral dimensionality. Journal of Mathematical Physics, 18(4):577–585, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p577_s1. See [1128].

Das:1977:SGF

[839] A. Das and S. Kloster. Stationary gravitational fields of a dually charged perfect fluid. Journal of Mathematical Physics, 18(4):586–589, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p586_s1.

Baker:1977:RGS

[840] George A. Baker, Jr. and Samuel Krinsky. Renormalization group structure for translationally invariant ferromagnets. Journal of Mathematical Physics, 18(4):590–607, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i4/p590_s1.

Jordan:1977:IVO

[841] Thomas F. Jordan. Identification of the velocity operator for an irreducible unitary representation of the Poincaré group. Journal of Mathematical Physics, 18(4):608–610, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i4/p608_s1.

Wu:1977:ATM

[842] F. Y. Wu. Ashkin–Teller model as a vertex problem. Journal of Mathematical Physics, 18(4):611–613, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i4/p611_s1.

Kiehn:1977:PMQ

[843] R. M. Kiehn. Periods on manifolds, quantization, and gauge. Journal of Mathematical Physics, 18(4):614–624, April 1977. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v18/i4/p614_s1. See comments

[1065].

Winnberg:1977:SES

[844] J. O. Winnberg. Superfields as an extension of the spin representation of the orthogonal group. Journal of Mathematical Physics, 18(4):625– 628, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p625_s1.

Carey:1977:AFQ

[845] A. L. Carey, J. M. Gaffney, and C. A. Hurst. A C∗-algebra formulation of the quantization of the electromagnetic field. Journal of Mathematical Physics, 18(4):629–640, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i4/p629_s1. See [1005].

Sohn:1977:GEE

[846] R. Sohn and K. Haller. Gauge equivalence of the electrodynamics of charged bosons. Journal of Mathematical Physics, 18(4):641–647, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p641_s1.

Raduta:1977:SDC

[847] A. A. Raduta, M. Badea, and E. Badralexe. A simple derivation of a closed formula for Bogoliubov boson transformations. Journal of Mathematical Physics, 18(4):648–650, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i4/p648_s1.

Corcoran:1977:MTA

[848] C. T. Corcoran and P. W. Langhoff. Moment-theory approximations for nonnegative spectral densities. Journal of Mathematical Physics, 18(4): 651–657, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p651_s1.

Cuer:1977:SET

[849] Michel Cuer. Some examples of transparent potentials in the classical approximation. Journal of Mathematical Physics, 18(4):658–661, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p658_s1.

Henry:1977:FEF

[850] Leighton L. Henry. Free electromagnetic fields on a compact Lie group manifold. Journal of Mathematical Physics, 18(4):662–663, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/
p662_s1.

Einstein:1977:CSD

[851] S. Einstein and R. Finkelstein. A class of solutions of the Dirac equation in the Kerr–Newman space. Journal of Mathematical Physics, 18(4): 664–671, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p664_s1.

Wainwright:1977:CSI

[852] J. Wainwright. Characterization of the Szekeres inhomogeneous cosmologies as algebraically special spacetimes. Journal of Mathematical Physics, 18(4):672–675, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i4/p672_s1.

Pardee:1977:AEP

[853] W. J. Pardee. Acoustic emission and the plate Green’s function. Journal of Mathematical Physics, 18(4):676–686, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p676_s1. See [1210].

Matsuno:1977:ESL

[854] Koichiro Matsuno. Ergodic state with least mean recurrence time in Markovian kinetics. Journal of Mathematical Physics, 18(4):687–689, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p687_s1.

Chiu:1977:GES

[855] S.-C. Chiu and J. F. Ladik. Generating exactly soluble nonlinear discrete evolution equations by a generalized Wronskian technique. Journal of Mathematical Physics, 18(4):690–700, April 1977. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p690_s1.

Ladik:1977:SNN

[856] John F. Ladik and Shiu-Chu Chiu. Solutions of nonlinear network equations by the inverse scattering method. Journal of Mathematical Physics, 18(4):701–704, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i4/p701_s1.

Esposito:1977:GRM

[857] F. Paul Esposito and E. N. Glass. General relativistic magnetofluids. Journal of Mathematical Physics, 18(4):705–707, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/p705_s1.

Glass:1977:RSS

[858] E. N. Glass. Rotating steady state configurations in general relativity. Journal of Mathematical Physics, 18(4):708–711, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/p708_s1.

Khandekar:1977:PIO

[859] D. C. Khandekar and S. V. Lawande. Path integration of a onedimensional three-body problem with three-body forces. Journal of Mathematical Physics, 18(4):712–716, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p712_s1.

Znojil:1977:GMR

[860] M. Znojil. Generalized method of a resolvent operator expansion. Journal of Mathematical Physics, 18(4):717–719, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p717_s1.

Ruijsenaars:1977:CPE

[861] S. N. M. Ruijsenaars. Charged particles in external fields. I. Classical theory. Journal of Mathematical Physics, 18(4):720–737, April 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/p720_
s1.

Asgekar:1977:CMS

[862] G. G. Asgekar and T. H. Date. Collineations and motions in selfgravitating magnetofluids. Journal of Mathematical Physics, 18(4):738– 742, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p738_s1.

Basu:1977:NRS

[863] Debabrata Basu. Note on the representation spaces of O(2, 1). Journal of Mathematical Physics, 18(4):743–745, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p743_s1.

Cheng:1977:MTS

[864] Kuo-Shung Cheng. Monopole theory without string. Journal of Mathematical Physics, 18(4):746–749, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i4/p746_s1.

Barston:1977:EPL

[865] E. M. Barston. Eigenvalue problem for Lagrangian systems. VIII. Eigenvalue bounds with application to fluid and plasma flows. Journal of Mathematical Physics, 18(4):750–755, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i4/p750_s1.

Misra:1977:ZPQ

[866] B. Misra and E. C. G. Sudarshan. The Zeno’s paradox in quantum theory. Journal of Mathematical Physics, 18(4):756–763, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/p756_s1.

Kelley:1977:CEO

[867] C. T. Kelley. Convolution and H-equations for operator-valued functions with applications to neutron transport theory. Journal of Mathematical Physics, 18(4):764–769, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i4/p764_s1.

Corum:1977:RRA

[868] James F. Corum. Relativistic rotation and the anholonomic object. Journal of Mathematical Physics, 18(4):770–776, April 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i4/p770_s1.

Demiralp:1977:DSS

[869] Metin Demiralp and Erdogan Suhubi. Discrete spectrum for a system of electrically charged particles. Journal of Mathematical Physics, 18(4): 777–785, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p777_s1.

Mizrahi:1977:SEQ

[870] Maurice M. Mizrahi. On the semiclassical expansion in quantum mechanics for arbitrary Hamiltonians. Journal of Mathematical Physics, 18(4): 786–790, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p786_s1.

Gilbert:1977:SQR

[871] J. Dennis Gilbert. Second-quantized representation for a model system with composite particles. Journal of Mathematical Physics, 18(4):791– 805, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p791_s1.

Gutzwiller:1977:BST

[872] Martin C. Gutzwiller. Bernoulli sequences and trajectories in the anisotropic Kepler problem. Journal of Mathematical Physics, 18(4): 806–823, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p806_s1.

Eris:1977:FAH

[873] Ahmet Eris. Further applications of harmonic mappings of Riemannian manifolds to gravitational fields. Journal of Mathematical Physics, 18(4): 824–829, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p824_s1.

Arms:1977:LSE

[874] Judith M. Arms. Linearization stability of the Einstein–Maxwell system. Journal of Mathematical Physics, 18(4):830–833, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i4/p830_s1.

Meiman:1977:TOD

[875] N. N. Meiman. The theory of one-dimensional Schrödinger operators with a periodic potential. Journal of Mathematical Physics, 18(4):834– 848, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i4/p834_s1.

Gruber:1977:MME

[876] B. Gruber and T. S. Santhanam. Multiquark models and the embedding of groups. Journal of Mathematical Physics, 18(4):849–853, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p849_s1.

Steiger:1977:EEE

[877] Arno D. Steiger. Erratum: “Electromagnetic eigenmode perturbations caused by deformation of a plasma core in a spherical cavity resonator”. Journal of Mathematical Physics, 18(4):853, April 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i4/p853_s1. See [800].

Sudarshan:1977:CRR

[878] E. C. G. Sudarshan, M. A. K. Khalil, and W. J. Hurley. A criterion for reducibility of a relativistic wave equation. Journal of Mathematical Physics, 18(5):855–857, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i5/p855_s1.

Shahinpoor:1977:AGT

[879] M. Shahinpoor. Asymmetric gas thermodynamics. Journal of Mathematical Physics, 18(5):858–861, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i5/p858_s1. See [1275].

Pule:1977:CSI

[880] J. V. Pulè and A. Verbeure. Classical systems of infinitely many noninteracting particles. Journal of Mathematical Physics, 18(5):862–867, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i5/p862_s1.

Whitman:1977:FIS

[881] Patrick G. Whitman. On a family of interior solutions for relativistic fluid spheres with possible applications to highly collapsed stellar objects. Journal of Mathematical Physics, 18(5):868–869, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p868_s1.

Chacon:1977:GTC

[882] E. Chacón and M. Moshinsky. Group theory of the collective model of the nucleus. Journal of Mathematical Physics, 18(5):870–880, May 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p870_
s1.

Ray:1977:SIP

[883] J. R. Ray and John Charles Zimmerman. A systematic investigation of the Petrov G_{4} types. Journal of Mathematical Physics, 18(5):881–884, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i5/p881_s1.

Lauten:1977:IST

[884] W. T. Lauten III and J. R. Ray. Investigations of space–times with four-parameter groups of motions acting on null hypersurfaces. Journal of Mathematical Physics, 18(5):885–888, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p885_s1.

Krause:1977:LTS

[885] Jorge Krause. Lorentz transformations as space–time reflections. Journal of Mathematical Physics, 18(5):889–893, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p889_s1.

Rosen:1977:CYF

[886] Lon Rosen. Construction of the Yukawa_{2} field theory with a large external field. Journal of Mathematical Physics, 18(5):894–897, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/
p894_s1.

Johnson:1977:BBS

[887] Russell A. Johnson. The bundle boundary in some special cases. Journal of Mathematical Physics, 18(5):898–902, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p898_s1.

Froman:1977:PIC

[888] Nanny Fröman and Per Olof Fröman. Phase-integral calculation of quantal matrix elements without the use of wavefunctions. Journal of Mathematical Physics, 18(5):903–906, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i5/p903_s1.

Albeverio:1977:EFH

[889] Sergio Albeverio, Raphael Høegh-Krohn, and Ludwig Streit. Energy forms, Hamiltonians, and distorted Brownian paths. Journal of Mathematical Physics, 18(5):907–917, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i5/p907_s1.

Avron:1977:ICS

[890] J. E. Avron, J. Zak, A. Grossmann, and L. Gunther. Instability of the continuous spectrum. The N-band Stark ladder. Journal of Mathematical Physics, 18(5):918–921, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i5/p918_s1.

Newell:1977:SS

[891] Alan C. Newell. Synchronized solitons. Journal of Mathematical Physics, 18(5):922–926, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i5/p922_s1.

vanHaeringen:1977:MER

[892] H. van Haeringen. T matrix and effective range function for Coulomb plus rational separable potentials especially for l = 1. Journal of Mathematical Physics, 18(5):927–940, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i5/p927_s1.

vanHaeringen:1977:NBS

[893] H. van Haeringen, C. V. M. van der Mee, and R. van Wageningen. The number of bound states of the Coulomb plus Yamaguchi potential. Journal of Mathematical Physics, 18(5):941–943, May 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i5/p941_s1.

Ferreira:1977:RTV

[894] Erasmo M. Ferreira, Ricardo Merí, and Javier Sesma. Regge trajectories for a velocity dependent potential. Journal of Mathematical Physics, 18(5):944–951, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i5/p944_s1.

Kaiser:1977:PSA

[895] Gerald Kaiser. Phase-space approach to relativistic quantum mechanics. I. Coherent-state representation for massive scalar particles. Journal of Mathematical Physics, 18(5):952–959, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p952_s1.

Thorpe:1977:CIS

[896] John A. Thorpe. Curvature invariants and space–time singularities. Journal of Mathematical Physics, 18(5):960–964, May 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i5/p960_s1.

Luehr:1977:NVC

[897] C. P. Luehr, M. Rosenbaum, M. P. Ryan, Jr., and L. C. Shepley. Nonstandard vector connections given by nonstandard spinor connections. Journal of Mathematical Physics, 18(5):965–970, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p965_s1.

Kapoor:1977:MSD

[898] A. K. Kapoor. Matrix superpropagators with derivatives. Journal of Mathematical Physics, 18(5):971–983, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p971_s1.

Molinet:1977:EUP

[899] Frédéric A. Molinet. Existence, uniqueness and properties of the solutions of the Boltzmann kinetic equation for a weakly ionized gas. I. Journal of Mathematical Physics, 18(5):984–996, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p984_s1.

Robbins:1977:PUI

[900] Steven Robbins. Paraboson uniqueness for infinitely many degrees of freedom. Journal of Mathematical Physics, 18(5):997–1005, May 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p997_
s1.

Fahnline:1977:CCF

[901] Donald E. Fahnline. Covariant canonical formulation and centers of mass and motion for a relativistic two-body problem. Journal of Mathematical Physics, 18(5):1006–1011, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i5/p1006_s1.

Manoukian:1977:ABF

[902] Edward B. Manoukian. Asymptotic behavior of Feynman integrals with zero mass particles. Journal of Mathematical Physics, 18(5):1012–1014, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i5/p1012_s1.

Krips:1977:QTM

[903] H. Krips. Quantum theory and measures on Hilbert space. Journal of Mathematical Physics, 18(5):1015–1021, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p1015_s1.

Boyer:1977:HTI

[904] Charles P. Boyer and Jerzy F. Plebañski. Heavens and their integral manifolds. Journal of Mathematical Physics, 18(5):1022–1031, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/
p1022_s1.

Boyer:1977:SHJ

[905] C. P. Boyer and E. G. Kalnins. Symmetries of the Hamilton–Jacobi equation. Journal of Mathematical Physics, 18(5):1032–1045, May 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p1032_
s1.

Wolf:1977:SRF

[906] Kurt Bernardo Wolf. On self-reciprocal functions under a class of integral transforms. Journal of Mathematical Physics, 18(5):1046–1051, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i5/p1046_s1.

Sutton:1977:SCA

[907] G. R. Sutton and J. J. McCoy. Spatial coherence of acoustic signals in randomly inhomogeneous waveguides — a multiple-scatter theory. Journal of Mathematical Physics, 18(5):1052–1057, May 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i5/p1052_s1.

McCoy:1977:PFT

[908] Barry M. McCoy, Craig A. Tracy, and Tai Tsun Wu. Painlevé functions of the third kind. Journal of Mathematical Physics, 18(5):1058– 1092, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i5/p1058_s1.

Hsieh:1977:VMN

[909] Din-Yu Y. Hsieh. Variation method and nonlinear stability problems. Journal of Mathematical Physics, 18(5):1093–1104, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p1093_s1.

Naito:1977:RJM

[910] Seichi Naito. Rederivation of the Jin–Martin lower bound. Journal of Mathematical Physics, 18(5):1105–1106, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p1105_s1.

Buchholz:1977:DI

[911] Detlev Buchholz and Klaus Fredenhagen. Dilations and interaction. Journal of Mathematical Physics, 18(5):1107–1111, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i5/p1107_s1.

Zorbas:1977:RES

[912] J. Zorbas. Renormalized off-energy-shell Coulomb scattering. Journal of Mathematical Physics, 18(5):1112–1120, May 1977. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i5/p1112_s1.

Buhring:1977:SEY

[913] Wolfgang Bühring. Schrödinger equation with Yukawa potential, a differential equation with two singular points. Journal of Mathematical Physics, 18(5):1121–1136, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i5/p1121_s1.

Haberman:1977:INC

[914] Richard Haberman. An infinite number of conservation laws for coupled nonlinear evolution equations. Journal of Mathematical Physics, 18(5): 1137–1139, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i5/p1137_s1.

Schwartz:1977:ACP

[915] Charles Schwartz. Addendum: “A classical perturbation theory [J. Math. Phys. 18, 110 (1977)]”. Journal of Mathematical Physics, 18(5):1139, May 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i5/p1139_s1. See [761].

Bargmann:1977:SAF

[916] V. Bargmann and I. T. Todorov. Spaces of analytic functions on a complex cone as carriers for the symmetric tensor representations of SO(n). Journal of Mathematical Physics, 18(6):1141–1148, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1141_s1.

Barucchi:1977:CPC

[917] G. Barucchi and T. Regge. Conformal properties of a class of exactly solvable N-body problems in space dimension one. Journal of Mathematical Physics, 18(6):1149–1153, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i6/p1149_s1.

Cycon:1977:CEG

[918] H. Cycon and K.-E. Hellwig. Conditional expectations in generalized probability theory. Journal of Mathematical Physics, 18(6):1154–1161, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i6/p1154_s1.

Laskar:1977:HWS

[919] W. Laskar. Highest weights of semisimple Lie algebras. Journal of Mathematical Physics, 18(6):1162–1177, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i6/p1162_s1. See erratum [1091].

Lafon:1977:TPA

[920] J.-P. J. Lafon. On a topological problem arising in physics. Journal of Mathematical Physics, 18(6):1178–1187, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i6/p1178_s1.

Cox:1977:KTK

[921] D. Cox. Kerr’s theorem and the Kerr–Schild congruences. Journal of Mathematical Physics, 18(6):1188–1190, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i6/p1188_s1.

McQuistan:1977:CNN

[922] Richmond B. McQuistan. Composite next nearest neighbor degeneracy. Journal of Mathematical Physics, 18(6):1191–1199, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1191_s1.

McQuistan:1977:ENN

[923] R. B. McQuistan. Exact nearest neighbor statistics for λ-bell particles on a one-dimensional lattice. Journal of Mathematical Physics, 18(6):1200– 1205, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i6/p1200_s1.

Hansen:1977:KIR

[924] R. O. Hansen and Jeffrey Winicour. Killing inequalities for relativistically rotating fluids. II. Journal of Mathematical Physics, 18(6):1206– 1209, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i6/p1206_s1.

Quesne:1977:ISC

[925] C. Quesne. Inequivalent sets of commuting missing label operators for SU(4) ⊆ SU(2) × SU(2). Journal of Mathematical Physics, 18(6):1210– 1211, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i6/p1210_s1.

Olver:1977:EEP

[926] Peter J. Olver. Evolution equations possessing infinitely many symmetries. Journal of Mathematical Physics, 18(6):1212–1215, June 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1212_
s1.

DeFacio:1977:SUQ

[927] B. De Facio and C. L. Hammer. Symmetries of ultralocal quantum field theories. Journal of Mathematical Physics, 18(6):1216–1223, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i6/p1216_s1.

Stillinger:1977:ABS

[928] Frank H. Stillinger. Axiomatic basis for spaces with noninteger dimension. Journal of Mathematical Physics, 18(6):1224–1234, June 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1224_
s1.

Lehr:1977:SDK

[929] William J. Lehr and James L. Park. A stochastic derivation of the Klein–Gordon equation. Journal of Mathematical Physics, 18(6):1235– 1240, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i6/p1235_s1.

Crawford:1977:VMC

[930] Oakley H. Crawford. Variational methods for chemical and nuclear reactions. Journal of Mathematical Physics, 18(6):1241–1255, June 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1241_
s1.

Macki:1977:ITD

[931] Jack W. Macki. Invariants for the time-dependent harmonic oscillator. Journal of Mathematical Physics, 18(6):1256–1258, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1256_s1.

Vazquez:1977:LSNa

[932] Luis Vazquez. Localized solutions of a nonlinear electromagnetic field. Journal of Mathematical Physics, 18(6):1259–1263, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1259_s1.

Woo:1977:PCT

[933] G. Woo. Pseudoparticle configurations in two-dimensional ferromagnets. Journal of Mathematical Physics, 18(6):1264–1266, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1264_s1.

Katzin:1977:TDDb

[934] Gerald H. Katzin and Jack Levine. Time-dependent dynamical symmetries and constants of motion III. Time-dependent harmonic oscillator. Journal of Mathematical Physics, 18(6):1267–1274, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1267_s1.

Baltes:1977:EDO

[935] H. P. Baltes and B. Steinle. Eigenfrequency density oscillations and Walfisz lattice sums. Journal of Mathematical Physics, 18(6):1275–1276, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i6/p1275_s1.

McKean:1977:PBP

[936] H. P. McKean. −Δ plus a bad potential. Journal of Mathematical Physics, 18(6):1277–1279, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i6/p1277_s1.

Vishveshwara:1977:RRD

[937] C. V. Vishveshwara and J. Winicour. Relativistically rotating dust cylinders. Journal of Mathematical Physics, 18(6):1280–1284, June 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1280_
s1.

Olsson:1977:IDE

[938] P. O. M. Olsson. On the integration of the differential equations of fiveparametric double-hypergeometric functions of second order. Journal of Mathematical Physics, 18(6):1285–1294, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i6/p1285_s1.

Weaver:1977:ASS

[939] D. L. Weaver. Algebraic solutions to some spin-1 problems. Journal of Mathematical Physics, 18(6):1295–1296, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i6/p1295_s1.

Deutsch:1977:EOC

[940] C. Deutsch. Erratum: “One-component plasma in 2 + ϵ dimensions”. Journal of Mathematical Physics, 18(6):1297, June 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i6/p1297_

s1. See [632].

Forte:1977:REF

[941] B. Forte and C. C. A. Sastri. Representation of the entropy functional for a grand canonical ensemble in classical statistical mechanics. Journal of Mathematical Physics, 18(7):1299–1302, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i7/p1299_s1.

Eris:1977:SAS

[942] A. Eris and M. Gürses. Stationary axially symmetric solutions of Einstein–Maxwell massless scalar field equations. Journal of Mathematical Physics, 18(7):1303–1304, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i7/p1303_s1.

Norris:1977:FST

[943] L. K. Norris, L. H. Green, and W. R. Davis. Fluid space–times including electromagnetic fields admitting symmetry mappings belonging to the family of contracted Ricci collineations. Journal of Mathematical Physics, 18(7):1305–1311, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i7/p1305_s1.

Rund:1977:EMT

[944] Hanno Rund. Energy-momentum tensors in the theory of electromagnetic fields admitting electric and magnetic charge distributions. Journal of Mathematical Physics, 18(7):1312–1315, July 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1312_s1.

Keizer:1977:MED

[945] Joel Keizer. On the macroscopic equivalence of descriptions of fluctuations for chemical reactions. Journal of Mathematical Physics, 18(7): 1316–1321, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1316_s1.

DeCanniere:1977:DPO

[946] J. De Cannière, J. Pulè, P. Vanheuverzwijn, and A. Verbeure. Dispersive properties and observables at infinity for classical KMS systems. Journal of Mathematical Physics, 18(7):1322–1326, July 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1322_s1.

Gzyl:1977:QSS

[947] Henryk Gzyl. Quantum system subject to random pulses. Journal of Mathematical Physics, 18(7):1327–1329, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v18/i7/p1327_s1. See comments

[1451].

Spero:1977:ASG

[948] Alan Spero and Ralph Baierlein. Approximate symmetry groups of inhomogeneous metrics. Journal of Mathematical Physics, 18(7):1330–1340, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i7/p1330_s1.

Vazquez:1977:LSNb

[949] Luis Vazquez. Localized solutions of a nonlinear scalar field with a scalar potential. Journal of Mathematical Physics, 18(7):1341–1342, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i7/p1341_s1.

Vazquez:1977:SLS

[950] Luis Vazquez. Stationary localized solutions in nonlinear classical fields. Journal of Mathematical Physics, 18(7):1343–1347, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1343_s1.

Newton:1977:NPG

[951] Roger G. Newton. Noncentral potentials: The generalized Levinson theorem and the structure of the spectrum. Journal of Mathematical Physics, 18(7):1348–1357, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i7/p1348_s1.

Drechsler:1977:GCF

[952] W. Drechsler. Group contraction in a fiber bundle with Cartan connection. Journal of Mathematical Physics, 18(7):1358–1366, July 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1358_
s1.

Ahner:1977:CIP

[953] Henry F. Ahner and Allan E. Moose. Covariant inverse problem of the calculus of variations. Journal of Mathematical Physics, 18(7):1367– 1373, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1367_s1. See corrigendum [1414].

Collins:1977:SSS

[954] C. B. Collins. Static stars : Some mathematical curiosities. Journal of Mathematical Physics, 18(7):1374–1377, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i7/p1374_s1.

Brans:1977:CIC

[955] Carl H. Brans. Complete integrability conditions of the Einstein–Petrov equations, type I. Journal of Mathematical Physics, 18(7):1378–1381, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i7/p1378_s1.

Daumens:1977:ORG

[956] M. Daumens and M. Perroud. Orbits of the rotation group on the density matrices of spin-1 particles. Journal of Mathematical Physics, 18(7): 1382–1386, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1382_s1.

Nwachuku:1977:GFE

[957] C. O. Nwachuku and M. A. Rashid. Generating functions for the eigenvalues of the Casimir operators of the orthogonal and symplectic groups. Journal of Mathematical Physics, 18(7):1387–1391, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1387_s1.

Gates:1977:HDD

[958] D. J. Gates. Hydrostatic density distribution of fluids in steep field gradients and near critical points. Journal of Mathematical Physics, 18(7): 1392–1398, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1392_s1.

Malament:1977:CCT

[959] David B. Malament. The class of continuous timelike curves determines the topology of spacetime. Journal of Mathematical Physics, 18(7):1399– 1404, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1399_s1.

Magri:1977:ETN

[960] Franco Magri. Equivalence transformations for nonlinear evolution equations. Journal of Mathematical Physics, 18(7):1405–1411, July 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1405_
s1.

Chen:1977:IAS

[961] F. C. Chen and K. Young. Inclusions of arbitrary shape in an elastic medium. Journal of Mathematical Physics, 18(7):1412–1416, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i7/p1412_s1.

Eichinger:1977:GCS

[962] B. E. Eichinger. Geometrization of configuration space. Journal of Mathematical Physics, 18(7):1417–1428, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i7/p1417_s1.

Actor:1977:PTD

[963] Alfred Actor. Polarization theorem for diffractive excitation at large momentum transfer. Journal of Mathematical Physics, 18(7):1429–1433, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i7/p1429_s1.

Bohm:1977:MEV

[964] A. Böhm and R. B. Teese. Matrix elements of vector and axial vector operators in an algebraic formulation of baryon decays. Journal of Mathematical Physics, 18(7):1434–1448, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i7/p1434_s1.

Patera:1977:SRT

[965] J. Patera and P. Winternitz. Subalgebras of real three- and fourdimensional Lie algebras. Journal of Mathematical Physics, 18(7):1449– 1455, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1449_s1.

Jaric:1977:NAM

[966] Marko V. Jarić and Joseph L. Birman. New algorithms for the Molien function. Journal of Mathematical Physics, 18(7):1456–1458, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/
p1456_s1.

Jaric:1977:CMG

[967] Marko V. Jarić and Joseph L. Birman. Calculation of the Molien generating function for invariants of space groups. Journal of Mathematical Physics, 18(7):1459–1465, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i7/p1459_s1. See [1066].

Singh:1977:ARR

[968] S. R. Singh. On approximating the resolvent of a rotated Hamiltonian in the scattering region. Journal of Mathematical Physics, 18(7):1466– 1469, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1466_s1.

Singh:1977:EBS

[969] S. R. Singh and G. Turchetti. Error bounds on some approximate solutions of the Fredholm equations and applications to potential scattering and bound state problems. Journal of Mathematical Physics, 18(7): 1470–1475, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1470_s1.

Gervois:1977:BLT

[970] A. Gervois and M. L. Mehta. Broken linear transformations. Journal of Mathematical Physics, 18(7):1476–1479, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i7/p1476_s1.

Wanders:1977:BLG

[971] G. Wanders and E. Morenzoni. Bounds for ladder graph scattering amplitudes on the boundary of their cut. Journal of Mathematical Physics, 18(7):1480–1484, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i7/p1480_s1.

Motoyama:1977:KDP

[972] Akiko Motoyama and Haruo Hosoya. King and domino polynomials for polyomino graphs. Journal of Mathematical Physics, 18(7):1485– 1490, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i7/p1485_s1.

Lovelock:1977:BFT

[973] David Lovelock. Bivector field theories, divergence-free vectors and the Einstein–Maxwell field equations. Journal of Mathematical Physics, 18(7):1491–1498, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i7/p1491_s1.

Truman:1977:CAN

[974] Aubrey Truman. The classical action in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 18(7):1499–1509, July 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1499_
s1.

Bongaarts:1977:MEA

[975] P. J. M. Bongaarts. Maxwell’s equations in axiomatic quantum field theory. I. Field tensor and potentials. Journal of Mathematical Physics, 18(7):1510–1516, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i7/p1510_s1.

Hacyan:1977:KSP

[976] S. Hacyan and J. F. Plebański. D(1, 0) Killing structures and ϵ potentials. Journal of Mathematical Physics, 18(7):1517–1519, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i7/p1517_s1.

Patera:1977:EHI

[977] J. Patera, R. T. Sharp, and P. Winternitz. Erratum: “Higher indices of group representations”. Journal of Mathematical Physics, 18(7):1519, July 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i7/p1519_s1. See [709].

Miklavc:1977:QAR

[978] Adolf Miklavc. Quantum axiomatics, representation theorem, and communicability of observations in physics. Journal of Mathematical Physics, 18(8):1521–1525, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i8/p1521_s1.

Miklavc:1977:FQM

[979] Adolf Miklavc. On the field of the quantum mechanical Hilbert space. Journal of Mathematical Physics, 18(8):1526–1528, August 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1526_
s1.

Kinnersley:1977:SSEa

[980] William Kinnersley. Symmetries of the stationary Einstein–Maxwell field equations. I. Journal of Mathematical Physics, 18(8):1529–1537, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i8/p1529_s1.

Kinnersley:1977:SSEb

[981] William Kinnersley and D. M. Chitre. Symmetries of the stationary Einstein–Maxwell field equations. II. Journal of Mathematical Physics, 18(8):1538–1542, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i8/p1538_s1.

Gordon:1977:SSC

[982] Paul Gordon. Stability of solutions of the compressible Navier–Stokes equations. Journal of Mathematical Physics, 18(8):1543–1552, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i8/p1543_s1.

Carey:1977:FMQ

[983] A. L. Carey and C. A. Hurst. The Fermi method of quantizing the electromagnetic field as a model for quantum field theory. Journal of Mathematical Physics, 18(8):1553–1561, August 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1553_s1.

Weber:1977:WSP

[984] T. A. Weber and C. L. Hammer. Wavepacket scattering in potential theory. Journal of Mathematical Physics, 18(8):1562–1567, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/
p1562_s1.

Tipler:1977:CSS

[985] Frank J. Tipler. Causally symmetric spacetimes. Journal of Mathematical Physics, 18(8):1568–1573, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i8/p1568_s1.

Ling:1977:ABI

[986] Rina Ling. Asymptotic behavior of integral and integrodifferential equations. Journal of Mathematical Physics, 18(8):1574–1576, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/
p1574_s1.

Alonso:1977:GTF

[987] L. Martinez Alonso. Group theoretical foundations of classical and quantum mechanics. I. Observables associated with Lie algebras. Journal of Mathematical Physics, 18(8):1577–1581, August 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1577_s1.

Newton:1977:NIG

[988] Roger G. Newton. Nonlocal interactions: The generalized Levinson theorem and the structure of the spectrum. Journal of Mathematical Physics, 18(8):1582–1588, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i8/p1582_s1.

Kuchar:1977:GTS

[989] Karel Kuchař. Geometrodynamics with tensor sources. IV. Journal of Mathematical Physics, 18(8):1589–1597, August 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1589_s1.

Dollard:1977:PIS

[990] John D. Dollard and Charles N. Friedman. Product integrals and the Schrödinger equation. Journal of Mathematical Physics, 18(8):1598– 1607, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i8/p1598_s1.

Leach:1977:TTD

[991] P. G. L. Leach. On the theory of time-dependent linear canonical transformations as applied to Hamiltonians of the harmonic oscillator type. Journal of Mathematical Physics, 18(8):1608–1611, August 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1608_
s1.

delaPena-Auerbach:1977:DQM

[992] L. de la Peña-Auerbach and A. M. Cetto. Derivation of quantum mechanics from stochastic electrodynamics. Journal of Mathematical Physics, 18(8):1612–1622, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i8/p1612_s1.

Martin:1977:IZE

[993] André Martin and Pierre C. Sabatier. Impedance, zero energy wavefunction, and bound states. Journal of Mathematical Physics, 18(8):1623– 1626, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i8/p1623_s1.

Schorr:1977:VSB

[994] Amir Schorr. Varieties of symmetry breaking in a class of gauge theories. Journal of Mathematical Physics, 18(8):1627–1639, August 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1627_
s1.

Schorr:1977:SSB

[995] Amir Schorr. Spontaneous symmetry breaking of U_{1}(M) ⊗ U_{2}(M) ⊗ U_{3}(M) with colored Higgs particles. Journal of Mathematical Physics, 18(8):1640–1649, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i8/p1640_s1.

Kac:1977:CFS

[996] Mark Kac and Colin J. Thompson. Correlation functions in the spherical and mean spherical models. Journal of Mathematical Physics, 18(8):1650–1653, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i8/p1650_s1.

Lamb:1977:SMS

[997] G. L. Lamb, Jr. Solitons on moving space curves. Journal of Mathematical Physics, 18(8):1654–1661, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i8/p1654_s1.

Finley:1977:SSE

[998] J. D. Finley III and J. F. Plebański. Spinorial structures and electromagnetic hyperheavens. Journal of Mathematical Physics, 18(8):1662– 1667, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i8/p1662_s1.

Szafron:1977:CIP

[999] D. A. Szafron and J. Wainwright. A class of inhomogeneous perfect fluid cosmologies. Journal of Mathematical Physics, 18(8):1668–1672, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i8/p1668_s1.

Szafron:1977:ICN

[1000] D. A. Szafron. Inhomogeneous cosmologies: New exact solutions and their evolution. Journal of Mathematical Physics, 18(8):1673–1677, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i8/p1673_s1.

Schindler:1977:DTP

[1001] Susan Schindler and R. Mirman. The decomposition of the tensor product of representations of the symmetric group. Journal of Mathematical Physics, 18(8):1678–1696, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i8/p1678_s1.

Schindler:1977:CGC

[1002] Susan Schindler and R. Mirman. The Clebsch–Gordan coefficients of S_{n}. Journal of Mathematical Physics, 18(8):1697–1704, August 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1697_

s1. See addendum [1556].

Humi:1977:ISC

[1003] Mayer Humi. Invariant solutions for a class of diffusion equations. Journal of Mathematical Physics, 18(8):1705–1708, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i8/p1705_s1.

Baur:1977:EZG

[1004] M. E. Baur and R. M. Sinder. Erratum: “Zeros of the grand canonical partition function: Generalization of a result of Lee and Yang”. Journal of Mathematical Physics, 18(8):1708, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i8/p1708_s1. See [733].

Carey:1977:ECA

[1005] A. L. Carey, J. M. Gaffney, and C. A. Hurst. Erratum: “A C∗-algebra formulation of the quantization of the electromagnetic field”. Journal of Mathematical Physics, 18(8):1708, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i8/p1708_s2. See [845].

Klauder:1977:AFO

[1006] John R. Klauder. Affine fields and the operator formulation of augmented scalar fields. Journal of Mathematical Physics, 18(9):1711–1719, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1711_s1.

Rodrigues:1977:GFG

[1007] Paulo R. Rodrigues. On generating forms of K-generalized Lagrangian and Hamiltonian systems. Journal of Mathematical Physics, 18(9):1720– 1723, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1720_s1.

Patil:1977:DSE

[1008] K. R. Patil and N. K. Thakare. On dual series equations involving Konhauser biorthogonal polynomials. Journal of Mathematical Physics, 18 (9):1724–1726, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i9/p1724_s1.

Dadhich:1977:ECG

[1009] Naresh Dadhich and P. P. Kale. Equatorial circular geodesics in the Kerr– Newman geometry. Journal of Mathematical Physics, 18(9):1727–1728, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1727_s1.

Grewe:1977:KSEa

[1010] N. Grewe and W. Klein. The Kirkwood–Salsburg equations for a bounded stable Kac potential. I. General theory and asymptotic solutions. Journal of Mathematical Physics, 18(9):1729–1734, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1729_s1.

Grewe:1977:KSEb

[1011] N. Grewe and W. Klein. The Kirkwood–Salsburg equations for a bounded stable Kac potential. II. Instability and phase transitions. Journal of Mathematical Physics, 18(9):1735–1740, September 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1735_s1.

Kalnins:1977:LTWc

[1012] E. G. Kalnins and W. Miller, Jr. Lie theory and the wave equation in space–time. 5. R-separable solutions of the wave equation ψ_{tt} −Δ_{3}ψ = 0. Journal of Mathematical Physics, 18(9):1741–1751, September 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1741_
s1.

Lee:1977:CCI

[1013] L. L. Lee. Continuum calculus. II. The heterogeneous continuous functional differentiation applied to the Feynman path integral. Journal of Mathematical Physics, 18(9):1752–1755, September 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1752_s1.

Woo:1977:TEH

[1014] G. Woo. The topology of Euclidean Higgs fields. Journal of Mathematical Physics, 18(9):1756–1758, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i9/p1756_s1.

Yoneya:1977:TEY

[1015] Tamiaki Yoneya. Topology of Euclidean Yang–Mills fields: instantons and monopoles. Journal of Mathematical Physics, 18(9):1759–1762, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1759_s1.

Gonzalez-Gascon:1977:NSS

[1016] F. Gonzalez-Gascon. Notes on the symmetries of systems of differential equations. Journal of Mathematical Physics, 18(9):1763–1767, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i9/p1763_s1.

Wong:1977:BME

[1017] M. K. F. Wong and Hsin-Yang Yeh. Boost matrix elements and Clebsch– Gordan coefficients of the homogeneous Lorentz group. Journal of Mathematical Physics, 18(9):1768–1781, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i9/p1768_s1.

Weir:1977:CKT

[1018] G. J. Weir. Conformal Killing tensors in reducible spaces. Journal of Mathematical Physics, 18(9):1782–1787, September 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1782_s1.

Malin:1977:MEE

[1019] S. Malin. Maxwell’s equations in an expanding universe. Journal of Mathematical Physics, 18(9):1788–1790, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i9/p1788_s1.

Maduemezia:1977:FBS

[1020] Awele Maduemezia. Fibre-bundle structure of thermodynamic states. Journal of Mathematical Physics, 18(9):1791–1793, September 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1791_
s1.

Glassey:1977:BSC

[1021] R. T. Glassey. On the blowing up of solutions to the Cauchy problem for nonlinear Schrödinger equations. Journal of Mathematical Physics, 18(9):1794–1797, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i9/p1794_s1.

Gowdy:1977:ICS

[1022] Robert H. Gowdy. Instantaneous Cauchy surfaces, topology change, and exploding black holes. Journal of Mathematical Physics, 18(9):1798– 1801, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1798_s1.

Richardson:1977:PLL

[1023] R. W. Richardson. Pairing in the limit of a large number of particles. Journal of Mathematical Physics, 18(9):1802–1811, September 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1802_
s1.

Lerner:1977:TIR

[1024] David E. Lerner. Twistors and induced representations of SU(2, 2). Journal of Mathematical Physics, 18(9):1812–1817, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i9/p1812_s1.

Barker:1977:PNT

[1025] B. M. Barker and R. F. O’Connell. Post-Newtonian two-body and n-body problems with electric charge in general relativity. Journal of Mathematical Physics, 18(9):1818–1824, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i9/p1818_s1. See [1329].

Ziff:1977:BDF

[1026] Robert M. Ziff. On the bulk distribution functions and fluctuation theorems. Journal of Mathematical Physics, 18(9):1825–1831, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i9/p1825_s1.

Chen:1977:MSI

[1027] Etang Chen. Markov systems of imprimitivity. Journal of Mathematical Physics, 18(9):1832–1837, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i9/p1832_s1.

Phares:1977:GTL

[1028] Alain J. Phares. A general theory of linear finite-difference equations with a few applications. Journal of Mathematical Physics, 18(9):1838–1848, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1838_s1.

Fackerell:1977:SWA

[1029] Edward D. Fackerell and Robert G. Crossman. Spin-weighted angular spheroidal functions. Journal of Mathematical Physics, 18(9):1849–1854, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i9/p1849_s1.

Cornille:1977:GIE

[1030] Henri Cornille. Generalization of the inversion equations and application to nonlinear partial differential equations. I. Journal of Mathematical Physics, 18(9):1855–1869, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i9/p1855_s1.

Bincer:1977:LO

[1031] Adam M. Bincer. U(n) lowering operators. Journal of Mathematical Physics, 18(9):1870, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i9/p1870_s1.

Luban:1977:ESL

[1032] Marshall Luban and Batya Nudler-Blum. Exact solution for the linear response of a hydrogenic atom to an external electromagnetic field. I. Frequencies below photoelectric threshold. Journal of Mathematical Physics, 18(9):1871–1881, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i9/p1871_s1.

Luban:1977:CSV

[1033] Marshall Luban and Moshe Deutsch. Classes of solvable Volterra integral equations. Journal of Mathematical Physics, 18(9):1882, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i9/p1882_s1.

Lohe:1977:SPS

[1034] M. A. Lohe, L. C. Biedenharn, and J. D. Louck. Structural properties of the self-conjugate SU(3) tensor operators. Journal of Mathematical Physics, 18(9):1883–1891, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i9/p1883_s1.

Hai:1977:RCQ

[1035] Yao Te Hai. A remark on the construction of quantum fields from Markov fields. Journal of Mathematical Physics, 18(9):1892–1897, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i9/p1892_s1.

Yao:1977:RCQ

[1036] T. H. Yao. A remark on the construction of quantum fields from Markov fields. Journal of Mathematical Physics, 18(9):1892–1897, September 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ray:1977:SMF

[1037] Dipankar Ray. Scalar meson fields in a conformally flat space. Journal of Mathematical Physics, 18(10):1899–1901, October 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i10/p1899_s1.

Leach:1977:IWS

[1038] P. G. L. Leach. Invariants and wavefunctions for some time-dependent harmonic oscillator-type Hamiltonians. Journal of Mathematical Physics, 18(10):1902–1907, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i10/p1902_s1.

Lavine:1977:GSP

[1039] Richard Lavine and Michael O’Carroll. Ground state properties and lower bounds for energy levels of a particle in a uniform magnetic field and external potential. Journal of Mathematical Physics, 18(10):1908– 1912, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i10/p1908_s1.

Challifour:1977:SFP

[1040] John L. Challifour and Steven P. Slinker. On Schwinger functionalspositive extensions, moment problems, and representations. Journal of Mathematical Physics, 18(10):1913–1917, October 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i10/p1913_s1.

Coll:1977:EEK

[1041] Bartolomé Coll. On the evolution equations for Killing fields. Journal of Mathematical Physics, 18(10):1918–1922, October 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i10/p1918_s1.

Lloyd-Evans:1977:GAS

[1042] D. J. R. Lloyd-Evans. Geometric aspects of supergauge theory. Journal of Mathematical Physics, 18(10):1923–1927, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/p1923_s1.

Ramanathan:1977:CNP

[1043] G. V. Ramanathan. Correlations near phase transition in a simple fluid. Journal of Mathematical Physics, 18(10):1928–1934, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/
p1928_s1.

Wu:1977:SAC

[1044] Alfred C. T. Wu. Structure of the Azzarelli–Collas representation for the scattering amplitude and generalization to the Rice representation and the Euler–Pochhammer representation. Journal of Mathematical Physics, 18(10):1935–1937, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i10/p1935_s1.

Major:1977:QMR

[1045] M. E. Major. The quantum mechanical representations of the anisotropic harmonic oscillator group. Journal of Mathematical Physics, 18(10): 1938–1943, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i10/p1938_s1.

Major:1977:ADA

[1046] M. E. Major. On the accidental degeneracy of the anisotropic harmonic oscillator. I. Journal of Mathematical Physics, 18(10):1944–1951, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i10/p1944_s1.

Major:1977:NGD

[1047] M. E. Major. A noninvariance group for the n-dimensional isotropic harmonic oscillator. Journal of Mathematical Physics, 18(10):1952– 1957, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i10/p1952_s1.

Major:1977:ADD

[1048] M. E. Major. On the accidental degeneracy of the n-dimensional anisotropic harmonic oscillator. II. Journal of Mathematical Physics, 18(10):1958–1965, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i10/p1958_s1.

Stelzer:1977:FQE

[1049] M. Stelzer, J. Mühlstein, and V. Ernst. A finite quantum electrodynamical base for many photon collective phenomena. Journal of Mathematical Physics, 18(10):1966–1977, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i10/p1966_s1.

Kohler:1977:MSS

[1050] Werner E. Kohler. A multiple-scales space–time analysis of a randomly perturbed one-dimensional wave equation. Journal of Mathematical Physics, 18(10):1978–1984, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i10/p1978_s1.

Greenberg:1977:IFL

[1051] William Greenberg. The interaction function and lattice duals. Journal of Mathematical Physics, 18(10):1985–1986, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/p1985_s1.

Tilgner:1977:ELG

[1052] Hans Tilgner. Extensions of Lie-graded algebras. Journal of Mathematical Physics, 18(10):1987–1991, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i10/p1987_s1.

Kannenberg:1977:IIN

[1053] L. Kannenberg. Invariant integrals in nonmetric supersymmetry spaces. Journal of Mathematical Physics, 18(10):1992–1995, October 1977.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/
p1992_s1.

Chui:1977:EOD

[1054] S. T. Chui and H. L. Frisch. The equivalence of a one-dimensional turbulence problem and the one-dimensional Coulomb gas. Journal of Mathematical Physics, 18(10):1996–1997, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i10/p1996_s1.

Gibbs:1977:NTP

[1055] Alan G. Gibbs and Edward W. Larsen. Neutron transport in plane geometry with general anisotropic, energy-dependent scattering. Journal of Mathematical Physics, 18(10):1998–2007, October 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i10/p1998_s1.

Keener:1977:GFL

[1056] J. P. Keener and D. W. McLaughlin. A Green’s function for a linear equation associated with solitons. Journal of Mathematical Physics, 18(10):2008–2013, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i10/p2008_s1.

Ehlers:1977:PSM

[1057] Jürgen Ehlers and Egon Köhler. Path structures on manifolds. Journal of Mathematical Physics, 18(10):2014–2018, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/p2014_s1.

Ling:1977:UVS

[1058] Rina Ling. Uniformly valid solutions of Volterra integral equations. Journal of Mathematical Physics, 18(10):2019–2025, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/p2019_s1.

Hamilton:1977:CDE

[1059] M. J. Hamilton and A. Das. On the combined Dirac–Einstein–Maxwell field equations. Journal of Mathematical Physics, 18(10):2026–2030, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i10/p2026_s1.

Scarfone:1977:QAB

[1060] Leonard M. Scarfone and James A. Burns. The N-quantum approximation in the Bronzan–Lee extended-source static model. Journal of Mathematical Physics, 18(10):2031–2043, October 1977. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v18/i10/p2031_s1.

Case:1977:SRW

[1061] K. M. Case and S. C. Chiu. Some remarks on the Wronskian technique and the inverse scattering transform. Journal of Mathematical Physics, 18(10):2044–2052, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i10/p2044_s1.

Bonora:1977:SPC

[1062] L. Bonora and P. Pasti. Spectral properties of completely monotonic operators. Journal of Mathematical Physics, 18(10):2053–2064, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i10/p2053_s1.

Dirl:1977:IPR

[1063] R. Dirl. Induced projective representations. Journal of Mathematical Physics, 18(10):2065–2080, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i10/p2065_s1.

Hongoh:1977:CSA

[1064] M. Hongoh. Coherent states associated with the continuous spectrum of noncompact groups. Journal of Mathematical Physics, 18(10):2081– 2084, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i10/p2081_s1.

Post:1977:CPM

[1065] E. J. Post. Comments on “Periods on manifolds, quantization, and gauge” by R. M. Kiehn [J. Math. Phys. 18, 614 (1977)]. Journal of Mathematical Physics, 18(10):2084, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i10/p2084_s1. See [843].

Jaric:1977:ECM

[1066] Marko V. Jarić and Joseph L. Birman. Erratum: “Calculation of the Molien generating function for invariants of space groups”. Journal of Mathematical Physics, 18(10):2085, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i10/p2085_s1. See [967].

Fabec:1977:FPS

[1067] Raymond C. Fabec. Free particle systems and their dynamics in de Sitter space. Journal of Mathematical Physics, 18(11):2087–2091, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2087_s1.

Srivastava:1977:TDU

[1068] Sushil K. Srivastava. A tachyon dust universe. Journal of Mathematical Physics, 18(11):2092–2096, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i11/p2092_s1.

Talukdar:1977:MWS

[1069] B. Talukdar and M. N. Sinha Ray. K-matrix for the Woods–Saxon potential. Journal of Mathematical Physics, 18(11):2097–2098, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2097_s1.

Pulvirenti:1977:SEK

[1070] M. Pulvirenti. Stability, equilibrium and KMS for an infinite classical system. Journal of Mathematical Physics, 18(11):2099–2103, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2099_s1.

Ekhaguere:1977:NMP

[1071] G. O. S. Ekhaguere. On notions of Markov property. Journal of Mathematical Physics, 18(11):2104–2107, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i11/p2104_s1.

Berryman:1977:ESP

[1072] James G. Berryman. Evolution of a stable profile for a class of nonlinear diffusion equations with fixed boundaries. Journal of Mathematical Physics, 18(11):2108–2115, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i11/p2108_s1.

Collins:1977:GSK

[1073] C. B. Collins. Global structure of the “Kantowski–Sachs” cosmological models. Journal of Mathematical Physics, 18(11):2116–2124, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2116_s1.

Goldoni:1977:NCV

[1074] Riccardo Goldoni. A necessary condition for the validity of Huygens’ principle on a curved space–time. Journal of Mathematical Physics, 18 (11):2125–2128, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i11/p2125_s1.

Fulinski:1977:PSP

[1075] A. Fuliński and C. Jedrzejek. Perturbation solution of the Percus– Yevick equation for the square-mound potential. Journal of Mathematical Physics, 18(11):2129–2132, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i11/p2129_s1.

Haba:1977:FEE

[1076] Z. Haba. Functional equations for extended hadrons. Journal of Mathematical Physics, 18(11):2133–2137, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i11/p2133_s1.

Srinivas:1977:QCP

[1077] M. D. Srinivas. Quantum counting processes. Journal of Mathematical Physics, 18(11):2138–2145, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i11/p2138_s1.

Nackoney:1977:GIB

[1078] Raymond W. Nackoney. The gravitational influence of a beam of light of variable flux. Journal of Mathematical Physics, 18(11):2146–2152, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i11/p2146_s1.

Moreno:1977:SPN

[1079] Carlos Moreno. Spaces of positive and negative frequency solutions of field equations in curved space–times. I. The Klein–Gordon equation in stationary space–times. Journal of Mathematical Physics, 18(11):2153– 2161, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i11/p2153_s1.

Napiorkowski:1977:NRA

[1080] Marek Napiórkowski. A note on the range of the applicability of the Ornstein–Zernike theory in the van der Waals model. Journal of Mathematical Physics, 18(11):2162–2165, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i11/p2162_s1.

Kraus:1977:GID

[1081] K. Kraus, L. Polley, and G. Reents. Generators of infinite direct products of unitary groups. Journal of Mathematical Physics, 18(11):2166– 2171, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i11/p2166_s1.

Gross:1977:STP

[1082] Eugene P. Gross. Sequences of transformations for particle-field Hamiltonians. Journal of Mathematical Physics, 18(11):2172–2176, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2172_s1.

Brihaye:1977:MMG

[1083] Y. Brihaye and J. Nuyts. Magnetic monopoles in SU(4) gauge theories. Journal of Mathematical Physics, 18(11):2177–2190, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2177_s1.

Kloster:1977:SGF

[1084] S. Kloster and A. Das. Stationary gravitational fields of a charged perfect fluid. Journal of Mathematical Physics, 18(11):2191–2193, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2191_s1.

Raju:1977:LMP

[1085] Cvavb. Chandra Raju. Lee model with 3V particles. Journal of Mathematical Physics, 18(11):2194–2203, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i11/p2194_s1.

Dixon:1977:FBI

[1086] Geoffrey Dixon. Fermi–Bose and internal symmetries with universal Clifford algebras. Journal of Mathematical Physics, 18(11):2204–2206, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i11/p2204_s1.

Corones:1977:LGF

[1087] James Corones, B. L. Markovski, and V. A. Rizov. A Lie group framework for soliton equations. I. Path independent case. Journal of Mathematical Physics, 18(11):2207–2213, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i11/p2207_s1.

Antippa:1977:DPA

[1088] Adel F. Antippa. Discrete path approach to linear recursion relations. Journal of Mathematical Physics, 18(11):2214–2222, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i11/
p2214_s1.

Coz:1977:TKD

[1089] Marcel Coz and Pierre Rochus. The translation kernel in the ndimensional scattering problem. Journal of Mathematical Physics, 18 (11):2223–2231, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i11/p2223_s1.

Coz:1977:TKV

[1090] M. Coz and P. Rochus. Translation kernels for velocity dependent interactions. Journal of Mathematical Physics, 18(11):2232–2240, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i11/p2232_s1.

Laskar:1977:EHW

[1091] W. Laskar. Erratum: “Higher [i.e., highest] weights of semisimple Lie algebras”, [J. Math. Phys. 18, 1162 (1977)]. Journal of Mathematical Physics, 18(11):2240, November 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v18/i11/p2240_s1. See [919].

Moses:1977:GGL

[1092] H. E. Moses. A generalization of the Gel’fand–Levitan equation for the one-dimensional Schrödinger equation. Journal of Mathematical Physics, 18(12):2243–2250, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2243_s1.

Joseph:1977:MSL

[1093] R. I. Joseph. Multi-soliton-like solutions to the Benjamin–Ono equation. Journal of Mathematical Physics, 18(12):2251–2258, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i12/p2251_s1.

Patera:1977:CSG

[1094] J. Patera, R. T. Sharp, P. Winternitz, and H. Zassenhaus. Continuous sub-group of the fundamental groups of physics. III. The de Sitter groups. Journal of Mathematical Physics, 18(12):2259–2288, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i12/p2259_s1.

Benioff:1977:FIM

[1095] Paul A. Benioff. Finite and infinite measurement sequences in quantum mechanics and randomness: The Everett interpretation. Journal of Mathematical Physics, 18(12):2289–2295, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/p2289_s1.

Hostler:1977:VSH

[1096] Levere Hostler. Vector spherical harmonics of the unit hyperboloid in Minkowski space. Journal of Mathematical Physics, 18(12):2296–2307, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2296_s1.

Truman:1977:CMD

[1097] Aubrey Truman. Classical mechanics, the diffusion (heat) equation, and the Schrödinger equation. Journal of Mathematical Physics, 18(12):2308– 2315, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2308_s1.

Hill:1977:PHI

[1098] Robert Nyden Hill. Proof that the H^{−} ion has only one bound state. Details and extensions to finite nuclear mass. Journal of Mathematical Physics, 18(12):2316–2330, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i12/p2316_s1.

Fox:1977:GLE

[1099] Ronald Forrest Fox. The generalized Langevin equation with Gaussian fluctuations. Journal of Mathematical Physics, 18(12):2331–2335, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2331_s1.

Chandler:1977:BQS

[1100] Colston Chandler and A. G. Gibson. N-body quantum scattering theory in two Hilbert spaces. I. The basic equations. Journal of Mathematical Physics, 18(12):2336–2347, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i12/p2336_s1.

Detrich:1977:AEI

[1101] John Detrich and Robert W. Conn. Analytic evaluation of an important integral in collision theory. Journal of Mathematical Physics, 18 (12):2348–2351, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2348_s1.

Lindblom:1977:MPN

[1102] Lee Lindblom. Mirror planes in Newtonian stars with stratified flows. Journal of Mathematical Physics, 18(12):2352–2355, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/
p2352_s1.

Gurses:1977:SSS

[1103] Metin Gürses. Some solutions of stationary, axially-symmetric gravitational field equations. Journal of Mathematical Physics, 18(12):2356– 2359, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2356_s1.

Wu:1977:USF

[1104] Alfred C. T. Wu. Universal singular functions in local field theory. Journal of Mathematical Physics, 18(12):2360–2363, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/p2360_s1.

Pulvirenti:1977:KMC

[1105] M. Pulvirenti and G. Riela. K.M.S. condition for stable states of infinite classical systems. Journal of Mathematical Physics, 18(12):2364–2367, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2364_s1.

Stachel:1977:CPS

[1106] John Stachel and Jerzy Plebański. Classical particles with spin. I. The WKBJ approximation. Journal of Mathematical Physics, 18(12):2368– 2374, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2368_s1.

Ishikawa:1977:PCC

[1107] Hirohisa Ishikawa. Properties of causally continuous closed universes. Journal of Mathematical Physics, 18(12):2375–2376, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/
p2375_s1.

Harris:1977:CAT

[1108] Frank E. Harris. Convergence acceleration technique for lattice sums arising in electronic-structure studies of crystalline solids. Journal of Mathematical Physics, 18(12):2377–2381, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i12/p2377_s1.

Okubo:1977:CIV

[1109] Susumu Okubo. Casimir invariants and vector operators in simple and classical Lie algebras. Journal of Mathematical Physics, 18(12):2382– 2394, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2382_s1.

Mansouri:1977:NCS

[1110] Freydoon Mansouri. A new class of superalgebras and local gauge groups in superspace. Journal of Mathematical Physics, 18(12):2395– 2396, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2395_s1.

Chen:1977:DSP

[1111] Pocheng Chen and Yih-Hsing Pao. The diffraction of sound pulses by a circular cylinder. Journal of Mathematical Physics, 18(12):2397–2406, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2397_s1.

Atkinson:1977:RRE

[1112] D. Atkinson, T. P. Pool, and H. A. Slim. Regularization of the Roy equations with a smooth cutoff. Journal of Mathematical Physics, 18 (12):2407–2413, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2407_s1.

Krause:1977:PTG

[1113] F. Krause. Perturbation theory for Green’s functions as an effective mass formalism. Journal of Mathematical Physics, 18(12):2414–2422, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2414_s1.

Park:1977:MQS

[1114] Yong Moon Park. Massless quantum sine-Gordon equation in two space– time dimensions. Correlation inequalities and infinite volume limit. Journal of Mathematical Physics, 18(12):2423–2426, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/p2423_s1.

Ranada:1977:KGK

[1115] Antonio F. Rañada and Manuel F. Rañada. Klein–Gordon kinks with fourth order derivative self-coupling. Journal of Mathematical Physics, 18(12):2427–2431, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2427_s1.

Cary:1977:TDC

[1116] John R. Cary. Time dependent canonical transformations and the symmetry-equals-invariant theorem. Journal of Mathematical Physics, 18(12):2432–2435, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2432_s1.

Cacuci:1977:NST

[1117] Dan G. Cacuci and Herbert Goldstein. Neutron slowing down and transport in a medium of constant cross-section. Journal of Mathematical Physics, 18(12):2436–2444, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i12/p2436_s1.

Kanal:1977:VPG

[1118] M. Kanal and H. E. Moses. A variational principle for the Gel’fand– Levitan equation and the Korteweg–de Vries equation. Journal of Mathematical Physics, 18(12):2445–2447, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v18/i12/p2445_s1.

Boffi:1977:ITR

[1119] V. C. Boffi and G. Spiga. Integral theory of radiative heat transfer with anisotropic scattering and general boundary conditions. Journal of Mathematical Physics, 18(12):2448–2455, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/p2448_s1.

Milgram:1977:PCP

[1120] Michael S. Milgram. On the properties of collision probability integrals in annular geometry. I. analysis. Journal of Mathematical Physics, 18 (12):2456–2467, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2456_s1.

Twersky:1977:CSF

[1121] Victor Twersky. Coherent scalar field in pair-correlated random distributions of aligned scatterers. Journal of Mathematical Physics, 18 (12):2468–2486, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2468_s1.

Jancewicz:1977:ODC

[1122] Bernard Jancewicz. Operator density current and relativistic localization problem. Journal of Mathematical Physics, 18(12):2487–2494, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i12/p2487_s1.

Sloan:1977:ANS

[1123] Alan Sloan. An application of the non-standard Trotter product formula. Journal of Mathematical Physics, 18(12):2495–2496, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i12/p2495_s1.

McQuistan:1977:ODN

[1124] R. B. McQuistan. One-dimensional, neighbor-pair degeneracies for dumbbells with distinguishable ends. Journal of Mathematical Physics, 18(12):2497–2501, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i12/p2497_s1.

Yamazaki:1977:KTS

[1125] Masatoshi Yamazaki. On the Kerr–Tomimatsu–Sato family of spinning mass solutions. Journal of Mathematical Physics, 18(12):2502–2508, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i12/p2502_s1.

Siewert:1977:ECC

[1126] C. E. Siewert. On an exceptional case concerning plasma oscillations. Journal of Mathematical Physics, 18(12):2509–2510, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/
p2509_s1.

Plebanski:1977:SES

[1127] Jerzy F. Plebański. On the separation of Einsteinian sub-structures. Journal of Mathematical Physics, 18(12):2511–2520, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/
p2511_s1.

Dhar:1977:ELE

[1128] Deepak Dhar. Erratum: “Lattices of effectively nonintegral dimensionality”. Journal of Mathematical Physics, 18(12):2520, December 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i12/

p2520_s1. See [838].

CastilhoAlcaras:1978:FAH

[1129] J. A. Castilho Alcarás. On the first approximation of the K-harmonics method. Journal of Mathematical Physics, 19(1):1–4, January 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p1_s1.

Dhar:1978:SAR

[1130] Deepak Dhar. Self-avoiding random walks: Some exactly soluble cases. Journal of Mathematical Physics, 19(1):5–11, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p5_s1.

Iwai:1978:PSD

[1131] Toshihiro Iwai. On projective symmetries of dynamical systems. Journal of Mathematical Physics, 19(1):12–18, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p12_s1.

Miodek:1978:SNE

[1132] I. Miodek. ‘IST-solvable’ nonlinear evolution equations and existence — An extension of Lax’s method. Journal of Mathematical Physics, 19(1): 19–31, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p19_s1.

Gouedard:1978:DEG

[1133] C. Gouedard and C. Deutsch. Dense electron-gas response at any degeneracy. Journal of Mathematical Physics, 19(1):32–38, January 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p32_s1.

See [1941].

Battle:1978:DPT

[1134] Guy A. Battle. Dynamics and phase transitions for a continuous system of quantum particles in a box. Journal of Mathematical Physics, 19(1): 39–51, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p39_s1.

Taylor:1978:IGT

[1135] William J. Taylor. Inverse Gaussian transforms: General properties and application to Slater-type orbitals with noninteger n in the coordinate and momentum representations. Journal of Mathematical Physics, 19(1): 52–58, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p52_s1. See [1453].

Garber:1978:NCCa

[1136] W. D. Garber and H. Reeh. Nontranslationally covariant currents and associated symmetry generators. Journal of Mathematical Physics, 19(1): 59–66, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p59_s1.

Ohnuki:1978:WMR

[1137] Y. Ohnuki and S. Kamefuchi. On the wave-mechanical representation of a Bose-like oscillator. Journal of Mathematical Physics, 19(1):67– 78, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p67_s1.

Filter:1978:TDC

[1138] E. Filter and E. O. Steinborn. The three-dimensional convolution of reduced Bessel functions and other functions of physical interest. Journal of Mathematical Physics, 19(1):79–84, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p79_s1.

Morris:1978:GPS

[1139] H. C. Morris. A generalized prolongation structure and the Bäcklund transformation of the anticommuting massive Thirring model. Journal of Mathematical Physics, 19(1):85–87, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p85_s1.

Weaver:1978:ES

[1140] D. L. Weaver. Eigenvalues of S · π for spins 1/2, 1, and 3/2. Journal of Mathematical Physics, 19(1):88–91, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p88_s1.

Moreno:1978:SPN

[1141] Carlos Moreno. Spaces of positive and negative frequency solutions of field equations in curved space–times. II. The massive vector field equations in static space–times. Journal of Mathematical Physics, 19(1):92– 99, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p92_s1.

Ray:1978:VDB

[1142] John R. Ray. A variational derivation of the Bach–Lanczos identity. Journal of Mathematical Physics, 19(1):100–102, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p100_s1.

Hammer:1978:UCB

[1143] C. L. Hammer and B. Defacio. Uniqueness connection between charge conjugation and statistics. Journal of Mathematical Physics, 19(1):103– 108, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p103_s1.

Dokmeci:1978:TVC

[1144] M. Cengiz Dökmeci. Theory of vibrations of coated, thermopiezoelectric laminae. Journal of Mathematical Physics, 19(1):109–126, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i1/p109_s1.

Fox:1978:FTG

[1145] Ronald Forrest Fox. Fluctuation theories and Gaussian stochastic processes. Journal of Mathematical Physics, 19(1):127–130, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/
p127_s1.

Wannier:1978:TGW

[1146] Gregory H. Wannier. Transforming Gaussians into Wannier functions. Journal of Mathematical Physics, 19(1):131–134, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p131_s1.

Menikoff:1978:DVE

[1147] R. Menikoff and D. H. Sharp. A derivation of the virial expansion with application to Euclidean quantum field theory. Journal of Mathematical Physics, 19(1):135–150, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i1/p135_s1.

Ibanez:1978:MSSa

[1148] J. L. Ibañez and M. G. Velarde. Multiple steady states in a simple reaction-diffusion model with Michaelis–Menten (first-order Hinshelwood–Langmuir) saturation law: The limit of large separation in the two diffusion constants. Journal of Mathematical Physics, 19(1):151– 156, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p151_s1.

Lenard:1978:CLT

[1149] A. Lenard. A characterization of Lorentz transformations. Journal of Mathematical Physics, 19(1):157, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p157_s1.

Fennelly:1978:ESS

[1150] A. J. Fennelly. Expanding shearfree spatially homogeneous universes with a nonsynchronous time coordinate and anisotropy of the universe. Journal of Mathematical Physics, 19(1):158–163, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p158_s1.

Edwards:1978:NAE

[1151] S. A. Edwards. A new approach to the eigenvalues of the Gel’fand invariants for the unitary, orthogonal, and symplectic groups. Journal of Mathematical Physics, 19(1):164–167, January 1978. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p164_s1.

Chiu:1978:SIT

[1152] S. C. Chiu. On the self-induced transparency effect of the three-wave resonance process. Journal of Mathematical Physics, 19(1):168–176, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i1/p168_s1.

Zorbas:1978:PTC

[1153] J. Zorbas. Perturbation theory for Coulomb scattering. Journal of Mathematical Physics, 19(1):177–186, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i1/p177_s1.

Hauser:1978:FAS

[1154] I. Hauser and R. J. Malhiot. Forms of all spacetime metrics which admit [(11) (11)] Killing tensors with nonconstant eigenvalues. Journal of Mathematical Physics, 19(1):187–194, January 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i1/p187_s1.

Kumei:1978:RBC

[1155] Sukeyuki Kumei. On the relationship between conservation laws and invariance groups of nonlinear field equations in hamilton’s canonical form. Journal of Mathematical Physics, 19(1):195–199, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p195_s1.

Boyer:1978:SSV

[1156] C. P. Boyer, E. G. Kalnins, and W. Miller. Symmetry and separation of variables for the Hamilton–Jacobi equation W_{t}^{2} − W_{x}^{2} − W_{y}^{2} = 0. Journal of Mathematical Physics, 19(1):200–211, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p200_s1.

Ibanez:1978:MSSb

[1157] J. L. Ibanez and M. G. Velarde. Multiple steady states in a simple reaction-diffusion model with Michaelis–Jacobi equation W_{t}^{2} − W_{x}^{2}W_{y}^{2} = 0. Journal of Mathematical Physics, 19(1):200–211, January 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Bergmann:1978:SGT

[1158] Peter G. Bergmann and Edward J. Flaherty, Jr. Symmetries in gauge theories. Journal of Mathematical Physics, 19(1):212–214, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/
p212_s1.

Twersky:1978:CEW

[1159] Victor Twersky. Coherent electromagnetic waves in pair-correlated random distributions of aligned scatterers. Journal of Mathematical Physics, 19(1):215–230, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i1/p215_s1.

Mutze:1978:NGT

[1160] U. Mutze. A no-go theorem concerning the cluster decomposition property of direct-interaction scattering theories. Journal of Mathematical Physics, 19(1):231–236, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i1/p231_s1.

Cazzola:1978:IEE

[1161] P. Cazzola, L. Lucaroni, and R. Semenzato. Integral equations for even parity perturbations of hot perfect fluid nonrotating neutron stars. Journal of Mathematical Physics, 19(1):237–246, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p237_s1.

Jordan:1978:IVO

[1162] Thomas F. Jordan. Identification of the velocity operator in an irreducible unitary representation of the Poincaré group for imaginary mass or zero mass and variable helicity. Journal of Mathematical Physics, 19(1):247–248, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i1/p247_s1.

Garbanati:1978:USL

[1163] L. Freede Garbanati, William Greenberg, and P. F. Zweifel. Uniqueness of solutions to the linearized Boltzmann equation. Journal of Mathematical Physics, 19(1):249–252, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i1/p249_s1.

Hooshyar:1978:ISP

[1164] M. A. Hooshyar. On the inverse scattering problem for a class of spinorbit and central potentials. Journal of Mathematical Physics, 19(1):253– 263, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p253_s1.

Hillion:1978:SHE

[1165] Pierre Hillion. The spinor Helmholtz equation. Journal of Mathematical Physics, 19(1):264–269, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i1/p264_s1.

Tikochinsky:1978:TBG

[1166] Yoel Tikochinsky. Transformation brackets for generalized Bogolyubovboson transformations. Journal of Mathematical Physics, 19(1):270– 276, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p270_s1.

Tsallis:1978:DMG

[1167] Constantino Tsallis. Diagonalization methods for the general bilinear Hamiltonian of an assembly of bosons. Journal of Mathematical Physics, 19(1):277–286, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i1/p277_s1.

Chernavskii:1978:RAB

[1168] D. S. Chernavskii and R. Kerner. Remarks on the asymptotic behaviour of the ’t Hooft’s magnetic monopole. Journal of Mathematical Physics, 19(1):287–291, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i1/p287_s1.

Klein:1978:WAB

[1169] Abraham Klein. WKB approximation for bound states by Heisenberg matrix mechanics. Journal of Mathematical Physics, 19(1):292–297, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i1/p292_s1.

Mizrahi:1978:PSP

[1170] Maurice M. Mizrahi. Phase space path integrals, without limiting procedure. Journal of Mathematical Physics, 19(1):298–307, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/

p298_s1. See erratum [1946].

Antippa:1978:LPS

[1171] Adel F. Antippa and Alain J. Phares. The linear potential: a solution in terms of combinatorics functions. Journal of Mathematical Physics, 19(1):308–319, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i1/p308_s1.

Yang:1978:GDM

[1172] Chen Ning Yang. Generalization of Dirac’s monopole to su_{2} gauge fields. Journal of Mathematical Physics, 19(1):320–328, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p320_s1.

Antoine:1978:OSN

[1173] J. P. Antoine and A. Grossmann. Orthocomplemented subspaces of nondegenerate partial inner product spaces. Journal of Mathematical Physics, 19(1):329–335, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i1/p329_s1.

Ginsburg:1978:ANI

[1174] Charles A. Ginsburg and Elliott W. Montroll. Application of a novel interpolative perturbation scheme to the determination of anharmonic oscillator wavefunctions. Journal of Mathematical Physics, 19(1):336– 346, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i1/p336_s1.

Post:1978:GVA

[1175] E. J. Post. The Gaussian version of Ampère’s law. Journal of Mathematical Physics, 19(1):347, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i1/p347_s1.

Furutani:1978:EBS

[1176] Y. Furutani and C. Deutsch. Erratum: “The n-bubble series in the theory of the classical one-component plasma”. Journal of Mathematical Physics, 19(1):348, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i1/p348_s1. See [796].

Tariq:1978:NBR

[1177] Nessim Tariq and Raymond G. McLenaghan. Note on the Bertotti– Robinson electromagnetic universe. Journal of Mathematical Physics, 19(2):349–351, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p349_s1.

Weiszmann:1978:CFH

[1178] Andrei N. Weiszmann. Correlation functions and higher order coherence in inelastically scattered quantum radiation. Journal of Mathematical Physics, 19(2):352–353, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i2/p352_s1.

Weiszmann:1978:RFD

[1179] Andrei N. Weiszmann. The reduced field density operator for an inelastically scattered quantum radiation. Journal of Mathematical Physics, 19(2):354–356, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p354_s1.

Lonngren:1978:CIB

[1180] K. E. Lonngren and H. C. S. Hsuan. A consequence of the invariance of Biot’s variational principle in thermal conduction. Journal of Mathematical Physics, 19(2):357–358, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i2/p357_s1.

Besieris:1978:KFT

[1181] Ioannis M. Besieris, Wojciech B. Stasiak, and Frederick D. Tappert. A kinetic formulation of the three-dimensional quantum mechanical harmonic oscillator under a random perturbation. Journal of Mathematical Physics, 19(2):359–369, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i2/p359_s1.

Krause:1978:LTS

[1182] Jorge Krause. Lorentz transformations as space–time reflections. II. Time like reflections. Journal of Mathematical Physics, 19(2):370–374, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i2/p370_s1.

Milton:1978:SGI

[1183] Kimball A. Milton and Lester L. DeRaad, Jr. Strings and gauge invariance. Journal of Mathematical Physics, 19(2):375–382, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i2/p375_s1.

Pao:1978:SSE

[1184] C. V. Pao. Subcriticality and supercriticality of energy dependent neutron transport in slab geometry. Journal of Mathematical Physics, 19(2):383–386, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p383_s1.

Vazquez:1978:ISL

[1185] Luis Vázquez. Interaction and stability of localized solutions in a classical non-linear scalar field theory. Journal of Mathematical Physics, 19(2):387–389, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p387_s1.

Kuchar:1978:EPH

[1186] Karel Kuchar. On equivalence of parabolic and hyperbolic super-Hamiltonians. Journal of Mathematical Physics, 19(2):390–400, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i2/p390_s1.

Vasholz:1978:RLS

[1187] David P. Vasholz. Random linear systems and temporal homogeneity. Journal of Mathematical Physics, 19(2):401–408, February 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/p401_
s1.

Uchiyama:1978:LBT

[1188] Tadashi Uchiyama. Lower bounds on the total cross section and the slope parameter for some measurable sequences of s → ∞. Journal of Mathematical Physics, 19(2):409–417, February 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i2/p409_s1.

Ovrut:1978:ISH

[1189] Burt A. Ovrut. Isotropy subgroups of SO(3) and Higgs potentials. Journal of Mathematical Physics, 19(2):418–425, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/p418_s1.

Weilenmann:1978:AAN

[1190] Jürg Weilenmann. An alternative approach to the normal frequencies of a randomly disordered linear chain. Journal of Mathematical Physics, 19(2):426–433, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p426_s1.

Siewert:1978:ERQ

[1191] C. E. Siewert. Explicit results for the quantum-mechanical energy states basic to a finite square-well potential. Journal of Mathematical Physics, 19(2):434–435, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p434_s1.

Maciejko:1978:SCN

[1192] R. Maciejko. Symmetry conditions and non-Abelian gauge fields. Journal of Mathematical Physics, 19(2):436–441, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/p436_s1.

Goswami:1978:SSN

[1193] G. K. Goswami. Some static and non-static solutions of Brans–Dicke theory of gravitation. Journal of Mathematical Physics, 19(2):442–445, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i2/p442_s1.

Leach:1978:QHQ

[1194] P. G. L. Leach. Quadratic Hamiltonians, quadratic invariants and the symmetry group SU(n). Journal of Mathematical Physics, 19(2):446– 451, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i2/p446_s1.

Cattaneo:1978:SRG

[1195] U. Cattaneo. Splitting and representation groups for Polish groups. Journal of Mathematical Physics, 19(2):452–460, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/p452_s1.

Faris:1978:IUP

[1196] William G. Faris. Inequalities and uncertainty principles. Journal of Mathematical Physics, 19(2):461–466, February 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i2/p461_s1.

Raynal:1978:DPG

[1197] Jacques Raynal. On the definition and properties of generalized 3-j symbols. Journal of Mathematical Physics, 19(2):467–476, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/
p467_s1.

Bountis:1978:ESM

[1198] Tassos Bountis and Robert H. G. Helleman. Exact statistical mechanics of some classical 1D systems. Journal of Mathematical Physics, 19(2):477–481, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p477_s1.

Havas:1978:CEG

[1199] Peter Havas and Jerzy Plebański. Conformal extensions of the Galilei group and their relation to the Schrödinger group. Journal of Mathematical Physics, 19(2):482–488, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i2/p482_s1.

Ernst:1978:CDL

[1200] Frederick J. Ernst. Coping with different languages in the null tetrad formulation of general relativity. Journal of Mathematical Physics, 19(2):489–493, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p489_s1. See [1887].

Kelley:1978:ACO

[1201] C. T. Kelley. Analytic continuation of an operator-valued H-function with applications to neutron transport theory. Journal of Mathematical Physics, 19(2):494–499, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i2/p494_s1.

Kelley:1978:SIE

[1202] C. T. Kelley and T. W. Mullikin. Solution by iteration of H-equations in multigroup neutron transport. Journal of Mathematical Physics, 19(2):500–501, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p500_s1.

Kaiser:1978:PSA

[1203] Gerald R. Kaiser. Phase-space approach to relativistic quantum mechanics. II. Geometrical aspects. Journal of Mathematical Physics, 19(2):502– 507, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i2/p502_s1.

Guler:1978:UAR

[1204] Y. Güler. Unitary analytic representations of SL(3, R) and hadronic Regge sequences. Journal of Mathematical Physics, 19(2):508–510, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i2/p508_s1.

Partensky:1978:C

[1205] A. Partensky and C. Maguin. The SU(4) ⊃ SU(2)⊗SU(2) chain. Journal of Mathematical Physics, 19(2):511–525, February 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i2/p511_s1.

Elizalde:1978:PSG

[1206] Emili E. Elizalde. Poincaré is a subgroup of Galilei in one space dimension more. Journal of Mathematical Physics, 19(2):526–528, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i2/p526_s1.

Wybourne:1978:SI

[1207] B. G. Wybourne. su_{6} × su_{3}^{c} scalars in E_{7} irreps. Journal of Mathematical Physics, 19(2):529–530, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i2/p529_s1.

Murray:1978:PFW

[1208] J. C. Murray. Poisson’s formulas for wave propagation in a superfluid. Journal of Mathematical Physics, 19(2):531–534, February 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/p531_
s1.

Zenk:1978:ASS

[1209] L. G. Zenk and A. Das. An algebraically special subclass of vacuum metrics admitting a Killing motion. Journal of Mathematical Physics, 19(2):535–538, February 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i2/p535_s1.

Pardee:1978:EAE

[1210] W. J. Pardee. Erratum: “Acoustic emission and the plate Green’s function”. Journal of Mathematical Physics, 19(2):538, February 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i2/p538_

s1. See [853].

Hoenselaers:1978:SEM

[1211] C. Hoenselaers. On the stationary Einstein–Maxwell equations. Journal of Mathematical Physics, 19(3):539–541, March 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i3/p539_s1.

Nickel:1978:ESF

[1212] Bernie G. Nickel. Evaluation of simple Feynman graphs. Journal of Mathematical Physics, 19(3):542–548, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i3/p542_s1.

Sommers:1978:GGF

[1213] Paul Sommers. The geometry of the gravitational field at spacelike infinity. Journal of Mathematical Physics, 19(3):549–554, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p549_
s1.

Tsamparlis:1978:PMV

[1214] Michael Tsamparlis. On the Palatini method of variation. Journal of Mathematical Physics, 19(3):555–557, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i3/p555_s1.

Fannes:1978:GTS

[1215] M. Fannes and A. Verbeure. Global thermodynamical stability and correlation inequalities. Journal of Mathematical Physics, 19(3):558– 560, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i3/p558_s1.

Dhurandhar:1978:TSW

[1216] S. V. Dhurandhar. Tachyonic scalar waves in the Schwarzschild space– time. Journal of Mathematical Physics, 19(3):561–567, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p561_
s1.

Naito:1978:SHI

[1217] Seichi Naito. Slavnov-’t Hooft identities in Mandelstam’s formalism. Journal of Mathematical Physics, 19(3):568–573, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p568_s1.

Kopsky:1978:ISB

[1218] Vojtech Kopský. On independent sets of basis functions for irreducible representations of finite groups. Journal of Mathematical Physics, 19(3): 574–576, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i3/p574_s1.

Zorbas:1978:STS

[1219] J. Zorbas. Scattering theory for Stark Hamiltonians involving longrange potentials. Journal of Mathematical Physics, 19(3):577–580, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i3/p577_s1.

Smith:1978:MEE

[1220] George B. Smith. Matrix element expansion of a spin wavefunction. Journal of Mathematical Physics, 19(3):581–585, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p581_s1.

Lovelock:1978:IGE

[1221] David Lovelock. The interaction of the gravitational and electromagnetic fields. Journal of Mathematical Physics, 19(3):586–592, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p586_
s1.

Lim:1978:MWR

[1222] Chee-Seng Lim. Multidimensional wave radiation from a source with Gaussian time variation and Gaussian-approximated distribution about a spherical sheet. Journal of Mathematical Physics, 19(3):593–607, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i3/p593_s1.

Hanckowiak:1978:IPP

[1223] Jerzy Hańćkowiak. Invariant properties of n-point functions and n-point functionals connected with the translational invariance of the formal measure. Journal of Mathematical Physics, 19(3):608–612, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p608_
s1.

Shurtleff:1978:GTC

[1224] Richard Shurtleff. A geometric theory of charge and mass. Journal of Mathematical Physics, 19(3):613–618, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i3/p613_s1.

Benza:1978:LTB

[1225] V. Benza and E. Montaldi. Large time behavior of the superfluorescent decay. Journal of Mathematical Physics, 19(3):619–623, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p619_

s1. See [1452].

Camenzind:1978:HIW

[1226] M. Camenzind. Homogeneous and isotropic world models in the Yang– Mills dynamics of gravity. The structure of the adiabats. Journal of Mathematical Physics, 19(3):624–634, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i3/p624_s1.

Johnston:1978:IRP

[1227] George L. Johnston. Integral representations of particular integrals of a class of inhomogeneous linear ordinary differential equations. Journal of Mathematical Physics, 19(3):635–641, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i3/p635_s1.

Garbaczewski:1978:QSF

[1228] Piotr Garbaczewski. Quantization of spinor fields. Journal of Mathematical Physics, 19(3):642–652, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i3/p642_s1.

Benfatto:1978:LNA

[1229] G. Benfatto and F. Nicolò. The local von Neumann algebras for the massless scalar free field and the free electromagnetic field. Journal of Mathematical Physics, 19(3):653–660, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i3/p653_s1.

Hauser:1978:TGF

[1230] I. Hauser. Type N gravitational field with twist. II. Journal of Mathematical Physics, 19(3):661–667, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i3/p661_s1.

Horndeski:1978:BTM

[1231] Gregory Walter Horndeski. Birkhoff’s theorem and magnetic monopole solutions for a system of generalized Einstein–Maxwell field equations. Journal of Mathematical Physics, 19(3):668–674, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p668_s1.

Karaki:1978:RRF

[1232] Yukihiko Karaki. Rigorous results on ferromagnetic lattice spin systems. I. Journal of Mathematical Physics, 19(3):675–682, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p675_s1.

Kramer:1978:NCT

[1233] P. Kramer, M. Moshinsky, and T. H. Seligman. Nonbijective canonical transformations and their representation in quantum mechanics. Journal of Mathematical Physics, 19(3):683–693, March 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i3/p683_s1.

Robinson:1978:NVB

[1234] Peter D. Robinson. New variational bounds on generalized polarizabilities. Journal of Mathematical Physics, 19(3):694–699, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p694_
s1.

Aldersley:1978:CCM

[1235] S. J. Aldersley. Charge conservation in metric-torsion gravitational theories. Journal of Mathematical Physics, 19(3):700–702, March 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/p700_
s1.

Wong:1978:WC

[1236] M. K. F. Wong. On the Weyl coefficients of SO(n). Journal of Mathematical Physics, 19(3):703–708, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i3/p703_s1.

Rittenberg:1978:ECG

[1237] V. Rittenberg and M. Scheunert. Elementary construction of graded Lie groups. Journal of Mathematical Physics, 19(3):709–713, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i3/p709_s1.

Geroch:1978:ASS

[1238] Robert Geroch and Basilis C. Xanthopoulos. Asymptotic simplicity is stable. Journal of Mathematical Physics, 19(3):714–719, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i3/
p714_s1.

Braunschweig:1978:BRS

[1239] D. Braunschweig and K. T. Hecht. Branching rules for the subgroups of the unitary group. Journal of Mathematical Physics, 19(3):720– 722, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i3/p720_s1.

Ray:1978:SGI

[1240] Dipankar Ray. Self-gravitating irrotational barotropic fluid in a conformally flat space. Journal of Mathematical Physics, 19(3):723–726, March 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i3/p723_s1.

Gaskell:1978:GFP

[1241] R. Gaskell, A. Peccia, and R. T. Sharp. Generating functions for polynomial irreducible tensors. Journal of Mathematical Physics, 19(4):727– 733, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p727_s1.

Gazeau:1978:GLP

[1242] J.-P. Gazeau, M.-Cl. Dumont-Lepage, and A. Ronveaux. Gel’fand lattice polynomials and irreducible representations of U(n). Journal of Mathematical Physics, 19(4):734–748, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i4/p734_s1.

Duval:1978:CIM

[1243] C. Duval and H. H. Fliche. A conformal invariant model of localized spinning test particles. Journal of Mathematical Physics, 19(4):749– 752, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p749_s1.

Hoffmann:1978:MMD

[1244] Patrick Hoffmann. The moments of the multiple distribution functions. Journal of Mathematical Physics, 19(4):753–757, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p753_s1.

Cahill:1978:SVI

[1245] K. Cahill and A. Comtet. Scalar-vector instantons in n-dimensions: surface terms. Journal of Mathematical Physics, 19(4):758–759, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i4/p758_s1.

Finley:1978:KVP

[1246] J. D. Finley III and J. F. Plebański. Killing vectors in plane HH spaces. Journal of Mathematical Physics, 19(4):760–766, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p760_s1.

Cattaneo:1978:QMP

[1247] U. Cattaneo. The quantum mechanical Poincaré and Galilei groups. Journal of Mathematical Physics, 19(4):767–773, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p767_s1.

Tenti:1978:SCE

[1248] G. Tenti and W. H. Hui. Some classes of exact solutions of the nonlinear Boltzmann equation. Journal of Mathematical Physics, 19(4): 774–779, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p774_s1.

Martin:1978:NIT

[1249] J. Martin and J. L. Sanz. No-interaction theorem of Currie, Jordan, and Sudarshan: Expansions in c^{−1}. Journal of Mathematical Physics, 19(4): 780–788, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p780_s1.

Bosanac:1978:MCR

[1250] S. Bosanac. A method for calculation of Regge poles in atomic collisions. Journal of Mathematical Physics, 19(4):789–797, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p789_s1.

Kaup:1978:ESD

[1251] David J. Kaup and Alan C. Newell. An exact solution for a derivative nonlinear Schrödinger equation. Journal of Mathematical Physics, 19(4):

798–801, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p798_s1.

Romerio:1978:AOC

[1252] M. Romerio. Absence of ordering in a class of real liquids-application to nematic liquid crystals. Journal of Mathematical Physics, 19(4):802– 805, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p802_s1.

Dollard:1978:TDS

[1253] John D. Dollard. Time-dependent scattering theory for infinite delta function potentials in one dimension. Journal of Mathematical Physics, 19(4):806–807, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p806_s1.

Kracklauer:1978:TET

[1254] Aloysius F. Kracklauer. A theory of the electromagnetic two-body interaction. Journal of Mathematical Physics, 19(4):808–811, April 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p808_
s1.

McCartor:1978:EBC

[1255] Gary McCartor. Error bounds for the complex-coordinate method. Journal of Mathematical Physics, 19(4):812–815, April 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i4/p812_s1.

Yu:1978:LHS

[1256] M. Y. Yu. Lower hybrid solitary waves. Journal of Mathematical Physics, 19(4):816–818, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p816_s1.

Kesarwani:1978:HPD

[1257] R. N. Kesarwani and Y. P. Varshni. High-precision determination of the critical screening length for the static screened Coulomb potential. Journal of Mathematical Physics, 19(4):819–820, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p819_s1.

Garibotti:1978:SPW

[1258] C. R. Garibotti and F. F. Grinstein. Summation of partial wave expansions in the scattering by long range potentials. I. Journal of Mathematical Physics, 19(4):821–829, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i4/p821_s1.

Lee:1978:PTB

[1259] Chien er Lee and Jung-Kai Hong. Pseudoscalar transition between a spin-J and a spin-5/2 baryon. Journal of Mathematical Physics, 19(4): 830–832, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i4/p830_s1.

Arenstorf:1978:ERN

[1260] Richard F. Arenstorf, Jeffrey M. Cohen, and Lawrence S. Kegeles. Electromagnetic radiation near black holes and neutron stars. Journal of Mathematical Physics, 19(4):833–837, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i4/p833_s1.

Ventura:1978:SQF

[1261] I. Ventura and G. C. Marques. Semiclassical quantization of a field theoretic model in any number of spatial dimensions. Journal of Mathematical Physics, 19(4):838–847, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p838_s1.

Davidson:1978:LTI

[1262] A. Davidson and A. Dar. Lorentz-type internal symmetry. Journal of Mathematical Physics, 19(4):848–855, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i4/p848_s1.

Glass:1978:RSS

[1263] E. N. Glass and S. P. Goldman. Relativistic spherical stars reformulated. Journal of Mathematical Physics, 19(4):856–859, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p856_s1.

Lieb:1978:SDF

[1264] Elliott H. Lieb and Manuel de Llano. Solitons and the delta function fermion gas in Hartree–Fock theory. Journal of Mathematical Physics, 19(4):860–868, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p860_s1.

Howe:1978:LSDa

[1265] P. S. Howe and R. W. Tucker. Local supersymmetry in (2 + 1) dimensions. I. Supergravity and differential forms. Journal of Mathematical Physics, 19(4):869–873, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p869_s1.

Curtis:1978:CLB

[1266] W. D. Curtis and D. E. Lerner. Complex line bundles in relativity. Journal of Mathematical Physics, 19(4):874–877, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p874_s1.

Bloore:1978:WIP

[1267] F. J. Bloore and S. J. Swarbrick. Wavefunctions of identical particles. Journal of Mathematical Physics, 19(4):878–879, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p878_s1.

Atanackovic:1978:GMS

[1268] T. M. Atanackovic. Geometry of motion of a single elastic body point. Journal of Mathematical Physics, 19(4):880–882, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i4/p880_s1.

Pomraning:1978:VI

[1269] G. C. Pomraning. Variational interpolation. Journal of Mathematical Physics, 19(4):883–887, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p883_s1.

Tikochinsky:1978:EPQ

[1270] Yoel Tikochinsky. Exact propagators for quadratic Hamiltonians. Journal of Mathematical Physics, 19(4):888–891, April 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i4/p888_s1.

Rajan:1978:MDC

[1271] V. T. Rajan, Chia-Wei Woo, and F. Y. Wu. Multiple density correlation functions for inhomogeneous systems. Journal of Mathematical Physics, 19(4):892–897, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i4/p892_s1.

Pereira:1978:SIH

[1272] Nino R. Pereira. Scaling invariance of helical curve motion and soliton equations. Journal of Mathematical Physics, 19(4):898–900, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i4/p898_s1.

Cazzola:1978:AED

[1273] P. Cazzola, L. Lucaroni, and R. Semenzato. Asymptotic eigenfrequency distribution for even-parity perturbations of hot perfect-fluid relativistic neutron stars. Journal of Mathematical Physics, 19(4):901–910, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i4/p901_s1.

Otsuki:1978:DTP

[1274] Takuro Otsuki. Diffraction by two parallel slits in a plane. Journal of Mathematical Physics, 19(4):911–915, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i4/p911_s1.

Shahinpoor:1978:EAG

[1275] M. Shahinpoor. Erratum: “Asymmetric gas thermodynamics”. Journal of Mathematical Physics, 19(4):915, April 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i4/p915_s1. See [879].

Manoukian:1978:HEB

[1276] Edward B. Manoukian. High-energy behavior of renormalized Feynman amplitudes. Journal of Mathematical Physics, 19(5):917–923, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p917_s1.

Chen:1978:STA

[1277] K.-H. Chen and C. C. Yang. Spectral theory for the acoustic wave equations with generalized Neumann boundary conditions in exterior domain. Journal of Mathematical Physics, 19(5):924–929, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p924_s1.

Larsen:1978:USG

[1278] M. L. Larsen and W. W. Repko. The use of the symmetric group in the construction of multispinor Lagrangians. Journal of Mathematical Physics, 19(5):930–934, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p930_s1.

Sidhu:1978:DPS

[1279] S. S. Sidhu and R. H. Good, Jr. Definition of polarization of a spin 1/2 particle in an external electromagnetic field. Journal of Mathematical Physics, 19(5):935–936, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p935_s1.

Chehab:1978:SOC

[1280] R. Chehab. A second order calculation of the adiabatic invariant of a charged particle spiraling in a longitudinal magnetic field. Journal of Mathematical Physics, 19(5):937–941, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i5/p937_s1.

Gursey:1978:GLG

[1281] Feza Gürsey and Louis Marchildon. The graded Lie groups SU(2, 2/1) and OSp(1/4). Journal of Mathematical Physics, 19(5):942–951, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p942_s1.

Brooke:1978:GFS

[1282] J. A. Brooke. A Galileian formulation of spin. I. Clifford algebras and Spin groups. Journal of Mathematical Physics, 19(5):952–959, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/
p952_s1.

Leibbrandt:1978:SSE

[1283] G. Leibbrandt. Solitonlike solutions of the elliptic sine-cosine equation by means of harmonic functions. Journal of Mathematical Physics, 19(5): 960–966, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p960_s1.

Ohya:1978:SWI

[1284] Masanori Ohya. Stability of Weiss Ising model. Journal of Mathematical Physics, 19(5):967–971, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p967_s1.

Ruhl:1978:ARP

[1285] W. Rühl and B. C. Yunn. The asymptotic reduction of products of representations of the universal covering group of SU(1, 1). Journal of Mathematical Physics, 19(5):972–980, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i5/p972_s1.

Howe:1978:LSDb

[1286] P. S. Howe and R. W. Tucker. Local supersymmetry in (2 + 1) dimensions. II. An action for a spinning membrane. Journal of Mathematical Physics, 19(5):981–984, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p981_s1.

Garber:1978:NCCb

[1287] W. D. Garber and H. Reeh. Nontranslationally covariant currents and symmetries of the S matrix. Journal of Mathematical Physics, 19(5): 985–986, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p985_s1.

Johri:1978:ESB

[1288] V. B. Johri and G. K. Goswami. Electromagnetic solutions of Brans– Dicke theory of gravitation from Einstein theory. Journal of Mathematical Physics, 19(5):987–990, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i5/p987_s1.

Magg:1978:SCF

[1289] M. Magg. Some constraints on finite energy solutions in non-Abelian gauge theories. Journal of Mathematical Physics, 19(5):991–993, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p991_s1.

McCormick:1978:IPT

[1290] N. J. McCormick and J. A. R. Veeder. On the inverse problem of transport theory with azimuthal dependence. Journal of Mathematical Physics, 19(5):994–998, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p994_s1. See addendum [1590].

Weingarten:1978:ABG

[1291] Don Weingarten. Asymptotic behavior of group integrals in the limit of infinite rank. Journal of Mathematical Physics, 19(5):999–1001, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p999_s1.

Graffi:1978:BSI

[1292] S. Graffi and V. Grecchi. Borel summability and indeterminacy of the Stieltjes moment problem: Application to the anharmonic oscillators. Journal of Mathematical Physics, 19(5):1002–1006, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1002_s1.

Zivkovic:1978:ACB

[1293] Tomislav P. Zivković and Hendrik J. Monkhorst. Analytic connection between configuration-interaction and coupled-cluster solutions. Journal of Mathematical Physics, 19(5):1007–1022, May 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1007_s1.

Ascoli:1978:RLO

[1294] R. Ascoli, P. Gabbi, and G. Palleschi. On the representation of linear operators in L^{2} spaces by means of “generalized matrices”. Journal of Mathematical Physics, 19(5):1023–1027, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i5/p1023_s1.

Lubkin:1978:ESC

[1295] Elihu Lubkin. Entropy of an n-system from its correlation with a Kreservoir. Journal of Mathematical Physics, 19(5):1028–1031, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/
p1028_s1.

Boisseau:1978:QLM

[1296] B. Boisseau and C. Barrabes. Quantization of the Liouville mechanics for systems with singular Lagrangians. Journal of Mathematical Physics, 19(5):1032–1036, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1032_s1.

Chen:1978:ZEC

[1297] Augustine C. Chen. The zero-energy Coulomb problem. Journal of Mathematical Physics, 19(5):1037–1040, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i5/p1037_s1.

Gazeau:1978:FEC

[1298] J. P. Gazeau. Four Euclidean conformal group in atomic calculations: Exact analytical expressions for the bound-bound two-photon transition matrix elements in the H atom. Journal of Mathematical Physics, 19(5): 1041–1048, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1041_s1. See erratum [1943].

Sarlet:1978:ICL

[1299] Willy Sarlet. Invariance and conservation laws for Lagrangian systems with one degree of freedom. Journal of Mathematical Physics, 19(5): 1049–1054, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1049_s1. See addendum [1726].

McQuistan:1978:NOD

[1300] R. B. McQuistan. On the normalization of one-dimensional lattice statistics. Journal of Mathematical Physics, 19(5):1055–1056, May 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1055_
s1.

Mori:1978:LAP

[1301] Masao Mori. On Lie algebraic properties of the step operators acting on P or confluent P functions. Journal of Mathematical Physics, 19(5): 1057–1063, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1057_s1.

Hioe:1978:GFC

[1302] F. T. Hioe. A Green’s function for a cubic lattice. Journal of Mathematical Physics, 19(5):1064–1067, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i5/p1064_s1.

Newton:1978:TDS

[1303] Roger G. Newton. Three-dimensional solitons. Journal of Mathematical Physics, 19(5):1068–1073, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1068_s1.

Davidson:1978:RQS

[1304] Russell Davidson and John J. Kozak. On the relaxation to quantum statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. VIII. Emission in an infinite system in the presence of an extra photon. Journal of Mathematical Physics, 19(5):1074–1086, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p1074_s1.

Mishima:1978:TEL

[1305] Nobuhiko Mishima and Tomio Yamakoshi Petrosky. On time evolution for large quantum systems in superspace. Journal of Mathematical Physics, 19(5):1087–1099, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i5/p1087_s1.

Nowakowski:1978:NCS

[1306] Jerzy Nowakowski and Andrzej Trautman. Natural connections on Stiefel bundles are sourceless gauge fields. Journal of Mathematical Physics, 19(5):1100–1103, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1100_s1.

Beig:1978:RDU

[1307] R. Beig. Radiation (damping) in a universe with topologically closed space sections. Journal of Mathematical Physics, 19(5):1104–1109, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p1104_s1.

Lonke:1978:CSS

[1308] A. Lonke. Coherent states and stochastic formulation of quantum mechanics. Journal of Mathematical Physics, 19(5):1110–1117, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/
p1110_s1.

Tilgner:1978:GSP

[1309] Hans Tilgner. The group structure of pseudo-Riemannian curvature spaces. Journal of Mathematical Physics, 19(5):1118–1125, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/
p1118_s1.

Newell:1978:NT

[1310] Alan C. Newell. Nonlinear tunnelling. Journal of Mathematical Physics, 19(5):1126–1133, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p1126_s1.

Challifour:1978:CEL

[1311] J. L. Challifour and D. Weingarten. Cluster expansion for lattice gauge theories with fermions. Journal of Mathematical Physics, 19(5):1134– 1136, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1134_s1.

Ling:1978:UAI

[1312] Rina Ling. Uniform approximations to integral and integrodifferential equations. Journal of Mathematical Physics, 19(5):1137–1140, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/
p1137_s1.

Froman:1978:CPD

[1313] Nanny Fröman. On the calculation of the probability density at the origin of an s-state. Journal of Mathematical Physics, 19(5):1141–1146, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p1141_s1.

Hsieh:1978:VMM

[1314] D. Y. Hsieh. Variational method and Mathieu equation. Journal of Mathematical Physics, 19(5):1147–1151, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i5/p1147_s1.

Jang:1978:PEG

[1315] Pong Soo Jang. On the positivity of energy in general relativity. Journal of Mathematical Physics, 19(5):1152–1155, May 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1152_s1. See er-

ratum [1591].

Magri:1978:SMI

[1316] Franco Magri. A simple model of the integrable Hamiltonian equation. Journal of Mathematical Physics, 19(5):1156–1162, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1156_s1.

Jantzen:1978:THS

[1317] Robert T. Jantzen. Tensor harmonics on the 3-sphere. Journal of Mathematical Physics, 19(5):1163–1172, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i5/p1163_s1.

Bincer:1978:RLO

[1318] Adam M. Bincer. Raising and lowering operators for the orthogonal groups. Journal of Mathematical Physics, 19(5):1173–1178, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/
p1173_s1.

Bincer:1978:NCS

[1319] Adam M. Bincer. Normalization coefficients for the O(n) shift operators. Journal of Mathematical Physics, 19(5):1179–1183, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1179_s1.

Haley:1978:ERA

[1320] Stephen B. Haley. I. exact relations among correlation functions of inherently nondynamic interacting systems. Journal of Mathematical Physics, 19(5):1184–1186, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1184_s1.

Haley:1978:IES

[1321] Stephen B. Haley. II. exact solutions of disordered Ising spin chains in a magnetic field. Journal of Mathematical Physics, 19(5):1187–1191, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i5/p1187_s1.

Bacry:1978:ECL

[1322] H. Bacry. Eigenstates of complex linear combinations of J_{1}, J_{2}, J_{3} for any representation of SU(2). Journal of Mathematical Physics, 19(5): 1192–1195, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1192_s1.

Bacry:1978:PLA

[1323] H. Bacry. The projective Lie algebra of the Lorentz group and homographic transformations. Journal of Mathematical Physics, 19(5):1196– 1197, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1196_s1.

Maugin:1978:CERa

[1324] G. A. Maugin. On the covariant equations of the relativistic electrodynamics of continua. I. General equations. Journal of Mathematical Physics, 19(5):1198–1205, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1198_s1.

Maugin:1978:CERb

[1325] G. A. Maugin. On the covariant equations of the relativistic electrodynamics of continua. II. Fluids. Journal of Mathematical Physics, 19(5):

1206–1211, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1206_s1.

Maugin:1978:CERc

[1326] G. A. Maugin. On the covariant equations of the relativistic electrodynamics of continua. III. Elastic solids. Journal of Mathematical Physics, 19(5):1212–1219, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1212_s1.

Maugin:1978:CERd

[1327] G. A. Maugin. On the covariant equations of the relativistic electrodynamics of continua. IV. Media with spin. Journal of Mathematical Physics, 19(5):1220–1226, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i5/p1220_s1.

Spohn:1978:EPQ

[1328] Herbert Spohn. Entropy production for quantum dynamical semigroups. Journal of Mathematical Physics, 19(5):1227–1230, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1227_s1.

Barker:1978:EPN

[1329] B. M. Barker and R. F. O’Connell. Erratum: “Post-Newtonian two-body and n-body problems with electric charge in general relativity”. Journal of Mathematical Physics, 19(5):1231, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i5/p1231_s2. See [1025].

Lee:1978:ECC

[1330] L. L. Lee. Erratum: “Continuum calculus and Feynman’s path integrals”. Journal of Mathematical Physics, 19(5):1231, May 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i5/p1231_

s1. See [711].

Kalnins:1978:LTWa

[1331] E. G. Kalnins and W. Miller, Jr. Lie theory and the wave equation in space–time. 4. the Klein–Gordon equation and the Poincaré group. Journal of Mathematical Physics, 19(6):1233–1246, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1233_s1.

Kalnins:1978:LTWb

[1332] E. G. Kalnins and W. Miller, Jr. Lie theory and the wave equation in space–time. 5. R-separable solutions of the wave equation ψ_{tt} −Δ_{3}ψ = 0. Journal of Mathematical Physics, 19(6):1247–1257, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1247_s1.

Kanal:1978:VPT

[1333] Madhoo Kanal and Harry E. Moses. A variational principle for transport theory. Journal of Mathematical Physics, 19(6):1258–1264, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1258_s1.

Perroud:1978:NFS

[1334] M. Perroud. A new formula for the spinor norm. Journal of Mathematical Physics, 19(6):1265–1266, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i6/p1265_s1.

Lehnigk:1978:BCS

[1335] Siegfried H. Lehnigk. Boundary condition solutions of the generalized Feller equation. Journal of Mathematical Physics, 19(6):1267–1275, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1267_s1.

LHuillier:1978:SMB

[1336] M. L’Huillier, Edward F. Redish, and P. C. Tandy. The structure of the many-body wavefunction for scattering. Journal of Mathematical Physics, 19(6):1276–1285, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i6/p1276_s1.

Wang:1978:CMN

[1337] P. K. C. Wang. A class of multidimensional nonlinear Langmuir waves. Journal of Mathematical Physics, 19(6):1286–1292, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1286_s1.

Sesma:1978:OAR

[1338] J. Sesma and V. Vento. Optical analysis of resonances in a velocitydependent potential. Journal of Mathematical Physics, 19(6):1293–1299, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1293_s1.

Geroch:1978:NIG

[1339] Robert Geroch. Null infinity is not a good initial-data surface. Journal of Mathematical Physics, 19(6):1300–1303, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i6/p1300_s1.

Enz:1978:PMB

[1340] U. Enz. A particle model based on stringlike solitons. Journal of Mathematical Physics, 19(6):1304–1306, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i6/p1304_s1. See [1686].

Hioe:1978:CML

[1341] F. T. Hioe. Correlations in an N-mode laser. Journal of Mathematical Physics, 19(6):1307–1315, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i6/p1307_s1.

Hauser:1978:GNS

[1342] Isidore Hauser and Frederick J. Ernst. On the generation of new solutions of the Einstein–Maxwell field equations from electrovac spacetimes with isometries. Journal of Mathematical Physics, 19(6):1316– 1323, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i6/p1316_s1.

Spero:1978:ASG

[1343] Alan Spero and Ralph Baierlein. Approximate symmetry groups of inhomogeneous metrics: Examples. Journal of Mathematical Physics, 19(6): 1324–1334, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i6/p1324_s1.

Dalton:1978:NRD

[1344] B. J. Dalton. Nonlinear realizations of the direct product of two Lorentz groups on a skew-symmetric tensor space. Journal of Mathematical Physics, 19(6):1335–1339, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i6/p1335_s1.

Mason:1978:DFS

[1345] D. P. Mason and D. K. McIlroy. The distribution function for stretched dipoles in an applied electric field. Journal of Mathematical Physics, 19(6):1340–1344, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i6/p1340_s1.

vanVliet:1978:LRT

[1346] K. M. van Vliet. Linear response theory revisited. I. The many-body van Hove limit. Journal of Mathematical Physics, 19(6):1345–1370, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1345_s1.

Goldstein:1978:NGT

[1347] Gary R. Goldstein and Michael J. Moravcsik. A no-go theorem for polarization structure. Journal of Mathematical Physics, 19(6):1371–1375, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1371_s1.

Yamazaki:1978:CKT

[1348] Masatoshi Yamazaki. On the charged Kerr–Tomimatsu–Sato family of solutions. Journal of Mathematical Physics, 19(6):1376–1378, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1376_s1.

vanHaeringen:1978:SAE

[1349] H. van Haeringen. Simple analytic expressions for the Coulomb off-shell Jost functions. Journal of Mathematical Physics, 19(6):1379–1380, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1379_s1.

Cahill:1978:TC

[1350] K. Cahill and N. Stoltzfus. Topological cohesion. Journal of Mathematical Physics, 19(6):1381, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i6/p1381_s1.

Jordan:1978:SPN

[1351] Thomas F. Jordan. Simple proof of no position operator for quanta with zero mass and nonzero helicity. Journal of Mathematical Physics, 19(6): 1382–1385, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i6/p1382_s1.

Hagelstein:1978:SML

[1352] P. L. Hagelstein. Solution of a modified lamé equation with an integral term. Journal of Mathematical Physics, 19(6):1386–1390, June 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1386_
s1.

Rashid:1978:ISG

[1353] M. A. Rashid. The intelligent states. I. Group-theoretic study and the computation of matrix elements. Journal of Mathematical Physics, 19(6): 1391–1396, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i6/p1391_s1.

Rashid:1978:ISI

[1354] M. A. Rashid. The intelligent states. II. The computation of the Clebsch– Gordan coefficients. Journal of Mathematical Physics, 19(6):1397–1402, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1397_s1.

Boyd:1978:TFW

[1355] T. J. M. Boyd and J. G. Turner. Three- and four-wave interactions in plasmas. Journal of Mathematical Physics, 19(6):1403–1413, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1403_s1.

Lee:1978:PIS

[1356] S. W. Lee. Path integrals for solving some electromagnetic edge diffraction problems. Journal of Mathematical Physics, 19(6):1414–1422, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1414_s1.

Bandyopadhyay:1978:CSD

[1357] Nikhilendu Bandyopadhyay. Cylindrically symmetric dust distributions in rigid rotation in Brans–Dicke theory. Journal of Mathematical Physics, 19(6):1423–1425, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i6/p1423_s1.

Naundorf:1978:MCS

[1358] Friedrich F. Naundorf. A method for computing scattering phase shifts and eigenvalues of the Schrödinger equation with singular potentials. Journal of Mathematical Physics, 19(6):1426–1432, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1426_s1.

Kress:1978:TPH

[1359] R. Kress and G. F. Roach. Transmission problems for the Helmholtz equation. Journal of Mathematical Physics, 19(6):1433–1437, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/
p1433_s1.

Samuel:1978:FIA

[1360] Stuart Samuel. Functional integrals over anticommuting variables and the one-dimensional scattering problem. Journal of Mathematical Physics, 19(6):1438–1444, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i6/p1438_s1.

Boyd:1978:CPM

[1361] John P. Boyd. A Chebyshev polynomial method for computing analytic solutions to eigenvalue problems with application to the anharmonic oscillator. Journal of Mathematical Physics, 19(6):1445–1456, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1445_s1.

Mermin:1978:HGC

[1362] N. D. Mermin. The homotopy groups of condensed matter physics. Journal of Mathematical Physics, 19(6):1457–1462, June 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1457_s1.

Cornille:1978:IIE

[1363] Henri Cornille. An inversionlike integral equation in the multidimensional case. Journal of Mathematical Physics, 19(6):1463–1466, June 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i6/p1463_
s1.

Hosokawa:1978:ASS

[1364] I. Hosokawa. Asymptotic steady-state solution in the functional randomwalk model. Journal of Mathematical Physics, 19(6):1467–1470, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i6/p1467_s1.

Staunton:1978:LSA

[1365] L. P. Staunton. Lorentz subgroup analysis of the Lie algebra of Sp(8, R) and a null-plane, boson realization. Journal of Mathematical Physics, 19(6):1471–1476, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i6/p1471_s1.

Domokos:1978:AC

[1366] G. Domokos and S. Kövesi-Domokos. The algebra of color. Journal of Mathematical Physics, 19(6):1477–1481, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i6/p1477_s1.

Zecca:1978:CL

[1367] Antonio Zecca. On the coupling of logics. Journal of Mathematical Physics, 19(6):1482–1485, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i6/p1482_s1.

Bednar:1978:FDC

[1368] Miroslav Bednář and Vladimír Šachl. Fourth degree Casimir operator of the semisimple graded Lie algebra (Sp(2N); 2N). Journal of Mathematical Physics, 19(7):1487–1492, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i7/p1487_s1.

Jagannathan:1978:NRP

[1369] R. Jagannathan and R. Vasudevan. A note on representation of para-Fermi algebra. Journal of Mathematical Physics, 19(7):1493–1494, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i7/p1493_s1.

Bergere:1978:RPB

[1370] M. C. Bergère and C. Gilain. Regge pole behavior in ϕ^{3} field theory. Journal of Mathematical Physics, 19(7):1495–1512, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i7/p1495_s1.

Millener:1978:NRC

[1371] D. J. Millener. A note on recoupling coefficients for SU(3). Journal of Mathematical Physics, 19(7):1513–1514, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i7/p1513_s1.

Taubes:1978:SAK

[1372] Clifford Henry Taubes. Solution of the almost-Killing equation and conformal almost- Killing equation in the Kerr spacetime. Journal of Mathematical Physics, 19(7):1515–1525, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i7/p1515_s1.

Devaney:1978:NIS

[1373] A. J. Devaney. Nonuniqueness in the inverse scattering problem. Journal of Mathematical Physics, 19(7):1526–1531, July 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i7/p1526_s1.

Dupre:1978:CRV

[1374] Maurice J. Dupré and Steven I. Rosencrans. Classical and relativistic vorticity in a semi-Riemannian manifold. Journal of Mathematical Physics, 19(7):1532–1535, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i7/p1532_s1.

Spero:1978:SCF

[1375] Alan Spero and Duane A. Szafron. Spatial conformal flatness in homogeneous and inhomogeneous cosmologies. Journal of Mathematical Physics, 19(7):1536–1541, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i7/p1536_s1.

Ashtekar:1978:UTN

[1376] Abhay Ashtekar and R. O. Hansen. A unified treatment of null and spatial infinity in general relativity. I. Universal structure, asymptotic symmetries, and conserved quantities at spatial infinity. Journal of Mathematical Physics, 19(7):1542–1566, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i7/p1542_s1.

Ashtekar:1978:TAS

[1377] Abhay Ashtekar and Anne Magnon-Ashtekar. A technique for analyzing the structure of isometries. Journal of Mathematical Physics, 19(7):1567– 1572, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i7/p1567_s1.

Osborne:1978:SSG

[1378] Anthony Osborne and Allan E. G. Stuart. On the separability of the sine-Gordon equation and similar quasilinear partial differential equations. Journal of Mathematical Physics, 19(7):1573–1579, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i7/p1573_s1.

Budic:1978:EIC

[1379] Robert Budic, Mark J. Gotay, and Robert H. Gowdy. Existence of instantaneous Cauchy surfaces. Journal of Mathematical Physics, 19(7): 1580–1583, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i7/p1580_s1.

Kornacker:1978:DFN

[1380] Karl Kornacker. Discrete finite nilpotent Lie analogs: New models for unified gauge field theory. Journal of Mathematical Physics, 19(7):1584– 1586, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i7/p1584_s1. See erratum [1557].

Siewert:1978:PEE

[1381] C. E. Siewert. On a possible experiment to evaluate the validity of the one-speed or constant cross section model of the neutron-transport equation. Journal of Mathematical Physics, 19(7):1587–1588, July 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i7/p1587_
s1.

Haller:1978:LTO

[1382] K. Haller and R. B. Sohn. Lorentz transformations of observable and ghost particle states in quantum electrodynamics and in a massive gauge theory. Journal of Mathematical Physics, 19(7):1589–1593, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i7/
p1589_s1.

Hagelstein:1978:SSO

[1383] P. L. Hagelstein. Solution of a second-order integro-differential equation which occurs in laser modelocking. Journal of Mathematical Physics, 19(7):1594–1599, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i7/p1594_s1.

Druhl:1978:TCL

[1384] Kai Drühl. A theory of classical limit for quantum theories which are defined by real Lie algebras. Journal of Mathematical Physics, 19(7): 1600–1606, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i7/p1600_s1.

Xanthopoulos:1978:EVS

[1385] Basilis C. Xanthopoulos. Exact vacuum solutions of Einstein’s equation from linearised solutions. Journal of Mathematical Physics, 19(7):1607– 1609, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i7/p1607_s1.

Chandler:1978:BQS

[1386] Colston Chandler and A. G. Gibson. N-body quantum scattering theory in two Hilbert spaces. II. Some asymptotic limits. Journal of Mathematical Physics, 19(7):1610–1616, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i7/p1610_s1.

Carlson:1978:ISK

[1387] George T. Carlson, Jr. and John L. Safko. An investigation of some of the kinematical aspects of plane symmetric space–times. Journal of Mathematical Physics, 19(7):1617–1624, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i7/p1617_s1.

Chitre:1978:CCS

[1388] D. M. Chitre. Characterization of certain stationary solutions of Einstein’s equations. Journal of Mathematical Physics, 19(7):1625–1626, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i7/p1625_s1.

Rideau:1978:NRR

[1389] G. Rideau. Noncompletely reducible representations of the Poincaré group associated with the generalized Lorentz gauge. Journal of Mathematical Physics, 19(7):1627–1634, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i7/p1627_s1.

Wong:1978:MFS

[1390] M. K. F. Wong. Multiplicity-free 6-j symbols and Weyl coefficients of U(n): Explicit evaluation. Journal of Mathematical Physics, 19(7):1635– 1643, July 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i7/p1635_s1.

Huddleston:1978:IWC

[1391] P. L. Huddleston. Inönü-Wigner contractions of the real four-dimensional Lie algebras. Journal of Mathematical Physics, 19(8):1645–1649, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i8/p1645_s1.

Doria:1978:MTF

[1392] Francisco Antonio Doria and Sergio Murilo Abrahão. Mesonic test fields and spacetime cohomology. Journal of Mathematical Physics, 19(8):1650–1653, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1650_s1.

Talukdar:1978:CSW

[1393] B. Talukdar, R. N. Chaudhuri, U. Das, and P. Banerjee. Continuum S state wavefunctions for the Debye potential by Ecker–Weizel approximations. Journal of Mathematical Physics, 19(8):1654–1655, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i8/p1654_s1.

Cheng:1978:CLD

[1394] Kuo shung Cheng. Constraints of the Lorentz–Dirac equation. Journal of Mathematical Physics, 19(8):1656–1657, August 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i8/p1656_s1.

Morgan:1978:ABB

[1395] John D. Morgan III. The asymptotic behavior of bound eigenfunctions of Hamiltonians for single-variable systems. Journal of Mathematical Physics, 19(8):1658–1661, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i8/p1658_s1. See erratum [1605].

Sarma:1978:PBS

[1396] C. R. Sarma and K. V. Dinesha. Projected basis set for the irreducible representation {2^{N/2−S}, 1^{2S}} of U(n). Journal of Mathematical Physics, 19(8):1662–1666, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1662_s1.

Basu:1978:LGO

[1397] Debabrata Basu. The Lorentz group in the oscillator realization. I. The group SO(2, 1) and the transformation matrices connecting the SO(2) and SO(1, 1) bases. Journal of Mathematical Physics, 19(8):1667– 1670, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i8/p1667_s1.

Yasue:1978:SDO

[1398] Kunio Yasue. A simple derivation of the Onsager–Machlup formula for one-dimensional nonlinear diffusion process. Journal of Mathematical Physics, 19(8):1671–1673, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v19/i8/p1671_s1.

Sullivan:1978:RCS

[1399] John J. Sullivan. Recoupling coefficients of the symmetric group involving outer plethysms. Journal of Mathematical Physics, 19(8):1674– 1680, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i8/p1674_s1.

Sullivan:1978:RCG

[1400] John J. Sullivan. Recoupling coefficients of the general linear group in bases adapted to shell theories. Journal of Mathematical Physics, 19(8):1681–1687, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1681_s1.

Saenz:1978:UEM

[1401] A. W. Sáenz and W. W. Zachary. Unitarily equivalent multiparticle Hamiltonian systems yielding equal scattering for corresponding states. Journal of Mathematical Physics, 19(8):1688–1698, August 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i8/p1688_
s1.

Babbitt:1978:GSR

[1402] Donald Babbitt and Lawrence Thomas. Ground state representation of the infinite one-dimensional Heisenberg ferromagnet. III. Scattering theory. Journal of Mathematical Physics, 19(8):1699–1704, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i8/
p1699_s1.

Srinivas:1978:CPS

[1403] M. D. Srinivas. Conditional probabilities and statistical independence in quantum theory. Journal of Mathematical Physics, 19(8):1705–1710, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i8/p1705_s1.

Dhar:1978:EPT

[1404] Deepak Dhar. Entropy and phase transitions in partially ordered sets. Journal of Mathematical Physics, 19(8):1711–1713, August 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i8/p1711_
s1.

CastilhoAlcaras:1978:HAH

[1405] J. A. Castilho Alcarás. On the higher approximations of the K-harmonics method. Journal of Mathematical Physics, 19(8):1714–1719, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i8/p1714_s1.

Sattinger:1978:BRI

[1406] D. H. Sattinger. Bifurcation from rotationally invariant states. Journal of Mathematical Physics, 19(8):1720–1732, August 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i8/p1720_s1.

Ladde:1978:SDI

[1407] G. S. Ladde. Systems of differential inequalities and stochastic differential equations. IV. Journal of Mathematical Physics, 19(8):1733–1741, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i8/p1733_s1.

Truman:1978:FMW

[1408] Aubrey Truman. The Feynman maps and the Wiener integral. Journal of Mathematical Physics, 19(8):1742–1750, August 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i8/p1742_s1. See er-

ratum [1814].

Ekhaguere:1978:TSR

[1409] G. O. S. Ekhaguere. The theory of superselection rules. I. A class of inequivalent, irreducible ^{∗}-representations of the canonical commutation relations of the free electromagnetic field. Journal of Mathematical Physics, 19(8):1751–1757, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i8/p1751_s1.

Burdet:1978:OGS

[1410] G. Burdet, J. Patera, M. Perrin, and P. Winternitz. The optical group and its subgroups. Journal of Mathematical Physics, 19(8):1758– 1780, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i8/p1758_s1.

Castanos:1978:GFC

[1411] O. Castaños, A. Frank, and M. Moshinsky. The gradient formula for the O(5) ⊆ O(3) chain of groups. Journal of Mathematical Physics, 19(8):1781–1789, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1781_s1.

Elizalde:1978:GPG

[1412] E. Elizalde and J. Gomis. The groups of Poincaré and Galilei in arbitrary dimensional spaces. Journal of Mathematical Physics, 19(8):1790– 1792, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i8/p1790_s1.

Kanal:1978:DIPa

[1413] Madhoo Kanal and Harry E. Moses. Direct-inverse problems in transport theory. 1. The inverse problem. Journal of Mathematical Physics, 19(8):1793–1798, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1793_s1.

Ahner:1978:CCI

[1414] Henry F. Ahner and Allan E. Moose. Corrigendum to “Covariant inverse problem of the calculus of variations”. Journal of Mathematical Physics, 19(8):1799–1800, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1799_s1. See [953].

Sorkin:1978:EEF

[1415] Rafael Sorkin. Erratum: “The electromagnetic field on a simplicial net”. Journal of Mathematical Physics, 19(8):1800, August 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i8/p1800_

s1. See [389].

Verri:1978:QDS

[1416] Maurizio Verri and Vittorio Gorini. Quantum dynamical semigroups and multipole relaxation of a spin in isotropic surroundings. Journal of Mathematical Physics, 19(9):1803–1807, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1803_s1.

Tanabe:1978:GSA

[1417] Yukio Tanabe. Generation of stationary axisymmetric Einstein–Maxwell fields. Journal of Mathematical Physics, 19(9):1808–1811, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p1808_s1.

Fisher:1978:TRG

[1418] Kurt Fisher and Henry F. Ahner. Theorem on the representations of SO(n) groups. Journal of Mathematical Physics, 19(9):1812–1815, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i9/p1812_s1.

Ernst:1978:FEI

[1419] F. J. Ernst and I. Hauser. Field equations and integrability conditions for special type N twisting gravitational fields. Journal of Mathematical Physics, 19(9):1816–1822, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i9/p1816_s1.

Froman:1978:AGQ

[1420] Nanny Fröman and Per Olof Fröman. On the application of the generalized quantal Bohr–Sommerfeld quantization condition to single-well potentials with very steep walls. Journal of Mathematical Physics, 19 (9):1823–1829, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1823_s1.

Froman:1978:PIQ

[1421] Nanny Fröman and Per Olof Fröman. On phase-integral quantization conditions for bound states in one-dimensional smooth single-well potentials. Journal of Mathematical Physics, 19(9):1830–1837, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p1830_s1.

Cohen:1978:TVT

[1422] Leon Cohen. The tensor virial theorem in quantum mechanics. Journal of Mathematical Physics, 19(9):1838–1840, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1838_s1.

Marlow:1978:QTH

[1423] A. R. Marlow. Quantum theory and Hilbert space. Journal of Mathematical Physics, 19(9):1841–1846, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i9/p1841_s1.

Yamazaki:1978:KTS

[1424] Masatoshi Yamazaki. On the Kerr–Tomimatsu–Sato family of solutions with nonintegral distortion parameter. Journal of Mathematical Physics, 19(9):1847–1849, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1847_s1.

Onofri:1978:SPH

[1425] E. Onofri and M. Pauri. Search for periodic Hamiltonian flows: A generalized Bertrand’s theorem. Journal of Mathematical Physics, 19 (9):1850–1858, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1850_s1.

Lowenstein:1978:LCF

[1426] John H. Lowenstein and Eugene R. Speer. Light-cone finite normal products. Journal of Mathematical Physics, 19(9):1859–1865, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p1859_s1.

Gersten:1978:TDG

[1427] Joel I. Gersten. A time dependent generalization of the multiple scattering formalism. Journal of Mathematical Physics, 19(9):1866–1868, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i9/p1866_s1.

Aldaya:1978:VPO

[1428] Víctor Aldaya and José A. de Azcárraga. Variational principles on rth order jets of fibre bundles in field theory. Journal of Mathematical Physics, 19(9):1869–1875, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1869_s1.

Aldaya:1978:VBO

[1429] V. Aldaya and J. A. de Azcárraga. Vector bundles, rth order Noether invariants and canonical symmetries in Lagrangian field theory. Journal of Mathematical Physics, 19(9):1876–1880, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1876_s1.

Arnal:1978:NSS

[1430] D. Arnal. Nuclearity of some spaces of C^{∞} vectors in induced representations. Journal of Mathematical Physics, 19(9):1881–1886, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p1881_s1.

Martin:1978:SMA

[1431] J. Martin and J. L. Sanz. Slow motion approximation in predictive relativistic mechanics. I. Approximated dynamics up to order c^{−4}. Journal of Mathematical Physics, 19(9):1887–1891, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1887_s1.

Yasue:1978:SQW

[1432] Kunio Yasue. Stochastic quantization of wave fields and its application to dissipatively interacting fields. Journal of Mathematical Physics, 19 (9):1892–1897, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1892_s1.

Letelier:1978:UTD

[1433] Patricio S. Letelier. Unified theory of direct interaction between particles, strings, and membranes. Journal of Mathematical Physics, 19(9):1898– 1908, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i9/p1898_s1.

Bencze:1978:GAT

[1434] Gy. Bencze and Edward F. Redish. General algebraic theory of identical particle scattering. Journal of Mathematical Physics, 19(9):1909–1917, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i9/p1909_s1.

Katzin:1978:EMT

[1435] Gerald H. Katzin and Jack Levine. Energy-momentum tensor symmetries and concomitant conservation laws. I. Einstein-massless-scalar (meson) field. Journal of Mathematical Physics, 19(9):1918–1925, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p1918_s1.

Kinnersley:1978:SSEa

[1436] William Kinnersley and D. M. Chitre. Symmetries of the stationary Einstein–Maxwell field equations. III. Journal of Mathematical Physics, 19(9):1926–1931, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1926_s1.

Underhill:1978:QMC

[1437] J. Underhill. Quantization on a manifold with connection. Journal of Mathematical Physics, 19(9):1932–1935, September 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1932_s1.

Cornille:1978:GII

[1438] Henri Cornille. Generalization of the inversionlike integral equations and application to nonlinear partial differential equations. II. Journal of Mathematical Physics, 19(9):1936–1945, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1936_s1.

Dewar:1978:RPW

[1439] R. L. Dewar. Random phase wave: a soluble non-Markovian system. Journal of Mathematical Physics, 19(9):1946–1951, September 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1946_
s1.

Srinivas:1978:QGK

[1440] M. D. Srinivas. Quantum generalization of Kolmogorov entropy. Journal of Mathematical Physics, 19(9):1952–1961, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1952_s1.

Rowe:1978:SDF

[1441] E. G. Peter Rowe. Spherical delta functions and multipole expansions. Journal of Mathematical Physics, 19(9):1962–1968, September 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1962_
s1.

Hall:1978:LEB

[1442] Richard L. Hall. Lower energy bounds for translation-invariant atomlike systems. Journal of Mathematical Physics, 19(9):1969–1972, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p1969_s1.

Venkatesh:1978:SCa

[1443] K. Venkatesh. Symmetries of the 6j coefficient. Journal of Mathematical Physics, 19(9):1973–1974, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i9/p1973_s1.

Davidson:1978:EQM

[1444] Mark Davidson. On the equivalence of quantum mechanics and a certain class of Markov processes. Journal of Mathematical Physics, 19(9):1975– 1978, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i9/p1975_s1.

Wierman:1978:CPP

[1445] John C. Wierman. On critical probabilities in percolation theory. Journal of Mathematical Physics, 19(9):1979–1982, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1979_s1.

Radin:1978:PET

[1446] Charles Radin. Pointwise ergodic theory on operator algebras. Journal of Mathematical Physics, 19(9):1983–1985, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1983_s1.

Ernst:1978:GM

[1447] Frederick J. Ernst. Generalized C-metric. Journal of Mathematical Physics, 19(9):1986–1987, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i9/p1986_s1. See erratum [1888].

Chapman:1978:IVM

[1448] David M. F. Chapman. Ideal vortex motion in two dimensions: Symmetries and conservation laws. Journal of Mathematical Physics, 19 (9):1988–1992, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i9/p1988_s1.

Fox:1978:HFT

[1449] Ronald Forrest Fox. Hydrodynamic fluctuation theories. Journal of Mathematical Physics, 19(9):1993–1999, September 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i9/p1993_s1.

Srivastava:1978:TST

[1450] Sushil K. Srivastava and Mahendra P. Pathak. A time-symmetric tachyon universe. Journal of Mathematical Physics, 19(9):2000–2003, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i9/p2000_s1.

Blume:1978:CQS

[1451] M. Blume. Comments on: “Quantum system subject to random pulses”, [J. Math. Phys. 18 (1977), no. 7, 1327–1329; MR 81e:81038b]. Journal of Mathematical Physics, 19(9):2004, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i9/p2004_s1. See [947].

Benza:1978:ELT

[1452] V. Benza and E. Montaldi. Erratum: “Large time behavior of the superfluorescent decay”. Journal of Mathematical Physics, 19(9):2005, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i9/p2005_s2. See [1225].

Taylor:1978:EIG

[1453] William J. Taylor. Erratum: “Inverse Gaussian transforms: General properties and application to Slater-type orbitals with noninteger and integer n in the coordinate and momentum representations”. Journal of Mathematical Physics, 19(9):2005, September 1978. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i9/p2005_s1. See [1135].

Gruber:1978:MFF

[1454] Bruno Gruber and A. U. Klimyk. Multiplicity free and finite multiplicity indecomposable representations of the algebra su(1,1). Journal of Mathematical Physics, 19(10):2009–2017, October 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2009_s1.

Mitra:1978:IT

[1455] A. K. Mitra. On the interaction of the type λx^{2}/(1 + gx^{2}). Journal of Mathematical Physics, 19(10):2018–2022, October 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2018_s1.

March:1978:AFF

[1456] N. H. March. Average force and force correlation formulas for momentum transfer cross section. Journal of Mathematical Physics, 19(10): 2023, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i10/p2023_s1.

Curtis:1978:CWN

[1457] W. D. Curtis, D. E. Lerner, and F. R. Miller. Complex pp waves and the nonlinear graviton construction. Journal of Mathematical Physics, 19(10):2024–2027, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2024_s1.

Maekawa:1978:CFR

[1458] Takayoshi Maekawa. On the computation formulas of the SO(n − 1, 1) representation matrix elements. Journal of Mathematical Physics, 19 (10):2028–2036, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2028_s1.

Kinnersley:1978:SSEb

[1459] William Kinnersley and D. M. Chitre. Symmetries of the stationary Einstein–Maxwell equations. IV. Transformations which preserve asymptotic flatness. Journal of Mathematical Physics, 19(10):2037–2042, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2037_s1.

Creutz:1978:II

[1460] Michael Creutz. On invariant integration over SU(N). Journal of Mathematical Physics, 19(10):2043–2046, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i10/p2043_s1.

Catenacci:1978:ASS

[1461] Roberto Catenacci and Franco Salmistraro. Almost-structures and structures in Lorentzian manifolds. I. Almost-Hermite- and almost-product- (2 × 2)-structures. Journal of Mathematical Physics, 19(10):2047–2051, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i10/p2047_s1.

Harris:1978:IDL

[1462] R. A. Harris and J. D. Zund. An investigation of Dubourdieu’s list of space–times which admit holonomy groups. Journal of Mathematical Physics, 19(10):2052–2054, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2052_s1.

Lee:1978:CCI

[1463] L. L. Lee. Continuum calculus. III. Skorohod’s weak distributions in the evaluation of a class of Feynman path integrals. Journal of Mathematical Physics, 19(10):2055–2059, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2055_s1.

Venkatesh:1978:SCb

[1464] K. Venkatesh. Symmetries of the 3j coefficient. Journal of Mathematical Physics, 19(10):2060–2063, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2060_s1.

Brydges:1978:NFM

[1465] David Brydges and Paul Federbush. A new form of the Mayer expansion in classical statistical mechanics. Journal of Mathematical Physics, 19(10):2064–2067, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2064_s1.

Hughes:1978:SOG

[1466] J. W. B. Hughes and J. Yadegar. O(3) shift operators: the general analysis. Journal of Mathematical Physics, 19(10):2068–2078, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i10/
p2068_s1.

Ullah:1978:PEH

[1467] Nazakat Ullah. On the polynomial expansion of the Hill–Wheeler integral for the rotational energies. Journal of Mathematical Physics, 19 (10):2079–2081, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2079_s1.

Mitchell:1978:VCT

[1468] Thomas P. Mitchell. Vector constants and their algebras for classical Hamiltonians. Journal of Mathematical Physics, 19(10):2082–2084, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2082_s1.

Smrz:1978:RDL

[1469] P. K. Smrz. Relativity and deformed Lie groups. Journal of Mathematical Physics, 19(10):2085–2088, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2085_s1.

Sharma:1978:PBC

[1470] J. K. Sharma, C. L. Mehta, and E. C. G. Sudarshan. Para-Bose coherent states. Journal of Mathematical Physics, 19(10):2089–2093, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2089_s1.

Hostler:1978:CLQ

[1471] Levere Hostler. Coulomb-like quantization of the electromagnetic field on spacelike hyperboloids. Journal of Mathematical Physics, 19(10):2094– 2099, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i10/p2094_s1.

Portinari:1978:IST

[1472] João C. Portinari. An inverse scattering transform for potentials of compact support. Journal of Mathematical Physics, 19(10):2100–2102, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2100_s1.

Dixon:1978:SCA

[1473] Geoffrey Dixon. Super Clifford algebra. Journal of Mathematical Physics, 19(10):2103–2107, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2103_s1.

Angelopoulos:1978:LRB

[1474] E. Angelopoulos. The linear representations of \overline{SL}(3, R) on a Banach space. Journal of Mathematical Physics, 19(10):2108–2120, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2108_s1.

Leitman:1978:SDD

[1475] M. Leitman, N. Schwartz, and M. J. Beran. Scattering in the depth direction for an anisotropic random medium. Journal of Mathematical Physics, 19(10):2121–2123, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2121_s1.

Sakmar:1978:CUS

[1476] I. A. Sakmar. Conditions on the uniqueness of the solution of the elastic unitarity equation. Journal of Mathematical Physics, 19(10):2124– 2125, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i10/p2124_s1.

Beckers:1978:TFI

[1477] J. Beckers, J. Harnad, M. Perroud, and P. Winternitz. Tensor fields invariant under subgroups of the conformal group of space–time. Journal of Mathematical Physics, 19(10):2126–2153, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2126_s1.

McNamara:1978:SCA

[1478] Brendan McNamara. Super-convergent adiabatic invariants with resonant denominators by Lie transforms. Journal of Mathematical Physics, 19(10):2154–2164, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2154_s1.

vanHaeringen:1978:BSS

[1479] H. van Haeringen. The bound states for the symmetric shifted Coulomb potential. Journal of Mathematical Physics, 19(10):2165–2170, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2165_s1.

vanHaeringen:1978:BSL

[1480] H. van Haeringen. Bound states for r^{−2}-like potentials in one and three dimensions. Journal of Mathematical Physics, 19(10):2171–2179, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i10/p2171_s1.

Ablowitz:1978:SRS

[1481] M. J. Ablowitz and J. Satsuma. Solitons and rational solutions of nonlinear evolution equations. Journal of Mathematical Physics, 19(10):2180– 2186, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i10/p2180_s1.

Majumdar:1978:STM

[1482] Chanchal K. Majumdar and Indrani Bose. Solutions of the three-magnon bound state equation. III. The physical eigenstate. Journal of Mathematical Physics, 19(10):2187–2192, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2187_s1.

Rittenberg:1978:SGL

[1483] V. Rittenberg and D. Wyler. Sequences of Z_{2} ⊗ Z_{2} graded Lie algebras and superalgebras. Journal of Mathematical Physics, 19(10):2193– 2200, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i10/p2193_s1.

Liu:1978:IWT

[1484] Kuang Chi Liu. On the inversion of the Weyl transform. Journal of Mathematical Physics, 19(10):2201–2202, October 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2201_s1.

Kannenberg:1978:KVG

[1485] L. Kannenberg. Killing vectors in gauge supersymmetry. Journal of Mathematical Physics, 19(10):2203–2206, October 1978. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2203_s1.

Davies:1978:GED

[1486] K. T. R. Davies and Richard L. Becker. General expressions for divergence relations and multipole expansions for arbitrary scalar functions. Journal of Mathematical Physics, 19(10):2207–2215, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i10/
p2207_s1.

Ashtekar:1978:ICA

[1487] Abhay Ashtekar and Basilis C. Xanthopoulos. Isometries compatible with asymptotic flatness at null infinity: a complete description. Journal of Mathematical Physics, 19(10):2216–2222, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2216_s1.

vanOost:1978:GRM

[1488] E. van Oost and H. N. W. Lekkerkerker. Geometry of the roots of matrices with Onsager–Casimir symmetry. Journal of Mathematical Physics, 19(10):2223–2228, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i10/p2223_s1.

Schaposnik:1978:FBA

[1489] F. A. Schaposnik and J. E. Solomin. Fibre bundle analysis of topological charges in spontaneously broken gauge theories. Journal of Mathematical Physics, 19(10):2229–2231, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i10/p2229_s1.

Uzes:1978:MRI

[1490] C. A. Uzes. Mechanical response and the initial value problem. Journal of Mathematical Physics, 19(10):2232–2238, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i10/p2232_s1.

Phares:1978:EEE

[1491] Alain J. Phares. The energy eigenvalue equation for the linear potential. Journal of Mathematical Physics, 19(10):2239–2244, October 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i10/
p2239_s1.

Jha:1978:GCE

[1492] Dhirendra Kumar Jha. A general commutator equation in the algebra of spin operators. Journal of Mathematical Physics, 19(11):2247–2252, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2247_s1.

Okolowski:1978:NSF

[1493] J. A. Okolowski. A nonlinear scalar field theory in isotropic homogeneous space–time. Journal of Mathematical Physics, 19(11):2253–2255, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2253_s1. See erratum [1815].

Demoen:1978:ESS

[1494] Bart Demoen, Paul Vanheuverzwijn, and André Verbeure. Energetically stable systems. Journal of Mathematical Physics, 19(11):2256–2259, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2256_s1.

Prugovecki:1978:RQK

[1495] Eduard Prugovečki. Relativistic quantum kinematics on stochastic phase spaces for massive particles. Journal of Mathematical Physics, 19(11):2260–2270, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i11/p2260_s1.

Prugovecki:1978:SPS

[1496] Eduard Prugovečki. Stochastic phase space kinematics of the photon. Journal of Mathematical Physics, 19(11):2271–2277, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i11/
p2271_s1.

Snellman:1978:RWL

[1497] Håkan Snellman. Representations of the Weyl Lie algebra as models of elementary particles. Journal of Mathematical Physics, 19(11):2278– 2282, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2278_s1.

Wesson:1978:ESE

[1498] Paul S. Wesson. An exact solution to Einstein’s equations with a stiff equation of state. Journal of Mathematical Physics, 19(11):2283–2284, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2283_s1.

McGuinness:1978:CDK

[1499] Mark J. McGuinness. The conserved densities of the Korteweg–de Vries equation. Journal of Mathematical Physics, 19(11):2285–2288, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i11/p2285_s1.

Lapedes:1978:BTP

[1500] A. S. Lapedes. Bogoliubov transformations, propagators, and the Hawking effect. Journal of Mathematical Physics, 19(11):2289–2293, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i11/p2289_s1.

Gowdy:1978:GSP

[1501] Robert H. Gowdy. Geometrical spacetime perturbation theory: Regular first-order structures. Journal of Mathematical Physics, 19(11):2294– 2304, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2294_s1.

Chinea:1978:CHS

[1502] F. J. Chinea. Causality in homogeneous spaces. Journal of Mathematical Physics, 19(11):2305–2307, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i11/p2305_s1.

Boon:1978:CSL

[1503] M. Boon and J. Zak. Coherent states and lattice sums. Journal of Mathematical Physics, 19(11):2308–2311, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i11/p2308_s1.

Shadwick:1978:BPE

[1504] W. F. Shadwick. The Bäcklund problem for the equation ∂^{2}z/∂x^{1}∂x^{2} = f(z). Journal of Mathematical Physics, 19(11):2312–2317, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i11/p2312_s1.

Inomata:1978:PIP

[1505] Akira Inomata and Vijay A. Singh. Path integrals with a periodic constraint: Entangled strings. Journal of Mathematical Physics, 19 (11):2318–2323, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i11/p2318_s1.

Palmer:1978:CEG

[1506] T. N. Palmer. Conservation equations and the gravitational symplectic form. Journal of Mathematical Physics, 19(11):2324–2331, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i11/p2324_s1.

Farrukh:1978:CQF

[1507] M. O. Farrukh. Constructing quantum fields in a Fock space using a new picture of quantum mechanics. Journal of Mathematical Physics, 19(11):2332–2340, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i11/p2332_s1.

Deliyannis:1978:CSI

[1508] Platon C. Deliyannis. Conditioning of states. II. Journal of Mathematical Physics, 19(11):2341–2345, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i11/p2341_s1.

Deans:1978:URI

[1509] Stanley R. Deans. A unified Radon inversion formula. Journal of Mathematical Physics, 19(11):2346–2349, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i11/p2346_s1.

Plebanski:1978:EGH

[1510] Jerzy F. Plebański and Ivor Robinson. Electromagnetic and gravitational Hertz potentials. Journal of Mathematical Physics, 19(11):2350– 2358, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2350_s1.

Constantinescu:1978:QFT

[1511] Florin Constantinescu and Herbert M. Ruck. Quantum field theory Potts model. Journal of Mathematical Physics, 19(11):2359–2361, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i11/p2359_s1.

Patera:1978:PIT

[1512] J. Patera, R. T. Sharp, and P. Winternitz. Polynomial irreducible tensors for point groups. Journal of Mathematical Physics, 19(11):2362–2376, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i11/p2362_s1.

Baumeister:1978:GHJ

[1513] Richard Baumeister. Generalized Hamilton–Jacobi theories. Journal of Mathematical Physics, 19(11):2377–2387, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i11/p2377_s1.

Gotay:1978:PMD

[1514] Mark J. Gotay, James M. Nester, and George Hinds. Presymplectic manifolds and the Dirac–Bergmann theory of constraints. Journal of Mathematical Physics, 19(11):2388–2399, November 1978. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i11/p2388_s1.

Backhouse:1978:CFD

[1515] Nigel Backhouse. A classification of four-dimensional Lie superalgebras. Journal of Mathematical Physics, 19(11):2400–2402, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i11/p2400_s1.

Garibotti:1978:PPA

[1516] C. R. Garibotti and F. F. Grinstein. Punctual Padé approximants as a regularization procedure for divergent and oscillatory partial wave expansions of the scattering amplitude. Journal of Mathematical Physics, 19(12):2405–2409, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i12/p2405_s1.

Sabatier:1978:SSI

[1517] Pierre C. Sabatier. Spectral and scattering inverse problems. Journal of Mathematical Physics, 19(12):2410–2425, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2410_s1.

Zorbas:1978:RPW

[1518] J. Zorbas. Renormalized plane wave-projected and Coulomb-projected T - matrices. Journal of Mathematical Physics, 19(12):2426–2430, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2426_s1.

Corones:1978:ILG

[1519] James Corones. An illustration of the Lie group framework for soliton equations: Generalizations of the Lund–Regge model. Journal of Mathematical Physics, 19(12):2431–2436, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i12/p2431_s1.

Nikolic:1978:TRN

[1520] Zlatiborka A. Nikolić. Time-reversal noninvariance of the quantummechanical kinetic equation of Kadanoff and Baym. Journal of Mathematical Physics, 19(12):2437–2440, December 1978. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i12/p2437_s1.

Sandberg:1978:TSH

[1521] Vernon D. Sandberg. Tensor spherical harmonics on S^{2} and S^{3} as eigenvalue problems. Journal of Mathematical Physics, 19(12):2441–2446, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i12/p2441_s1.

dInverno:1978:CTS

[1522] R. A. d’Inverno and John Stachel. Conformal two-structure as the gravitational degrees of freedom in general relativity. Journal of Mathematical Physics, 19(12):2447–2460, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2447_s1.

Incoul:1978:ELS

[1523] R. Incoul. Effective Lagrangian in spontaneously broken gauge theories. Journal of Mathematical Physics, 19(12):2461–2468, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2461_s1.

Schlup:1978:RJE

[1524] W. A. Schlup. The retarded Josephson equation. Journal of Mathematical Physics, 19(12):2469–2475, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i12/p2469_s1.

Berryman:1978:ESP

[1525] James G. Berryman and Charles J. Holland. Evolution of a stable profile for a class of nonlinear diffusion equations. II. Journal of Mathematical Physics, 19(12):2476–2480, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2476_s1.

Banerjee:1978:ESS

[1526] S. N. Banerjee and S. N. Chakravorty. On the exact scattering solution of the Schrödinger and Dirac equations with a short range potential. Journal of Mathematical Physics, 19(12):2481–2484, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2481_s1.

Sud:1978:ACL

[1527] A. R. Sud and K. K. Sud. Analytic continuations of the Lauricella function. Journal of Mathematical Physics, 19(12):2485–2490, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2485_s1.

Maciejko:1978:YME

[1528] R. Maciejko. Yang–Mills equations in Maxwell form. Journal of Mathematical Physics, 19(12):2491–2493, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i12/p2491_s1.

Moreau:1978:NEP

[1529] M. Moreau. Note on the entropy production in a discrete Markov system. Journal of Mathematical Physics, 19(12):2494–2498, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2494_s1.

Pao:1978:CNT

[1530] C. V. Pao. Criticality of neutron transport in a slab with finite reflectors. Journal of Mathematical Physics, 19(12):2499–2505, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2499_s1.

Shukla:1978:ESI

[1531] P. K. Shukla and M. Y. Yu. Exact solitary ion acoustic waves in a magnetoplasma. Journal of Mathematical Physics, 19(12):2506–2508, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i12/p2506_s1.

Handelman:1978:GED

[1532] Michael Handelman. A generalized eigenvalue distribution. Journal of Mathematical Physics, 19(12):2509–2513, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2509_s1.

Tuong:1978:EDT

[1533] Ton That Tuong. Explicit decomposition of tensor products of certain analytic representations of symplectic groups. Journal of Mathematical Physics, 19(12):2514–2519, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2514_s1.

Bracken:1978:PFQ

[1534] A. J. Bracken. The point form of quantum dynamics and a 4-vector coordinate operator for a spinless particle. Journal of Mathematical Physics, 19(12):2520–2527, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i12/p2520_s1.

Brihaye:1978:PSD

[1535] Y. Brihaye, D. B. Fairlie, J. Nuyts, and R. G. Yates. Properties of the self-dual equations for an SU(n) gauge theory. Journal of Mathematical Physics, 19(12):2528–2532, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2528_s1.

Besieris:1978:BPF

[1536] Ioannis M. Besieris. Beam propagation in focusing media with randomaxis misalignments: Second- and higher-order moments. Journal of Mathematical Physics, 19(12):2533–2545, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2533_s1.

Ekong:1978:SRA

[1537] Samuel D. Ekong. Some results on automorphisms of affine algebraic groups. Journal of Mathematical Physics, 19(12):2546–2554, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2546_s1.

Bose:1978:GCA

[1538] S. K. Bose and D. N. Tripathy. The general coupled anharmonic oscillators and the coherent state representation. Journal of Mathematical Physics, 19(12):2555–2557, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2555_s1.

Bayman:1978:GSH

[1539] B. F. Bayman. A generalization of the spherical harmonic gradient formula. Journal of Mathematical Physics, 19(12):2558–2562, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2558_s1.

Sengupta:1978:ANS

[1540] A. Sengupta and H. Karnick. Analysis of the neutron slowing down equation. Journal of Mathematical Physics, 19(12):2563–2569, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2563_s1.

Anderson:1978:SDF

[1541] I. M. Anderson. On the structure of divergence-free tensors. Journal of Mathematical Physics, 19(12):2570–2575, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2570_s1.

Twersky:1978:CCD

[1542] Victor Twersky. Constraint on the compound depolarization factor of aligned ellipsoids. Journal of Mathematical Physics, 19(12):2576–2578, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i12/p2576_s1.

Bender:1978:AGC

[1543] Carl M. Bender and William E. Caswell. Asymptotic graph counting techniques in Ψ^{2N} field theory. Journal of Mathematical Physics, 19 (12):2579–2586, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i12/p2579_s1.

Lax:1978:ASM

[1544] M. Lax. Analytic solution of a model pulse propagation problem. Journal of Mathematical Physics, 19(12):2587–2590, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2587_s1.

Gibbs:1978:TEA

[1545] Alan G. Gibbs and Richard Seto. The transport equation with anisotropic scattering in finite slab geometry. Journal of Mathematical Physics, 19(12):2591–2604, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2591_s1.

Girardeau:1978:SQR

[1546] M. D. Girardeau. Second-quantization representation for a nonrelativistic system of composite particles. II. Kinematical properties of the multispecies Tani transformation. Journal of Mathematical Physics, 19 (12):2605–2613, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i12/p2605_s1.

Anandan:1978:PBY

[1547] J. Anandan and R. Roskies. Polynomial basis for Yang–Mills invariants. Journal of Mathematical Physics, 19(12):2614–2618, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2614_s1.

Siewert:1978:NAI

[1548] C. E. Siewert. A new approach to the inverse problem. Journal of Mathematical Physics, 19(12):2619–2621, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i12/p2619_s1.

Yang:1978:MH

[1549] Chen Ning Yang. SU_{2} monopole harmonics. Journal of Mathematical Physics, 19(12):2622–2627, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i12/p2622_s1.

Cazzola:1978:NSA

[1550] P. Cazzola and L. Lucaroni. Neutron stars: Analytical continuation of the ingoing gravitational wave amplitudes and asymptotic eigenfrequencies distribution for even modes. Journal of Mathematical Physics, 19 (12):2628–2633, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i12/p2628_s1.

Rao:1978:SSI

[1551] A. N. V. Rao, R. A. Tourgee, and C. P. Tsokos. Stochastic systems and integral inequalities. Journal of Mathematical Physics, 19(12):2634– 2640, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i12/p2634_s1.

Kanal:1978:DIPb

[1552] Madhoo Kanal and Harry E. Moses. Direct-inverse problems in transport theory, the inverse albedo problem for a finite medium. Journal of Mathematical Physics, 19(12):2641–2645, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2641_s1.

Saraswati:1978:STRa

[1553] Dipak K. Saraswati and H. Primas. A system theoretic representation of mechanical systems. I. Decomposition of a mechanical system into a hierarchy of orthogonal stationary linear dynamical systems. Journal of Mathematical Physics, 19(12):2646–2654, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2646_s1.

Saraswati:1978:STRb

[1554] Dipak K. Saraswati and H. Primas. A system theoretic representation of mechanical systems. II. Stochastic interpretation. Journal of Mathematical Physics, 19(12):2655–2658, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v19/i12/p2655_s1.

Lucquiaud:1978:RBE

[1555] J. C. Lucquiaud. On the relativistic Boltzmann equation. Journal of Mathematical Physics, 19(12):2659–2665, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v19/i12/p2659_s1.

Schindler:1978:ACG

[1556] Susan Schindler and R. Mirman. Addendum: “The Clebsch–Gordan coefficients of S_{n}”. Journal of Mathematical Physics, 19(12):2665, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i12/p2665_s1. See [1002].

Kornacker:1978:EDF

[1557] Karl Kornacker. Erratum: “Discrete finite nilpotent Lie analogs: new models for unified gauge field theory”, [J. Math. Phys. 19 (1978), no. 7, 1584–1586; MR 58 #14473]. Journal of Mathematical Physics, 19(12): 2666, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i12/p2666_s1. See [1380].

Dyer:1979:GHK

[1558] C. C. Dyer and E. Honig. Geometry of homothetic Killing trajectories and stationary limit surfaces. Journal of Mathematical Physics, 20(1): 1–5, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i1/p1_s1.

Zorbas:1979:IM

[1559] John Zorbas. The ionization T matrix. Journal of Mathematical Physics, 20(1):6–10, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p6_s1.

Sakmar:1979:ADL

[1560] I. A. Sakmar. An angle-dependent lower bound on the solution of the elastic unitarity integral and a new uniqueness condition resulting from it. Journal of Mathematical Physics, 20(1):11–12, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/p11_s1.

Poenaru:1979:TSG

[1561] V. Poénaru and G. Toulouse. Topological solitons and graded Lie algebras. Journal of Mathematical Physics, 20(1):13–19, January 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/p13_s1.

Nouri-Moghadam:1979:AAC

[1562] M. Nouri-Moghadam and Tetz Yoshimura. Application of Altman’s contractor techniques to nonlinear integral equations for Green’s functions of augmented quantum field theory. Journal of Mathematical Physics, 20(1):20–22, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p20_s1.

Reddy:1979:SSS

[1563] D. R. K. Reddy. Spherically symmetric static conformally flat solutions in Brans–Dicke and Sen–Dunn theories of gravitation. Journal of Mathematical Physics, 20(1):23–24, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i1/p23_s1.

Martin:1979:SMA

[1564] J. Martin and J. L. Sanz. Slow motion approximation in predictive relativistic mechanics. II. A noninteraction theorem for interactions derived from the classical field theory. Journal of Mathematical Physics, 20(1): 25–34, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i1/p25_s1.

Castanos:1979:GTI

[1565] O. Castaños, E. Chacón, A. Frank, and M. Moshinsky. Group theory of the interacting boson model of the nucleus. Journal of Mathematical Physics, 20(1):35–44, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p35_s1.

Wodkiewicz:1979:ESS

[1566] K. Wódkiewicz. Exact solutions of some multiplicative stochastic processes. Journal of Mathematical Physics, 20(1):45–48, January 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/p45_s1.

See [1939].

Horvathy:1979:VFS

[1567] P. Horváthy. Variational formalism for spinning particles. Journal of Mathematical Physics, 20(1):49–52, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i1/p49_s1.

Weston:1979:NAI

[1568] V. H. Weston. Nonlinear approach to inverse scattering. Journal of Mathematical Physics, 20(1):53–59, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i1/p53_s1.

Adler:1979:CIS

[1569] M. Adler. Completely integrable systems and symplectic actions. Journal of Mathematical Physics, 20(1):60–67, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i1/p60_s1.

Lloyd-Evans:1979:FQB

[1570] D. J. R. Lloyd-Evans. Formal quark binding and geometric strings. Journal of Mathematical Physics, 20(1):68–73, January 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i1/p68_s1.

Desmier:1979:PTD

[1571] P. E. Desmier and R. T. Sharp. Polynomial tensors for double point groups. Journal of Mathematical Physics, 20(1):74–82, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/
p74_s1.

Floyd:1979:HOM

[1572] Edward R. Floyd. Higher order modified potentials for the effective phase integral approximation. Journal of Mathematical Physics, 20(1): 83–85, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i1/p83_s1. See comments [1952].

Leach:1979:NTT

[1573] P. G. L. Leach. On nonlinear transformations for time-dependent polynomial Hamiltonians. Journal of Mathematical Physics, 20(1):86–95, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i1/p86_s1.

Leach:1979:TIT

[1574] P. G. L. Leach. Towards an invariant for the time-dependent anharmonic oscillator. Journal of Mathematical Physics, 20(1):96–100, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i1/p96_s1.

Arrowsmith:1979:PTM

[1575] D. K. Arrowsmith. Percolation theory on multirooted directed graphs. Journal of Mathematical Physics, 20(1):101–103, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/p101_s1.

Hirondel:1979:PRM

[1576] D. Hirondel. Predictive relativistic mechanics. Journal of Mathematical Physics, 20(1):104–113, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i1/p104_s1.

Hudson:1979:CFT

[1577] R. L. Hudson and S. N. Peck. Canonical Fourier transforms. Journal of Mathematical Physics, 20(1):114–119, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i1/p114_s1.

Halil:1979:CPG

[1578] Mustafa Halil. Colliding plane gravitational waves. Journal of Mathematical Physics, 20(1):120–125, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i1/p120_s1.

Cruz:1979:CDE

[1579] A. Cruz and J. Sesma. Convergence of the Debye expansion for the S matrix. Journal of Mathematical Physics, 20(1):126–131, January 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/p126_
s1.

Snellman:1979:RSM

[1580] Håkan Snellman and Göran Tengstrand. Relativistic scattering of model particles. Journal of Mathematical Physics, 20(1):132–140, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/
p132_s1.

Garibotti:1979:SPW

[1581] C. R. Garibotti and F. F. Grinstein. Summation of partial wave expansions in the scattering by long range potentials. II. Numerical applications. Journal of Mathematical Physics, 20(1):141–147, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i1/p141_s1.

Epifanio:1979:MRC

[1582] G. Epifanio and C. Trapani. On the matrix representation of closed linear operators: an extension of von Neumann’s theory. Journal of Mathematical Physics, 20(1):148–152, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i1/p148_s1.

So:1979:WC

[1583] S. I. So and D. Strottman. Wigner coefficients for SU(6) ⊇ SU(3)⊗SU(2). Journal of Mathematical Physics, 20(1):153–176, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i1/p153_s1.

Anile:1979:FGO

[1584] A. M. Anile and P. Pantano. Foundation of geometrical optics in general relativistic dispersive media. Journal of Mathematical Physics, 20(1):177–183, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p177_s1.

Miyazaki:1979:EOS

[1585] Eizo Miyazaki. Exact occupation statistics of two- and three-dimensional lattices of mixed single particles. Journal of Mathematical Physics, 20(1):184–187, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p184_s1.

Strinati:1979:MWP

[1586] G. Strinati. Multipole wavefunctions for photoelectrons in crystals. IV. The irregular functions and the matching to an impurity. Journal of Mathematical Physics, 20(1):188–194, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i1/p188_s1.

Herrera:1979:NFA

[1587] L. Herrera and J. Jiménez. Neutrino fields in axially and reflection symmetric space–times. Journal of Mathematical Physics, 20(1):195– 198, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i1/p195_s1.

Cornille:1979:SGN

[1588] Henri Cornille. Solutions of the generalized nonlinear Schrödinger equation in two spatial dimensions. Journal of Mathematical Physics, 20(1):199–209, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p199_s1.

Campanino:1979:IFE

[1589] M. Campanino and G. Del Grosso. Independence of the free energy for one-dimensional systems of fermions. Journal of Mathematical Physics, 20(1):210–215, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i1/p210_s1.

McCormick:1979:AIP

[1590] N. J. McCormick and J. A. R. Veeder. Addendum: “On the inverse problem of transport theory with azimuthal dependence [J. Math. Phys. 19, 994 (1978)]”. Journal of Mathematical Physics, 20(1):216, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i1/p216_s1. See [1290].

Jang:1979:EPE

[1591] Pong Soo Jang. Erratum: “On the positivity of energy in general relativity”, (J. Math. Phys. 19 (1978), no. 5, 1152–1155). Journal of Mathematical Physics, 20(1):217, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i1/p217_s1. See [1315].

MartinezAlonso:1979:GTF

[1592] L. Martínez Alonso. Group-theoretical foundations of classical and quantum mechanics. II. Elementary systems. Journal of Mathematical Physics, 20(2):219–230, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i2/p219_s1.

Petry:1979:ESS

[1593] Herbert Rainer Petry. Exotic spinors in superconductivity. Journal of Mathematical Physics, 20(2):231–240, February 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i2/p231_s1.

Perez:1979:EPP

[1594] R. B. Perez. On the expansion of the propagator in power series of the coupling constant. Journal of Mathematical Physics, 20(2):241– 244, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i2/p241_s1.

Raychaudhuri:1979:CFS

[1595] A. K. Raychaudhuri and S. R. Maiti. Conformal flatness and the Schwarzschild interior solution. Journal of Mathematical Physics, 20 (2):245–246, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i2/p245_s1.

Wong:1979:IOT

[1596] M. K. F. Wong and Hsin Yang Yeh. Invariant operators of IU(n) and IO(n) and their eigenvalues. Journal of Mathematical Physics, 20(2):247– 250, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i2/p247_s1.

Harvey:1979:ABM

[1597] Alex Harvey. Automorphisms of the Bianchi model Lie groups. Journal of Mathematical Physics, 20(2):251–253, February 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i2/p251_s1.

Chen:1979:PSP

[1598] M. Chen. Perturbative solution of the Percus–Yevick integral equation for a general class of intermolecular potential. Journal of Mathematical Physics, 20(2):254–259, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i2/p254_s1.

Anandan:1979:CGF

[1599] J. Anandan. Classification of gauge fields. Journal of Mathematical Physics, 20(2):260–268, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i2/p260_s1.

Horwitz:1979:EGG

[1600] L. P. Horwitz and L. C. Biedenharn. Exceptional gauge groups and quantum theory. Journal of Mathematical Physics, 20(2):269–298, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i2/p269_s1.

Geronimo:1979:STP

[1601] J. S. Geronimo and K. M. Case. Scattering theory and polynomials orthogonal on the unit circle. Journal of Mathematical Physics, 20(2):299– 310, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i2/p299_s1.

Dudley:1979:CPW

[1602] Alan L. Dudley and J. D. Finley III. Covariant perturbed wave equations in arbitrary type-D backgrounds. Journal of Mathematical Physics, 20(2):311–328, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i2/p311_s1.

Hoenders:1979:CNM

[1603] B. J. Hoenders. On the completeness of the natural modes for quantum mechanical potential scattering. Journal of Mathematical Physics, 20(2):329–335, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i2/p329_s1.

Peck:1979:CSS

[1604] S. N. Peck. Canonical Schwartz spaces and generalized operators. Journal of Mathematical Physics, 20(2):336–343, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i2/p336_s1.

Morgan:1979:EAB

[1605] John D. Morgan III. Erratum: “The asymptotic behavior of bound eigenfunctions of Hamiltonians for single variable systems”, [J. Math. Phys. 19, 1658 (1978)]. Journal of Mathematical Physics, 20(2):344, February 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i2/p344_s1. See [1395].

Davies:1979:PBI

[1606] E. B. Davies. Particle-boson interactions and the weak coupling limit. Journal of Mathematical Physics, 20(3):345–351, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/p345_s1.

Reichl:1979:MTSa

[1607] L. E. Reichl. Microscopic theory of superfluid Fermi systems. I. Binary expansion. Journal of Mathematical Physics, 20(3):352–360, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i3/p352_s1.

Reichl:1979:MTSb

[1608] L. E. Reichl. Microscopic theory of superfluid Fermi systems. II. Application to low density hard sphere system. Journal of Mathematical Physics, 20(3):361–366, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i3/p361_s1.

Bednar:1979:CCF

[1609] Miroslav Bednář and Vladimír F. Šachl. Classification and construction of finite-dimensional irreducible representations of the graded algebras; application to the (Sp(2n); 2n) algebra. Journal of Mathematical Physics, 20(3):367–371, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p367_s1.

Talukdar:1979:ESRa

[1610] B. Talukdar, U. Das, and S. Mukhopadhyaya. Off-energy-shell results for scattering by a nonlocal potential. I. Journal of Mathematical Physics, 20(3):372–377, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p372_s1.

Talukdar:1979:ESRb

[1611] B. Talukdar, U. Das, N. Mallick, and M. Sen Gupta. Off-energy-shell results for scattering by a nonlocal potential. II. Journal of Mathematical Physics, 20(3):378–381, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p378_s1.

Mitchell:1979:LAA

[1612] Thomas P. Mitchell. Lie algebras associated with motion in axisymmetric electromagnetic fields. Journal of Mathematical Physics, 20(3): 382–384, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i3/p382_s1.

Christensen:1979:ETO

[1613] Mark J. Christensen. Extension theorems for operator-valued measures. Journal of Mathematical Physics, 20(3):385–389, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/p385_s1.

Jagannathan:1979:RPB

[1614] R. Jagannathan and R. Vasudevan. Representations of a para-Bose algebra using only a single Bose field. Journal of Mathematical Physics, 20(3): 390–392, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i3/p390_s1. See erratum [1889].

Sage:1979:CCC

[1615] Martin L. Sage and Kameshwar C. Wali. Computation of a class of 3j-coefficients. Journal of Mathematical Physics, 20(3):393–395, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i3/p393_s1.

Bosanac:1979:SPE

[1616] S. Bosanac. Semiclassical perturbation expansion of the multichannel scattering matrix. Journal of Mathematical Physics, 20(3):396– 405, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i3/p396_s1.

Tikochinsky:1979:DGQ

[1617] Yoel Tikochinsky. On the diagonalization of the general quadratic Hamiltonian for coupled harmonic oscillators. Journal of Mathematical Physics, 20(3):406–408, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p406_s1.

Dyer:1979:CKH

[1618] C. C. Dyer and E. Honig. Conformal Killing horizons. Journal of Mathematical Physics, 20(3):409–412, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i3/p409_s1.

Lewis:1979:LAI

[1619] H. Ralph Lewis and Keith R. Symon. Linearized analysis of inhomogeneous plasma equilibria: General theory. Journal of Mathematical Physics, 20(3):413–436, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i3/p413_s1. See errata [1890, 1940].

Abellanas:1979:IEA

[1620] L. Abellanas and L. Martínez Alonso. Invariants in enveloping algebras under the action of Lie algebras of derivations. Journal of Mathematical Physics, 20(3):437–440, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p437_s1.

Rosen:1979:CHF

[1621] Gerald Rosen, Oleg Jakubowicz, and Vahe Tatoian. Constrained Hamiltonian formulation for interacting fields: Stable particlelike solutions. Journal of Mathematical Physics, 20(3):441–442, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/p441_s1.

Arms:1979:LSG

[1622] Judith M. Arms. Linearization stability of gravitational and gauge fields. Journal of Mathematical Physics, 20(3):443–453, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/p443_s1.

Glassey:1979:CLC

[1623] R. T. Glassey and W. A. Strauss. Conservation laws for the classical Maxwell–Dirac and Klein–Gordon–Dirac equations. Journal of Mathematical Physics, 20(3):454–458, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i3/p454_s1.

Smallwood:1979:CPG

[1624] J. Smallwood. The Cauchy problem in general relativity. III. On locally imbedding a family of null hypersurfaces. Journal of Mathematical Physics, 20(3):459–464, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p459_s1.

Rowe:1979:GDK

[1625] D. J. Rowe and G. Rosensteel. Geometric derivation of the kinetic energy in collective models. Journal of Mathematical Physics, 20(3):465– 468, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i3/p465_s1.

delaPena:1979:QHO

[1626] L. de la Peña and A. M. Cetto. The quantum harmonic oscillator revisited: a new look from stochastic electrodynamics. Journal of Mathematical Physics, 20(3):469–483, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i3/p469_s1.

Geronimo:1979:STH

[1627] J. S. Geronimo. Szegő’s theorem for Hankel determinants. Journal of Mathematical Physics, 20(3):484–491, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i3/p484_s1.

Basu:1979:MAF

[1628] Debabrata Basu and S. Datta Majumdar. The master analytic function and the Lorentz group. III. Coupling of continuous representations of O(2, 1). Journal of Mathematical Physics, 20(3):492–498, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/
p492_s1.

Langguth:1979:RHP

[1629] Wolfgang Langguth and Akira Inomata. Remarks on the Hamiltonian path integral in polar coordinates. Journal of Mathematical Physics, 20(3):499–504, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p499_s1.

Akyeampong:1979:THH

[1630] D. A. Akyeampong. Tensor harmonics in the helicity basis. Journal of Mathematical Physics, 20(3):505–508, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i3/p505_s1.

Bertero:1979:POR

[1631] M. Bertero, C. De Mol, and G. A. Viano. On the problems of object restoration and image extrapolation in optics. Journal of Mathematical Physics, 20(3):509–521, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i3/p509_s1.

Aldersley:1979:HEO

[1632] S. J. Aldersley. Higher Euler operators and some of their applications. Journal of Mathematical Physics, 20(3):522–531, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i3/p522_s1. See

erratum [1935].

Barnsley:1979:DMP

[1633] M. Barnsley, D. Bessis, and P. Moussa. The Diophantine moment problem and the analytic structure in the activity of the ferromagnetic Ising model. Journal of Mathematical Physics, 20(4):535–546, April 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i4/p535_
s1.

Cornwell:1979:DDL

[1634] J. F. Cornwell. Direct determination of the Langlands decompositions for the parabolic subalgebras of noncompact semisimple real Lie algebras. Journal of Mathematical Physics, 20(4):547–555, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i4/p547_s1.

Arnal:1979:GTC

[1635] Didier Arnal and Jean-Claude Cortet. Geometrical theory of contractions of groups and representations. Journal of Mathematical Physics, 20(4): 556–563, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i4/p556_s1.

Lewis:1979:FLV

[1636] H. Ralph Lewis and Leaf Turner. Formulation of the linearized Vlasovfluid model for a sharp-boundary screw pinch. Journal of Mathematical Physics, 20(4):564–571, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p564_s1.

Klein:1979:SQN

[1637] Abraham Klein and Ching teh Li. Semiclassical quantization of nonseparable systems. Journal of Mathematical Physics, 20(4):572–578, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i4/p572_s1.

Moncrief:1979:GDC

[1638] Vincent Moncrief. Gribov degeneracies: Coulomb gauge conditions and initial value constraints. Journal of Mathematical Physics, 20(4):579– 585, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i4/p579_s1.

Okubo:1979:QTI

[1639] Susumu Okubo. Quartic trace identity for exceptional Lie algebras. Journal of Mathematical Physics, 20(4):586–593, April 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i4/p586_s1.

Ben-Artzi:1979:SST

[1640] Matania Ben-Artzi and Allen Devinatz. Spectral and scattering theory for the adiabatic oscillator and related potentials. Journal of Mathematical Physics, 20(4):594–607, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i4/p594_s1. See erratum [1948].

Bowden:1979:NIC

[1641] R. L. Bowden. A note on the iteration of the Chandrasekhar nonlinear H-equation. Journal of Mathematical Physics, 20(4):608–610, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i4/p608_s1.

Combe:1979:QSS

[1642] Ph. Combe, R. Rodriguez, M. Sirugue-Collin, and M. Sirugue. On the quantization of spin systems and Fermi systems. Journal of Mathematical Physics, 20(4):611–616, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p611_s1.

Furutsu:1979:ATP

[1643] K. Furutsu. An analytical theory of pulse wave propagation in turbulent media. Journal of Mathematical Physics, 20(4):617–628, April 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i4/p617_
s1.

Wong:1979:LRO

[1644] M. K. F. Wong and Hsin Yang Yeh. Lowering and raising operators of IU(n) and IO(n) and their normalization constants. Journal of Mathematical Physics, 20(4):629–633, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i4/p629_s1.

Gordon:1979:QPE

[1645] Paul Gordon. Quasisteady primitive equations with associated upper boundary conditions. Journal of Mathematical Physics, 20(4):634–658, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i4/p634_s1.

Dirl:1979:CGCa

[1646] R. Dirl. Clebsch–Gordan coefficients: General theory. Journal of Mathematical Physics, 20(4):659–663, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i4/p659_s1.

Dirl:1979:MSG

[1647] R. Dirl. Multiplicities for space group representations. Journal of Mathematical Physics, 20(4):664–670, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i4/p664_s1.

Dirl:1979:CGCb

[1648] R. Dirl. Clebsch–Gordan coefficients for space groups. Journal of Mathematical Physics, 20(4):671–678, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i4/p671_s1.

Dirl:1979:CGCc

[1649] R. Dirl. Clebsch–Gordan coefficients for P n3n. Journal of Mathematical Physics, 20(4):679–684, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p679_s1.

Graffi:1979:ROD

[1650] S. Graffi, V. Grecchi, S. Levoni, and M. Maioli. Resonances in onedimensional Stark effect and continued fractions. Journal of Mathematical Physics, 20(4):685–690, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i4/p685_s1.

Maekawa:1979:LRD

[1651] Takayoshi Maekawa. Linear representations of any dimensional Lorentz group and computation formulas for their matrix elements. Journal of Mathematical Physics, 20(4):691–711, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i4/p691_s1.

Scheunert:1979:GLA

[1652] M. Scheunert. Generalized Lie algebras. Journal of Mathematical Physics, 20(4):712–720, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p712_s1.

Symes:1979:RAG

[1653] W. Symes. Relations among generalized Korteweg–de Vries systems. Journal of Mathematical Physics, 20(4):721–725, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i4/p721_s1.

Horndeski:1979:NEF

[1654] Gregory Walter Horndeski. Null electromagnetic fields in the generalized Einstein–Maxwell field theory. Journal of Mathematical Physics, 20(4): 726–732, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i4/p726_s1.

Pule:1979:CLQ

[1655] J. V. Pulè and A. Verbeure. The classical limit of quantum dissipative generators. Journal of Mathematical Physics, 20(4):733–735, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i4/p733_s1.

Darewych:1979:FOE

[1656] J. W. Darewych and J. Sokoloff. First order estimates of the error in approximate calculations of scattering phase shifts. Journal of Mathematical Physics, 20(4):736–739, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i4/p736_s1.

Das:1979:SEM

[1657] A. Das. On the static Einstein–Maxwell field equations. Journal of Mathematical Physics, 20(4):740–743, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i4/p740_s1.

Rosenbaum:1979:STH

[1658] Marcos Rosenbaum, Michael P. Ryan, Jr., and L. C. Shepley. Space– time homogeneous models with torsion. Journal of Mathematical Physics, 20(4):744–751, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p744_s1.

Choudhury:1979:SPMa

[1659] M. H. Choudhury and D. B. Pearson. The spectral properties of manyelectron atomic Hamiltonians and the method of configuration interaction. I. Proof of convergence of the configuration interaction method. Journal of Mathematical Physics, 20(4):752–756, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i4/p752_s1.

Davidson:1979:CPM

[1660] A. Davidson. Covariant n^{2}-plet mass formulas. Journal of Mathematical Physics, 20(5):757–761, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i5/p757_s1.

Greenberger:1979:CMA

[1661] Daniel M. Greenberger. A critique of the major approaches to damping in quantum theory. Journal of Mathematical Physics, 20(5):762–770, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p762_s1.

Greenberger:1979:NAP

[1662] Daniel M. Greenberger. A new approach to the problem of dissipation in quantum mechanics. Journal of Mathematical Physics, 20(5): 771–780, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p771_s1.

Rochus:1979:UBM

[1663] Pierre Rochus. Upper bounds for the many-channel Marchenko transformation operator and its derivatives. Journal of Mathematical Physics, 20(5):781–792, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i5/p781_s1.

Ashtekar:1979:CQG

[1664] Abhay Ashtekar and Anne Magnon-Ashtekar. On conserved quantities in general relativity. Journal of Mathematical Physics, 20(5):793–800, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p793_s1.

Miller:1979:TPC

[1665] David E. Miller. Thermodynamical properties of a class of solvable statistical models for hadronic matter. Journal of Mathematical Physics, 20(5): 801–809, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p801_s1.

Carey:1979:AAQ

[1666] A. L. Carey and C. A. Hurst. Applications of an algebraic quantization of the electromagnetic field. Journal of Mathematical Physics, 20(5): 810–819, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p810_s1.

Lemire:1979:FAC

[1667] F. Lemire and J. Patera. Formal analytic continuation of Gel′fand’s finite-dimensional representations of gl(n, C). Journal of Mathematical Physics, 20(5):820–829, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i5/p820_s1.

Marchesin:1979:SLF

[1668] Dan Marchesin. The scaling limit of the φ^{2} field in the anharmonic oscillator. Journal of Mathematical Physics, 20(5):830–836, May 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i5/p830_
s1.

Gonzalez:1979:RKN

[1669] C. González, L. Herrera, and J. Jiménez. Radiating Kerr–Newman metric. Journal of Mathematical Physics, 20(5):837–843, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i5/p837_s1.

Mizrahi:1979:SEA

[1670] Maurice M. Mizrahi. The semiclassical expansion of the anharmonicoscillator propagator. Journal of Mathematical Physics, 20(5):844–855, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p844_s1.

Marmo:1979:GSR

[1671] G. Marmo, E. J. Saletan, and A. Simoni. A general setting for reduction of dynamical systems. Journal of Mathematical Physics, 20(5): 856–860, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p856_s1.

vanHaeringen:1979:ISE

[1672] H. van Haeringen. On an inhomogeneous Schrödinger equation and its solutions in scattering theory. Journal of Mathematical Physics, 20(5): 861–867, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p861_s1.

Miglietta:1979:EQP

[1673] Francesco Miglietta. Electromagnetic quadripotential for the pureradiation field generated by a classical charged point-particle. Journal of Mathematical Physics, 20(5):868–870, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i5/p868_s1.

Maison:1979:CIS

[1674] Dieter Maison. On the complete integrability of the stationary, axially symmetric Einstein equations. Journal of Mathematical Physics, 20(5): 871–877, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p871_s1.

Malo:1979:ASL

[1675] Otieno Malo. An attempt to separate the long and short range forces by Gaussian method. Journal of Mathematical Physics, 20(5):878–886, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p878_s1.

Das:1979:WSM

[1676] U. Das, S. Chakravarty, and B. Talukdar. S-wave off-shell T and K matrices for the Yukawa potential by Ecker–Weizel approximations. Journal of Mathematical Physics, 20(5):887–890, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i5/p887_s1.

Gilmore:1979:CLQ

[1677] R. Gilmore. The classical limit of quantum nonspin systems. Journal of Mathematical Physics, 20(5):891–893, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i5/p891_s1.

Dashen:1979:PIW

[1678] Roger Dashen. Path integrals for waves in random media. Journal of Mathematical Physics, 20(5):894–920, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i5/p894_s1.

Mallett:1979:HMI

[1679] Ronald L. Mallett. Higgs mechanism and the inverse Einstein–Infeld– Hoffman problem. Journal of Mathematical Physics, 20(5):921–922, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p921_s1.

Oliveira:1979:ZAG

[1680] Colber G. Oliveira and Marcos D. Maia. Zorn algebra in general relativity. Journal of Mathematical Physics, 20(5):923–930, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i5/p923_s1.

Harnad:1979:YMS

[1681] J. Harnad, S. Shnider, and Luc Vinet. The Yang–Mills system in compactified Minkowski space; invariance conditions and SU(2) invariant solutions. Journal of Mathematical Physics, 20(5):931–942, May 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i5/p931_
s1.

Choudhury:1979:SPMb

[1682] M. H. Choudhury. The spectral properties of many-electron atomic Hamiltonians and the method of configuration interaction. II. Compactness proof associated with an infinite system of linear equations for twoelectron atoms. Journal of Mathematical Physics, 20(5):943–952, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p943_s1.

Montesinos:1979:GSF

[1683] Angel Montesinos. Geometry of spacetime founded on spacelike metric. Journal of Mathematical Physics, 20(5):953–965, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i5/p953_s1.

Ringhofer:1979:KPR

[1684] Klaus Ringhofer. Koornwinder’s polynomials and representations of the conformal group. Journal of Mathematical Physics, 20(5):966–971, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i5/p966_s1.

Case:1979:PBO

[1685] K. M. Case. Properties of the Benjamin–Ono equation. Journal of Mathematical Physics, 20(5):972–977, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i5/p972_s1.

Enz:1979:EPM

[1686] U. Enz. Erratum: “A particle model based on stringlike solitons [J. Math. Phys. 19, 1304 (1978)]”. Journal of Mathematical Physics, 20 (5):978, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i5/p978_s1. See [1340].

Schwarzmeier:1979:VNE

[1687] James L. Schwarzmeier and Keith R. Symon. Variational nature of eigenvalue problems with a parameter. Journal of Mathematical Physics, 20(6):979–983, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i6/p979_s1.

Loupias:1979:EKS

[1688] G. Loupias and M. Mebkhout. Extension of KMS states and angular momentum. Journal of Mathematical Physics, 20(6):984–998, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p984_s1.

Ablowitz:1979:NMT

[1689] Mark J. Ablowitz, Martin Kruskal, and Harvey Segur. A note on Miura’s transformation. Journal of Mathematical Physics, 20(6):999–1003, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p999_s1.

Plebanski:1979:SEE

[1690] Jerzy F. Plebański and Shahen Hacyan. Some exceptional electrovac type D metrics with cosmological constant. Journal of Mathematical Physics, 20(6):1004–1010, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i6/p1004_s1.

Doebner:1979:GPS

[1691] H. D. Doebner and J.-E. Werth. Global properties of systems quantized via bundles. Journal of Mathematical Physics, 20(6):1011–1014, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1011_s1.

Anderson:1979:CGC

[1692] Ronald Anderson and G. C. Joshi. SU(4) Clebsch–Gordan coefficients for the formation of baryonium and exotic baryons. Journal of Mathematical Physics, 20(6):1015–1029, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i6/p1015_s1.

Shenoy:1979:IIE

[1693] S. K. Shenoy and A. J. Mockel. Invariant imbedding equations in general geometries: Numerical solution in spherical geometry. Journal of Mathematical Physics, 20(6):1030–1040, June 1979. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i6/p1030_s1.

Hauser:1979:GEM

[1694] I. Hauser and F. J. Ernst. SU(2, 1) generation of electrovacs from Minkowski space. Journal of Mathematical Physics, 20(6):1041–1055, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1041_s1.

Wald:1979:NSS

[1695] Robert M. Wald. Note on the stability of the Schwarzschild metric. Journal of Mathematical Physics, 20(6):1056–1058, June 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1056_s1. See er-

ratum [1938].

Osborn:1979:IBB

[1696] T. A. Osborn and D. Bollé. Integral bounds for N-body total cross sections. Journal of Mathematical Physics, 20(6):1059–1063, June 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1059_
s1.

Rasmussen:1979:AAC

[1697] W. O. Rasmussen and S. Salamó. An algebraic approach to Coulomb scattering in N dimensions. Journal of Mathematical Physics, 20(6): 1064–1067, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1064_s1.

Leon:1979:GCF

[1698] J. León, M. Quirós, and J. Ramirez Mittelbrunn. Group content of the Foldy–Wouthuysen transformation and the nonrelativistic limit for arbitrary spin. Journal of Mathematical Physics, 20(6):1068–1076, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1068_s1.

Daumens:1979:IGG

[1699] M. Daumens and M. Perroud. Internal Galilei group for a two-particle system. Journal of Mathematical Physics, 20(6):1077–1089, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/
p1077_s1.

Federbush:1979:PCC

[1700] Paul Federbush. Positivity conditions on correlation functions that imply Debye screening. Journal of Mathematical Physics, 20(6):1090–1092, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1090_s1.

Lee:1979:GSS

[1701] K. C. Lee and S. E. Trullinger. Generalized susceptibility of a solitary wave. Journal of Mathematical Physics, 20(6):1093–1098, June 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1093_
s1.

Bowden:1979:RIT

[1702] R. L. Bowden, William Greenberg, and P. F. Zweifel. Resolvent integration techniques for generalized transport equations. Journal of Mathematical Physics, 20(6):1099–1108, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i6/p1099_s1.

vanHaeringen:1979:CCL

[1703] H. van Haeringen. The Coulomb and Coulomb-like off-shell Jost functions. Journal of Mathematical Physics, 20(6):1109–1114, June 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1109_
s1.

Halford:1979:PTV

[1704] W. D. Halford. Petrov type N vacuum metrics and homothetic motions. Journal of Mathematical Physics, 20(6):1115–1117, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1115_s1.

Fahnline:1979:CMT

[1705] Donald E. Fahnline. Circular motion for a time-asymmetric relativistic two-body problem. Journal of Mathematical Physics, 20(6):1118– 1120, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1118_s1.

Bolle:1979:TDB

[1706] D. Bollé and T. A. Osborn. Time delay in N-body scattering. Journal of Mathematical Physics, 20(6):1121–1134, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i6/p1121_s1.

Barnsley:1979:CMB

[1707] M. Barnsley and D. Bessis. Constructive methods based on analytic characterizations and their application to nonlinear elliptic and parabolic differential equations. Journal of Mathematical Physics, 20(6):1135–1145, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1135_s1.

Novotny:1979:MPA

[1708] Mark A. Novotny. Matrix products with application to classical statistical mechanics. Journal of Mathematical Physics, 20(6):1146–1150, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1146_s1.

Moses:1979:JSG

[1709] Harry E. Moses. Jost solutions and Green’s functions for the threedimensional Schrödinger equation. Journal of Mathematical Physics, 20(6):1151–1156, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i6/p1151_s1.

Choudhury:1979:SPMc

[1710] M. H. Choudhury. The spectral properties of many-electron atomic Hamiltonians and the method of configuration interaction. III. Compactness proof associated with an infinite system of linear equations for n-electron atoms. Journal of Mathematical Physics, 20(6):1157–1169, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1157_s1.

Tsang:1979:MMT

[1711] L. Tsang and J. A. Kong. Modified modal theory of transient response in layered media. Journal of Mathematical Physics, 20(6):1170–1182, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i6/p1170_s1.

Greene:1979:MDS

[1712] John M. Greene. A method for determining a stochastic transition. Journal of Mathematical Physics, 20(6):1183–1201, June 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1183_s1.

Fulling:1979:AES

[1713] S. A. Fulling. Adiabatic expansions of solutions of coupled second-order linear differential equations. II. Journal of Mathematical Physics, 20(6): 1202–1209, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1202_s1.

Romo:1979:AFR

[1714] W. J. Romo. Asymptotic forms of radial wavefunctions and Jost functions for cutoff potentials. Journal of Mathematical Physics, 20(6):1210– 1224, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1210_s1. See erratum [1936].

Cox:1979:YAF

[1715] Dale Cox and William Kinnersley. Yet another formulation of the Einstein equations for stationary axisymmetry. Journal of Mathematical Physics, 20(6):1225–1228, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i6/p1225_s1.

deDormale:1979:RLM

[1716] Bernard M. de Dormale. A relativistic Lee model. Journal of Mathematical Physics, 20(6):1229–1231, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i6/p1229_s1.

Takahashi:1979:SSN

[1717] K. Takahashi. Soliton solutions of nonlinear Dirac equations. Journal of Mathematical Physics, 20(6):1232–1238, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i6/p1232_s1.

Abellanas:1979:CDN

[1718] L. Abellanas and A. Galindo. Conserved densities for nonlinear evolution equations. I. Even order case. Journal of Mathematical Physics, 20(6):

1239–1243, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1239_s1. See erratum [1942].

Bergere:1979:IRD

[1719] M. C. Bergère and F. David. Integral representation for the dimensionally regularized massive Feynman amplitude. Journal of Mathematical Physics, 20(6):1244–1255, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i6/p1244_s1.

Galindo:1979:ACE

[1720] A. Galindo. An algorithm to construct evolution equations with a given set of conserved densities. Journal of Mathematical Physics, 20(6):1256– 1259, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1256_s1.

Nwachuku:1979:NEE

[1721] C. O. Nwachuku. New expressions for the eigenvalues of the invariant operators of O(N) and Sp(2n). Journal of Mathematical Physics, 20(6): 1260–1266, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1260_s1.

Gielerak:1979:MEB

[1722] Roman Gielerak. Mayer equations for Bose field theoretical models. Journal of Mathematical Physics, 20(6):1267–1273, June 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1267_s1.

Folland:1979:FIS

[1723] N. O. Folland. Factored irreducible symmetry operators and space groups. Journal of Mathematical Physics, 20(6):1274–1280, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/
p1274_s1.

Flaherty:1979:ILG

[1724] Francis J. Flaherty and G. David Kerlick. Inhomogeneous local gauge transformations in spacetimes with torsion. Journal of Mathematical Physics, 20(6):1281–1284, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i6/p1281_s1.

Ibanez:1979:TBG

[1725] J. Ibañez and J. L. Sanz. Two-body gravidynamics. Journal of Mathematical Physics, 20(6):1285–1295, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i6/p1285_s1.

Sarlet:1979:AIC

[1726] Willy Sarlet. Addendum: “Invariance and conservation laws for Lagrangian systems with one degree of freedom [J. Math. Phys. 19, 1049 (1978)]”. Journal of Mathematical Physics, 20(6):1295, June 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i6/p1295_

s1. See [1299].

Marshall:1979:PIE

[1727] John T. Marshall and John L. Pell. Path-integral evaluation of the space– time propagator for quadratic Hamiltonian systems. Journal of Mathematical Physics, 20(7):1297–1302, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1297_s1.

Reina:1979:ISA

[1728] Cesare Reina. Internal symmetries of the axisymmetric gravitational fields. Journal of Mathematical Physics, 20(7):1303–1305, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1303_
s1.

Si:1979:RFE

[1729] Do Tan Si. Resolution of Fredholm equations with kernels K(z − z_{0}) by operational calculus. Journal of Mathematical Physics, 20(7):1306– 1307, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1306_s1.

Lapiedra:1979:CPEa

[1730] R. Lapiedra and A. Molina. Classical predictive electrodynamics of two charges with radiation: General framework. I. Journal of Mathematical Physics, 20(7):1308–1315, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1308_s1.

Lapiedra:1979:CPEb

[1731] R. Lapiedra, F. Marqués, and A. Molina. Classical predictive electrodynamics of two charges with radiation: Energy and 3-momentum balance and scattering cross sections. II. Journal of Mathematical Physics, 20(7): 1316–1320, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1316_s1.

Larsson:1979:CRAa

[1732] J. Larsson. Current responses of all orders in a collisionless plasma. I. General theory. Journal of Mathematical Physics, 20(7):1321–1330, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1321_s1.

Larsson:1979:CRAb

[1733] J. Larsson. Current responses of all orders in a collisionless plasma. II. Homogeneous plasma. Journal of Mathematical Physics, 20(7):1331– 1335, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1331_s1.

Emch:1979:WQN

[1734] Gérard G. Emch and Kalyan B. Sinha. Weak quantization in a nonperturbative model. Journal of Mathematical Physics, 20(7):1336–1340, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1336_s1.

Kim:1979:RPG

[1735] Y. S. Kim, Marilyn E. Noz, and S. H. Oh. Representations of the Poincaré group for relativistic extended hadrons. Journal of Mathematical Physics, 20(7):1341–1344, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1341_s1.

Miller:1979:BKH

[1736] J. G. Miller. Bifurcate Killing horizons. Journal of Mathematical Physics, 20(7):1345–1348, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1345_s1.

Messer:1979:DLC

[1737] Joachim Messer. Decay of local correlations and absence of phase transitions. Journal of Mathematical Physics, 20(7):1349–1351, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1349_
s1.

Burzlaff:1979:SAS

[1738] J. Burzlaff and D. Maison. On stationary, axially symmetric solutions of Jordan’s five-dimensional, unified theory. Journal of Mathematical Physics, 20(7):1352–1355, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1352_s1.

Miller:1979:RAT

[1739] Bonnie D. Miller. Repulsive and attractive timelike singularities in vacuum cosmologies. Journal of Mathematical Physics, 20(7):1356–1361, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1356_s1.

Ashtekar:1979:AMS

[1740] Abhay Ashtekar and Michael Streubel. On angular momentum of stationary gravitating systems. Journal of Mathematical Physics, 20(7): 1362–1365, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1362_s1.

Erdahl:1979:AGT

[1741] R. M. Erdahl, C. Garrod, B. Golli, and M. Rosina. The application of group theory to generate new representability conditions for rotationally invariant density matrices. Journal of Mathematical Physics, 20(7):1366– 1374, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1366_s1.

Gallone:1979:SQD

[1742] Franco Gallone and Antonio Sparzani. Segal quantization of dynamical systems. Journal of Mathematical Physics, 20(7):1375–1384, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1375_s1.

Ali:1979:SRP

[1743] S. Twareque Ali. On some representations of the Poincaré group on phase space. Journal of Mathematical Physics, 20(7):1385–1391, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1385_s1.

Rund:1979:RDG

[1744] Hanno Rund. Representations of the duals of gauge field tensors. Journal of Mathematical Physics, 20(7):1392–1397, July 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1392_s1.

Seznec:1979:SDS

[1745] R. Seznec and J. Zinn-Justin. Summation of divergent series by order dependent mappings: Application to the anharmonic oscillator and critical exponents in field theory. Journal of Mathematical Physics, 20(7): 1398–1408, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1398_s1.

Murray:1979:SFD

[1746] J. C. Murray. On the sound field due to a moving source in a superfluid. Journal of Mathematical Physics, 20(7):1409–1412, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1409_

s1. See erratum [1947].

Reddy:1979:SCF

[1747] D. R. K. Reddy. Static conformally flat solution in a scalar-tensor theory of gravitation. Journal of Mathematical Physics, 20(7):1413–1414, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1413_s1.

Plebanski:1979:AGH

[1748] Jerzy F. Plebañski. The algebra generated by Hodge’s star and external differential. Journal of Mathematical Physics, 20(7):1415–1422, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1415_s1.

Parker:1979:DG

[1749] Phillip E. Parker. Distributional geometry. Journal of Mathematical Physics, 20(7):1423–1426, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1423_s1.

Barker:1979:CIC

[1750] B. M. Barker and R. F. O’Connell. Center of inertia and coordinate transformations in the post- Newtonian charged n-body problem in gravitation. Journal of Mathematical Physics, 20(7):1427–1434, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1427_
s1.

Leung:1979:MDS

[1751] C. N. Leung and J. Rosner. Mass dependence of Schrödinger wavefunctions. Journal of Mathematical Physics, 20(7):1435–1438, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1435_
s1.

Costen:1979:EPF

[1752] Robert C. Costen. Equations for the propagation of finite local disturbances in vector fields of physics and mechanics. Journal of Mathematical Physics, 20(7):1439–1445, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1439_s1.

Aurilia:1979:TSMa

[1753] A. Aurilia and D. Christodoulou. Theory of strings and membranes in an external field. I. General formulation. Journal of Mathematical Physics, 20(7):1446–1452, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1446_s1.

Harter:1979:AEF

[1754] William G. Harter and Chris W. Patterson. Asymptotic eigensolutions of fourth and sixth rank octahedral tensor operators. Journal of Mathematical Physics, 20(7):1453–1459, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1453_s1.

Maekawa:1979:NLR

[1755] Takayoshi Maekawa. Note on the linear representations of any dimensional Lorentz group and their matrix elements. Journal of Mathematical Physics, 20(7):1460–1463, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1460_s1.

Doria:1979:NGP

[1756] Francisco Antonio Doria. Noncontinuous gauge potentials without magnetic monopoles. Journal of Mathematical Physics, 20(7):1464–1465, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1464_s1.

Darkhosh:1979:DGF

[1757] Teymour Darkhosh. Discontinuities of the Green’s functions across the characteristic surfaces of field equations. Journal of Mathematical Physics, 20(7):1466–1472, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1466_s1.

Li:1979:CIG

[1758] Wann-Quan Li and Wei-Tou Ni. Coupled inertial and gravitational effects in the proper reference frame of an accelerated, rotating observer. Journal of Mathematical Physics, 20(7):1473–1480, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1473_s1.

Peters:1979:TBH

[1759] P. C. Peters. Toroidal black holes? Journal of Mathematical Physics, 20(7):1481–1485, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i7/p1481_s1.

Tanabe:1979:GLS

[1760] Yukio Tanabe. Generation of the Lewis solutions from the Weyl solutions. Journal of Mathematical Physics, 20(7):1486–1489, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1486_s1.

Bellissard:1979:RAD

[1761] J. Béllissard. A remark about the duality for non-Abelian lattice fields. Journal of Mathematical Physics, 20(7):1490–1493, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1490_s1.

Bandyopadhyay:1979:DDC

[1762] N. Bandyopadhyay. Dust distribution (cylindrically symmetric) in nonrigid rotation in Brans–Dicke theory. Journal of Mathematical Physics, 20(7):1494–1495, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i7/p1494_s1.

Satsuma:1979:TDL

[1763] J. Satsuma and M. J. Ablowitz. Two-dimensional lumps in nonlinear dispersive systems. Journal of Mathematical Physics, 20(7):1496–1503, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1496_s1.

McCormick:1979:TSC

[1764] N. J. McCormick. Transport scattering coefficients from reflection and transmission measurements. Journal of Mathematical Physics, 20(7): 1504–1507, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i7/p1504_s1.

Glass:1979:SFG

[1765] E. N. Glass. Shear-free gravitational collapse. Journal of Mathematical Physics, 20(7):1508–1513, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1508_s1.

Rashid:1979:BME

[1766] M. A. Rashid. Boost matrix elements of the homogeneous Lorentz group. Journal of Mathematical Physics, 20(7):1514–1519, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1514_s1.

Dalton:1979:NNR

[1767] B. J. Dalton. New nonlinear realizations of SL(2, C). Journal of Mathematical Physics, 20(7):1520–1529, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1520_s1.

Rose:1979:TOM

[1768] Conrad M. Rose and Ioannis M. Besieris. Nth-order multifrequency coherence functions: A functional path integral approach. Journal of Mathematical Physics, 20(7):1530–1538, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i7/p1530_s1.

Gunther:1979:UDD

[1769] Neil J. Günther. Unitary declension of dynamical symmetries for the time-dependent harmonic oscillator. Journal of Mathematical Physics, 20(7):1539–1544, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i7/p1539_s1.

Backhouse:1979:PR

[1770] N. B. Backhouse and J. W. B. Hughes. Projective representations of SU(2)_{Λ}R^{4}. Journal of Mathematical Physics, 20(7):1545–1554, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i7/p1545_s1.

Dadhich:1979:SGN

[1771] Naresh Dadhich. Spacelike geodesics in the nordström geometry. Journal of Mathematical Physics, 20(7):1555–1558, July 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1555_s1.

Chowdhury:1979:PSL

[1772] A. Roy Chowdhury and T. Roy. Prolongation structure for Langmuir solitons. Journal of Mathematical Physics, 20(7):1559–1561, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1559_
s1.

Dirl:1979:CCG

[1773] R. Dirl. Coupling coefficients: General theory. Journal of Mathematical Physics, 20(7):1562–1565, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1562_s1.

Dirl:1979:CCSa

[1774] R. Dirl. Complex conjugation of space group representations. Journal of Mathematical Physics, 20(7):1566–1569, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i7/p1566_s1.

Dirl:1979:CCSb

[1775] R. Dirl. Coupling coefficients for space groups. Journal of Mathematical Physics, 20(7):1570–1574, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1570_s1.

Dirl:1979:CC

[1776] R. Dirl. Coupling coefficients for P n3n. Journal of Mathematical Physics, 20(7):1575–1578, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i7/p1575_s1.

Rudra:1979:GTA

[1777] P. Rudra and M. K. Sikdar. Group theoretical analysis of secondorder phase transitions in magnetic structures. Journal of Mathematical Physics, 20(7):1579–1592, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i7/p1579_s1.

Srinivas:1979:SOR

[1778] M. D. Srinivas. Successive observations in relativistic quantum theory. Journal of Mathematical Physics, 20(7):1593–1611, July 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1593_s1.

CastilhoAlcaras:1979:CSC

[1779] J. A. Castilho Alcarás and L. C. Biedenharn. On the coupling of selfconjugate systems with \overline{SL(3, R)} symmetry. Journal of Mathematical Physics, 20(8):1615–1621, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1615_s1. See erratum [1949].

Chatterjee:1979:CWS

[1780] R. Chatterjee and J. M. Dixon. Comment on the Wigner 9-j symbol. Journal of Mathematical Physics, 20(8):1622–1623, August 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1622_
s1.

Klimyk:1979:RME

[1781] A. U. Klimyk and A. M. Gavrilik. Representation matrix elements and Clebsch–Gordan coefficients of the semisimple Lie groups. Journal of Mathematical Physics, 20(8):1624–1642, August 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1624_s1.

Strottman:1979:EWC

[1782] D. Strottman. Evaluation of SU(6) ⊇ SU(3)⊗SU(2) Wigner coefficients. Journal of Mathematical Physics, 20(8):1643–1647, August 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1643_
s1.

Weir:1979:SSH

[1783] G. J. Weir. Singular spin harmonics. Journal of Mathematical Physics, 20(8):1648–1649, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1648_s1.

Bassanini:1979:ALE

[1784] Piero Bassanini. An application of Lax’s estimates to the determination of critical crystal thickness in nonlinear laser optics. Journal of Mathematical Physics, 20(8):1650–1652, August 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1650_s1.

Cornille:1979:SNW

[1785] H. Cornille. Solutions of the nonlinear 3-wave equations in three spatial dimensions. Journal of Mathematical Physics, 20(8):1653–1666, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v20/i8/p1653_s1. See erratum [1950].

Lakshmanan:1979:RBM

[1786] M. Lakshmanan. Rigid body motions, space curves, prolongation structures, fiber bundles, and solitons. Journal of Mathematical Physics, 20(8):1667–1672, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1667_s1. See erratum [1945].

Epifanio:1979:RTG

[1787] G. Epifanio and C. Trapani. Remarks on a theorem by G. Epifanio: “On the matrix representation of unbounded operators”, [J. Math. Phys. 17 (1976), no. 9, 1688–1691; MR 55 #8879]. Journal of Mathematical Physics, 20(8):1673–1675, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1673_s1.

Shalitin:1979:TBN

[1788] Dan Shalitin and Yoel Tikochinsky. Transformation between the normal and antinormal expansions of boson operators. Journal of Mathematical Physics, 20(8):1676–1678, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1676_s1.

Ekhaguere:1979:MFN

[1789] G. O. S. Ekhaguere. Markov fields in noncommutative probability theory on W^{∗} algebras. Journal of Mathematical Physics, 20(8):1679–1683, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i8/p1679_s1.

Steeb:1979:NLT

[1790] W.-H. Steeb. A note on the Lorentz transformation. Journal of Mathematical Physics, 20(8):1684–1686, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i8/p1684_s1.

Devaney:1979:IPR

[1791] A. J. Devaney. The inverse problem for random sources. Journal of Mathematical Physics, 20(8):1687–1691, August 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1687_s1.

Aurilia:1979:TSMb

[1792] A. Aurilia and D. Christodoulou. Theory of strings and membranes in an external field. II. The string. Journal of Mathematical Physics, 20(8):1692–1699, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1692_s1.

Beau:1979:LGF

[1793] Daniel Beau and Salah Horchani. A Lie group framework for composite particles and mass spectrum. Journal of Mathematical Physics, 20(8):1700–1711, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1700_s1.

Kay:1979:URS

[1794] Bernard S. Kay. A uniqueness result in the Segal–Weinless approach to linear Bose fields. Journal of Mathematical Physics, 20(8):1712– 1713, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i8/p1712_s1.

Gzyl:1979:BIS

[1795] Henryk Gzyl. On Blume’s integration of Schrödinger’s equation for a quantum system subject to random pulses. Journal of Mathematical Physics, 20(8):1714–1715, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1714_s1.

Froman:1979:SQH

[1796] Nanny Fröman and Karl-Erik Thylwe. Settling the question of the highenergy behavior of phase shifts produced by repulsive, strongly singular, inverse-power potentials. Journal of Mathematical Physics, 20(8):1716– 1719, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i8/p1716_s1.

Suranyi:1979:CCS

[1797] Peter Suranyi. Classical, cross-section generating solutions of field equations. Journal of Mathematical Physics, 20(8):1720–1726, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i8/
p1720_s1.

Dollard:1979:SBP

[1798] John D. Dollard and Charles N. Friedman. Scattering of a beam of particles by a potential. Journal of Mathematical Physics, 20(8):1727– 1730, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i8/p1727_s1.

Persides:1979:DAM

[1799] S. Persides. A definition of asymptotically Minkowskian space–times. Journal of Mathematical Physics, 20(8):1731–1740, August 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1731_
s1.

Cantor:1979:NSC

[1800] Murray Cantor. A necessary and sufficient condition for York data to specify an asymptotically flat spacetime. Journal of Mathematical Physics, 20(8):1741–1744, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1741_s1.

Horndeski:1979:CSC

[1801] Gregory Walter Horndeski. Characteristic surfaces and characteristic initial data for the generalized Einstein–Maxwell field equations. Journal of Mathematical Physics, 20(8):1745–1751, August 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1745_s1. See er-

ratum [1937].

Campanino:1979:LLT

[1802] Massimo Campanino, Gabriella Del Grosso, and Brunello Tirozzi. Local limit theorem for Gibbs random fields of particles and unbounded spins. Journal of Mathematical Physics, 20(8):1752–1758, August 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1752_
s1.

Rabinovitch:1979:SRG

[1803] A. Rabinovitch and R. N. Sen. Solution of renormalization group recursion relations for a finite Ising chain. Journal of Mathematical Physics, 20(8):1759–1760, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1759_s1.

Gallinar:1979:MMH

[1804] J.-P. Gallinar. Moment-method in a hard-sphere problem on a lattice, and physical applications. I. Uniform chain. Journal of Mathematical Physics, 20(8):1761–1768, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1761_s1.

Essam:1979:PMP

[1805] J. W. Essam. Potts models, percolation, and duality. Journal of Mathematical Physics, 20(8):1769–1773, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i8/p1769_s1.

Horiguchi:1979:HTM

[1806] Tsuyoshi Horiguchi. Husimi–Temperley model under a random field. Journal of Mathematical Physics, 20(8):1774–1775, August 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1774_
s1.

Larsen:1979:SMN

[1807] Edward W. Larsen. The spectrum of the multigroup neutron transport operator for bounded spatial domains. Journal of Mathematical Physics, 20(8):1776–1782, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1776_s1.

Chaichian:1979:PEQ

[1808] M. Chaichian, M. Hayashi, N. F. Nelipa, and A. E. Pukhov. On the problem of existence of quantum field theory. Journal of Mathematical Physics, 20(8):1783–1790, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1783_s1.

Dimock:1979:SQFa

[1809] J. Dimock. Scalar quantum field in an external gauge field. Journal of Mathematical Physics, 20(8):1791–1796, August 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i8/p1791_s1.

DeMeyer:1979:CPG

[1810] H. E. De Meyer. On a common property of gravitational and chiral exponential Green functions. Journal of Mathematical Physics, 20(8):1797– 1800, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i8/p1797_s1.

Sudarshan:1979:SRC

[1811] E. C. G. Sudarshan and N. Mukunda. Structure and representation of correlation functions and the density matrix for a statistical wave field in optics. Journal of Mathematical Physics, 20(8):1801–1810, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i8/p1801_s1.

Eimerl:1979:TTP

[1812] D. Eimerl. Theory of temporal pump stochasticity in stimulated Raman scattering in dispersionless media. Journal of Mathematical Physics, 20(8):1811–1823, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i8/p1811_s1.

Willemse:1979:ECP

[1813] M. W. M. Willemse and W. J. Caspers. Electrical conductivity of polycrystalline materials. Journal of Mathematical Physics, 20(8):1824– 1831, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i8/p1824_s1.

Truman:1979:EFM

[1814] A. Truman. Erratum: “The Feynman maps and the Wiener integral”, (J. Math. Phys. 19 (1978), no. 8, 1742–1750). Journal of Mathematical Physics, 20(8):1832–1833, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i8/p1832_s1. See [1408].

Okolowski:1979:ENS

[1815] J. A. Okolowski. Erratum: “A nonlinear scalar field theory in isotropic homogeneous space–time”, [J. Math. Phys. 19, 2253 (1978)]. Journal of Mathematical Physics, 20(8):1834, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i8/p1834_s1. See [1493].

Shebalin:1979:DAS

[1816] John V. Shebalin. A double angle sum formula for Gegenbauer polynomials. Journal of Mathematical Physics, 20(9):1837, September 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i9/p1837_
s1.

Laedke:1979:AVA

[1817] E. W. Laedke and K. H. Spatschek. On the applicability of the variation of action method to some one-field solitons. Journal of Mathematical Physics, 20(9):1838–1841, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i9/p1838_s1.

Ochs:1979:WDP

[1818] Wilhelm Ochs. When does a projective system of state operators have a projective limit? Journal of Mathematical Physics, 20(9):1842–1847, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i9/p1842_s1.

Thomas:1979:BSB

[1819] Lawrence E. Thomas. Birman–Schwinger bounds for the Laplacian with point interactions. Journal of Mathematical Physics, 20(9):1848–1850, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i9/p1848_s1.

Urbantke:1979:RPF

[1820] H. Urbantke. On rotating plane-fronted waves and their Poincaréinvariant differential geometry. Journal of Mathematical Physics, 20 (9):1851–1860, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i9/p1851_s1.

Yasue:1979:ROM

[1821] Kunio Yasue. The role of the Onsager–Machlup Lagrangian in the theory of stationary diffusion process. Journal of Mathematical Physics, 20 (9):1861–1864, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i9/p1861_s1.

Davidson:1979:MSO

[1822] Mark Davidson. A model for the stochastic origins of Schrödinger’s equation. Journal of Mathematical Physics, 20(9):1865–1869, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i9/p1865_s1.

Khandekar:1979:EST

[1823] D. C. Khandekar and S. V. Lawande. Exact solution of a time-dependent quantal harmonic oscillator with damping and a perturbative force. Journal of Mathematical Physics, 20(9):1870–1877, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i9/p1870_s1.

Muller-Kirsten:1979:ESW

[1824] H. J. W. Müller-Kirsten, G. E. Hite, and S. K. Bose. Explicit solution of the wave equation for arbitrary power potentials with application to charmonium spectroscopy. Journal of Mathematical Physics, 20(9):1878– 1890, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i9/p1878_s1.

Stillinger:1979:SQM

[1825] Frank H. Stillinger. Solution of a quantum mechanical eigenvalue problem with long range potentials. Journal of Mathematical Physics, 20 (9):1891–1895, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i9/p1891_s1.

Haba:1979:SLE

[1826] Z. Haba. A scale limit of φ^{2} in φ^{4} Euclidean field theory. I. Journal of Mathematical Physics, 20(9):1896–1904, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i9/p1896_s1.

Aldersley:1979:CCD

[1827] S. J. Aldersley. Comments on certain divergence-free tensor densities in a 3-space. Journal of Mathematical Physics, 20(9):1905–1907, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i9/p1905_s1.

Banerjee:1979:PSS

[1828] A. Banerjee and D. Bhattacharya. Plane symmetric static fields in Brans–Dicke theory of gravitation. Journal of Mathematical Physics, 20(9):1908–1910, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i9/p1908_s1.

Cattell:1979:NMS

[1829] David F. Cattell and M. A. Melvin. Neutrino, Maxwell, and scalar fields in the cylindrical magnetic or plasma universe. Journal of Mathematical Physics, 20(9):1911–1924, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i9/p1911_s1.

Li:1979:EAM

[1830] Wann-Quan Li and Wei-Tou Ni. Expansions of the affinity, metric and geodesic equations in Fermi normal coordinates about a geodesic. Journal of Mathematical Physics, 20(9):1925–1929, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i9/p1925_s1.

Lim:1979:EMS

[1831] S. C. Lim. Euclidean and Minkowski space formulations of linearized gravitational potential in various gauges. Journal of Mathematical Physics, 20(9):1930–1937, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i9/p1930_s1.

Finley:1979:CAS

[1832] J. D. Finley III and J. F. Plebański. The classification of all H spaces admitting a Killing vector. Journal of Mathematical Physics, 20(9):1938– 1945, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i9/p1938_s1.

Plebanski:1979:NNT

[1833] Jerzy F. Plebański. The nondiverging and nontwisting type D electrovac solutions with λ. Journal of Mathematical Physics, 20(9):1946–1962, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i9/p1946_s1.

Gudder:1979:BDR

[1834] Stanley Gudder, Jean-Paul Marchand, and Walter Wyss. Bures distance and relative entropy. Journal of Mathematical Physics, 20(9):1963–1966, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i9/p1963_s1.

Bollini:1979:SPA

[1835] C. G. Bollini, J. J. Giambiagi, and J. Tiomno. Singular potentials and analytic regularization in classical Yang–Mills equations. Journal of Mathematical Physics, 20(9):1967–1971, September 1979. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i9/p1967_s1.

Einstein:1979:SRS

[1836] S. Einstein and R. Finkelstein. Solutions of the Rarita–Schwinger equation in the Kerr–Newman space. Journal of Mathematical Physics, 20 (9):1972–1976, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i9/p1972_s1.

Bars:1979:CLA

[1837] Itzhak Bars and Murat Günaydin. Construction of Lie algebras and Lie superalgebras from ternary algebras. Journal of Mathematical Physics, 20(9):1977–1993, September 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i9/p1977_s1.

Klimyk:1979:MEIa

[1838] Anatoli U. Klimyk and Bruno Gruber. Matrix elements for infinitesimal operators of the groups U(p + q) and U(p, q) in a U(p) × U(q) basis. I. Journal of Mathematical Physics, 20(10):1995–2010, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/
p1995_s1.

Klimyk:1979:MEIb

[1839] A. U. Klimyk and Bruno Gruber. Matrix elements for infinitesimal operators of the groups U(p + q) and U(p, q) in a U(p) × U(q) basis. II. Journal of Mathematical Physics, 20(10):2011–2013, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/p2011_s1.

Partensky:1979:SRM

[1840] A. Partensky and C. Quesne. Sum rules for the matrices of the generators of SU(3) in an SO(3) basis. Journal of Mathematical Physics, 20(10):2014–2027, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i10/p2014_s1.

vandenBroek:1979:CGC

[1841] P. M. van den Broek. Clebsch–Gordan coefficients of finite magnetic groups. Journal of Mathematical Physics, 20(10):2028–2035, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i10/p2028_s1.

Hinton:1979:SMC

[1842] F. L. Hinton. Stokes multipliers for a class of ordinary differential equations. Journal of Mathematical Physics, 20(10):2036–2046, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/
p2036_s1.

Moses:1979:GLE

[1843] H. E. Moses. Gel’fand–Levitan equations with comparison measures and comparison potentials. Journal of Mathematical Physics, 20(10):2047– 2053, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i10/p2047_s1.

Ray:1979:NTT

[1844] John R. Ray and James L. Reid. Noether’s theorem, time-dependent invariants and nonlinear equations of motion. Journal of Mathematical Physics, 20(10):2054–2057, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i10/p2054_s1.

Boisseau:1979:GTH

[1845] B. Boisseau and C. Barrabès. Gauge theory in Hamiltonian classical mechanics : The electromagnetic and gravitational fields. Journal of Mathematical Physics, 20(10):2058–2062, October 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i10/p2058_s1.

Helstrom:1979:QDA

[1846] Carl W. Helstrom. Quasiprobability distributions and the analysis of the linear quantum channel with thermal noise. Journal of Mathematical Physics, 20(10):2063–2068, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i10/p2063_s1.

Papanicolaou:1979:GDB

[1847] N. Papanicolaou. Gravitational duality and Bäcklund transformations. Journal of Mathematical Physics, 20(10):2069–2077, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/
p2069_s1.

Letelier:1979:SGF

[1848] P. S. Letelier. Self-gravitating fluids with cylindrical symmetry. II. Journal of Mathematical Physics, 20(10):2078–2080, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/p2078_s1.

Lindblad:1979:GQS

[1849] Göran Lindblad. Gaussian quantum stochastic processes on the CCR algebra. Journal of Mathematical Physics, 20(10):2081–2087, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i10/p2081_s1.

Matsumoto:1979:QSC

[1850] H. Matsumoto, G. Oberlechner, M. Umezawa, and H. Umezawa. Quantum soliton and classical soliton. Journal of Mathematical Physics, 20 (10):2088–2096, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i10/p2088_s1.

Gidas:1979:DSP

[1851] Basilis Gidas. Deformations and spectral properties of merons. Journal of Mathematical Physics, 20(10):2097–2109, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/p2097_s1.

Schaposnik:1979:GFS

[1852] F. A. Schaposnik and J. E. Solomin. Gauge field singularities and noninteger topological charge. Journal of Mathematical Physics, 20(10):2110– 2114, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i10/p2110_s1.

Jarvis:1979:CIC

[1853] P. D. Jarvis and H. S. Green. Casimir invariants and characteristic identities for generators of the general linear, special linear and orthosymplectic graded Lie algebras. Journal of Mathematical Physics, 20(10):2115– 2122, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i10/p2115_s1.

Sastry:1979:NSD

[1854] V. V. S. S. Sastry and A. Chakrabarti. A note on Sommerfeld’s diffraction problem. Journal of Mathematical Physics, 20(10):2123–2125, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i10/p2123_s1.

Arthur:1979:ADF

[1855] M. D. Arthur, R. L. Bowden, and P. F. Zweifel. Analysis of the dispersion function for anisotropic longitudinal plasma waves. Journal of Mathematical Physics, 20(10):2126–2130, October 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v20/i10/p2126_s1.

Klimas:1979:AME

[1856] Alexander J. Klimas. An approximation method for electrostatic Vlasov turbulence. Journal of Mathematical Physics, 20(10):2131–2138, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i10/p2131_s1.

Fortini:1979:FSM

[1857] André Fortini, Alain Hairie, and Jean-Pierre Girard. First-stage magnetization and metastable migration field in a type I. superconducting slab. Journal of Mathematical Physics, 20(10):2139–2149, October 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i10/
p2139_s1.

Kim:1979:NMMa

[1858] Shoon K. Kim. A new method of matrix transformation. I. Matrix diagonalizations via involutional transformations. Journal of Mathematical Physics, 20(11):2153–2158, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i11/p2153_s1. See erratum [1951].

Kim:1979:NMMb

[1859] Shoon K. Kim. A new method of matrix transformations. II. General theory of matrix diagonalizations via reduced characteristic equations and its application to angular momentum coupling. Journal of Mathematical Physics, 20(11):2159–2167, November 1979. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i11/p2159_s1.

Carinena:1979:RGS

[1860] José F. Cariñena and Mariano Santander. Representation groups for semiunitary projective representations of finite groups. Journal of Mathematical Physics, 20(11):2168–2177, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i11/p2168_s1.

Agrawala:1979:MTI

[1861] Vishnu K. Agrawala. Micu-type invariants of simple Lie algebras. Journal of Mathematical Physics, 20(11):2178–2186, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/p2178_s1.

Christensen:1979:QDH

[1862] Mark J. Christensen and John Piepenbrink. Quantum dynamics of Hamiltonians perturbed by pulses. Journal of Mathematical Physics, 20(11):2187–2190, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i11/p2187_s1.

Maurone:1979:ABD

[1863] Philip A. Maurone and Alain J. Phares. On the asymptotic behavior of the derivatives of Airy functions. Journal of Mathematical Physics, 20(11):2191, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i11/p2191_s1.

Hanckowiak:1979:SE

[1864] Jerzy Hańćkowiak. Schwinger equations. Journal of Mathematical Physics, 20(11):2192–2203, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i11/p2192_s1.

Landgren:1979:AMP

[1865] J. J. Landgren and P. A. Rejto. An application of the maximum principle to the study of essential self-adjointness of Dirac operators. I. Journal of Mathematical Physics, 20(11):2204–2211, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/p2204_s1.

Behroozi:1979:CPP

[1866] F. Behroozi and R. Gagnon. Cyclic pursuit in a plane. Journal of Mathematical Physics, 20(11):2212–2216, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v20/i11/p2212_s1.

Moore:1979:SBR

[1867] Marion E. Moore. Stability and boundedness results of stochastic differential equations relative to a generalized norm. Journal of Mathematical Physics, 20(11):2217–2220, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i11/p2217_s1.

Maharatna:1979:SAT

[1868] A. Maharatna, Ranabir Dutt, and D. Chattarji. Some applications of time-dependent canonical transformations to nonlinear nonconservative classical systems. Journal of Mathematical Physics, 20(11):2221–2225, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i11/p2221_s1.

Fronsdal:1979:IPQ

[1869] Christian Fronsdal. Invariant ^{∗}-product quantization of the onedimensional Kepler problem. Journal of Mathematical Physics, 20 (11):2226–2232, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i11/p2226_s1.

Kaushal:1979:ENB

[1870] R. S. Kaushal and H. J. W. Müller-Kirsten. Explicit normalization of bound-state wave functions and the calculation of decay widths. Journal of Mathematical Physics, 20(11):2233–2243, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/p2233_s1.

Barut:1979:QTI

[1871] A. O. Barut, C. K. E. Schneider, and Raj Wilson. Quantum theory of infinite component fields. Journal of Mathematical Physics, 20(11):2244– 2256, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i11/p2244_s1.

Cattaneo:1979:BMB

[1872] U. Cattaneo. Borel multipliers for the Bondi–Metzner–Sachs group. Journal of Mathematical Physics, 20(11):2257–2263, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/
p2257_s1.

Fang:1979:DGG

[1873] J. Fang and C. Fronsdal. Deformations of gauge groups. Gravitation. Journal of Mathematical Physics, 20(11):2264–2271, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/
p2264_s1.

Farhoosh:1979:SCM

[1874] Hamid Farhoosh and Robert L. Zimmerman. Stationary charged Cmetric. Journal of Mathematical Physics, 20(11):2272–2279, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i11/p2272_s1.

Singh:1979:GSE

[1875] T. Singh and L. N. Rai. Generalized static electromagnetic fields in Brans–Dicke theory. Journal of Mathematical Physics, 20(11):2280–2285, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i11/p2280_s1.

Pule:1979:DOI

[1876] J. V. Pulè and A. Verbeure. Dissipative operators for infinite classical systems and equilibrium. Journal of Mathematical Physics, 20(11):2286– 2290, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i11/p2286_s1.

Vaidya:1979:OPR

[1877] H. G. Vaidya and C. A. Tracy. One particle reduced density matrix of impenetrable bosons in one dimension at zero temperature. Journal of Mathematical Physics, 20(11):2291–2312, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/p2291_s1.

ElWakil:1979:TRC

[1878] S. A. El Wakil, M. H. Haggag, and E. A. Saad. Transmission and reflection coefficients in the slowing down region. Journal of Mathematical Physics, 20(11):2313–2318, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i11/p2313_s1.

Naito:1979:URH

[1879] Seichi Naito. Unitarity and renormalized ’t Hooft identities. Journal of Mathematical Physics, 20(11):2319–2326, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/p2319_s1.

Asorey:1979:EMP

[1880] M. Asorey and L. J. Boya. Electromagnetism without monopoles is possible in nontrivial U(1)-fibre bundles. Journal of Mathematical Physics, 20(11):2327–2329, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i11/p2327_s1.

Znojil:1979:GMR

[1881] M. Znojil. Generalized method of a resolvent operator expansion. II. Journal of Mathematical Physics, 20(11):2330–2333, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/
p2330_s1.

Sanz:1979:LDE

[1882] J. L. Sanz. On the Lorentz–Dirac equation. Journal of Mathematical Physics, 20(11):2334–2338, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i11/p2334_s1.

Weinert:1979:MRK

[1883] Ulrich Weinert. Matrix representation of the kinetic theory differential operator. Journal of Mathematical Physics, 20(11):2339–2346, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i11/p2339_s1.

Collins:1979:NAIa

[1884] C. B. Collins and D. A. Szafron. A new approach to inhomogeneous cosmologies: Intrinsic symmetries. I. Journal of Mathematical Physics, 20(11):2347–2353, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i11/p2347_s1.

Szafron:1979:NAI

[1885] D. A. Szafron and C. B. Collins. A new approach to inhomogeneous cosmologies: intrinsic symmetries. II. Conformally flat slices and an invariant classification. Journal of Mathematical Physics, 20(11):2354–2361, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i11/p2354_s1.

Collins:1979:NAIb

[1886] C. B. Collins and D. A. Szafron. A new approach to inhomogeneous cosmologies: intrinsic symmetries. III. Conformally flat slices and their analysis. Journal of Mathematical Physics, 20(11):2362–2370, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i11/p2362_s1.

Ernst:1979:ECD

[1887] Frederick J. Ernst. Erratum: “Coping with different languages in the null tetrad formulation of general relativity”, [J. Math. Phys. 19, 489 (1978)]. Journal of Mathematical Physics, 20(11):2371, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i11/

p2371_s1. See [1200].

Ernst:1979:EGC

[1888] Frederick J. Ernst. Erratum: “Generalized C metric”, [J. Math. Phys. 19, 1986 (1978)]. Journal of Mathematical Physics, 20(11):2371, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i11/p2371_s2. See [1447].

Jagannathan:1979:ERP

[1889] R. Jagannathan and R. Vasudevan. Erratum: “Representations of a para-Bose algebra using only a single Bose field”, [J. Math. Phys. 20 (1979), no. 3, 390–392; MR 80h:81039a]. Journal of Mathematical Physics, 20(11):2372, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i11/p2372_s1. See [1614].

Lewis:1979:ELA

[1890] H. Ralph Lewis and Keith R. Symon. Erratum: “Linearized analysis of inhomogeneous plasma equilibria: general theory”, (J. Math. Phys. 20 (1979), no. 3, 413–436). Journal of Mathematical Physics, 20(11): 2372, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i11/p2372_s2. See [1619].

Antippa:1979:TSS

[1891] Adel F. Antippa and Nguyen Ky Toan. Topological solution for systems of simultaneous linear equations. Journal of Mathematical Physics, 20 (12):2375–2379, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2375_s1.

Schneider:1979:LOG

[1892] C. K. E. Schneider and Raj Wilson. Ladder operators of group matrix elements. Journal of Mathematical Physics, 20(12):2380–2390, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2380_s1.

Wong:1979:MFW

[1893] M. K. F. Wong. On the multiplicity-free Wigner and Racah coefficients of U(n). Journal of Mathematical Physics, 20(12):2391–2397, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2391_s1.

Raynal:1979:DPG

[1894] Jacques Raynal. On the definition and properties of generalized 6-j symbols. Journal of Mathematical Physics, 20(12):2398–2415, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2398_s1.

Chudnovsky:1979:MSN

[1895] D. V. Chudnovsky. Meromorphic solutions of nonlinear partial differential equations and many-particle completely integrable systems. Journal of Mathematical Physics, 20(12):2416–2422, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2416_s1.

Couch:1979:PSS

[1896] W. E. Couch and R. J. Torrence. A particular N-soliton solution and scalar wave equations. Journal of Mathematical Physics, 20(12):2423– 2426, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2423_s1.

Garczynski:1979:RFF

[1897] W. Garczyński and J. Hańćkowiak. Restoration of a functional from its functional derivatives. Journal of Mathematical Physics, 20(12):2427– 2430, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2427_s1.

Hellsten:1979:VGA

[1898] H. Hellsten. Visual geometry and the algebraic properties of spinors. Journal of Mathematical Physics, 20(12):2431–2437, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/
p2431_s1.

Zhang:1979:DTA

[1899] Hwe Ik Zhang and Jae Il Lee. Disequilibrium theory applied to two-spin Glauber model. Journal of Mathematical Physics, 20(12):2438–2444, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2438_s1.

Littlejohn:1979:GCH

[1900] Robert G. Littlejohn. A guiding center Hamiltonian: a new approach. Journal of Mathematical Physics, 20(12):2445–2458, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/
p2445_s1.

Salas:1979:WLI

[1901] Angel Salas. World-line invariance in predictive mechanics. Journal of Mathematical Physics, 20(12):2459–2466, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2459_s1.

Fox:1979:TOO

[1902] Ronald Forrest Fox. Time ordered operator cumulants: Statistical independence and noncommutativity. Journal of Mathematical Physics, 20 (12):2467–2470, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2467_s1.

Muller-Kirsten:1979:SWE

[1903] H. J. W. Müller-Kirsten and S. K. Bose. Solution of the wave equation for the logarithmic potential with application to particle spectroscopy. Journal of Mathematical Physics, 20(12):2471–2480, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2471_s1.

Hall:1979:CES

[1904] Richard L. Hall and B. Schwesinger. The complete exact solution to the translation-invariant N-body harmonic oscillator problem. Journal of Mathematical Physics, 20(12):2481–2483, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2481_s1.

Johnson:1979:RRC

[1905] Bruce R. Johnson and Joseph O. Hirschfelder. The radial reduced Coulomb Green’s function. Journal of Mathematical Physics, 20(12): 2484–2501, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2484_s1. See erratum [1954].

Narnhofer:1979:DTE

[1906] H. Narnhofer. Does there exist a scattering theory for time automorphism groups of C^{∗}-algebras corresponding to two-body interactions. Journal of Mathematical Physics, 20(12):2502–2505, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2502_s1.

Panangaden:1979:PNF

[1907] Prakash Panangaden. Positive and negative frequency decompositions in curved spacetime. Journal of Mathematical Physics, 20(12):2506–2510, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2506_s1.

Klink:1979:CPS

[1908] William H. Klink. Crossing properties of scattering operators. Journal of Mathematical Physics, 20(12):2511–2513, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2511_s1.

Klink:1979:RMS

[1909] W. H. Klink. A representation of multiparticle scattering operators satisfying unitarity and crossing properties. Journal of Mathematical Physics, 20(12):2514–2519, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2514_s1.

vanHaeringen:1979:SSC

[1910] H. van Haeringen. Off-shell scattering by Coulomb-like potentials. Journal of Mathematical Physics, 20(12):2520–2525, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2520_s1.

Hoenselaers:1979:SSEa

[1911] C. Hoenselaers. Symmetries of the stationary Einstein–Maxwell field equations. V. Journal of Mathematical Physics, 20(12):2526–2529, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2526_s1.

Hoenselaers:1979:SSEb

[1912] C. Hoenselaers, William Kinnersley, and Basilis C. Xanthopoulos. Symmetries of the stationary Einstein–Maxwell equations. VI. Transformations which generate asymptotically flat spacetimes with arbitrary multipole moments. Journal of Mathematical Physics, 20(12):2530–2536, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2530_s1.

Pant:1979:CFS

[1913] D. N. Pant and A. Sah. Charged fluid sphere in general relativity. Journal of Mathematical Physics, 20(12):2537–2539, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2537_s1.

Gerlach:1979:EPJ

[1914] U. H. Gerlach and U. K. Sengupta. Even parity junction conditions for perturbations on most general spherically symmetric space–times. Journal of Mathematical Physics, 20(12):2540–2546, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2540_s1.

Fischer:1979:EBT

[1915] Elliot Fischer. An extension of the Bonnor transformation. Journal of Mathematical Physics, 20(12):2547–2548, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2547_s1.

Dimock:1979:SQFb

[1916] J. Dimock. Scalar quantum field in an external gravitational field. Journal of Mathematical Physics, 20(12):2549–2555, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2549_s1.

Heidenreich:1979:CEE

[1917] W. Heidenreich. Conformal electrodynamics in Einstein spacetime. Journal of Mathematical Physics, 20(12):2556–2558, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2556_s1.

Zasada:1979:LTT

[1918] C. S. Zasada and R. K. Pathria. Low-temperature thermodynamics of a weakly interacting Bose gas confined to restricted geometries. Journal of Mathematical Physics, 20(12):2559–2566, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2559_s1.

Ferrario:1979:NMR

[1919] Mauro Ferrario and Paolo Grigolini. The non-Markovian relaxation process as a “contraction” of a multidimensional one of Markovian type. Journal of Mathematical Physics, 20(12):2567–2572, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/
p2567_s1.

vanVliet:1979:LRT

[1920] K. M. van Vliet. Linear response theory revisited. II. The master equation approach. Journal of Mathematical Physics, 20(12):2573–2595, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2573_s1.

Garber:1979:ASB

[1921] W. D. Garber. All spontaneously broken symmetries for noncovariant currents. Journal of Mathematical Physics, 20(12):2596–2604, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2596_s1.

Cho:1979:CGF

[1922] Y. M. Cho. Classification of gauge fields. Journal of Mathematical Physics, 20(12):2605–2612, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i12/p2605_s1. See erratum [1944].

Cognola:1979:CNA

[1923] G. Cognola, R. Soldati, L. Vanzo, and S. Zerbini. Classical non-Abelian gauge theories in the space of reference frames. Journal of Mathematical Physics, 20(12):2613–2618, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i12/p2613_s1.

Reina:1979:FOS

[1924] C. Reina, M. Martellini, and P. Sodano. A first order system of differential equations for covariant σ models. Journal of Mathematical Physics, 20(12):2619–2620, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2619_s1.

Daumens:1979:ILB

[1925] M. Daumens and M. Perroud. Internal Lorentz basis for two-particle states. Journal of Mathematical Physics, 20(12):2621–2627, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2621_s1.

Pohlmeyer:1979:RTD

[1926] K. Pohlmeyer and K.-H. Rehren. Reduction of the two-dimensional O(n) nonlinear σ-model. Journal of Mathematical Physics, 20(12):2628–2632, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2628_s1.

Kanal:1979:APR

[1927] Madhoo Kanal. Albedo problem of radiative transfer in inhomogeneous isotropically scattering atmospheres: Application of the maximum variational principle. Journal of Mathematical Physics, 20(12):2633–2638, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2633_s1.

Lucena:1979:COE

[1928] S. J. Lucena, T. A. Minelli, and M. Pusterla. Classical oscillators excited parametrically and their applications in plasma physics. Journal of Mathematical Physics, 20(12):2639–2648, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2639_s1.

Dragt:1979:NFM

[1929] Alex J. Dragt and John M. Finn. Normal form for mirror machine Hamiltonians. Journal of Mathematical Physics, 20(12):2649–2660, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2649_s1.

Runnels:1979:ZGP

[1930] L. K. Runnels and John Runnels. Zeros of the grand partition function: One-dimensional Widom–Rowlinson lattice gas. Journal of Mathematical Physics, 20(12):2661–2664, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i12/p2661_s1.

Vanheuverzwijn:1979:MSI

[1931] Paul Vanheuverzwijn. Metastable states in the infinite Ising model. Journal of Mathematical Physics, 20(12):2665–2670, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v20/i12/p2665_s1.

Eringen:1979:CTN

[1932] A. Cemal Eringen. Continuum theory of nematic liquid crystals subject to electromagnetic fields. Journal of Mathematical Physics, 20(12):2671– 2681, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i12/p2671_s1.

McCarthy:1979:GWD

[1933] M. F. McCarthy and H. F. Tiersten. The growth of wave discontinuities in piezoelectric semiconductors. Journal of Mathematical Physics, 20 (12):2682–2691, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2682_s1.

Bonilla:1979:SPA

[1934] L. L. Bonilla and M. G. Velarde. Singular perturbations approach to the limit cycle and global patterns in a nonlinear diffusion-reaction problem with autocatalysis and saturation law. Journal of Mathematical Physics, 20(12):2692–2703, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2692_s1.

Aldersley:1979:EHE

[1935] S. J. Aldersley. Erratum: “Higher Euler operators and some of their applications [J. Math. Phys. 20, 522 (1979)]”. Journal of Mathematical Physics, 20(12):2704, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i12/p2704_s1. See [1632].

Romo:1979:EAF

[1936] W. J. Romo. Erratum: “Asymptotic forms of radial wavefunctions and Jost functions for cutoff potentials”, [J. Math. Phys. 20, 1210(1979)]. Journal of Mathematical Physics, 20(12):2704, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v20/i12/p2704_s2. See [1714].

Horndeski:1980:ECS

[1937] Gregory Walter Horndeski. Erratum: “Characteristic surfaces and characteristic initial data for the generalized Einstein–Maxwell field equations”. Journal of Mathematical Physics, 21(1):218, January 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1801].

Wald:1980:ENS

[1938] Robert M. Wald. Erratum: “Note on the stability of the Schwarzschild metric”, (J. Math. Phys. 20 (1979), no. 6, 1056–1058). Journal of Mathematical Physics, 21(1):218, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1695].

Wodkiewicz:1980:EES

[1939] K. Wódkiewicz. Erratum: “Exact solutions of some multiplicative stochastic processes”, (J. Math. Phys. 20 (1979), no. 1, 45–48). Journal of Mathematical Physics, 21(1):218, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1566].

Lewis:1980:ELA

[1940] H. Ralph Lewis and Keith R. Symon. Erratum: “Linearized analysis of inhomogeneous plasma equilibria: general theory”, [J. Math. Phys. 20 (1979), no. 3, 413–436; MR 80b:82045a]. Journal of Mathematical Physics, 21(3):599, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i3/p599_s1. See [1619].

Gouedard:1980:EDE

[1941] C. Gouedard and C. Deutsch. Erratum: “Dense electron-gas response at any degeneracy”, [J. Math. Phys. 19, 32 (1978)]. Journal of Mathematical Physics, 21(4):949, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p949_s1. See erratum [1133].

Abellanas:1980:ECD

[1942] L. Abellanas and A. Galindo. Erratum: “Conserved densities for nonlinear evolution equations. I. Even order case”, [J. Math. Phys. 20, 1239 (1979)]. Journal of Mathematical Physics, 21(5):1267, May 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1267_

s2. See [1718].

Gazeau:1980:EFE

[1943] J. P. Gazeau. Erratum:“Four Euclidean conformal groups in atomic calculations: Exact analytical expressions for the bound-bound two photon transition matrix elements in the H atom”, [J. Math. Phys. 19, 1041 (1978)]. Journal of Mathematical Physics, 21(5):1267, May 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1267_

s1. See [1298].

Cho:1980:ECG

[1944] Yongmin Cho. Erratum: “Classification of gauge fields”, [J. Math. Phys. 20 (1979), no. 12, 2605–2612; MR 80k:81086]. Journal of Mathematical Physics, 21(6):1552, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i6/p1552_s3. See [1922].

Lakshmanan:1980:ERB

[1945] M. Lakshmanan. Erratum: “Rigid body motions, space curves, prolongation structures, fiber bundles, and solitons”, [J. Math. Phys. 20 (1979), no. 8, 1667–1672; MR 80g:58003]. Journal of Mathematical Physics, 21 (6):1552, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1786].

Mizrahi:1980:EPS

[1946] Maurice M. Mizrahi. Erratum: “Phase space path integrals, without limiting procedure”, [J. Math. Phys. 19 (1978), no. 1, 298–307; MR 56 #17627]. Journal of Mathematical Physics, 21(7):1965, July 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1170].

Murray:1980:ESF

[1947] J. C. Murray. Erratum: “On the sound field due to a moving source in a superfluid”, [J. Math. Phys. 20, 1409 (1979)]. Journal of Mathematical Physics, 21(8):2313, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2313_s1. See [1746].

Ben-Artzi:1980:ESS

[1948] Matania Ben-Artzi and Allen Devinatz. Erratum: “Spectral and scattering theory for the adiabatic oscillator and related potentials”. Journal of Mathematical Physics, 21(9):2471, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1640].

CastilhoAlcaras:1980:ECS

[1949] J. A. Castilho Alcarás and L. C. Beidenharn. Erratum: “On the coupling of self-conjugate systems with \overline{SL(3, R)} symmetry”, [J. Math. Phys. 20 (1979), no. 8, 1615–1621; MR 80i:81069]. Journal of Mathematical Physics, 21(12):2852, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1779].

Cornille:1981:ESN

[1950] H. Cornille. Erratum: “Solutions of the nonlinear 3-wave equations in three spatial dimensions”, [J. Math. Phys. 20 (1979), no. 8, 1653–1666; MR 81d:35069]. Journal of Mathematical Physics, 22(1):219, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1785].

Kim:1981:ENM

[1951] Shoon K. Kim. Erratum: “A new method of matrix transformations. II. General theory of matrix diagonalizations via reduced characteristic equations and its application to angular momentum coupling”, [J. Math. Phys. 20 (1979), no. 11, 2159–2167; MR 81j:15015b]. Journal of Mathematical Physics, 22(4):916, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1858].

Froman:1981:CHO

[1952] Nanny Fröman and Per Olof Fröman. Comments on: “Higher order modified potentials for the effective phase integral approximation”, [J. Math. Phys. 20 (1979), no. 1, 83–85; MR 80f:81031 ] by Edward R. Floyd. Journal of Mathematical Physics, 22(6):1190–1191, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1572].

Bleistein:1982:ENI

[1953] Norman Bleistein and Jack K. Cohen. Erratum: “Nonuniqueness in the inverse source problem in acoustics and electromagnetics”, [J. Math. Phys. 18 (1977), no. 2, 194–201; MR 55 #11812]. Journal of Mathematical Physics, 23(5):915, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p915_s2. See [778].

Johnson:2021:ERR

[1954] Bruce R. Johnson. Erratum: “The radial reduced Coulomb Green’s function” [j. math. phys. 20, 2484 (1979)]. Journal of Mathematical Physics, 62(2):029901, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1905].